A Bibliography of Papers in Lecture Notes in Computer Science (2002) (Part 4 of 4)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

02 May 2020 Version 1.11

Title word cross-reference

(2^{m}) [378]. 2 [1470, 574, 631, 1132, 1467, 1502, 1164, 643, 1158]. 2^{m} [377]. 3 [574, 9, 537, 1519, 659, 612, 1280, 1356, 973, 462]. ^{2} [165, 1009]. ^{TM} [487, 1514]. _{A} [329]. _{d} [1288]. flow [703]. Δ [681]. E [693]. H [1468]. K [990, 1089, 1090, 996, 1093, 632, 1474, 953, 1501]. MAX − MIN [1161]. U [1333]. X [1333]. μ [1338, 1317]. ω [1466, 1508]. Σ_{3} [1336].

-Calculus [1317, 1338]. -Chinese [1090]. -Closest-Pairs [996]. -Colorability [1470]. -Colourings [1468]. -Connected [632, 1158]. -D [537]. -Dominating [1474]. -Dynamic [703]. -Letter [1502]. -Line [1089]. -Medoids-Based [990]. -Ontology [612]. -Query [1467]. -Regular [1466]. -Round [631]. -SAT [1280]. -service [693]. -Splittable [1093]. -testable [1501]. -Trees [681]. -Valued [1356].

1-Connected [632]. 1-Type [1062]. 1.375-Approximation [1101].

2 [575, 265]. 2SAT [1277].

3 [1288]. 3GPP [359].

802.11a [575].

A* [812]. A.I [664]. A/D [848]. Abductive [680]. Abelian [637, 641]. Abstract [934, 963, 186, 1383, 1327, 1387, 625, 1386, 1362, 1339, 644, 1141, 1385, 1331, 1366, 1317, 1384, 638, 1156, 1381]. Abstracted [245]. AC [1288]. AC-3 [1288]. Academic [1312]. Accelerate [1303]. Accelerating [1245]. Accelerators [598]. Access [1112, 918, 364, 926, 310, 1532, 411, 143, 1432, 1038, 363, 779, 980, 318, 402, 1441]. Accesses [1034]. Accessible [602, 422]. Accessing [824, 273]. Accommodating [423]. Account [1347]. Accumulator [384]. Accumulators [624]. Accuracy [995]. Accurate [1147]. Achieving [774]. ACO [1195, 1175, 1162, 1191, 1170, 1173, 1193]. Acquire [73]. Acquiring [737]. Acquisition [20, 1542, 1426]. across [895]. ACS [1188]. ACS-TSP [1188]. Action [668, 945, 59, 453, 665, 82, 58]. Actions [122, 57]. Active [630, 312, 83, 1450, 1539]. Activities [1524, 1049, 127, 102, 881]. Activity [874, 1524, 499]. Activity-Centered [1524]. Actors [1412]. Acts [920]. Actual [610]. Acyclic [67]. Ad [649, 992, 701]. Ad-Hoc [701, 649]. Adapt [73, 31]. Adaptability [318]. Adaptable [931]. Adaptation [3, 348, 345, 8, 187, 40, 23, 26, 300, 232]. Adapting [303, 685, 301]. Adaptive [36, 668, 1400, 1265, 537, 621, 472, 67, 906, 444, 241, 953, 666, 617, 464, 930, 106, 520, 988, 177]. Adders [852, 842]. Adding [1006, 588]. Additional [1191, 286]. Additive [392]. Additively [1132]. Additives [1325]. Address [876]. Address-Event-Representation [876]. Addressed [948]. ADEES [931]. Adiabatic [852]. adjoint [680]. Adjustable [101]. Administration [157]. Administrative [280]. Admon [157]. Admon-Time [157]. Advanced [702, 919, 141]. Advances [808]. Adversaries [636]. Advice [1480]. Advisory [906, 45]. Aerial [88]. AES [549, 620]. Affective [110]. Affine [1482]. against [699, 630, 751, 362, 636, 386, 385].

Age [464]. Agent

[1041, 110, 748, 677, 1394, 1028, 676, 733, 1031, 741, 1424, 1435, 731, 750, 1163, 746, 113, 115, 1434, 65, 87, 1205, 70, 64, 101, 86, 905, 678, 734, 747, 85, 1453, 1029]. Agent-Based [1028, 1031, 741, 731, 1163, 113, 115, 87]. Agents [1019, 726, 745, 751, 936, 1201, 1049, 736, 111, 403, 1416, 727, 18, 752, 405, 739, 1313, 1030, 1171, 1436, 929, 112, 1454]. Aggregating [32]. Agile [1437]. Agreement [1256]. Ahead [852]. AI [686]. Aimed [1178]. Aircraft [1303]. AlfaNum [803]. Algebra [1070, 194, 335, 662, 208, 417]. Algebraic [1070, 1358, 186, 205, 59, 1287, 555, 620]. Algebraically [1324]. Algebras [1406]. Algorithm [856, 811, 1101, 677, 491, 710, 857, 990, 758, 1178, 1119, 1124, 162, 1191, 541, 1129, 1433, 1131, 1043, 1004, 105, 1238, 79, 1002, 1189, 1140, 1141, 752, 1293, 1146, 812, 307, 454, 1249, 1286, 1164, 994, 66, 1158, 989, 1288, 597]. Algorithmic [1063, 1482]. Algorithms [1089, 1087, 538, 1479, 642, 1166, 1097, 1021, 1103, 1194, 238, 590, 832, 1108, 1109, 408, 574, 1274, 1407, 1137, 96, 909, 1297, 1188, 393, 818]. Alias [1361]. Alignment [967, 1515]. Alignment-Based [1515]. Allocation [1163, 1190]. Allocations [1240]. Almost [774]. Alphabetic [1094]. Alternating [1413]. Alternating-Time [1413]. Alternative [800, 899, 1216]. Alternatives [737, 875]. Always [1481]. Ambiguities [97, 124]. among [1268, 1454]. Amount [737]. Amplification [1239]. Analog [508, 888]. Analogical [1220]. Analogies [1199]. Analogy [3]. Analyse [414]. Analyser [782]. Analyses [1227, 1374]. Analysing [1510, 1378, 785]. Analysis [1373, 781, 247, 272, 242, 269, 1095, 807, 640, 234, 238, 1388, 629, 883, 303, 857, 1180, 262, 914, 1358, 1376, 361, 207, 837, 716, 1072, 1131, 814, 816, 71, 1377, 1389, 1423, 271, 1371, 1367, 387, 470, 301, 1026, 633, 1440, 992, 1369, 526, 1356, 214, 976, 942, 270, 1073, 547, 1361, 220, 433, 166, 1370]. Analytical [477]. Analyzes [1385]. Analyzing [206]. Ancient [220]. Annotation [159, 1212]. Anonymity [368]. Anonymous [624, 374]. Answer [128]. Answering [367, 1211, 1212]. Ant [1166, 1176, 1194, 1174, 1180, 1178, 1165, 1168, 1183, 1189, 1179, 1192, 1185, 1182, 1187, 1196, 1164, 1167, 1161, 1177, 1186, 1171]. Ant-Based [1179]. Ant-Like [1171]. Anti [1169]. Anti-pheromone [1169]. Ants [1181, 1184, 1172]. App [686]. Appendix [339, 338]. Application [646, 835, 36, 609, 551, 624, 475, 1117, 908, 40, 70, 762, 432, 917, 1416, 816, 440, 481, 802, 694, 106, 1196, 1401, 675, 734, 225]. Application-Specific [551]. Applications [1462, 349, 1112, 376, 649, 164, 1107, 618, 1119, 581, 1079, 952, 588, 282, 202, 1315, 472, 1134, 474, 1482, 1027, 777, 363, 645, 694, 49, 1501, 52, 283, 560, 1053, 872, 692, 1413, 894, 1029, 1236, 1075]. Applied [1162, 826, 827, 1241]. Apply [291]. Applying [1170, 459, 626, 2, 1427]. Appraisal [1400]. Appraisals [110]. Approach [92, 863, 944, 35, 133, 1028, 670, 490, 171, 969, 1358, 833, 1063, 163, 1404, 1127, 398, 280, 1242, 508, 40, 967, 317, 991, 940, 1020, 679, 156, 559, 86, 1273, 97, 678, 1013, 210, 680, 731, 301, 598, 663, 1163, 416, 24, 1302, 951, 102, 316, 1268, 1395, 746, 788, 21, 55, 724, 56, 606, 1029, 32, 400]. Approaches [248, 809, 943, 995, 829]. Approaching [1420]. Approximate [996, 408, 1530, 1378, 1399]. Approximating [1116, 1362]. Approximation [1089, 1087, 1101, 398, 1130, 1137, 1140, 1157]. Approximations [76]. Arabic [773]. Arbitrage [1024]. Arbitrary [637, 390]. Arbitrary-Sized [390]. Arc [1255, 61, 1286, 1288]. Arc-Consistency [1255, 61, 1286, 1288]. Architectural [1070, 324, 532]. Architecture [110, 609, 677, 552, 676, 471, 571, 186, 733, 729, 1444, 514, 377, 568, 444, 1448, 1025, 422, 671, 594, 323, 843, 106, 529, 405, 734, 509, 692, 561, 311, 507]. Architectures [885, 315, 1074, 478, 1430, 1435, 510, 359, 509, 547, 565]. Archives [926]. Arcs [1099, 1156]. Arctic [821]. Area [1077]. Areas [679]. Argumentation [56]. Argumentation-Theoretic [56]. Arguments [189]. Arithmetic [841]. Arithmetics [528]. Armies [1291]. Arrangements [1156]. Array [710, 1141, 957]. Arrays [508, 1361]. Art [53]. Article [674]. Artificial [743, 1513, 1417, 52]. Asian [1147]. ASP [396]. Aspect [297]. Aspects [327, 1534]. Assembly [1543, 1166]. Assessing [344, 37]. Assessment [1213, 901, 19, 1187]. assigned [1102]. Assignment [793, 1122, 1312, 1144, 794]. Assistance [734]. Assistant [37]. Assisted [184, 329, 566]. Associated [422]. Association [1486, 984, 986, 1198, 763, 762, 903, 1004, 761, 75, 457, 146]. Associative [548, 313]. Assurance [440, 316]. Asymptotic [1130]. Asynchronous [848, 851, 469, 605, 1440, 1289, 878, 196, 433]. ATM [1273]. Atomic [174]. Attack [388, 938, 385]. Attackers [621]. Attacks [630, 751, 362, 386, 385, 1378]. Attribute [28, 1504]. Attributes [1199, 225]. Auction [369, 1018]. Auctions [1021, 1114]. Audio [712, 581, 108, 821, 285, 107, 521]. Audio-to-Visual [108]. Audio-Visually [107]. Augmenting [1403]. Authenticated [644, 366]. Authentication [650]. Authenticity [1380]. Author [48, 137, 1355, 1391, 1418, 1463, 1485, 1516, 1546, 180, 203, 274, 330, 394, 427, 463, 599, 619, 687, 706, 725, 753, 771, 838, 889, 983, 1016, 1059, 1080, 1159, 1160, 1230, 1314, 671]. Authoring [459]. Authorization [993, 1432]. Authorizations [918]. Auto [873, 120]. Auto-Optimizing [873]. Automata [1492, 1495, 586, 1316, 182]. Automated [272, 1071, 38, 888, 903, 387, 261, 1363]. Automatic [811, 898, 267, 1217, 809, 1308, 187, 714, 257, 96, 256, 488, 470, 127, 685, 663, 224, 805, 229, 794, 893, 799, 793]. Automatically [16]. Automating [518]. Automation [39]. Automatizability [1352]. Automobile [1136]. Autonomous [676, 1302, 1229, 1313]. Autonomy [101]. Average [71, 72]. Aware [1517, 857, 1525, 1544]. Awareness [611, 1460, 816, 1535, 612, 1545]. Axiom [1408]. Axiomatic [398]. Axiomatization [1406]. Axioms [197]. Axis [1116].

B [339]. B2B [1022]. B2B/B2C [1022]. B2C [1022]. Back [78]. Background [765, 1]. Backhauls [1172]. BackPOLE [78]. Backpropagation [79]. Backtrack [1297]. Backward [1364, 214]. Badge [1539]. Balanced [1109]. Balanced-Replication [1109]. Balancing [1166, 908]. Ball [1410]. Bandwidth [1077]. Bank [222]. Bankcheck [223]. Base [74, 13, 1004, 16, 1001, 18, 26, 27, 651]. Based [34, 697, 435, 856, 538, 1452, 250, 578, 35, 36, 642, 801, 4, 502, 517, 993, 640, 609, 800, 234, 133, 1028, 1400, 497, 629, 592, 676, 164, 380, 1108, 447, 857, 990, 898, 922, 1379, 408, 37, 1200, 1458, 1266, 618, 892, 169, 1279, 9, 655, 162, 581, 460, 595, 436, 1022, 42, 1224, 1220, 10, 858, 1519, 736, 1170, 187, 310, 465, 696, 469, 1460, 459, 168, 1072, 116, 14, 762, 1530, 860, 1450, 940, 1221, 431, 432, 15, 492, 450, 257, 1031, 1042, 1532, 1018, 741, 298, 156, 448, 440, 143, 86, 172, 444, 135, 995, 404, 1050, 1448, 1423]. Based [437, 666, 1038, 434, 139, 1528, 1179, 1385, 795, 77, 907, 127, 63, 1278, 1435, 1225, 731, 1231, 1366, 740, 1527, 363, 791, 1046, 812, 694, 1163, 775, 453, 385, 307, 930, 134, 454, 972, 43, 22, 919, 1249, 401, 24, 25, 1421, 441, 526, 1172, 213, 102, 45, 861, 1374, 405, 46, 579, 1290, 1401, 1542, 28, 1395, 746, 113, 374, 142, 962, 1210, 458, 1512, 29, 723, 942, 21, 233, 103, 452, 115, 523, 78, 120, 965, 931, 318, 665, 433, 47, 439, 1434, 166, 1051, 1370, 1264, 31, 760, 1441, 104, 1036, 446, 153, 828, 704, 462, 99]. Based [1393, 806, 87, 455, 33, 170, 1455, 1456, 148, 989, 366, 443, 1301, 1515, 475, 589, 583, 527, 525, 468, 587, 513, 484, 577]. Bases [378]. Basic [862]. Basis [1099, 1350]. Batch [372, 1292]. Bayes [91, 93, 233, 1003]. Bayesian [90, 1397, 1219, 12, 89, 1223, 724]. BCH [512]. BDI [745]. Beam [77]. Beam-Wise [77]. Before [610]. Behavior [303, 738, 323, 112, 1187, 433]. Behavioral [1409, 1386]. Behaviors [107, 103]. Behaviour [736]. Behaviour-Based [736]. Behaviours [1019, 193]. Benchmarking [1069, 1068]. Benders [1270]. Benes [390]. Best [231, 1180, 837]. Best-Worst [1180]. Better [1195, 1169]. between [1309, 984, 736, 13, 1348, 1318, 1045, 222, 1036]. Beyond [1255, 1489, 1532]. Bi [1508]. Bi- [1508]. Bibliographic [261]. Bibliography [340]. BiCMOS [507]. Bid [369]. Bidimensional [1193]. Bigram [798]. Bijections [1318]. Bilingual [218, 217]. Binary [1252, 974]. Binary-Tree [974]. Binding [656, 1431]. Bioinformatics [101]. Biomedical [1227]. Biometrics [129]. Bipartite [1158]. Biped [84]. Birthday [638]. Bisection [1473]. Bisimilarity [197]. Bisimulation [198]. Bisimulations [200]. Bit [808, 487, 378, 647]. Bit-Serial [487]. Bitmap [976]. BizQuery [421]. Black [637]. Black-Box [640, 637]. Blended [422]. Blended-Query [422]. Blind [836, 1523]. Block [640, 477, 622]. Block-Cipher-Based [640]. Blocks [479, 888]. Blockwise [621]. Blockwise-Adaptive [621]. Bodies [109]. Book [428]. Boolean [1099, 1342, 654, 1201, 501, 980, 1244].

Boosting [1003]. Both [73]. Bottom [30]. Bottom-Up [30]. Bound [1103, 1124]. Boundaries [710]. Bounded [1104, 1251, 636, 1500, 1349]. Bounding [980]. Bounds [1090, 1063, 1332, 1467]. Box [640, 637]. Boxes [194]. BRAIN} [214]. Branch [1103, 1124]. Branch-And-Bound [1124, 1103]. Breadth [1143]. Breadth-First [1143]. Breaking [1261, 1258, 1259, 1307, 1291, 1260]. Breaks [385]. Bridging [1309]. Bringing [939]. Broadband [1537]. Broadcast [623, 404, 591, 406]. Broadcasting [581]. Broker [678]. Browse [731]. Browser [836]. Browsing [103, 159, 433]. Bucketing [163]. Buffer [868]. Buffers [1151, 861, 206]. Build [829]. Building [580, 616, 1025, 992, 27, 1177, 994, 47, 2]. Built [311]. Built-in [311]. Bulgarian [674]. Burrell [1034]. Bus [873]. Business [699, 691, 1049, 935, 700, 304, 695, 951, 1053, 1431, 703, 116, 688, 276, 277, 1058, 1054]. Business-to-Business [691]. Butterflies [1110].

C [552, 165, 872, 206]. Cache [172, 307, 144]. Cache-Coherence [307]. Caching [1125, 891]. CAI [460]. Cake [1157]. Calculations [1299]. Calculi [196]. Calculus [1338, 200, 191, 1317, 1454]. Call [1348, 1330]. Call-by-Name [1348]. Call-By-Value [1330, 1348]. Camera [1271]. Can [73]. Candidate [1182]. Can’t [759]. Capability [1195, 310]. Capability-Based [310]. Capacitances [850]. Capacitated [1153]. Capacity [1066, 589]. Capture [305]. Capturing [1520, 900]. Card [390]. Cardinal [1253]. Cardinality [1153]. Cards [1056, 201]. Care [1005]. CARROT [742]. Carry [852]. Case [34, 35, 36, 4, 1270, 1162, 5, 1517, 37, 9, 74, 908, 1220, 10, 11, 1017, 12, 13, 14, 15, 450, 16, 910, 18, 20, 627, 834, 43, 22, 44, 24, 25, 102, 45, 768, 26, 27, 1210, 21, 547, 47, 31, 33]. Case-Based [34, 36, 4, 37, 9, 1220, 10, 14, 15, 43, 22, 25, 102, 45, 1210, 21, 47, 31, 33]. Cases [317, 19, 29]. cash [368]. CASL [181]. Cast [41]. Catalogs [731, 24]. Catalyzes [1461]. Categorial [1512]. Categorical [1086, 758, 210]. Categories [1216, 789]. Categorization [1184, 96, 780]. Category [29]. Category-Based [29]. Causal [410, 11, 56]. Causality [193]. Cause [95]. Cayley [1471]. CBC [621]. CBCM [164]. CBR [38, 39, 41, 46]. CBR-Based [46]. Cell [90, 852, 164]. Cell-Based [164]. Cellular [586]. Center [1089]. Centered [1524]. Centralized [748]. Centric [466, 1078]. Cepstral [1225]. Certification [380]. Certified [371, 388]. Ceteris [1304]. Chains [1469, 335, 334, 336, 333, 115]. Challenges [504, 1357, 468, 727, 728]. Change [76, 1523, 134]. Change-Point [76]. Change1 [300]. Changes [145, 423, 73]. Changing [1531]. Channel [386, 1144, 385]. Channels [1112]. Chapter [339]. Chapters [338]. Character [221, 834, 236, 245, 228, 230]. Characteristic [643]. Characteristics [230, 845]. Characterization [887, 1478]. Characterizing [1282]. Characters [250, 109, 220]. Charge [861]. Charge-Based [861]. Charging [850]. Chart [272]. Chat [834]. Check [199, 481]. Checkers [118]. Checking [699, 1071, 1335, 191, 1383, 1387, 1386, 1364, 1384, 1253]. Checks [655]. Cheque [222]. Chinese [786, 1090]. Chip [575, 595]. Choosing [737, 231]. Chosen [648, 362]. Chosen-Ciphertext [362]. Chosen-Plaintext [648]. Cipher [640, 359]. Ciphers [652, 622, 653]. Ciphertext [362]. Circuit [572, 888, 886]. Circuits [863, 874, 851, 469, 505, 879, 881]. Cities [112]. Claim [1003]. Class [190, 654, 1404, 314, 213, 296]. Class-Based [213]. Classes [1491, 1060, 454, 1496].

Classical [3, 1095, 1345, 786]. Classification [822, 932, 267, 971, 714, 67, 1223, 44, 259, 966, 240, 457, 131, 1, 400]. Classifier [80, 98, 1225, 236, 1403, 232, 234, 235]. Classifiers [234, 93, 228, 230]. Classifying [1484, 1200]. Classroom [603, 438]. Clausal [1407, 1353]. Clauses [1308, 63, 1360]. Cleansing [937]. Client [1041, 157, 172, 323]. Client-Based [172]. Clinical [912]. Clique [1124, 1402]. Clock [859, 884]. Clock-Gating [884]. Clocked [853]. Clocking [497, 853]. Closed [1372, 1001]. Closest [996]. Closing [231]. Closure [559]. Cloud [1082]. Cluster [412, 908, 584, 307]. Cluster-Based [307]. Clustered [981, 865]. Clustering [92, 1214, 1028, 164, 1108, 163, 791, 94, 69, 978, 988, 989]. CMOS [874, 883, 857, 888, 860, 864, 846, 861, 879, 881]. CNF [1243]. Co [536, 84, 467, 944]. Co-evolving [84]. Co-operation [467]. Co-processor [536]. Coarse [568]. Coarse-Grain [568]. Code [181, 470, 296]. Codes [1112, 391, 1467]. Coding [808]. Coefficient [582, 1225]. Coercion [1376]. Coevolutionary [1194]. Coexisting [1291]. CoFD} [989]. Cognitive [1409, 733, 278]. Coherence [1219, 307, 144]. COLIBRI [7]. Collaboration [897, 320, 610, 1534]. Collaborative [34, 611, 1462, 609, 442, 1214, 613, 602, 1043, 257, 1042, 444, 938, 1207, 441, 1210, 21, 739]. Collecting [50]. Collections [821]. Colonies [1167]. Colony [1178, 1165, 1183, 1189, 1192, 1182, 1196, 1164, 1186]. Color [83, 270]. Colorability [1470]. Colored [1086]. Colour [250]. Colouring [675]. Colourings [1468]. Column [1261, 1303]. Combat [88]. Combination [1221, 236, 796]. Combinatorial [1476]. Combined [230]. Combiner [653]. Combining [1220, 749, 766, 723, 978, 225]. ComicDiary [1518]. Comics [1518].

Comments [373]. Commerce

[350, 281, 1443, 1057, 28, 1456, 1028, 1022, 1056, 363, 1053, 734, 894, 1029]. Commercial [1416, 1027, 49]. Commit [174]. Commitment [656]. Commitments [655, 700]. Committees [94]. committing [627]. Common [1398, 772, 679, 161, 413]. Communication [346, 1274, 1126, 432, 126, 325, 1461, 698, 1382, 746, 747]. Communications [486, 876, 53, 608, 513]. Communities [450]. Community [442, 732]. Commutativity [193]. Compact [861]. Compactly [1371]. Compactness [1321]. Companion [1125]. comparand [548]. Comparative [361, 96, 798, 871]. Compare [1216]. Comparing [1032, 1227]. Comparison [984, 899, 842, 809, 42, 714, 837, 910, 1150, 213, 1209, 796]. Comparison-Based [42]. Comparisons [1148]. Compatible [1142]. Competitive [83, 1114, 738]. Compilation [571]. Compiler [552, 1392, 503, 663, 553]. Compilers [1375]. Compiling [1394, 551]. Complementarity [234]. Complementarity-Based [234]. Complete [221, 1324]. Completion [229, 1155]. Complex [247, 253, 969, 265, 426, 888, 488, 295, 259, 756]. Complexity [3, 314, 1339, 1332, 1142, 1250, 627, 115, 270, 1288]. Component [609, 1450, 444, 322]. Component-Based [609, 444]. Components [1452, 1525, 187, 841, 951, 292, 321, 308]. Composable [656]. Composed [1173]. Composite [694]. Composition [703]. Compound [1198]. Compression [870, 775]. Computable [1315, 1536]. Computation [646, 632, 1274, 631, 586]. Computational [707, 1099, 1339, 114]. Computationally [382]. Computations [1368, 1441]. Computer [603, 574, 126, 591, 53, 104, 462]. Computer-Integrated [603]. Computers [481]. Computing [1398, 1517, 1524, 1525, 1082, 1118, 1472, 465, 1523, 584, 1530, 1536, 472, 559, 546, 1535, 174, 464, 1240, 1078, 558, 473, 532, 643, 1531, 534]. Concatenated [391]. Conceived [547]. Concept [147, 1178, 945, 992, 942]. Concepts [707, 133, 257]. Conceptual [1198, 434, 161]. Concise [761]. Conclusions [337]. Concurrency [404]. Concurrent [630, 1358, 193, 1302, 1382, 1363]. Condition [1336]. Conditional [814, 379]. Conditions [1390, 1496]. Conducting [606]. Cone [804]. Cone-Shaped [804]. Conference [1458]. Confidence [796]. Configurable [782, 510, 565, 544, 547]. Configuration [1328, 235, 282, 265, 667, 1046, 524]. Configurations [1488]. Confinement [1378]. Conflict [1112, 1279, 173]. Conflict-Based [1279]. Conflicts [1244]. Conformational [1120]. Conic [1099, 1156]. Connected [632, 1158, 908]. Connections [1273]. Connectivity [124]. Connectors [321]. Connects [1128]. Conscious [17]. Consciousness [1206]. Considerations [554]. Consistency [1255, 61, 151, 1286, 1253, 58, 1288]. Consistent [1490, 923]. Consolidated [920]. Constant [1466, 656, 582, 1240]. Constants [1359]. Constellation [245]. Constrained [1271, 1179, 381]. Constraint [697, 1232, 1235, 1270, 934, 1008, 1342, 1392, 1379, 1251, 1469, 1481, 1265, 1254, 935, 1411, 1312, 1303, 1276, 762, 1032, 1263, 1278, 1311, 1302, 1275, 1289, 1296, 1290, 1286, 1285, 760, 1256, 1284, 1294, 1236, 1298]. Constraint-Based [1379, 762, 1290, 760]. Constraint-Free [1008]. Constraints [1091, 1233, 916, 1299, 1255, 859, 745, 1237, 1258, 1346, 933, 1433, 1273, 63, 1250, 1275, 1300, 1247, 1305, 1304]. Constraints-Preserving [1433]. Construct [290]. Constructed [450, 114]. Constructing [1104, 74, 947].

Construction [407, 1446, 654, 1107, 1123, 89, 1141, 902, 645, 22, 103]. Constructions [640, 1112]. Constructive [23]. Constructively [1321]. Consumption [863, 1077, 877, 872]. Contactless [928]. Containment [1088]. Contemporary [1095]. Content [1045, 1042, 685, 401, 142, 158, 240, 153, 279]. Content- [142]. Content-Based [401]. Contents [977, 978]. Context [1486, 811, 1517, 1525, 1312, 945, 1544, 1325, 301, 1499, 778, 1542, 259, 557]. Context-Aware [1517]. Context-Free [1486, 778]. Context-Independent [811]. Continual [949]. Continuation [1347]. Continuator [718]. Continuity [1321]. Continuous [1319, 1178, 426, 403, 961, 1189, 485, 797, 1275, 1296, 968, 225]. Contract [1052]. Contracts [699, 1031, 913, 1054]. Contribution [319, 1270]. Control [81, 668, 1271, 1449, 364, 310, 404, 1432, 1038, 1331, 363, 721, 1369, 47]. Controlled [314, 891]. Controller [84, 589]. Controlling [1264]. Conventional [854]. Convergence [1116]. Conversational [11, 829]. Conversations [1394]. Conversion [848, 1045, 108]. Converters [562]. Convex [1372]. Convexity [1282, 1465]. Convolvers [582]. Cooperating [1415]. Cooperation [1413, 1438]. Cooperations [1439]. Cooperative [749, 70, 612, 113, 606, 1436, 443]. Coordinate [395]. Coordination [312, 729, 605, 85, 1454, 1453, 1455]. Coprocessor [538]. Coprocessors [563, 522]. COPSE [618]. COPSE-Web [618]. Copying [1227]. Copying-Identifier [1227]. Copyright [1050]. CORBA [321, 1456, 322]. CORBA-Based [1456]. Core [769, 474, 560]. Cores [494]. Corner [673]. Corpora [823]. Corporate [275, 743]. Corpus [971, 774, 791, 1403]. Corpus-Based [791]. Correct [774].

Correcting [237]. Correction [1233]. Cosine [515]. CoSMo [944]. Cost [1299, 160, 581, 1545, 63, 1231, 1007, 284, 877, 124]. Cost-Based [63, 1231]. Costs [924, 1154]. COTS [311]. Counter [389]. Countermeasure [385]. Counters [884, 1363]. Counting [1464, 1468]. Coupled [882]. Course [447, 1161]. Courseware [434, 441]. CovaTM [1451]. Cover [1103]. Coverage [699]. Coverage-Determination [699]. Covering [1138, 1185, 1177]. CPU [908, 561]. CPU-RFU [561]. Crawlers [726]. Create [317]. Creating [544, 438]. Creation [965]. Creator [682]. Credentials [624]. CREM [265]. Crime [914]. Criteria [990, 71, 605, 46]. Criterion [231]. Critical [439]. Criticality [1281]. Cross [1128, 1221, 779, 789, 1171]. Cross-Connects [1128]. Cross-Entropy [1171]. Cross-Language [1221, 779, 789]. Crosstalk [846]. CRT [635]. CRT-Exponent [635]. Cryptanalysis [652, 550, 635]. Cryptanalytic [543]. Cryptographic [654, 384, 541, 542, 379]. Cryptography [540, 1135]. Cryptology [641]. Cryptosystem [381]. Cryptosystems [642, 377]. CSAT [593]. CSCW [604, 1440, 614]. CSoC [533]. CSP [1252, 1309, 1274, 1287, 1249]. CSPs [1283]. Cubes [953]. Cubic [1473, 1140]. Cucumber [553]. Cumulative [1276]. cumulatives} [1235]. Curious [735]. Current [879]. Curve [1087, 540, 1135, 386, 879]. Curves [996, 643]. Custom [538, 580, 481, 473]. Customer [1049, 13]. Customers [1060]. Customisation [518]. Customising [1227]. Customization [289]. Cuts [1153, 1157]. CWM [161, 413]. Cyber [451, 1461]. Cyber-Hearth [1461]. Cyclic [1171]. Czech [793, 792, 790, 795, 798, 797, 805].

D [882, 848, 574, 9, 537, 1519, 659, 1132, 597, 1164, 973, 462]. D-algorithm [597]. D-Latch [882]. DAN [267]. DARP [521]. Data [890, 1086, 1092, 1464, 1012, 1096, 870, 1008, 1006, 963, 986, 1014, 1446, 979, 164, 1108, 147, 1009, 408, 958, 1111, 764, 1491, 1082, 1419, 163, 1015, 364, 770, 729, 926, 908, 176, 1495, 970, 426, 465, 1076, 572, 1072, 287, 125, 419, 741, 873, 411, 143, 901, 135, 1005, 910, 424, 1013, 953, 617, 150, 77, 585, 766, 902, 1145, 1026, 363, 925, 660, 768, 817, 1503, 872, 144, 406, 893, 962, 756, 1403, 1210, 1047, 1504, 976, 942, 867, 261, 270, 141, 965, 232, 1361, 968, 73, 1051, 379, 264, 402, 911, 153, 82, 1284, 1294, 152]. Data [988, 1298, 937, 1496]. Data-Driven [125, 817]. Data-Intensive [872]. Data-Processor [585]. Data-Stream-Based [465]. Database [412, 121, 670, 912, 946, 367, 928, 139, 795, 453, 787, 957, 939, 144, 982, 138, 956, 929, 455]. Databases [288, 948, 941, 729, 415, 173, 100, 422, 975, 283, 689, 924, 423, 456]. Dataflow [1373, 1374]. DataGrid [1079]. Datalog [416]. Datapath [841]. Dataset [760]. Datasets [1011]. Datatypes [1362]. DATR [674]. dbRouter [954]. DDH [657]. DDM [887]. Decathlon [1010]. Decidability [1349]. Decidable [190, 1335, 1359]. Decimal [201]. Decision [110, 1446, 914, 1396, 59, 995, 605, 823, 1023, 47]. Declarative [306, 9, 424]. Declination [793]. Decodable [1467]. Decoder [512]. Decomposing [667]. Decomposition [1270, 500]. Decompositions [1142]. Decompositon [249]. Dedicated [855]. Deep [511]. Defining [398, 276, 30]. Definite [674]. Definition [943]. Defragmentation [495]. Degradation [887, 216]. Degree [1104, 907]. Delay [882, 527, 525, 849, 887, 886, 861]. Delegation [379]. Deleting [27]. deling [944]. Delivered [448]. Delivery [279, 388, 158]. Demodulation [475]. Demodulator [476]. Dempster [670].

Dempster-Shafer [670]. Deniable [650]. Denial [751]. Denotational [213]. Density [1108]. Density-Based [1108]. Deontic [55]. Dependability [1071, 1072]. Dependence [1268]. Dependencies [1197, 770, 398]. Dependency [123, 782, 775]. Dependent [572, 1133, 144, 1292]. Deploying [1079]. Deployment [323]. Derandomization [1480]. Derivation [69, 915]. Derivational [3]. Deriving [934, 892]. Describing [1438]. Description [130, 263, 1070, 1204, 44, 939, 430, 1439]. Descriptions [1404, 1386, 58]. Design [435, 352, 1021, 1006, 609, 564, 303, 670, 1522, 291, 1458, 365, 476, 1425, 581, 733, 851, 293, 1220, 39, 40, 888, 432, 1020, 569, 1083, 567, 505, 568, 468, 477, 880, 1225, 301, 1145, 695, 1272, 43, 868, 813, 513, 1295, 1196, 46, 573, 326, 1153, 456, 1313, 31, 866, 1301]. Designated [157]. Designing [852, 280, 608, 1053]. Designs [497, 592, 1344, 1343, 479, 518, 10]. Desired [81]. Desk [351]. Detecting [1390, 755, 244]. Detection [248, 358, 506, 754, 426, 1000, 149, 132, 134, 768, 579, 246]. Detector [813]. Determination [699, 1021, 22, 1302]. Determining [1252, 757, 710, 1120]. Deterministic [1126, 1505]. D’etre [1461]. Developed [784]. Developing [449, 349, 618, 797, 1514]. Development [215, 1462, 502, 309, 306, 286, 910, 1027, 585, 717, 694, 211, 188, 1426, 47, 278]. Devices [517, 1538, 492, 544, 488, 731, 1229, 532, 1521]. DFG [1415, 1416, 1417]. DH [657]. Diagnosis [4, 1003, 577, 33]. Diagram [1116, 314]. Diagrams [1132, 59]. Dial [1139]. Dial-a-Ride [1139]. Dialog [1514]. Dialogs [28]. Dialogue [130, 830, 835, 831, 832, 659, 814, 823, 825]. Dialogues [1412, 824]. DIC [69]. Difference [160, 233]. Differences [327, 904, 216]. Different [809, 829, 28, 1286]. Differential [860, 911].

Difficulty [357]. Digest [965]. Digital [874, 254, 556, 554, 487, 6, 1117, 581, 885, 492, 468, 1040, 919, 879, 455, 177, 521]. Dimension [997]. Dimensional [1106, 165, 408, 1201, 418, 586, 962, 976, 989]. Dimensionality [1482]. Direct [668, 1526]. Directed [683, 67, 308]. Direction [930]. Direction-Based [930]. Directional [1318]. Directions [273, 1253]. Disambiguation [1198, 788]. disciplinary [1017]. Discontinuous [231]. Discourse [781]. Discourse-Semantic [781]. Discovering [712, 235, 1422]. Discovery [767, 711, 754, 967, 769, 762, 100, 998, 927, 77, 768, 402]. Discrete [539, 376, 515, 485, 558, 141, 69, 225, 871]. Discrete-Log [376]. Discriminant [816, 233]. Discrimination [792]. Discriminations [1404]. Disequality [1275]. Disjunctive [1296, 1305]. Disk [908]. Dismantleable [1469]. Dispatch [1030]. Dispatching [1136, 1424]. Display [1523]. Displays [1527, 1534]. Disputes [1344]. Disruptive [465]. Dissemination [741]. Dissociations [985]. Distance [800, 445, 428, 990, 408, 921, 1482, 429, 438, 1084, 457, 989]. Distance-Based [408]. Distillation [166]. Distinct [1464]. Distinguishers [154]. Distinguishing [222]. Distributed [1452, 1069, 748, 312, 491, 742, 1458, 382, 1039, 655, 1274, 743, 384, 1357, 61, 1447, 1448, 605, 1038, 1544, 1440, 924, 1290, 113, 954, 923, 1453, 1459, 929, 1256]. Distributed-Memory [491]. Distributing [667]. Distribution [859, 1009, 1109, 382, 892, 1050, 617, 233, 804]. Distributional [904]. Diverse [5]. Diversification [1162]. Diversity [17, 33]. Diversity-Conscious [17]. Division [517, 377, 1157]. DjVu [266]. Do [684, 157]. Document [319, 263, 269, 603, 262, 1184, 257, 268, 241, 271, 252, 266, 920, 245, 243, 270, 240, 249, 264, 227]. Document-Form [245]. Documentation [671]. Documents [92, 288, 255, 95, 918, 179, 267, 217, 964, 967, 1429, 900, 94, 259, 260, 220, 244, 375, 784]. Doll [1248]. DOM [139]. DOM-Based [139]. Domain [835, 898, 471, 658, 1425, 315, 588, 294, 788, 99, 1426]. Domain-Specific [471, 1425, 315]. Domains [807, 932, 1366, 28, 1211, 57, 895]. Dominant [850]. Dominating [1474]. Double [1114]. Down [226, 122, 30]. DP [559]. DPA [385]. Drawing [1002]. Drive [590]. Driven [262, 836, 125, 1025, 854, 470, 925, 817, 1301]. Driving [883, 273]. Drum [714]. DSMIO [307]. DSP [590, 474, 843]. DTD [1033]. Dual [854, 1257]. Duality [1348]. Dualization [1119]. Duplicated [616]. Duplication [1449]. during [1258, 716]. Dutch [1018]. DVB [475]. Dynamic [1394, 60, 624, 1415, 235, 1279, 1168, 163, 941, 1183, 770, 205, 908, 384, 1170, 1246, 582, 1132, 1428, 1429, 545, 1044, 688, 1163, 1268, 878, 1395, 482, 739, 891, 1431, 1190, 1064, 703, 57, 999]. Dynamically [571, 492, 570, 480].

E-business [116, 688, 276, 277, 1058, 1054]. E-cash [368]. E-Commerce [1456, 1028, 1022, 1056, 363, 1053, 734, 894, 1029]. E-content [279]. E-forms [280]. E-Learning [461, 103]. E-mails [98]. E-markets [1024]. E-science [1079]. E-service [696]. E-services [280, 702, 700, 689, 158]. Eager [1105]. Easy [1481, 460]. eBook [1045]. ECA [912, 1454]. ECDSA [573]. ecurity [944]. Edge [1269]. EDI [1036]. Editing [671]. Editors [1459]. EDM [595]. EDSM [1489]. Education [432, 446, 462, 455]. Educational [442]. Effect [605]. Effective [431, 93, 527, 823, 134]. Effects [526]. Efficiency [230, 400]. Efficient [646, 642, 1012, 1098, 1112, 624, 475, 960, 990, 171, 391, 858, 384, 167, 1129, 1246, 519, 409, 961, 903, 143, 378, 145, 880, 557, 802, 1145, 1026, 22, 868, 660, 879, 393, 535, 661, 390, 942, 1365, 1212, 138, 1370, 1441, 937, 140, 1288, 1092]. Efficiently [969, 1118, 664, 94]. Effort [842]. Eigenvalue [1483]. Elaborating [37]. Election [1440]. Electrical [590]. Electricity [921]. Electronic [350, 352, 1021, 281, 371, 1443, 1020, 928, 1050, 373, 266, 728, 840, 24, 46, 28]. Element [578, 558]. Element-Based [578]. Elements [1464, 853, 1148]. Elimination [990]. ELIZA [107]. ELIZA-Like [107]. Elliptic [540, 1135, 386, 393]. Embedded [870, 474, 1448, 587, 560, 1264]. Embedding [1051]. EMILE [1509]. Emotion [110, 175]. Emotions [1400]. Empirical [1267, 1209, 148]. Empty [959]. Emulation [597, 467]. EMV [1056]. Enable [545]. Enabled [1414, 952]. Enabling [1017, 450, 1057, 461]. Encoded [1429]. Encoding [899, 873, 875]. Encodings [167]. Encryption [623, 628, 621, 633, 627, 636]. End [1067]. End-to-End [1067]. Endangered [821]. Energy [863, 1077, 519, 865, 498, 868, 877, 867, 866]. Energy-Efficient [519, 868]. Enforcement [704, 1455]. Engine [154, 267, 1265, 1218, 472, 542, 260, 153]. Engineering [425, 265, 268, 1141, 1278, 1226, 1206]. Enhance [870, 177]. Enhanced [533, 1424]. Enhancement [819]. Enhancing [1493, 837]. Enough [396]. Enrolling [135]. Ensembles [32]. Ensure [199]. Ensuring [894]. Enterprise [306, 1437, 1432, 913, 300, 1438]. Enterprises [1022, 1017]. Entropy [28, 788, 1171]. Entropy-Based [28]. Envelope [1240]. Environment [604, 1041, 502, 1522, 749, 310, 450, 399, 1429, 404, 174, 605, 854, 1226, 20, 853, 1426, 1457, 994, 1453, 82, 1455, 443]. Environmental [73]. Environments [442, 396, 1028, 1415, 1460, 1523, 1535, 744, 1529, 925, 433]. Episodes [758, 769]. Episodic [70]. Epistemic [1413]. Equality [1275, 1374]. Equality-Based [1374]. Equational [197]. Equations [1238]. Equivalence [1342, 559]. Equivalent [1423]. Erasure [512]. EROICA [156]. Erroneous [1284, 1294]. Error [79]. Errors [78]. Escape [207]. Essential [620]. Establishment [913]. Estimating [1111, 987, 1500, 924]. Estimation [811, 522, 1113, 76, 885, 489, 802, 1007, 879, 872, 499]. Estimations [8, 1015]. Estuarine [43]. Eulerian [1142]. Eureka [927]. European [1079, 1147, 311]. European-Asian [1147]. Evaluate [414]. Evaluating [309, 613, 1224, 825, 307, 722]. Evaluation [412, 242, 1071, 748, 832, 1522, 1458, 1065, 826, 827, 961, 139, 498, 862, 1046, 775, 924, 456, 270, 665, 565, 818, 1244, 614]. Even [1503, 1158]. Event [1400, 943, 876, 134, 1540, 931]. Events [876]. Evidence [993]. Evolution [963, 941, 1444, 118, 441]. Evolutionary [1217, 112, 211]. Evolving [209, 84]. Exact [1146]. Examples [1514]. Exchange [629, 370, 705, 644, 366]. Exchanges [1020]. Exchanging [749]. Excluded [1320]. Exclusion [370]. Exclusion-Freeness [370]. Executable [1055]. Execution [1012, 1450, 509, 1054, 703, 62]. Exhaustive [421]. Expanded [1133]. Expanding [648]. Expansion [656, 257, 1061]. Experience [344, 349, 348, 342, 350, 352, 351, 347, 343, 421]. Experiences [449, 447, 1079, 448, 727, 1518]. Experiment [314, 1211]. Experimental [1095, 1165, 271, 1247, 295, 33]. Experiments [822, 1214, 10, 732, 805]. Expert [1397, 682, 982, 223]. Experts [73]. Explanation [1219, 32]. Exploitation [461]. Exploiting [11, 527, 271, 1243, 992, 789, 747, 845, 177]. Exploration [533, 1115, 1169, 573, 166]. Exploratory [1215]. Exploring [1524, 945].

Exponent [635]. Expressing [1445]. Expression [713, 716, 924, 140, 146]. Expressions [698]. Expressive [5, 708]. Expressiveness [1333]. Extended [963, 186, 1327, 783, 625, 1339, 644, 1141, 1317, 1023, 638, 1156]. Extending [1149, 204]. Extensible [1462, 931]. Extension [578, 73, 524]. External [1143]. External-Memory [1143]. Extract [152]. Extracted [935]. Extracting [1197, 98, 424, 150]. Extraction [984, 253, 254, 95, 218, 967, 168, 128, 251, 685, 261, 249, 264, 99, 804, 1353]. Extractors [636].

F2 [288]. F2/XML [288]. Fabric [478].

Fabric-Specific [478]. Face [1526, 105, 132, 1461, 1527, 106]. Face-Responsive [1526]. Face-to-Face [1461]. Facial [175]. Facilitator [1022]. Factor [656, 1057]. Factored [59]. Factors [561, 439]. Failed [664]. Failure [1072]. Fair [705]. Fairness [1334]. Fast [533, 1124, 595, 1402, 468, 855, 998, 360, 1147, 975, 513, 879, 467, 1502, 507]. Faster [1195, 1488]. Fault [1072]. Faults [4, 527, 525, 524]. Feasibility [1211]. Feature [218, 714, 409, 977, 233, 103]. Features [1215, 764, 216, 804]. Federation [821]. Federations [1007]. Feedback [1224]. Feedforward [1225]. Festival [806]. Few [1262, 1366]. Field [508, 1072, 1192, 473]. Field-Programmable [508, 473]. Fields [643, 999, 1529]. Fifth [460]. File [867, 866]. Files [911]. Filling [996]. Filter [653, 484]. Filter-Combiner [653]. Filtering [511, 176, 1043, 1042, 1238, 128, 740, 938, 1207, 401, 29, 1399, 21, 739, 1156, 760, 784]. Filters [554, 487, 580, 555]. Financial [1213, 87]. Find [684]. Finding [178, 1475, 1199, 1201, 1152, 724, 1171, 1158]. FindIT [1528]. Fine [545, 547]. Fine-Grain [545]. Fine-Grained [547]. Fingerprinting [391]. Finite [1406, 487, 1251, 1492, 1471, 643]. Finite-Variable [1251]. Finiteness [1389]. FIR [580]. First [822, 833, 1371, 1143, 886, 840, 543, 1275, 806]. First-Order [1371, 1275]. Fitting [820, 1405, 1076]. Fixed [1337]. Fixed-Point [1337]. Fixing [1153]. Fixpoint [1323]. Flashlight [1528]. Flattening [415]. Flaws [626]. Flexibility [549]. Flexible [155, 832, 447, 471, 568, 128, 453, 328, 496, 405, 520]. Flip [860]. Flip-Flops [860]. Floating [531, 518, 508, 528, 529, 1368]. Floating-Gate [508]. Floating-Point [531, 518, 528, 529]. Floor [1186]. Flops [860]. Flow [1373, 1093, 199, 659, 1133, 897, 320, 698, 1369]. Flow-Dependent [1133]. Flow-Insensitive [1373]. Flows [1530]. Fly [503]. Focusing [1009]. Folding [1164]. Forbidden [1488]. Force [1529]. Forecasting [38, 1202, 921]. Form [247, 1340, 1324, 460, 1060, 1206, 245, 1305, 1353]. Formal [1176, 715, 1055, 211, 942]. Formalism [1070, 291]. Forming [103]. forms [280]. Formula [1244]. Formulae [376]. Formulas [1320, 1383, 1243]. ‘ForSe [1529]. Forward [683]. Foundation [1176, 207, 249]. Founded [1405]. FPGA [539, 536, 504, 590, 475, 592, 490, 476, 515, 589, 486, 583, 505, 586, 550, 548, 468, 549, 470, 528, 576, 485, 598, 543, 529, 516, 500, 875, 523, 577, 524]. FPGA-based [475, 589, 583, 468, 577]. FPGA-DSP [590]. FPGAs [491, 556, 554, 487, 495, 593, 469, 582, 519, 569, 489, 567, 527, 525, 545, 501, 496, 526, 499, 553, 484, 507]. FPL [513]. FPL-based [513]. FPLD [562]. FPLDs [555]. Fractal [1076]. Fragmentation [1493]. Frame [1326]. Framework [3, 533, 1462, 1021, 309, 414, 832, 1511, 571, 1122, 952, 1519, 933, 696, 732, 1031, 325, 905, 612, 477, 1424, 1179, 276, 510, 1300, 295, 300, 296, 429, 1457, 430, 208, 954]. Fraud [1003]. Free [369, 1486, 1008, 669, 1499, 778]. Free-Text [669]. Freeness [370]. Frequencies [1102]. Frequency [1113, 555, 1144]. Frequent [767, 1001, 145, 402, 760, 152]. FSM [500]. FSMs [875]. Fully [187, 1130, 540, 1331]. Fun [1083]. Function [578, 640, 645, 1401, 47]. Functional [578, 398, 480, 188, 310, 322]. Functionality [690]. Functionals [1339]. Functions [189, 654, 648, 1178, 361, 1362, 849, 657, 1430, 456, 69, 643]. Furniture [1543]. Fused [41]. Fusion [696]. Future [727, 728]. Fuzziness [60]. Fuzzy [250, 1396, 903, 24, 26, 119, 69, 120]. FuzzyDrive [120].

G [1062, 1060, 607]. G-Networks [1060]. GA [1195, 559, 1228]. GA-DP [559]. GAACO [1195]. Gage [587]. Galleries [1174]. Games [1194, 1336, 1332, 1317, 195, 182]. Gap [1309]. Gaps [231]. Gate [874, 508]. Gates [882, 883]. Gathering [744]. Gating [884]. Gaussian [815]. Gaussians [819]. GDWFMS [179]. GEM [621]. Gene [596, 159, 146]. GeneAround [159]. General [709, 171, 1122, 13, 569, 1330, 585]. Generalises [1364]. Generalization [81, 296]. Generalizations [1340]. Generalized [1112, 1495, 509, 638, 264]. Generalizied [19]. Generate [860]. Generating [294, 1231, 441]. Generation [646, 776, 898, 471, 7, 1308, 84, 74, 1303, 783, 226, 304, 876, 671, 930, 1078, 1457, 119, 565, 99, 1298]. Generative [358, 123]. Generator [580]. Generators [578]. Generic [997, 298, 283, 1540, 1054]. Genes [1206]. Genetic [1004, 1206, 1208, 1188, 115, 66]. Genomic [975]. genre [685]. Geo [6]. Geo-spatial [6]. Geographic [410, 965]. Geometric [238, 1108, 1123, 933, 411, 501]. Geospatial [926]. German [820, 795]. Gesture [708]. Getting [763, 610]. GF} [378, 540, 377].

GHz [853]. Gibbsian [778]. GIDB [901]. Gigabit [506]. Gigabits [541]. GIS [143, 901]. Global [1299, 305, 1237, 89, 1238]. GnuPG [362]. Goal [683]. Goal-Directed [683]. Goes [601]. Good [1115, 1480, 834]. GPS [1196]. GQ [630]. Grain [568, 545]. Grained [547]. Gram [820]. Grammar [1509, 247, 123, 1500, 778, 723]. Grammar-Based [723]. Grammars [1486, 673, 1511, 1392, 1490, 1494, 1499, 1502, 1512, 1504, 1496]. Grammatical [1497, 1507]. Granidt [506]. granularity [934]. Granulometries [263]. Graph [247, 178, 215, 857, 1115, 1472, 167, 1539, 241, 77, 557, 1366, 675, 1257, 152]. Graph-Based [857, 77, 1366]. Graphical [1445]. Graphics [574, 1519, 239, 462]. Graphs [179, 1475, 1484, 1473, 1474, 1469, 67, 1133, 1140, 1471, 1142, 1145, 1084, 1158]. Gray [243]. Gray-Scale [243]. Greedy [1131]. Greek [806]. Greek-TTS [806]. Greibach [1324]. Grid [1086, 844, 446, 443, 277]. Grid-Based [446]. Ground [242]. Ground-Truthing [242]. GroundTruth [270]. Group [1162, 1112, 615, 945, 1460, 605, 374, 103, 1459]. Groups [611, 649, 1484, 637, 1258]. Groupware [618, 617, 600, 608, 607, 601, 614]. Groupwares [616]. Growing [428]. GSAT [1277]. Guarantee [1116]. Guaranteed [1375]. Guarded [209]. Guided [1486, 824, 1136, 44, 1171]. Guiding [1407]. GVR [1208].

HAGAR [557]. Hairetes [260]. Halftoning [1117]. Hamming [393]. Hand [731]. Hand-Held [731]. Handheld [604]. Handhelds [828]. Handling [1238, 524]. Handwriting [225, 227]. Handwritten [222]. Hard [916, 1490]. Harder [664]. Hardness [1267]. Hardware [502, 564, 551, 511, 537, 361, 588, 503, 569, 546, 570, 566, 529, 359, 579, 381, 524].

Hardware-Assisted [566]. Harmonisation [721]. HASCASL [188]. Hash [640, 658, 361, 383]. Hash-Function [640]. Haskell [666]. HBT [507]. HCI [460]. HCM [1438]. Heads [659]. Health [1005]. Hearth [1461]. Heating [859]. Heavy [1076]. Height [1354]. Height-Restrictions [1354]. Heights [1235]. Held [731]. Help [351, 1498]. Help-Desk [351]. Helping [731]. Here [464]. Heterarchical [1178]. Hetereogeneous [1362]. Heterogeneous [963, 946, 287, 325, 1447, 1305, 1036, 929]. Heuristic [1076, 1293, 813, 1285, 665]. Heuristics [1090, 160, 1239, 663]. Hidden [90, 51, 168, 64, 1203, 108, 647, 485, 719, 1506, 1264]. Hiding [656]. Hierarchical [1028, 979, 957, 789, 966]. Hierarchies [147, 329, 992]. Hierarchy [1479, 1098, 1096, 1338, 865, 567]. High [556, 554, 165, 408, 589, 486, 492, 505, 586, 1204, 555, 843, 316, 359, 962, 976, 1156, 493, 989]. High-Dimensional [408, 976]. High-Level [492, 1204, 316, 1156]. High-Performance [843]. High-Speed [359]. Higher [1346, 1339, 185, 1349]. Higher-Order [1346, 185, 1349]. Highly [794]. HIPERLAN [575]. HIPERLAN/2 [575]. Histograms [974, 170]. History [805]. Hit [757]. HMM [795, 134]. HMM-Based [795, 134]. Hoare [1322]. Hoc [992, 701, 649]. Holistic [226]. Homogeneous [1175, 1184, 1036]. Homologous [91]. Homology [975, 493]. Homomorphism [392]. Homotopic [1118]. Horn [1360, 1301]. Horn-Driven [1301]. HotSpot [588]. HP [1164]. HPIR [144]. HTML [833, 900, 244]. Human [250, 269, 1524, 53]. Human-Computer [53]. Humanoid [84]. Hungarian [781]. Hybrid [1195, 190, 38, 1022, 1404, 1519, 1242, 991, 1536, 559, 1273, 501, 25, 561, 87]. Hybridised [1228]. Hyper [969, 636]. Hyper-encryption [636]. Hyper-sphere [969]. Hypercube [1477]. Hyperelliptic [643]. Hypergraphs [1470]. Hyperlink [892]. Hyperlink-Based [892]. Hyperlinked [1008]. Hypermedia [425, 606]. Hypertubes [1271]. Hypothetical [63].

I.P [855]. I/O [1092, 533, 955, 1143]. I/O-efficient [1092]. IACBC [621]. ICs [846]. ID [1532]. ID-Based [1532]. IDEA [541]. Identifiability [1493]. Identifiable [1491]. Identification [630, 1215, 217, 391, 1490, 129, 106, 245, 227]. Identifier [1536, 1227]. Identify [1496]. Identifying [1024, 97]. Identity [380]. IEEE [575]. If [759]. II [517, 742, 583, 378]. IKE [629]. Illumination [105]. ILP [669]. Image [538, 269, 6, 514, 226, 591, 1192, 44, 216, 966, 1051, 131]. Images [250, 133, 253, 258, 241, 256, 252, 222, 243, 131, 227]. Imbalanced [366]. Immediate [173]. Impact [844, 293, 228]. Impersonation [630]. Implement [483]. Implementation [539, 533, 1095, 512, 554, 537, 541, 1129, 362, 486, 1451, 432, 505, 550, 548, 549, 849, 1447, 422, 576, 1145, 543, 513, 558, 579, 516]. Implementations [590, 361, 568, 360, 570]. Implemented [582]. Implementing [1092, 121, 487, 515, 593, 469, 1430, 24, 562, 1515]. Implications [1345]. Implicit [1339]. Implied [1308]. Imprint [223]. Improve [567, 400]. Improved [1379, 169, 1015, 841, 1203, 1043, 488, 480]. Improvement [79, 44]. Improvements [633]. Improving [1299, 690, 1124, 659, 1221, 93, 257, 135, 995, 18, 1374, 1277, 977, 21, 230]. Impulse [487]. Imputation [1011]. InBASE [776]. Incentive [1032]. Inclusion [1370]. Inclusion-Based [1370]. Incomplete [1011, 1294]. Incompleteor [1284]. Inconsistent [1233]. Incorporating [765]. Increasing [234, 1018, 33]. Incremental [969, 1123, 1396, 902, 1499, 69, 1427]. Independent [811, 864, 1117]. Index [48, 137, 1355, 1391, 1418, 1463, 1485, 1516, 1546, 180, 203, 274, 330, 394, 427, 463, 599, 619, 687, 706, 725, 753, 771, 838, 889, 983, 1016, 1059, 1080, 1159, 1160, 1230, 1314, 355, 839, 981, 990, 1217, 140]. Indexes [165, 960]. Indexical [1266]. Indexical-Based [1266]. Indexing [1197, 1198, 979, 981, 1429, 998, 256, 139, 127, 453, 975, 458]. Indices [974, 976]. Individual [1518]. Indonesian [124]. Induced [844, 1475]. Induction [1509, 77]. Inductive [4, 1511, 716, 762]. Industry [936, 689]. Inequality [990]. Inequations [1238]. Inexpensive [1539]. Inference [1486, 1411, 1497, 117, 1364, 1421, 1507, 1365, 703]. Inferring [1488, 1263, 1504]. Infinite [1337]. Inflected [826, 827, 794]. Influenced [28]. Influencing [561]. InfoFrax [41]. Informal [1461]. Informatics [278]. Information [737, 309, 199, 414, 83, 742, 1205, 410, 305, 926, 749, 783, 964, 696, 967, 659, 168, 1523, 1530, 1221, 432, 738, 1532, 741, 399, 1222, 1432, 304, 1139, 150, 744, 825, 727, 907, 766, 752, 972, 1023, 789, 397, 1395, 1512, 1399, 115, 318, 978, 1036, 1427, 1436, 99, 1256]. Infrastructure [618, 365, 608, 397, 931]. Inheritance [691]. Initial [805]. Injection [908, 1306, 1072]. Innovative [451]. Input [830, 669, 1445, 886]. Insensitive [1373, 849]. Insertion [673, 1172]. Inspecting [1447]. Instance [661]. Instances [1252, 755, 1281, 62]. Institutions [728]. Instruction [865, 847, 520]. Instructional [435, 102]. Instructions [1543]. Instrument [1225]. Instrumentation [847]. Integer [1091, 1232, 1234, 555]. Integers [214]. Integral [668]. Integral-Plus-State [668]. Integrated [603, 490, 671, 151, 188, 220]. Integrating [1232, 742, 149, 325, 411, 744, 501, 25, 1289, 46, 1401, 1]. Integration [963, 1014, 564, 691, 696, 116, 1450, 101, 287, 901, 1420, 53, 1037, 566, 1431, 799]. Intellectual [1051]. Intelligence [743, 1513, 52]. Intelligent [611, 1446, 88, 1416, 399, 727, 1229, 99, 87]. Intellimedia [52]. Intensification [1162]. Intensity [1107]. Intensity-Modulated [1107]. Intensive [872]. Intention [1459]. Inter [946, 970, 947, 775]. Inter-object [970]. Inter-Phrase [775]. Inter-relational [947]. Inter-system [946]. Interact [729]. Interacting [718]. Interaction [733, 1519, 945]. Interactive [1412, 269, 154, 1218, 1350, 335, 334, 336, 332, 583, 927, 1529, 1226, 1293, 1534, 438, 195, 462]. Interception [1410]. Interchangeability [1283, 1311]. Interconnection [545]. Interest [1484]. Interesting [755]. Interests [739]. Interface [584, 787, 229]. Interfaces [1526, 727]. Interfacing [828]. Interference [1331]. Intermediary [1022]. Internet [1458, 1113, 448, 1055, 1078, 405, 1455]. Internet-Based [1458]. Interpretation [707, 1387, 1386, 1362, 128, 1385, 672, 1381]. Interpretation-Based [1385]. Intersection [52]. Interval [1015, 404, 383, 662]. Intonation [717]. Intra [412]. Intra-Query [412]. Intrinsic [177]. Introducing [534]. Introduction [341, 331]. Intruder [356]. Intrusion [358, 506, 134, 651]. Intrusion-Resilient [651]. Inventory [86]. Inversion [377]. Inversion/Division [377]. Investigating [1417]. Investigation [1541]. Investigations [1217]. Investment [87]. INVISIP [926]. IP [46]. iPACE [472]. iPACE-V1 [472]. Irregular [251]. Irreversibility [177]. ISA [969]. Isolated [719]. Isomorphism [1342, 572]. Issues [890, 449, 1010, 422, 1521, 1075]. Item [1216]. Itemsets [1001, 402]. Iterated [8]. Iteration [1348]. Iterative [81, 490]. ITR [45].

J [932]. J-Pruning [932]. Japanese [775, 228, 230]. Java [201, 310, 296].

JavaBeans [306]. JavaCard [184]. JBits [487, 596, 481]. JECKO [301]. JML [204]. Join [1008, 997, 962]. Joins [948, 959]. Joint [987, 393]. JPEG [516]. JSP [441]. Juridical [1052].

Kanji [228]. Kannada [219]. KAON [1048]. KASUMI [359]. kd} [1085]. kd}-Trees [1085]. Kernel [816, 804]. Key [629, 710, 382, 377, 1050, 644, 724, 381, 366]. Key-Exchange [629]. Key-Finding [724]. Keys [380, 357, 1001]. Keyword [810, 917, 138]. Keywords [245]. KGCL [443]. Killer [686]. Kinetic [1085]. Kirkman [1262]. Knapsack [1134]. Knowledge [435, 1409, 345, 376, 6, 898, 1449, 51, 74, 743, 293, 149, 50, 459, 14, 1428, 1004, 1420, 1430, 927, 77, 1226, 1046, 1243, 682, 295, 1426, 765, 292, 73, 2, 1413, 1427, 828, 99, 1, 461, 443, 1212]. Knowledge-Based [435, 898, 1046]. Knowledge-Grid-Based [443]. Known [648]. Known-Plaintext [648]. Korean [229].

Lab [445, 614]. Labeled [766]. Labeling [1102, 241]. Labels [1126]. Laboratories [440]. Ladder [863]. Lagrangian [1153]. Lambek [1494]. Land [47]. Language [697, 781, 130, 1070, 709, 5, 1392, 436, 826, 827, 824, 820, 1221, 126, 1513, 777, 950, 798, 797, 779, 919, 721, 805, 787, 817, 1382, 789, 803, 794, 1512, 1504, 1212]. Languages [984, 1488, 414, 1510, 1466, 1200, 1491, 1498, 741, 821, 1501, 1503, 1505, 1508, 1370]. Large [449, 1048, 163, 1079, 1126, 282, 89, 732, 59, 1527, 112, 957, 817, 1534, 228, 1011, 784]. Large-Scale [449, 89, 957]. Larger [654]. Latch [882]. Latches [860]. Latency [29]. Latent [1198]. Latest [278]. Latin [1232]. Lattice [1476, 812]. Lattices [1119]. Law [1320, 373, 1483]. Layer [783, 1225, 845]. Layered [324, 1190, 116]. Layers [1106].

Layout [242, 1098, 238, 255, 241, 243]. Lazy [32]. LC [71, 72]. LC-Learning [71, 72]. Leaf [1140]. Leakage [856, 367]. Learn [1510, 739]. Learnability [1505, 1506]. Learners [32]. Learning [81, 4, 1487, 1397, 133, 445, 428, 147, 1511, 613, 1266, 1492, 602, 436, 1404, 1411, 50, 1460, 459, 70, 431, 432, 1203, 1498, 257, 71, 72, 448, 444, 96, 135, 68, 1267, 241, 1513, 1226, 663, 1499, 662, 718, 49, 44, 660, 1247, 1395, 1502, 1503, 429, 438, 458, 1512, 430, 1211, 233, 452, 985, 1508, 78, 1304, 433, 457, 31, 264, 82, 148, 461, 443, 1515, 905, 103, 71, 72]. Learning-Based [148]. Least [1398]. Lecture [453, 456]. LeedsCQ} [949]. Left [673]. Legacy [935]. Legal [906, 920]. Length [76, 1204]. less [484]. Lessons [468]. Letter [1502]. Level [874, 564, 522, 696, 492, 855, 849, 1204, 847, 316, 879, 1156]. Levelized [143]. Lexical [794]. Lexicographic [1237]. Lexicon [790]. Libraries [919, 1369]. Library [1372, 531, 852, 1135, 864, 455]. Life [109, 1257]. Life-Like [109]. LIGHT} [1392]. Lightweight [1141]. Like [1076, 815, 107, 109, 1171, 65]. Likelihood [800, 226]. Limit [1491, 1490, 1494, 1315, 1496]. Limit-Computable [1315]. Limited [807, 1113, 705]. Line [1089, 1166, 315, 1139, 636, 1121, 495, 1155]. Linear [1087, 1091, 1233, 1341, 652, 655, 1345, 1326, 550, 1467, 416, 1421, 1503, 1157, 1441, 170, 1158, 1301]. Linear-Logic [1326]. Linear-Order [1441]. Linear-Spline [170]. Lines [231, 883, 324]. Linguistic [672]. Link [978]. Linkage [178]. Linked [977, 1363]. Lip [799]. Lip-Reading [799]. Liquid [435]. List [1097, 353]. Lists [1363]. Load [908, 1542]. Loan [374]. Local [1341, 1299, 83, 1077, 664, 593, 1246, 89, 1241, 15, 323]. Locality [196]. Locally [1467, 1503]. Localness [907]. Location [748, 922, 1519, 1537, 1538, 403, 1536, 930, 1540, 144]. Location-Based [922]. Location-Sensitive [1519]. Locking [454]. Locking-Based [454]. Log [376, 992]. Logarithmic [528]. Logic [882, 1335, 852, 466, 1388, 1320, 1350, 1345, 111, 1326, 192, 716, 1337, 1333, 474, 849, 854, 488, 1390, 1329, 1316, 1356, 480, 879, 1368, 55, 204, 1365, 1413]. Logic-Centric [466]. Logical [842, 1351, 241]. Logics [1251, 1322, 939, 681]. Logistic [1167]. Logs [1419, 453, 1422]. Long [865]. Look [852]. Look-Ahead [852]. Loops [1181, 475, 1367]. Low [848, 808, 1104, 857, 581, 851, 865, 873, 1545, 555, 884, 843, 877, 500, 875, 845, 866, 1288]. Low-Cost [1545, 877]. Low-Power [848, 851, 843, 875, 845]. Lower [1090, 1467]. LSD [623]. LTL [1383, 1333, 1332]. LUC [647]. LUM [537]. LUT [592, 469, 527, 525, 484]. LUT-based [527, 525, 484]. Lyrebird [1514].

M [1062]. M/G/1 [1062]. M/G/1-Type [1062]. Machine [588, 1221, 125, 219, 96, 663, 49, 222, 1211, 264, 1393, 1155]. Machines [810, 358, 1154, 473]. Macro [841, 565]. Macro-Model [841]. Made [460]. Mail [371]. mails [98]. Main [1430]. maintainable [911]. Maintaining [1096, 1097, 349, 151]. Maintenance [160, 13, 1013, 18, 27, 923]. Maintenance-Cost [160]. Maiorana [654]. Majority [389, 236]. Make [160]. Makespan [1154]. Making [110, 729, 605, 823, 1023, 1384]. MALACH [779, 805]. Malleable [1130]. Manage [144]. Management [603, 36, 349, 342, 350, 352, 351, 121, 1446, 155, 6, 1449, 365, 936, 1039, 941, 743, 282, 1326, 459, 156, 1420, 172, 1448, 1424, 897, 320, 1430, 322, 1026, 925, 682, 1421, 957, 1395, 1054, 2, 153, 1427, 1436]. Manager [215]. Managing [700, 920, 115]. Manipulation [1526]. Manual [893]. Manufacture [41]. Manufacturing [1437].

Many [418]. Many-Dimensional [418]. Map [83, 1428, 140, 1076]. Mapping [1433, 1047]. Maps [1115, 82]. Marginal [987]. Market [1217, 921]. markets [1024]. Markov [1469, 335, 334, 336, 333, 64, 1203, 108, 485, 1061, 1192, 719, 1062, 1506]. Markov-Modulated [1061]. Markovian [1076]. MAS [675]. Masking [652]. Massively [597]. Master [448]. Matching [596, 128, 241, 252, 245, 196, 1431]. Mathematics [1315]. Matrix [538, 1261, 519, 1310]. Max [1090, 1249]. Max-CSP [1249]. Maxima [1106]. Maximize [1088]. Maximum [1121, 1124, 1402, 1140, 788]. MAXSAT [1295]. May [196]. MAYA [215]. McFarland [654]. MDPs [59]. MDSP [843]. Meaning [785]. Measure [1198, 791]. Measurement [874, 293, 1072, 846]. Measurement-Based [1072]. Measures [398, 782, 13, 30, 796, 270]. Measuring [736]. Mechanism [699, 83, 1083, 173, 422, 1431, 1051, 1453, 704]. Mechanisms [748, 899, 434, 387]. Media [450]. Medial [1116]. Medians [1085]. Medical [835, 577, 455]. Medium [1022]. Medoid [990]. Medoids [990]. MEDWRAP [923]. Meeting [604]. Meetings [606]. Mel [1225]. Melodic [722]. Melodies [720]. Membership [69]. Memories [2]. Memory [1098, 1096, 870, 491, 865, 998, 1143, 1500, 543, 723]. Memory-Based [723]. Memoryless [653]. Mental [390, 65]. Merchandises [1032]. Merge [539]. Merging [136]. Meshed [1171]. Message [76]. Messages [372]. Messaging [325]. Messor [1174]. Meta [154, 290, 420, 306, 1212]. Meta-knowledge [1212]. Meta-patterns [290]. Metacomputation [535]. Metadata [430, 455]. Metaheuristic [1175, 1295]. Metal [845]. Metamodel [415, 282, 161, 413]. Metamodelling [1035]. Method [487, 164, 751, 1425, 74, 202, 100, 71, 72, 411, 135, 79, 129, 1189, 953, 386, 775, 1147, 453, 385, 328, 558, 977, 233, 73, 249, 565, 1441, 988, 937]. Methodological [316, 300]. Methodologies [910, 626]. Methodology [1213, 851, 858, 888, 1046, 695, 675]. Methods [309, 826, 827, 265, 714, 1221, 93, 19, 813, 75, 28, 980]. Metric [892]. Metrical [317]. Metrics [314, 316]. Microfilm [256]. Microprocessor [1528]. Microscopic [92]. Middle [1320]. Middleware [1450, 325, 1457, 1436]. Mimetic [717]. Min [1090]. Min-Max [1090]. Mind [1206]. Minds [109]. Minimal [1129, 1228]. Minimizing [1121, 1154, 881, 1155]. Minimum [76, 1204, 1286, 124]. Mining [986, 164, 147, 922, 764, 908, 1495, 991, 762, 903, 1042, 901, 1005, 768, 259, 756, 1210, 765, 942, 457, 760, 175, 146, 145]. Mixed [590, 1234]. Mixers [556]. Mixing [1476, 1469]. Mixtures [819]. MKL [979]. MKL-Tree [979]. MLP [135]. MLP-Based [135]. Mo [944]. Mobile [1415, 751, 1049, 1519, 1074, 945, 1538, 403, 1057, 174, 1040, 752, 701, 405, 144, 893, 1457, 482, 158, 1030, 1434, 1436, 929]. Mobility [750]. Modal [1338, 191]. Mode [79]. Model [1071, 801, 1335, 709, 1400, 83, 1034, 420, 9, 364, 1383, 1387, 265, 841, 226, 1536, 1450, 431, 432, 569, 419, 298, 1287, 1432, 1139, 322, 825, 373, 1278, 1331, 887, 902, 114, 1037, 43, 496, 720, 951, 1384, 939, 861, 597, 653, 1442, 509, 141, 739, 692, 1030, 1054, 1264, 1190, 225, 704, 1438, 1455, 866]. Model-Based [1278, 1264]. Model-Checking [1335]. Model-Transformers [692]. Modeling [863, 882, 319, 1271, 1065, 418, 1312, 1357, 1184, 1437, 1076, 1010, 64, 1530, 313, 1377, 693, 886, 766, 847, 134, 1007, 1300, 607, 1073, 299, 433, 1011]. Modelling [395, 1213, 883, 522, 1034, 425, 193, 1417, 1173, 765, 65]. Models [1069, 358, 1217, 1261, 297, 820, 1306, 116, 1203, 125, 1310, 741, 294, 108, 304, 1423, 485, 798, 972, 276, 719, 1150, 1421, 817, 284, 723, 1506, 1422]. Modern [529]. Modes [621, 605]. Modifiable [503]. Modified [100]. Modular [676, 478, 183, 1369, 1408]. Modulated [1107, 1061]. Module [776]. Modules [531]. Modulo [646]. Modulus [1051]. Modulus-Based [1051]. Monadic [1351]. Monitor [595]. Monitoring [890, 403, 157, 1031]. Monitors [505]. monotone [376]. monotonic [1323]. Morphable [530]. Morphological [817, 124]. Morphology [90, 254, 84]. MORPHOSYS [574]. Motif [100]. Motivation [431]. Motivation-Supporting [431]. Move [730]. Moving [922, 961, 142, 956]. Mu [191]. Mu-Calculus [191]. Multi [1019, 263, 1235, 934, 234, 1205, 235, 371, 370, 1495, 1357, 1017, 70, 64, 116, 101, 548, 411, 86, 16, 905, 678, 1544, 545, 685, 680, 557, 853, 112, 1426, 734, 747, 223, 85, 923, 1454, 1453, 1459, 1029]. Multi-adjoint [680]. Multi-agent [1205, 70, 64, 101, 86, 905, 678, 734, 747, 85, 1453, 1029]. Multi-agents [112, 1454]. Multi-Case-Base [16]. Multi-classifier [234, 235]. Multi-comparand [548]. Multi-context [557]. Multi-disciplinary [1017]. Multi-domain [1426]. Multi-expert [223]. Multi-genre [685]. Multi-GHz [853]. Multi-granularity [934]. Multi-layered [116]. Multi-party [371, 370]. Multi-phased [1019]. Multi-relation [923]. Multi-relational [1495]. Multi-resource [1235]. Multi-scale [263, 411]. Multi-search [548]. Multi-sensor [1544]. Multi-stakeholder [1357]. Multi-tasking [545]. Multi-versioning [1459]. Multicast [1150]. Multiclass [67]. Multicontext [569]. Multidimensional [979, 981, 970, 1010, 161, 413, 662, 957, 423, 1011]. Multilayered [783]. Multilevel [1098]. Multilingual [776, 829, 777]. Multimedia [319, 459, 448, 498, 53, 52, 455, 461, 1075]. Multimedia-Oriented [498]. Multiminima [1178]. Multimodal [130, 1519, 659, 273]. Multinomial [1497]. Multiparty [631]. Multiple [372, 81, 4, 1112, 1397, 95, 958, 914, 1060, 80, 1032, 605, 236, 577, 895, 681]. Multiple-Cause [95]. Multiple-Incentive [1032]. Multiple-Valued [681]. Multiplication [517, 519, 378, 386, 393]. Multiplicative [392]. Multiplier [517, 568, 484]. Multiplier-Based [517]. Multiplier-less [484]. Multipliers [530]. Multistage [850]. Multitasking [546]. Multithreading [466]. Multiuser [579]. Multivariate [770]. Multiversion [406]. Mumble [801]. Museum [1520]. Music [707, 709, 711, 1222, 768]. Musical [712, 713, 716, 1225, 718, 719]. MUX [882]. MWM [893]. MXY [255].

N [820]. N-Gram [820]. ‘NAIL’ [1513]. Naive [91, 93, 1003]. Name [1348, 196]. Natural [824, 126, 1513, 777, 787, 1504, 1212]. Natural-Language [1212]. Navigation [434, 1047]. nceptual [944]. Near [1087, 1014]. Near-Linear [1087]. Nearest [171, 921, 1]. Nearest-Neighbor [1]. Need [289, 157]. Neem [1462]. Negation [672]. Negative [1235]. Negotiating [736]. Negotiation [1457, 1456]. Negotiations [1019]. Negotiator [735]. Neighbor [171, 1]. Neighborhoods [1100]. Neighbours [921]. Nest [1187]. Nested [1333, 1190]. Net [1423, 1442]. Nets [1510, 693, 1421]. Network [859, 511, 581, 589, 506, 584, 505, 68, 434, 1223, 1225, 1272, 570, 1153, 390, 119, 1441, 1436]. Networked [1538]. Networking [1533]. Networks [90, 36, 1095, 934, 632, 1397, 1213, 1174, 95, 1077, 1110, 1183, 1126, 1060, 1219, 783, 1000, 89, 1402, 118, 1273, 921, 545, 1144, 930, 1196, 579, 1171, 366]. Neural [90, 668, 1213, 1510, 1215, 1000, 118, 995, 921, 1225, 680, 570, 579, 1403, 119]. Neural-Based [995]. Neuro [1202]. Neuro-symbolic [1202]. Neuromanifold [68]. News [254]. Newspaper [256]. Next [930, 1078, 1457]. NI [584]. NL [776]. NLP [50]. NN [132]. No [196]. Node [589]. Nodes [90]. Noise [844, 176, 802, 884, 819]. Noisy [932, 227]. Non [1479, 376, 1510, 1323, 194, 1117, 310, 322, 627, 107, 232, 989]. Non-committing [627]. Non-distance [989]. Non-functional [310, 322]. Non-independent [1117]. Non-monotone [376]. Non-monotonic [1323]. Non-regular [1510]. Non-representative [232]. Non-safe [194]. Non-Speech [813]. Non-uniform [1479]. Non-verbal [107]. Nonblocking [1128]. Nondeterminism [1322]. Nonterminal [1487]. Nonuniform [1361]. Norma [920]. Norma-System [920]. Normal [1324, 460, 378, 887, 1305, 1353]. Normalizable [1360]. Normalization [1319, 105]. Notification [698]. Noun [1198, 674]. Nouns [793, 790]. Novel [592, 460, 1537, 529, 69]. NP [1255, 1490, 1478]. NP-hard [1490]. NRDR [121]. NTRU [633]. NULL [999]. Nullspace [1375]. Number [1252, 234, 1466, 647, 528, 230, 1157]. Numbers [784]. Numerical [1367, 1366]. Nutritional [1545]. Nyberg [372]. Nyberg-Rueppel [372].

O [533, 955, 1143]. O-efficient [1092]. Object [1088, 288, 303, 941, 293, 280, 415, 168, 616, 325, 419, 901, 304, 320, 285, 1461, 323, 328, 213, 283, 939, 316, 924, 26, 326, 302, 142, 973, 1408, 417, 970]. Object-Based [213]. Object-Extraction-Based [168]. Object-Oriented [941, 280, 325, 419, 901, 285, 328, 283, 316]. Objective [78]. Objects [312, 281, 289, 922, 1123, 1404, 933, 940, 961, 1289, 212, 430, 965, 956, 1533]. Oblivious [389]. Observable [82]. Observation [82]. OCOA [676]. OCR [218, 260]. ODBMS [307]. ODMG [417]. OEM [423]. Offers [37]. Offline [870, 368, 228, 230]. Offtalk [822]. OLAP [412, 162, 953, 1007]. OLAP-XML [1007]. On-Line [1139, 636, 1121, 495, 1155]. Ones [987]. Online [445, 1125, 79, 1151, 230]. Only [212]. Ontological [50]. Ontologies [294, 747]. Ontology [676, 1425, 293, 612, 1426, 915]. OO [121]. OODBMS [121, 282]. Open [676, 1254, 1350, 1357, 1010, 744, 1055, 441, 229, 429, 286]. Operand [867]. Operating [883, 534]. operation [467]. Operations [1099, 173, 136]. Operators [517, 1333, 1287]. Opinion [740]. Opinion-Based [740]. Opportunistic [1248]. Opportunities [504, 1024]. Optical [221, 589]. Optically [1528]. Optimal [1233, 737, 997, 1107, 637, 1115, 218, 84, 955, 71, 378, 849, 383, 1332, 1390, 1467, 1478]. Optimisation [947, 1182, 1209, 675]. Optimised [390]. Optimistic [404, 672]. Optimization [859, 1077, 1178, 1183, 421, 959, 1170, 1004, 1137, 68, 1267, 1189, 139, 752, 19, 416, 1164, 119, 954, 1375, 1186, 246, 1064]. Optimizations [529]. Optimize [416, 1188]. Optimizers [1167]. Optimizing [80, 873, 905]. Optimum [479]. Options [1147]. OQL [284]. Oracle [664, 627]. Oracles [1084]. Oral [805]. Orchestration [693]. Order [1097, 997, 1346, 1371, 185, 886, 385, 1275, 1349, 208, 1441]. Order-Sorted [208]. Ordered [1134]. Ordering [1105]. Orderings [1237]. Organisational [746, 2]. Organisations [745]. Organizational [699, 156, 302, 439]. Organizations [365, 406]. Organized [1174]. Organizing [319]. Oriented [1006, 155, 186, 941, 293, 280, 325, 419, 98, 901, 320, 498, 285, 1146, 328, 283, 939, 316, 26]. Orthogonal [1232]. Out-Linked [977].

Outcomes [449]. Outlier [1000, 149]. Output [864, 862]. Outsourced [365]. Outsourcing [1054]. Overcome [830]. Overfitting [932]. Overhead [873].

P2P [379]. PA [857]. PA-ZSA [857]. PAC [1505]. Paced [447]. Packet [511, 1113, 589, 486]. PACT [552]. PACT-XPP [552]. Padding [634]. Paddings [633]. Page [892, 1034, 258, 679, 978, 457]. Pages [258, 977, 891]. Pairing [642]. Pairing-Based [642]. Pairs [996]. Pairwise [655]. Paper [267, 268]. Paradigm [897]. Paradoxes [55]. Parallel [539, 491, 955, 970, 1130, 1013, 576, 307, 1272, 1185, 597, 284, 1030, 62]. Parallel/Reconfigurable [597]. Parallelism [412, 1323, 970, 1281]. Parallelizable [386]. Parameter [811, 532]. Parameterizable [540]. Parameterized [531]. Parameters [1500]. Parametric [190]. PARBIT [483]. Paribus [1304]. ParleE [1400]. Parma [1372]. Parser [773, 673, 837, 124]. Parsing [709, 778, 1507]. Part [772, 786]. Part-of-Speech [772, 786]. Partial [1094, 990, 658, 483, 173, 1337, 1259, 1307]. Partial-Domain [658]. Partially [494, 495, 1134, 1305, 747, 82]. Partially-Ordered [1134]. Particle [1209]. Particularly [1340]. Partition [1487, 962]. Partition-Based [962]. Partitioning [552, 492, 488]. Partitions [1472, 1318, 909]. PartJoin [1012]. Parts [258]. party [371, 370]. Pass [862]. Passing [1347]. Password [644, 366]. Password-Authenticated [644]. Path [1095, 525, 924, 140]. Paths [800, 1118, 1171, 402]. PATI [97]. Pattern [757, 922, 914, 754, 967, 79, 768, 759, 47, 760, 152]. Patterns [793, 711, 291, 764, 770, 39, 426, 998, 1005, 145, 662, 719, 756, 175, 290]. Payment [1057]. PC [908]. PCA [132].

PCFG [1487]. PCS [930]. PDA [1519]. PDA-Based [1519]. Peaceably [1291]. Peer [1110, 729, 732]. Peer-to-Peer [1110, 732]. Pen [229]. Penalized [1203]. People [1216, 836]. Perception [250, 83, 720, 723]. Perceptive [1526]. Perceptual [1462, 1525]. Percolation [1471]. Perfect [656]. Perform [780]. Performance [412, 242, 1071, 563, 1006, 512, 556, 554, 522, 1239, 1067, 842, 1065, 1074, 716, 13, 1423, 139, 1046, 44, 843, 547, 561, 871, 1075, 1422]. Performance-Oriented [1006]. Performances [1412]. Permutation [1173, 390]. Personal [37, 1531]. Personalization [405]. Personalized [1027, 49, 429, 457]. Personalizing [1521]. Perspective [707, 1006, 410, 570]. Perspectives [1058]. Pervasive [925]. Pessimistic [672]. Petri [194, 1423, 693, 1421, 1442]. PGP [362]. PGV [640]. Phase [1076, 904, 723, 484]. Phased [71, 72, 1019]. pheromone [1169]. Phoneme [812]. Phonological [816]. Phrase [775, 1393]. Phrase-Based [1393]. Physical [670]. Pilot [120]. Pin&Play [1533]. Pins [1533]. Pipeline [1377]. Pipelined [497]. Pixel [880]. PKI [365]. Place [612]. Placement [1216, 490, 1376, 488, 598]. Placement/Routing [598]. Plaintext [648]. Plan [1409, 3, 1400, 15, 823]. Plan-Recognition [1409]. Planar [1088, 1341, 1144]. Plane [1104, 1158]. Planning [697, 1066, 1409, 3, 1269, 831, 1234, 122, 20, 102, 1401, 85, 665, 606, 87]. PlantCare [1541]. Platform [184, 1462, 952, 1432, 585, 542, 558, 818]. Platforms [533, 504, 1035]. Playing [1194]. PLM} [703]. Pluggable [279]. Plus [668]. Poetry [7]. Point [1088, 531, 1082, 76, 518, 1337, 528, 386, 529, 1368]. Pointer [1373]. Points [1494, 1370]. Points-To [1370]. Policies [699, 289, 1268, 1436]. Policy [156, 1455]. POLIS [533]. Poly [484]. Poly-phase [484]. Polygonalization [1193]. Polygons [1099]. Polyhedra [1372]. Polymatroid [1064]. Polymorphs [1120]. Polynomial [1130, 1389, 1359, 1496]. Polynomial-Time [1130]. Population [1170]. Portable [472]. Portal [1443]. Portals [49]. Portion [252]. PoS [774]. PoS-Tagged [774]. Positioning [1196]. Positive [1491, 1060, 1346, 1332, 1503]. Possibilistic [948]. Possible [1330]. Possibly [1372]. Postman [1090]. Potential [327, 610]. Power [848, 856, 497, 844, 857, 885, 851, 841, 873, 854, 880, 555, 1483, 847, 884, 843, 496, 872, 500, 875, 845, 871, 866, 507]. Power-Aware [857]. Power-Driven [854]. Power-Grid-Induced [844]. Power/Low [884]. PR [144]. PR-Tree [144]. Practical [554, 1541, 1040, 915]. Practice [1123, 336, 302]. Practices [452]. Pragmatics [183]. Pre [1102, 1424]. Pre-assigned [1102]. Pre-dispatching [1424]. Precise [1365]. Precision [1374]. Predator [1209]. Predicate [146]. Predicates [980]. Predict [1510]. Predicting [1034, 176]. Prediction [826, 827, 14, 112, 877]. Predictions [15]. Preemption [1154]. Preference [1224, 1042]. Preference-Based [1224]. PrefixSpan [100]. Preliminary [1234, 72]. Presence [1526]. Presentation [434, 893]. Preservation [1459]. Preserve [1482]. Preserving [1433, 1384]. Previous [990]. Prey [1209]. Price [921]. Pricing [1147, 1395]. Primal [1127]. Prime [646]. Principle [1333, 1204]. Principles [407]. Printed [217, 219, 222]. Priority [1415, 1416, 1417]. Privacy [396, 392, 1049, 367, 1530, 1535, 368]. Private [632, 732, 357, 1256]. Privilege [365, 134]. Proactive [1543]. Probabilistic [197, 1479, 1175, 147, 1492, 885, 1246, 1131, 825, 1507]. Probabilistic-Time [1479].

Probabilities [987, 1329]. Probability [1335, 1334, 246]. Probing [1241]. Problem [1090, 1232, 1398, 1093, 1262, 1175, 1103, 171, 1165, 1127, 572, 1402, 86, 647, 1140, 744, 1146, 1228, 1272, 1440, 1208, 1149, 1185, 326, 1164, 675, 1257, 1161, 29, 56, 638, 240]. Problem-Solving [744]. Problematic [830]. Problems [1341, 1095, 1282, 1107, 1081, 1279, 1122, 664, 1170, 1306, 715, 1137, 1267, 1189, 1139, 667, 1179, 1173, 1311, 1172, 1295, 1296, 675, 1304, 1284, 1294]. Procedures [1057, 1322, 62]. Process [1070, 691, 936, 595, 286, 910, 694, 114, 721, 765, 1454, 703]. Processes [35, 449, 199, 613, 1049, 1443, 332, 695, 1062, 1053, 1167]. Processing [538, 1008, 669, 960, 162, 885, 514, 917, 1033, 284, 1504, 954, 223, 1292, 968, 140]. Processor [540, 548, 576, 585, 594, 560, 523, 521, 536]. Processors [865, 1137, 557, 520, 871]. Product [5, 1060, 315, 324]. Product-Lines [324]. Production [35, 663]. Products [646, 1119, 1395]. Profile [925, 607]. Profiles [98, 994]. Profiling [870]. Profit [1083]. Program [1415, 1416, 448, 1417, 1356, 872]. Programmable [508, 478, 488, 473, 871]. Programming [1091, 1232, 1270, 1176, 445, 1217, 1234, 306, 205, 111, 1326, 716, 313, 906, 1263, 750, 1302, 1268, 509, 115, 1264, 146, 1301]. Programs [1388, 1390, 416, 188, 1368, 1365]. Progressive [318, 158]. Project [449, 1079, 468, 20, 834, 779, 805]. Projection [1215]. Projects [510]. Prolog [502]. Prolog-Based [502]. Prolongation [231]. Promising [1231]. Promoting [938]. Proof [1352, 658, 1326, 185, 1347, 1478, 211, 626]. Proofnets [1325]. Proofs [269, 1354, 630, 376, 1350, 339, 338, 627, 1353]. Propagation [882, 1303, 887, 886, 328, 861, 1285, 78, 56].

Properties [1472, 1386, 13, 1287, 1385, 63, 161, 413, 1280, 1375]. Property [1282, 1051]. Proposal [130, 1022]. Propositional [1352, 1354, 1320]. Prosodic [822]. Protect [1056]. Protecting [751, 357, 368]. Protection [1049, 1051]. Protein [1164]. Protocol [629, 1052, 371, 389, 126, 388, 1055]. Protocols [1380, 1379, 912, 370]. Prototyping [512, 468, 855, 585, 467]. Provably [392, 625, 621]. Provider [702]. Proving [1375]. PRS [65]. PRS-Like [65]. Pruning [932, 971, 68]. Pseudo [233]. Pseudoknots [1002]. Pseudorandom [648, 645]. Public [691, 380, 377, 1050, 1534, 381]. Public-Key [377]. Publishing [266]. Pull [1150]. Pure [1280]. Purpose [585]. Purposes [40]. Pursuit [1215]. Pushdown [1336]. Pyramid [251].

Q [70, 409, 905]. Q-Learning [70, 905]. Q-Tree [409]. QAM [476]. QAP [1180]. QPSK [475]. Quadratic [1238]. Quadrature [556]. Qualitative [1253, 1410]. Quality [892, 567, 440, 316, 46]. Quality-Criteria [46]. Quantification [185]. Quantified [1255, 1244]. Quantifying [367]. Quantitative [563]. Quantization [880]. Quantum [1477]. Quartz [840]. Queens [1291]. Queries [948, 960, 1466, 408, 969, 955, 192, 367, 403, 762, 961, 917, 949, 953, 968]. Query [412, 1012, 776, 984, 171, 162, 959, 970, 257, 1137, 909, 947, 139, 422, 950, 1467, 416, 1033, 919, 284, 1211, 954, 138]. Querying [163, 285, 787]. QueryLens} [1532]. Question [456, 1212]. Question-Answering [1212]. Queues [1061]. Quo [982].

R [407, 54, 408]. R-Tree [407]. R-Trees [408]. R-UNSEARCHMO [54]. RADAR [1199]. Radiation [1107]. Radio [1102, 1126]. Railway [825, 1290]. Raison [1461]. RAM [869]. Ramification [56]. Ramp [886]. Random [1470, 1473, 657, 639, 1250, 627, 1192, 1280, 1245]. Randomization [1239, 117, 1297]. Randomized [1117, 1472, 1137]. Randomness [1478]. Randomness-Optimal [1478]. Range [1086, 955, 961, 909, 953, 1249]. Range-Based [1249]. Ranking [169, 1231]. Rapid [512, 489, 468]. Rapidly [1469]. Rarely [1480]. Rarity [1111]. Rate [808, 757, 506, 228]. RBAC [1455]. RC [863]. RC4 [639]. RDR [74, 73]. Re [544, 145, 510, 547]. Re-configurable [544, 547]. Re-mining [145]. Reachability [214]. Reaching [1256]. Reactive [677, 560]. Read [258, 212]. Read-Only [212]. Reading [224, 799]. Real [536, 1014, 1458, 713, 1076, 356, 472, 1136, 132, 174, 591, 797, 878, 438, 1186, 175, 1459, 577]. Real-Time [536, 1014, 1458, 356, 1136, 132, 174, 591, 797, 438, 175, 1459, 577]. Real-World [1186]. Realistic [371, 1416]. Reality [1519]. Realizability [1320]. Realization [484, 866]. Realized [1018]. Realizing [107]. Reasoning [35, 36, 683, 9, 1220, 933, 16, 63, 680, 114, 43, 22, 25, 102, 1253, 1264]. Receipt [369]. Receipt-Free [369]. Receiver [513]. Recognise [90]. Recognition [221, 1409, 808, 226, 15, 105, 679, 219, 79, 823, 485, 802, 798, 812, 797, 719, 236, 817, 245, 228, 220, 799, 225, 818, 265]. Recognizable [1360]. Recognizing [91]. Recommendation [42, 1043, 738, 1210, 21, 457]. Recommendations [5, 1044, 938]. Recommender [34, 952, 1224, 1042, 18, 29]. Recommenders [1214]. ReConfigME [534]. Reconfigurability [527]. Reconfigurable [563, 575, 564, 522, 471, 571, 494, 537, 595, 495, 514, 483, 584, 492, 568, 546, 474, 498, 591, 542, 594, 570, 480, 597, 558, 532, 509, 534, 521].

Reconfiguration [481, 535, 482, 493]. Reconstruction [12]. Recorded [779]. Recordings [712]. Recovering [1204, 1243]. Recovery [894]. Rectangular [263]. Recurrent [1510, 579]. Recursion [1348]. Recursive [1322, 1361]. Red [1202]. Reduce [932, 29]. Reduced [1231, 393, 62]. Reducing [1310, 135, 877]. Reduction [856, 1214, 176, 1149, 867]. Reductions [1482]. Redundant [189]. References [354, 212]. Refinement [1397, 54, 1383]. Reflecting [716, 977]. Reflective [303, 734]. REFLIX [560]. Refractory [41]. Refusal [195]. Refutations [1354]. Region [883]. Register [867, 866]. Registration [591]. Registries [690]. Regression [914]. Regular [1488, 1466, 1491, 1474, 1508, 1510]. Reinforcement [1191, 61, 71, 72, 1395]. Rejection [801]. Rejuvenation [1073]. Related [1352]. Relating [216]. relation [923]. Relational [147, 162, 1404, 763, 1433, 1278, 1331, 1366, 660, 787, 138, 1495, 947]. Relations [1351, 193, 11, 13, 1360, 259]. Relationships [700]. Relaxation [1276, 1236]. Relaxed [174]. Relevant [258, 55]. Reliability [1221, 1073]. Reliable [489]. Rememberer [1520]. Remote [438]. Removing [189, 55]. Rendering [583, 153]. Reorder [868]. Reorganization [329]. Repair [1279]. Repetitive [956, 565]. Replaceability [308]. Replanning [677]. Replication [1109, 955, 356, 322]. Replicator [1000]. Report [1234]. Reports [1227]. Repository [116]. Representation [1409, 201, 711, 958, 1376, 51, 1002, 876, 864, 18, 1208, 393, 142]. Representations [1398, 715, 761, 136]. representative [232]. Represented [255, 1304]. Representing [343, 345, 935, 14, 1362, 1371, 161, 413, 1518, 57]. Reputation [1039]. Request [697]. Requirements [37, 262, 305, 13]. Requirements-Driven [262]. Requires [1148]. Rerouting [1273]. Research [1357, 1017, 1437, 271, 585, 52, 1058].

Reservation [1273]. Residual [1492]. Resilient [651]. Resistance [386]. Resolution [1352, 1354, 1112, 683, 1407, 173, 1250]. Resolving [97]. Resonant [850]. Resource [1452, 1327, 908, 1326, 1448, 671, 1240, 1235]. Resource-Based [1452]. Resources [271, 805]. Respiration [716]. Response [487, 698]. Responses [814]. Responsive [1526, 1528, 323]. Restart [1268]. Restarts [1155]. Restricted [1139, 681]. Restrictions [1354]. Result [422]. Result-Visualization [422]. Results [1388, 1060, 567, 72]. Retaining [435]. Retargetable [565]. Retraction [661]. Retrieval [726, 347, 5, 255, 1205, 10, 11, 964, 168, 1221, 258, 257, 399, 1222, 1429, 907, 17, 972, 24, 789, 893, 142, 458, 973, 966, 965, 308, 1408, 1427]. Retrieving [752]. Reuse [352, 1388, 425, 46, 867, 292]. Revealing [1256]. Revenue [1018]. Reversals [1101]. Reverse [265, 268]. Review [295]. Revision [1397]. Revisited [1321, 160, 1333, 1260, 239]. Revisiting [407, 621, 210, 236]. Revocation [624]. Revokable [368]. Revolutionary [438]. Reward [71, 72]. Rewriting [421, 1318]. RFU [561]. RHiNET [584]. RHiNET/NI [584]. Rhythm [1200]. Ride [1139]. Right [1206]. Rigid [1490, 1494, 1502]. Rijndael [542]. Ring [649, 650]. Ripple [122]. RNA [1002]. RNS [858, 513]. RNS-Based [858]. Road [1095]. RoboCup [1415]. Robot [133, 84]. Robotic [81, 676]. Robots [1415, 1177]. Robust [856, 685, 802, 1207, 1272, 106, 1030]. Role [297, 777, 363, 326, 994]. Role-Based [363]. Roles [286, 327, 940, 746]. Rough [75]. Round [631]. Rounding [1117]. Routability [489]. Route [752]. Router [1095]. Routes [1030]. Routing [491, 471, 1165, 1183, 86, 488, 598, 501, 1208, 1172, 1084, 845]. Row [1282, 1261]. RPNI} [1488]. RSA [634, 628, 635, 366]. Rueppel [372]. rUID [917]. Rule [773, 289, 762, 1450, 156, 25, 75, 120, 457, 1454, 703, 146]. Rule-Based [1450, 156, 25, 120]. Rules [1486, 984, 986, 932, 235, 912, 935, 763, 421, 903, 761, 122, 947, 75, 915]. Run [303, 495, 908, 483, 481, 520, 493]. Run-Time [303, 495, 908, 483, 481, 520, 493]. Running [447]. Runs [1268]. Runtime [315]. Russian [821, 1248].

S [944]. SACReD [894, 1029]. Safe [646, 194]. Safe-Prime [646]. Safety [1386]. Salesman [1175, 1081]. Saliency [83]. Salient [780]. Samples [1076]. Sampling [1468, 1015, 555]. SAN [908]. SAN-connected [908]. Sancta [1174]. SAT [856, 1309, 1282, 1308, 664, 1242, 1246, 1297, 1280, 1281, 62]. SAT-Based [856]. Satisfaction [697, 1342, 1251, 1481, 1254, 1311, 1296, 1256, 1244]. Satisfactionwith [1275]. Satisfiability [1131, 501, 1301]. Saving [1147, 507]. Saving-Asian [1147]. Savings [497]. SC2000 [360]. SCAE [1041]. Scalable [949, 558]. Scalar [393]. Scalars [393]. Scale [449, 1048, 1079, 89, 732, 957, 243, 263, 411]. Scaleable [954]. Scaled [1225]. Scaling [1246, 854, 1280]. Scanning [1096]. Scarcely [1171]. Scenarios [1416]. Scene [253]. SCEP [387]. Scheduling [1269, 1162, 1279, 1303, 1130, 1241, 1134, 678, 454, 1151, 878, 1186, 1064]. Schema [916, 418, 941, 1433, 285, 1037]. Scheme [382, 623, 1130, 873, 930, 579, 374, 142, 1157]. Schemes [630, 658, 634, 656, 626, 375]. Schenkerian [715]. Schizophrenia [326]. Schnorr [630]. School [449, 601]. Schoolgirl [1262]. Science [1227, 1079]. SCIL [1091]. Scouts [1187]. Script [217]. Scripting [1452, 109]. Sea [363, 553]. Sea-of-Data [363]. Seal [223]. Sealed [369]. Sealed-Bid [369]. Search [1195, 856, 1487, 1299, 757, 154, 990, 1239, 1258, 664, 593, 1326, 1246, 89, 409, 1032, 1297, 1143, 1248, 185, 1169, 1293, 501, 812, 1289, 26, 260, 138, 166, 493, 62, 148, 1301, 548]. Searchability [299]. Searching [1086, 252, 975]. Second [541, 385]. Second-Order [385]. Seconds [1262]. Secrecy [1380]. Secret [646, 637, 655]. Secure [1041, 382, 392, 631, 202, 625, 621, 1050, 1030, 1256]. Securing [1435, 1382]. Security [630, 1380, 199, 396, 629, 1379, 658, 648, 1039, 628, 621, 647, 387, 104, 704, 1381, 177]. See [759]. Segment [795]. Segmentation [250, 811, 267, 809, 1192, 722, 243, 227]. Segmenting [758]. Selecting [16]. Selection [160, 714, 837, 125, 46, 28]. Selections [948]. Selective [884]. Self [351, 1174, 859, 447, 356, 106, 535, 911, 988]. Self-Adaptive [106]. Self-Heating [859]. Self-maintainable [911]. Self-Organized [1174]. Self-Paced [447]. Self-Reconfiguration [535]. Self-Replication [356]. Self-Service [351]. Self-Tuning [988]. Semantic [781, 1198, 683, 187, 783, 149, 116, 287, 947, 136, 416, 972, 401, 259, 144, 142, 1512, 1054, 1048, 275, 1414, 686, 1428, 1399]. Semantic-Based [142]. Semantics [248, 690, 1405, 1330, 900, 1325, 1385, 1329]. Semi [1340, 793]. Semi-automatic [793]. Semi-terms [1340]. Semirandom [1475]. Semirings [1324]. Semistructured [150, 423]. Sense [1128, 792, 788]. Sensing [1545, 1542]. Sensitive [1519]. Sensor [498, 587, 1544]. Sensors [1529]. Sentence [251, 775]. Separating [627]. Separation [657, 239]. Sequence [100, 1292]. Sequences [767, 769]. Sequencing [524]. Sequential [1486, 527, 525, 998, 82]. Serbian [803]. Serial [487, 378]. Series [1213, 176, 1044]. Server [1069, 1531]. Service [351, 396, 922, 751, 690, 186, 1031, 1136, 1078, 1457, 1431, 1436, 704, 696, 693].

Service-Oriented [186]. Services [697, 1066, 899, 1067, 1414, 157, 279, 1025, 688, 695, 919, 701, 1442, 1439, 280, 702, 700, 689, 158]. Session [702, 1442]. Set [1127, 1346, 421, 525, 98, 847, 1185, 1182, 75]. Set-Oriented [98]. Set-up [847]. Sets [1474, 163, 559, 228]. Setup [1292]. SEU [526]. SHA [361, 523]. SHA-1 [361]. SHA-256 [523]. SHA-512 [361]. Shafer [670]. Shallow [1507]. Shape [973]. Shaped [804]. Shapes [1082, 1193]. Shared [646, 747, 1521]. Sharing [637, 655, 212, 379]. Shop [1162, 1186]. ShopFloor [1269]. Short [1480]. Shortest [1095, 1118]. Shuffles [639]. Shuffling [390]. SiBIR [651]. Side [386, 385]. Side-Channel [386]. SiGe [507]. Sign [781]. Signal [538, 885]. Signals [1060]. Signature [372, 629, 658, 1033, 374, 723, 626]. Signature-Based [629]. Signatures [649, 651, 657, 1040, 373]. Signed [375]. Signer [651]. Signer-Base [651]. Significance [775]. Signing [1052]. Similarities [327]. Similarity [997, 6, 255, 1111, 1120, 1184, 409, 19, 791, 22, 44, 401, 28, 962, 973, 30, 966, 166, 308, 999]. Similarity-Based [166]. Similarity-Influenced [28]. Simple [1340, 1165, 1124, 167, 105, 1227, 1147, 243, 1505, 152, 1158]. Simplification [1087]. Simplified [1097]. Simulated [1412]. Simulation [512, 526, 480, 1187, 467, 1290, 195, 66, 895]. Simulation-Based [526]. Simulator [88, 1540]. Simultaneous [393]. Single [1155]. Sink [1152]. Site [395]. Sites [780]. Size [1100]. Sized [909, 390]. Sizing [317]. Skeletons [1047]. Skew [237]. SLA [1054]. SLAAC1 [546]. SLAAC1-V [546]. Slack [857]. Slightly [1479]. Slovenian [817, 794]. Small [1398, 517, 1474, 1022, 635, 359]. Smart [544, 1227, 587, 429, 438, 201]. Smartcard [202, 928]. smartFIX} [262].

SmartMoveX [1539]. Smoother [537].

Smoothing [1246]. SOAP [1434].

SOAP-Based [1434]. SoC [581]. Social [112, 1534, 66]. Socially [88, 1530, 114]. Socially-Based [1530]. Socially-Constructed [114]. Societies [1417]. Socionics [1417]. Soft [1283, 1136, 1247, 1304]. SoftTOTEM [576]. Software [1510, 1358, 1220, 10, 39, 1074, 1443, 187, 1135, 357, 1513, 360, 1204, 731, 301, 316, 211, 1073, 308]. Soldier [834]. Solution [1519, 772, 40]. Solutions [1252, 1010, 1061]. Solve [675, 124]. Solved [1305]. Solver [1266, 1278, 1301]. Solving [1262, 1175, 1336, 1081, 171, 969, 1279, 1411, 192, 59, 1189, 744, 1272, 1289, 1247, 1304, 56, 1298]. Some [1491, 1494, 715, 1010, 621, 387, 815, 1152, 1058, 461]. SOPC [587]. SOPC-based [587]. Sort [27]. Sorted [208]. Sorting [1101, 1148, 1151]. SOS [183]. Sounds [714]. Source [882, 776, 963, 470]. Source-Coupled [882]. Sources [742, 287, 923]. Space [996, 1113, 1476, 59, 1189, 636, 1169, 662, 573, 1288]. Space-Bounded [636]. Space-Complexity [1288]. Spaces [408, 409, 1482, 397, 962, 989]. Spanish [830, 787]. Spanners [1104]. Spanning [1140]. Sparse [1475, 660]. Sparsity [1214]. Spatial [1411, 901, 1253, 980, 141, 6]. Spatio [410, 142]. Spatio-Temporal [410, 142]. Speaker [135]. Speakers [91]. Speaking [1229]. Special [1484, 764]. Specialization [1248]. Specialized [511]. Speciated [118]. Specific [551, 471, 1425, 315, 588, 478]. Specification [184, 414, 297, 1055, 951, 188]. Specifications [181, 1445, 187, 209, 204]. Specifying [201]. Spectral [1061]. Speech [811, 807, 808, 1200, 809, 1519, 772, 659, 786, 795, 485, 802, 812, 797, 779, 819, 813, 817, 803, 799, 828, 806, 818, 804]. Speech/Non [813]. Speech/Non-Speech [813]. Speed [486, 505, 586, 135, 555, 359, 493]. Speeding [1373, 1123, 976]. Speedup [971]. sphere [969]. Spi [200, 1360]. Spin [1476]. Spiral [710]. Spline [170]. Split [539]. Splittable [1093]. Spoken [835, 798, 1514]. Spotting [810, 258]. SPP [1415, 1416, 1417]. SPP-1077 [1417]. SPP-1083 [1416]. SPP-1125 [1415]. Springy [816]. Spurious [881]. SQL [24]. Squares [1232]. SRAM [526]. SRAM-based [526]. st} [1105]. st}-Ordering [1105]. Stability [1367]. Stable [1127]. Stack [940]. Stack-Based [940]. Stacking [1213]. stakeholder [1357]. STaM [1201]. Stamping [383]. Standard [852, 475, 317, 413, 359, 516, 417]. Standard-Cell [852]. Standardized [459, 1057]. Standards [1045, 441]. Starting [1121]. State [668, 1492, 1401, 1264]. State-Based [1401]. Statements [1304]. States [1387, 65]. Static [1358, 1371, 1367, 1385, 1145, 1356]. Statistical [892, 786, 96, 129, 823, 813, 216, 1393]. Steganography [625]. Steiner [1228, 1149, 1146]. Stemmer [773]. Step [318]. Stepwise [1240]. Stepwise-Constant [1240]. Stereotype [29]. Stereotypes [1047]. Stitch [1452]. Stochastic [1070, 1063, 1495, 1497, 1423, 1501, 225, 62]. Stock [904, 905]. Storage [1012, 162, 1330, 139, 853, 957]. Storing [288]. Story [1218]. Strain [587]. Strategies [1019, 830, 1344, 16, 96, 797, 1182]. Strategy [118, 1456]. Stratified [1405]. Stream [1464, 652, 1111, 465, 653]. Streams [1113, 426, 968]. Stress [794]. Strictly [1370]. Strictly-Typed [1370]. String [206]. Strings [1490, 1502]. Striving [1014]. Stroke [230]. Stroke-Number-Characteristics [230]. Strong [636, 55]. Strongly [1179, 1360, 1384]. Structural [314, 1243, 140]. Structure [435, 707, 979, 1199, 712, 715, 917, 1002, 1204, 1223, 864, 620, 862, 119, 249]. Structured [420, 1429, 77, 1281, 1504]. Structures [1092, 178, 981, 364, 1337, 1371, 1361]. Structuring [324]. Student [1227, 510, 103, 433]. Study [412, 1270, 1162, 654, 1522, 1165, 908, 450, 1045, 126, 96, 910, 798, 1247, 240, 1434, 871, 33, 148]. Studying [1311]. Stuttering [1333]. Style [324, 1347, 1518]. Subcircuits [527, 525]. Subclasses [1488]. Subcube [1472]. Subgraph [572]. Subgraphs [1475]. Subject [839]. Sublinear [1143]. Submicronic [861]. Submitting [179]. Submodularity [1465]. Subproblems [1231]. Subsumers [1398]. Subsystems [533, 471]. Success [439]. Successive [512]. Sufficient [1496]. Suffix [1141]. Sum [1236]. Superimposed [1112]. Supersingular [641]. Supervectors [1225]. Supervised [1225, 1403]. Supply [854, 115]. Support [611, 810, 358, 351, 609, 1446, 912, 914, 186, 588, 970, 1444, 399, 1424, 897, 320, 617, 1545, 957, 1210, 456, 292, 47, 455]. Supporting [745, 1524, 420, 306, 729, 282, 286, 1460, 431, 605, 671, 902, 454, 46, 452, 509, 138, 1439]. Supports [32]. Surfaces [1538]. Survey [1027]. Suspensible [702]. Svejk [834]. Swarm [1209]. Switch [479, 477]. Switching [874, 881]. Symbolic [1091, 1202]. Symbols [353, 225]. Symmetric [578, 1244]. Symmetries [1261]. Symmetry [1258, 1310, 1259, 1307, 1291, 1260]. Synchronization [858]. Synchronizing [1519]. Synchronous [617, 884, 653]. Synergetic [327]. Syntactic [1197, 1331, 136]. Synthesis [807, 808, 554, 659, 478, 855, 854, 470, 795, 803, 182, 806]. Synthesizing [553]. Synthetic [876]. System [435, 272, 221, 1409, 830, 121, 575, 776, 564, 1446, 155, 522, 380, 179, 1009, 1392, 1180, 1379, 1458, 171, 262, 1205, 1165, 1039, 38, 1202, 595, 436, 908, 1444, 964, 1537, 1326, 580, 80, 1072, 583, 101, 949, 904, 738, 1042, 816, 1222, 546, 1535, 172, 135, 304, 1447, 437, 320, 1038, 1544, 1430, 528, 685, 1225, 273, 1527, 1046, 797, 453, 698, 1192, 718, 920, 134, 44, 112, 682, 368, 441, 236, 1185, 787, 45, 1196, 803, 326, 1154, 438, 374, 1161, 966, 292, 103, 159, 261, 120, 223, 965, 299, 220, 47, 1408, 534, 104, 153, 929, 577]. System [87, 946]. System-Level [564, 522]. System-of-Systems [1444]. System-on-Chip [575]. System-on-Chip-Based [595]. SystemC [566]. Systems [190, 34, 1069, 81, 1352, 870, 309, 1397, 234, 466, 831, 414, 832, 235, 912, 1476, 410, 305, 1168, 935, 1224, 858, 1443, 946, 1357, 1444, 483, 659, 888, 64, 616, 492, 325, 906, 156, 1541, 1238, 905, 304, 1448, 1423, 1424, 210, 1204, 727, 829, 608, 1440, 1023, 1421, 957, 52, 878, 1478, 113, 482, 29, 1514, 982, 318, 138, 439, 1264, 1036, 1436]. Systolic [377].

Table [248, 247, 246, 249]. Tableaux [1327]. Tables [244]. Tachograph [272]. Tackling [664]. Tag [1326]. Tag-Frame [1326]. Tagged [774]. Tagging [772, 786, 1528]. Tailed [1076]. Tailorable [279]. Tailoring [615]. Takes [1402, 1152]. Talking [659]. Talmudic [450]. Tamil [221]. Target [125, 525]. Targeted [1044]. Task [1312, 70, 825, 1163, 1190]. tasking [545]. Tasks [1130, 1424, 1025, 1223]. Taxonomic [11]. Taxonomy [473]. TDR [491]. TEA [784]. Teaching [451, 431, 432, 816, 448, 510, 458]. Teaching-Learning [431, 432]. Team [155, 897, 320]. Teams [1415]. Technical [671]. Technique [801, 997, 497, 1446, 592, 981, 860, 143, 470, 846, 69, 308]. Techniques [808, 870, 669, 1522, 1279, 926, 286, 714, 459, 1318, 448, 1430, 22, 1149, 27, 1434, 2, 785, 760, 507, 461]. Technologies [600, 159, 177]. Technology [835, 844, 506, 869, 432, 1428, 481, 302, 1427, 704].

Telecom [689]. Temperature [481]. Templates [264]. Temporal [934, 552, 922, 958, 1234, 410, 1407, 941, 749, 426, 192, 14, 1005, 1385, 950, 1421, 1247, 142, 204, 1304, 1413]. TEMPPLET [1425]. Term [257, 127, 791]. Term-Based [257]. Terminals [893]. Termination [1388, 1390, 1368]. Terms [92, 1350, 1340]. Terrain [1107, 1177]. Terrain-Covering [1177]. Test [1103, 1517, 1445, 1298]. Test-Data [1298]. Testable [1466, 1503, 1501]. Testbed [1079]. Tested [28]. Testing [312, 1112, 527, 1465, 196]. Tests [1497, 311]. Text [767, 1197, 248, 669, 253, 254, 95, 154, 764, 971, 93, 219, 96, 823, 600, 222, 261, 239, 1, 400, 784]. Text/Graphics [239]. Textual [741, 766]. Their [578, 531, 1482, 63, 1142, 977, 103]. Theorem [1479]. Theoretic [627, 56]. Theoretical [1347, 249]. Theories [1328]. Theory [1323, 302, 196]. Therapy [1107]. There [610, 1206, 759]. Thesauri [789]. Things [1138]. Think [1531]. Third [157]. Thoughts [1058]. Three [1106, 586]. Three-Dimensional [1106]. Threshold [578, 649, 145, 644, 752, 645, 1190]. Thresholding [96]. Tickets [928, 224]. Tides [1202]. Time [1087, 536, 395, 1479, 1341, 1213, 1014, 303, 1458, 713, 1476, 1165, 1121, 1015, 495, 908, 483, 356, 1130, 472, 157, 1133, 1136, 549, 1389, 132, 383, 174, 481, 864, 591, 797, 543, 1172, 878, 1152, 520, 438, 723, 1413, 175, 1459, 493, 577, 1158, 1496, 1155]. Time-Expanded [1133]. Time-Interval [1015]. Time-Memory [543]. Time-Stamping [383]. Timed [1421, 1422]. Times [176, 1133, 1292]. Timeseries [758]. Timestamp [404]. Timetabling [1161]. Timing [883, 859, 857, 1377]. TLS [628]. Together [939]. Tokenization [772]. Tonal [715, 720]. Tool [435, 242, 184, 291, 1520, 1357, 459, 927, 1169, 1512, 499, 784]. Tool-Assisted [184]. Toolbox [1509]. Toolkit [902]. Tools [1069, 793, 901, 1027].

Top [226, 953, 30]. Top- [953]. Top-Down [226, 30]. Topic [95, 166]. Topological [82]. Total [1155]. TOTEM [576]. Tourism [1023, 1036]. Tourists [730]. TPIE [1092]. Trace [647]. Traceability [375]. Traces [1387]. Tracing [356, 1050]. Tracker [1539]. Tracking [536, 713, 1519, 132, 1527, 175]. Tractate [450]. Tradeoff [549, 543]. Tradeoffs [1313]. Trading [1217, 1334, 904, 905, 1055, 1313]. Traffic [1076, 224]. Trail [952]. Trail-Enabled [952]. TrailTRECer [952]. Training [971, 1203, 1403, 232, 228, 439]. Traitor [1050]. Trajectories [81, 142]. Trajectory [1302]. Transaction [1450, 454, 988]. Transactional [1039, 280, 896, 1013]. Transactions [281, 896, 1056, 174, 701, 929, 894, 1029]. Transcription [805]. Transfer [1362]. Transform [539, 515, 486, 871]. Transformation [1353]. Transformations [420, 298, 470, 63, 1142, 211]. Transformed [918]. Transformers [692]. Transforming [199]. TRANSIMS [1095]. Transistor [855, 862]. Transit [1133]. Transition [864, 302]. Transitions [209]. Transitive [559]. Transitive-Closure [559]. Translate [296]. Translating [1088, 918]. Translation [1221, 125, 1257, 1393]. Transmission [883, 873]. Transparent [617]. Transport [581]. Travel [37, 45]. Traveling [1175, 1081]. Travelling [1343]. Traversal [383]. Traversals [1096]. Traversing [1096]. Treatment [1244]. Tree [1098, 407, 974, 673, 364, 1495, 967, 409, 383, 1140, 1146, 1228, 1149, 1501, 144, 1301, 979]. Trees [1085, 1094, 255, 1109, 408, 1396, 995, 151, 681]. Treewidth [1251]. Trends [465, 869, 688, 728]. Triangular [990]. Triangulation [1129]. Triangulations [1158]. Triggered [1528, 107]. Triggering [946]. Trim [165]. Truncation [880]. Trust [993, 396, 281, 736, 705, 740, 994, 704]. Trust-Based [704]. Truthful [1114].

Truthing [242]. TSP [1168, 1188, 1100]. TTS [806]. Tuning [35, 1297, 988]. Tutor [666]. Tutorial [1062]. Tutoring [113, 104]. TV [468]. TVQL [950]. Tweakable [622]. Two [1097, 238, 1216, 1201, 1481, 943, 1318, 904, 1137, 849, 63, 1225, 1478, 1211]. Two-Dimensional [1201]. Two-Directional [1318]. Two-Layer [1225]. Two-Level [849]. Two-Phase [904]. Type [1076, 157, 378, 1339, 1364, 1062, 1365]. Typed [1388, 1511, 1392, 1370]. Typed-Unification [1511, 1392]. Types [1351, 1263, 1145]. Typically [1480]. Typings [1381].

Ubicomp [1522]. Ubiquitous

[1523, 1530, 1536, 1541, 1535, 1521, 1531]. UDDI [690]. Ultra [857]. Ultra-Low [857]. Ultralow [869]. Ultralow-Voltage [869]. Ultrasonic [1537, 514, 591, 542]. UML [297, 298, 607, 296, 1047, 299]. UML-G [607]. Unbalanced [635]. Unbounded [1322]. Uncertain [117, 141]. UNCITRAL [373]. Understanding [563, 776, 262, 300]. Unguided [1136]. Unification [198, 1511, 1392, 1349]. Uniform [9, 1432, 1378, 1154, 818, 1479]. Uniformly [1146]. Uniformly-Oriented [1146]. Union [1123]. Unique [657]. Units [1136, 529]. Universal [3, 634, 479]. Universally [656]. University [451, 1161]. Unknown [794]. Unlabeled [766, 1403]. Unlimited [968]. Unlinkability [1049]. Unpredictable [956]. UNSEARCHMO [54]. Unstructured [784]. Unsupervised [758]. Update [173, 930]. upon [145]. Usage [169, 926, 991]. Usage-Based [169]. Use [531, 9, 783, 317, 259, 47]. Used [154]. Usefulness [1197]. User [346, 776, 1445, 1522, 37, 696, 1042, 98, 1025, 29, 734, 994, 739]. User-Driven [1025]. User-Level [696]. User-Role [994]. Users [731]. Using [1373, 247, 1092, 810, 263, 1071, 110, 90, 808, 121, 442, 1213, 932, 5, 1214, 1198, 254, 1488, 487, 745, 95, 960, 83, 990, 1215, 408, 1199, 842, 365, 1217, 290, 1034, 1178, 912, 1049, 8, 76, 515, 163, 941, 297, 39, 664, 1303, 596, 1000, 1538, 483, 659, 50, 91, 580, 403, 1072, 567, 1203, 1056, 59, 904, 125, 1428, 257, 126, 679, 906, 98, 122, 143, 378, 559, 108, 132, 700, 241, 1189, 251, 252, 1204, 1545, 880, 795, 161, 591, 485, 555, 266, 791, 1046, 1500, 453, 662, 1192, 1033, 328, 719, 1188, 819, 102, 817, 1534, 245, 75]. Using [144, 535, 756, 1403, 1514, 423, 1277, 966, 159, 1507, 780, 484, 230, 47, 457, 1051, 308, 1454, 911, 493, 99, 866, 804, 524, 1301, 1155]. Utility [344, 1213]. Utterance [800].

V [546]. V1 [472]. VA [165]. Vacuity [191]. vadis [982]. Vague [171, 969]. Valency [790]. Validated [284]. Validating [314]. Validation [1052, 1074, 173]. Value [1348, 1330, 877]. Valued [1356, 681]. Values [999]. Variability [425, 315]. Variable [1338, 1251, 909, 1153, 237]. Variables [1481, 51]. Variant [299]. Variants [1320, 1180, 1494]. Variations [236]. Varieties [641]. Varying [1100]. VC [552]. Vector [810, 358, 1122, 880, 977]. Vectors [233]. Vega [446]. Vehicle [1165, 1208, 1172]. Vehicles [1302]. Velocity [1410]. verbal [107]. Verbs [790]. Verifiable [657]. Verification [372, 184, 800, 1379, 135, 129, 566, 261, 223, 1363]. Verifying [201, 892]. Version [1215, 941]. Versioned [950]. versioning [1459]. versus [632, 921, 213]. Vertical [711]. Very [808, 163, 1179]. VHDL [1386]. via [654, 1183, 959, 1025, 1317, 1356, 877, 390, 115, 246, 1064]. Video [254, 514, 591, 453, 142, 458, 965]. View [160, 1476, 923]. View-Selection [160]. Viewer [178]. Views [934, 328, 911]. Virtex [517, 497, 494, 515, 516]. Virtex-Based [497]. Virtex-II [517]. Virtual [611, 1412, 442, 445, 745, 588, 1017, 732, 448, 897, 320, 834, 913, 606, 1438]. Vision [1409, 729, 1527]. Vision-Based [1527]. Visits [1520]. Visopt [1269]. Visual [536, 1215, 958, 108, 972, 606, 1408]. Visualisation [713, 9, 785]. Visualization [926, 679, 422, 462]. Visualizing [1428]. Visually [107]. VIZARD [583]. VLSI [857, 842, 858, 860]. Vocabulary [817, 229]. Voice [836, 834]. Voice-Driven [836]. VoiceXML [833]. Voltage [869, 854, 878]. Volume [583]. Volumes [1482]. Voronoi [1116, 1132]. Voting [236]. VPC [732]. vs [1387, 1411, 672, 28, 30, 788]. VSS [655].

W [1009]. Wakeup [1528]. Walk [1280]. Walking [84]. Walks [1477, 1245]. Walkthrough [915]. Want [684]. Warehouse [1006, 914, 1419, 419, 910, 1013, 161, 413, 902, 911]. Warehouses [1012, 1014, 1015]. Warehousing [1446]. WARMeR [763]. Watch [840]. Watermark [1051]. Watermarking [592, 177]. Waveform [862]. Wavelength [1183]. Wavelet [539, 486, 871]. Way [424, 1531]. WDM [1183, 589, 1128]. Weak [1320, 193]. Wearables [1544]. WearNET [1544]. Web [890, 697, 435, 1066, 1069, 250, 178, 395, 1041, 726, 1008, 396, 1048, 899, 447, 1484, 1067, 618, 892, 1034, 836, 690, 1414, 602, 460, 436, 1022, 824, 686, 167, 286, 967, 1184, 1460, 459, 168, 991, 431, 432, 1428, 450, 1018, 679, 440, 143, 172, 444, 702, 437, 666, 1025, 434, 422, 907, 1027, 688, 1046, 453, 695, 454, 992, 458, 1047, 765, 1399, 977, 452, 780, 692, 318, 433, 978, 439, 166, 104, 1036, 131, 704, 462, 894, 455, 148]. Web-Accessible [602, 422]. Web-Based [435, 447, 618, 460, 436, 1022, 459, 431, 432, 450, 440, 143, 444, 437, 666, 1046, 453, 454, 458, 452, 318, 433, 439, 104, 1036, 462, 455]. Web-Cache [172]. Web-Linkage [178]. Web-Server [1069]. Web-Services [697]. WebD [1009]. WebDAV [441]. WebMath [437]. Webs [275]. Weight [1104, 393, 1286].

Weighted [1132, 921]. Weighted-Distance [921]. Well [1405]. Well-Founded [1405]. Where [258]. Wide [1041, 1128]. Wide-Sense [1128]. Window [385]. Window-Method [385]. Windows [1111, 1165, 1136, 1172]. Winner [1021, 1402]. Winner-Takes-All [1402]. Winning [1336]. Wireless [36, 1077, 498, 49, 405, 366]. Wise [77]. within [380, 620]. without [1115, 1256]. Word [767, 226, 865, 125, 1004, 792, 251, 252, 788]. Wordlength [532]. WordNet [125]. Words [229, 794, 780]. Work [611, 155, 615]. Workbench [602]. Workflow [155, 179, 898, 1419, 420, 305, 896, 157, 156, 1420, 1448, 1423, 1424, 1421, 895, 1422]. Workflows [420, 896]. Workload [1065]. World [544, 1330, 1186, 1041]. Worst [1180]. Wrapper [133, 99]. Wrapper-Based [133]. Wrong [1480]. WSD [777]. WWW [399, 271, 891].

Xilinx [494, 515]. XMI [298]. XMI-Based [298]. XML [288, 916, 986, 918, 964, 1433, 459, 917, 173, 287, 268, 1429, 424, 1038, 139, 1026, 1316, 151, 1033, 1037, 1007, 919, 441, 1047, 159, 138, 153, 99, 140]. XML-Based [1038, 919, 153, 99]. XML-Encoded [1429]. XML-Schema [1433]. XML/JSP/WebDAV [441]. XOR [882]. XPath [960]. XPP [552]. XPP-VC [552]. XQuery [421]. XRL [919]. XSL [424]. XTR [647].

Yield [1395]. Yield-Management [1395].

Zero [857]. Zero-Slack [857]. Zeta [643]. Zone [240]. ZSA [857].

References

Zelikovitz:2002:IBK

[1] Sarah Zelikovitz and Haym Hirsh. Integrating background knowledge into nearest-neighbor text classification. Lecture Notes in Computer Science, 2416:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                              springer-ny.com/link/service/series/
                                              0558/bibs/2416/24160001.htm;   http:
                                              //link.springer-ny.com/link/service/
                                              series/0558/papers/2416/24160001.
                                              pdf.

Watson:2002:AKM

[2] Ian Watson. Applying knowledge management: Techniques for building organisational memories. Lecture Notes in Computer Science, 2416:6– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                              tronic). URL http://link.springer-
                                              ny.com/link/service/series/0558/
                                              bibs/2416/24160006.htm;   http:/
                                              /link.springer-ny.com/link/service/
                                              series/0558/papers/2416/24160006.
                                              pdf.

Au:2002:CPA

[3] Tsz-Chiu Au, Héctor Muñoz-Avila, and Dana S. Nau. On the complexity of plan adaptation by derivational analogy in a universal classical planning framework. Lecture Notes in Computer Science, 2416:13– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                              tronic). URL http://link.springer-
                                              ny.com/link/service/series/0558/
                                              bibs/2416/24160013.htm;   http:/
                                              /link.springer-ny.com/link/service/

series/0558/papers/2416/24160013.
pdf.

Baumeister:2002:ILC

[4] Joachim Baumeister, Martin Atzmüller, and Frank Puppe. Inductive learning for case-based diagnosis with multiple faults. Lecture Notes in Computer Science, 2416:28–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160028.
pdf.

Bridge:2002:DPR

[5] Derek Bridge and Alex Ferguson. Diverse product recommendations using an expressive language for case retrieval. Lecture Notes in Computer Science, 2416:43–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160043.
pdf.

Carswell:2002:DIS

[6] James D. Carswell, David C. Wilson, and Michela Bertolotto. Digital image similarity for geo-spatial knowledge management. Lecture Notes in Computer Science, 2416:58– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2416/24160058.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160058.
                                        pdf.

Diaz-Agudo:2002:PGC

[7] Belén Díaz-Agudo, Pablo Gervás, and Pedro A. González-Calero. Poetry generation in COLIBRI. Lecture Notes in Computer Science, 2416:73– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160073.
                                        pdf.

Falkman:2002:AUI

[8] Göran Falkman. Adaptation using iterated estimations. Lecture Notes in Computer Science, 2416:88– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160088.
                                        pdf.

Falkman:2002:UUD

[9] Göran Falkman. The use of a uniform declarative model in 3D visualisation for case-based reasoning. Lecture Notes in Computer Science, 2416: 103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160103.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2416/24160103.
pdf.

Gomes:2002:ECB

[10] Paulo Gomes, Francisco C. Pereira, Paulo Paiva, Nuno Seco, Paulo Carreiro, José L. Ferreira, and Carlos Bento. Experiments on case-based retrieval of software designs. Lecture Notes in Computer Science, 2416:118– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160118.
pdf.

Gupta:2002:ETC

[11] Kalyan Moy Gupta, David W. Aha, and Nabil Sandhu. Exploiting taxonomic and causal relations in conversational case retrieval. Lecture Notes in Computer Science, 2416:133– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160133.
pdf.

Hennessy:2002:BCR

[12] Daniel N. Hennessy, Bruce G. Buchanan, and John M. Rosenberg. Bayesian case reconstruction. Lecture Notes in Computer Science, 2416:148– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2416/24160148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160148.
                                        pdf.

Iglezakis:2002:RBC

[13] Ioannis Iglezakis and Thomas Reinartz. Relations between customer requirements, performance measures, and general case properties for case base maintenance. Lecture Notes in Computer Science, 2416:159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160159.
                                        pdf.

Jaere:2002:RTK

[14] Martha Dørum Jære, Agnar Aamodt, and Pål Skalle. Representing temporal knowledge for case-based prediction. Lecture Notes in Computer Science, 2416:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2416/24160174.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160174.
                                        pdf.

Kerkez:2002:LPC

[15] Boris Kerkez and Michael T. Cox. Local predictions for case-based plan recognition. Lecture Notes in Computer Science, 2416:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2416/24160189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160189.
pdf.

Leake:2002:ASS

[16] David B. Leake and Raja Sooriamurthi. Automatically selecting strategies for multi-case-base reasoning. Lecture Notes in Computer Science, 2416:204– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160204.
pdf.

McSherry:2002:DCR

[17] David McSherry. Diversity-conscious retrieval. Lecture Notes in Computer Science, 2416:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2416/24160219.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2416/24160219.pdf.

Montaner:2002:ICR

[18] Miquel Montaner, Beatriz López, and Josep Lluís de la Rosa. Improving case representation and case base maintenance in recommender agents. Lecture Notes in Computer Science, 2416:234– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160234.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160234.
                                        pdf.

Mougouie:2002:SAG

[19] Babak Mougouie and Ralph Bergmann. Similarity assessment for generalizied cases by optimization methods. Lecture Notes in Computer Science, 2416: 249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160249.
                                        pdf.

Mukkamalla:2002:CAP

[20] Sasidhar Mukkamalla and Héctor Muñoz-Avila. Case acquisition in a project planning environment. Lecture Notes in Computer Science, 2416:264– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160264.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160264.
                                        pdf.

Sullivan:2002:ICB

[21] Derry O’ Sullivan, David Wilson, and Barry Smyth. Improving casebased recommendation: A collaborative filtering approach. Lecture Notes in Computer Science, 2416:278– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160278.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2416/24160278.
pdf.

Patterson:2002:ESD

[22] David W. Patterson, Niall Rooney, and Mykola Galushka. Efficient similarity determination and case construction techniques for casebased reasoning. Lecture Notes in Computer Science, 2416:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160292.
pdf.

Plaza:2002:CA

[23] Enric Plaza and Josep-Lluís Arcos. Constructive adaptation. Lecture Notes in Computer Science, 2416:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160306.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160306.
pdf.

Portinale:2002:FCR

[24] Luigi Portinale and Stefania Montani. A fuzzy case retrieval approach based on SQL for implementing electronic catalogs. Lecture Notes in Computer Science, 2416:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2416/24160321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160321.
                                        pdf.

Prentzas:2002:IHR

[25] Jim Prentzas and Ioannis Hatzilygeroudis. Integrating hybrid rule-based with case-based reasoning. Lecture Notes in Computer Science, 2416:336– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160336.
                                        pdf.

Ruet:2002:SAF

[26] Magali Ruet and Laurent Geneste. Search and adaptation in a fuzzy object oriented case base. Lecture Notes in Computer Science, 2416:350– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160350.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160350.
                                        pdf.

Salamo:2002:DBS

[27] Maria Salamó and Elisabet Golobardes. Deleting and building sort out techniques for case base maintenance. Lecture Notes in Computer Science, 2416:365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2416/24160365.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2416/24160365.
pdf.

Schmitt:2002:EBV

[28] Sascha Schmitt, Philipp Dopichaj, and Patricia Domínguez-Marín. Entropybased vs. similarity-influenced: Attribute selection methods for dialogs tested on different electronic commerce domains. Lecture Notes in Computer Science, 2416:380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160380.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160380.
pdf.

Sollenborn:2002:CBF

[29] Mikael Sollenborn and Peter Funk. Category-based filtering and user stereotype cases to reduce the latency problem in recommender systems. Lecture Notes in Computer Science, 2416: 395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160395.
pdf.

Stahl:2002:DSM

[30] Armin Stahl. Defining similarity measures: Top-down vs. bottom-up. Lecture Notes in Computer Science, 2416: 406–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160406.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160406.
                                        pdf.

Wiratunga:2002:LAC

[31] Nirmalie Wiratunga, Susan Craw, and Ray Rowe. Learning to adapt for case-based design. Lecture Notes in Computer Science, 2416:421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160421.
                                        pdf.

Zenobi:2002:AAE

[32] Gabriele Zenobi and Pádraig Cunningham. An approach to aggregating ensembles of lazy learners that supports explanation. Lecture Notes in Computer Science, 2416:436–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160436.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160436.
                                        pdf.

Zhang:2002:ESI

[33] Lu Zhang, Frans Coenen, and Paul Leng. An experimental study of increasing diversity for casebased diagnosis. Lecture Notes in Computer Science, 2416:448–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160448.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160448.
pdf.

Aguzzoli:2002:CCB

[34] Stefano Aguzzoli, Paolo Avesani, and Paolo Massa. Collaborative casebased recommender systems. Lecture Notes in Computer Science, 2416:460– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160460.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160460.
pdf.

Bandini:2002:TPP

[35] Stefania Bandini, Sara Manzoni, and Carla Simone. Tuning production processes through a case based reasoning approach. Lecture Notes in Computer Science, 2416:475–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160475.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160475.
pdf.

Barbera:2002:ACB

[36] Massimo Barbera, Cristina Barbero, Paola Dal Zovo, Fernanda Farinaccio, Evangelos Gkroustiotis, Sofoklis Kyriazakos, Ivan Mura, and Gianluca Previti. An application of casebased reasoning to the adaptive management of wireless networks. Lecture Notes in Computer Science, 2416:490– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160490.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160490.
pdf.

Coyle:2002:CBP

[37] Lorcan Coyle, Pádraig Cunningham, and Conor Hayes. A casebased personal travel assistant for elaborating user requirements and assessing offers. Lecture Notes in Computer Science, 2416:505–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160505.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160505.
pdf.

Fdez-Riverola:2002:AHC

[38] Florentino Fdez-Riverola, Juan M. Corchado, and Jesús M. Torres. An automated hybrid CBR system for forecasting. Lecture Notes in Computer Science, 2416:519–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160519.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160519.
pdf.

Gomes:2002:UCA

[39] Paulo Gomes, Francisco C. Pereira, Paulo Paiva, Nuno Seco, Paulo Carreiro, José L. Ferreira, and Carlos Bento. Using CBR for automation of software design patterns. Lecture Notes in Computer Science, 2416:534– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160534.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160534.
pdf.

Hejazi:2002:NAS

[40] Mahmoudreza Hejazi and Kambiz Badie. A new approach to solution adaptation and its application for design purposes. Lecture Notes in Computer Science, 2416:549–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160549.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160549.
pdf.

Khemani:2002:ICF

[41] Deepak Khemani, Radhika B. Selvamani, Ananda Rabi Dhar, and S. M. Michael. InfoFrax : CBR in fused cast refractory manufacture. Lecture Notes in Computer Science, 2416:560– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160560.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2416/24160560.
                                        pdf.

Ginty:2002:CBR

[42] Lorraine Mc Ginty and Barry Smyth. Comparison-based recommendation. Lecture Notes in Computer Science, 2416:575–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2416/24160575.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160575.
                                        pdf.

Passone:2002:CBR

[43] Sara Passone, Paul W. H. Chung, and Vahid Nassehi. Case-based reasoning for estuarine model design. Lecture Notes in Computer Science, 2416:590– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160590.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160590.
                                        pdf.

Perner:2002:SGL

[44] Petra Perner, Horst Perner, and Bernd Müller. Similarity guided learning of the case description and improvement of the system performance in an image classification system. Lecture Notes in Computer Science, 2416:604– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160604.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2416/24160604.
pdf.

Ricci:2002:ICB

[45] Francesco Ricci, Bora Arslan, Nader Mirzadeh, and Adriano Venturini. ITR: A case-based travel advisory system. Lecture Notes in Computer Science, 2416:613–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2416/24160613.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2416/24160613.
pdf.

Schaaf:2002:SED

[46] Martin Schaaf, Rainer Maximini, Ralph Bergmann, Carsten Tautz, and Ralph Traphöner. Supporting electronic design reuse by integrating quality-criteria into CBRbased IP selection. Lecture Notes in Computer Science, 2416:628–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2416/24160628.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2416/24160628.
pdf.

Wang:2002:BCB

[47] Xingwen Wang and Anthony G. O. Yeh. Building a case-based decision support system for land development control using land use function pattern. Lecture Notes in Computer Science, 2416:642–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2416/24160642.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2416/24160642.
                                        pdf.

Anonymous:2002:AIa

[48] Anonymous. Author index. Lecture Notes in Computer Science, 2416:655– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2416/2416auth.pdf.

Pazzani:2002:CAM

[49] Michael J. Pazzani. Commercial applications of machine learning for personalized wireless portals. Lecture Notes in Computer Science, 2417:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170001.
                                        pdf.

Hovy:2002:LCU

[50] Eduard Hovy. Learning, collecting, and using ontological knowledge for NLP. Lecture Notes in Computer Science, 2417:6–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170006.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170006.
                                        pdf.

Foo:2002:HVK

[51] Norman Y. Foo. Hidden variables in knowledge representation. Lecture Notes in Computer Science, 2417:7– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170007.
pdf.

Rist:2002:ISR

[52] Thomas Rist. Intellimedia systems: Research and applications at the intersection of multimedia and artificial intelligence. Lecture Notes in Computer Science, 2417:9– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170009.
pdf.

Nakatsu:2002:IMA

[53] Ryohei Nakatsu. Integration of multimedia and art for new humancomputer communications. Lecture Notes in Computer Science, 2417:19– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170019.
pdf.

Chao:2002:RUR

[54] Yuyan Chao, Norimitsu Kawana, Lifeng He, Tsuyoshi Nakamura, and Hidenori Itoh. R-UNSEARCHMO: A refinement on UNSEARCHMO. Lecture Notes in Computer Science, 2417: 29–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170029.
pdf.

Tagawa:2002:DRL

[55] Takahiro Tagawa and Jingde Cheng. Deontic relevant logic: A strong relevant logic approach to removing paradoxes from deontic logic. Lecture Notes in Computer Science, 2417:39– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170039.
pdf.

Vo:2002:SRP

[56] Quoc Bao Vo and Norman Y. Foo. Solving the ramification problem: Causal propagation in an argumentation-theoretic approach. Lecture Notes in Computer Science, 2417:49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2417/24170049.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/2417/24170049.
pdf.

Zhang:2002:RAD

[57] Yan Zhang and Norman Y. Foo. Representing actions over dynamic domains. Lecture Notes in Computer Science, 2417:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2417/24170060.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2417/24170060.
pdf.

Zhang:2002:CAD

[58] Dongmo Zhang, Samir Chopra, and Norman Y. Foo. Consistency of action descriptions. Lecture Notes in Computer Science, 2417:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170070.
pdf.

Kim:2002:SFM

[59] Kee-Eung Kim and Thomas Dean. Solving factored MDPs with large action space using algebraic decision diagrams. Lecture Notes in Computer Science, 2417:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170080.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2417/24170080.
pdf.

Buller:2002:DF

[60] Andrzej Buller. Dynamic fuzziness. Lecture Notes in Computer Science, 2417:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2417/24170090.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2417/24170090.
pdf.

Hassine:2002:DRA

[61] Ahlem Ben Hassine and Khaled Ghédira. Distributed reinforcement of arc-consistency. Lecture Notes in Computer Science, 2417:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170097.
pdf.

Zhang:2002:PES

[62] Wenhui Zhang, Zhuo Huang, and Jian Zhang. Parallel execution of stochastic search procedures on reduced SAT instances. Lecture Notes in Computer Science, 2417:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170108.
pdf.

Matsuo:2002:TTC

[63] Yutaka Matsuo and Mitsuru Ishizuka. Two transformations of clauses into constraints and their properties for cost-based hypothetical reasoning. Lecture Notes in Computer Science, 2417: 118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170118.
pdf.

Itsuki:2002:HMM

[64] Noda Itsuki. Hidden Markov modeling for multi-agent systems. Lecture Notes in Computer Science, 2417:128– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170128.
pdf.

Wobcke:2002:MPL

[65] Wayne Wobcke. Modelling PRSlike agents’ mental states. Lecture Notes in Computer Science, 2417:138– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170138.
pdf.

Winoto:2002:GAS

[66] Pinata Winoto. Genetic algorithm and social simulation. Lecture Notes in Computer Science, 2417:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170148.
pdf.

Kijsirikul:2002:ADA

[67] Boonserm Kijsirikul, Nitiwut Ussivakul, and Surapant Meknavin. Adaptive directed acyclic graphs for multiclass classification. Lecture Notes in Computer Science, 2417:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170158.
pdf.

Lee:2002:NOT

[68] Hyunjin Lee, Hyeyoung Park, and Yillbyung Lee. Network optimization through learning and pruning in neuromanifold. Lecture Notes in Computer Science, 2417:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170169.
pdf.

Tung:2002:DND

[69] W. L. Tung and C. Quek. DIC: A novel discrete incremental clustering technique for the derivation of fuzzy membership functions. Lecture Notes in Computer Science, 2417:178– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170178.
pdf.

Ito:2002:AEQ

[70] Akira Ito. Application of episodic Q-learning to a multi-agent cooperative task. Lecture Notes in Computer Science, 2417:188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170188.
pdf.

Konda:2002:LLPa

[71] Taro Konda and Tomohiro Yamaguchi. LC-Learning: Phased method for average reward reinforcement learning — analysis of optimal criteria -. Lecture Notes in Computer Science, 2417:198– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170198.
pdf.

Konda:2002:LLPb

[72] Taro Konda, Shinjiro Tensyo, and Tomohiro Yamaguchi. LC-Learning: Phased method for average reward reinforcement learning — preliminary results. Lecture Notes in Computer Science, 2417:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170208.
                                        pdf.

Wada:2002:ERM

[73] Takuya Wada, Tetsuya Yoshida, Hiroshi Motoda, and Takashi Washio. Extension of the RDR method that can adapt to environmental changes and acquire knowledge from both experts and data. Lecture Notes in Computer Science, 2417:218–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170218.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170218.
                                        pdf.

Fujiwara:2002:CGM

[74] Keisei Fujiwara, Tetsuya Yoshida, Hiroshi Motoda, and Takashi Washio. Case generation method for constructing an RDR knowledge base. Lecture Notes in Computer Science, 2417:228– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2417/24170228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170228.
pdf.

Sarjon:2002:ARU

[75] Defit Sarjon and Noor Md Sap Mohd. Association rules using rough set and association rule methods. Lecture Notes in Computer Science, 2417:238– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170238.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170238.
pdf.

Fitzgibbon:2002:CPE

[76] Leigh J. Fitzgibbon, David L. Dowe, and Lloyd Allison. Change-point estimation using new minimum message length approximations. Lecture Notes in Computer Science, 2417:244– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170244.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170244.
pdf.

Matsuda:2002:KDS

[77] Takashi Matsuda, Hiroshi Motoda, Tetsuya Yoshida, and Takashi Washio. Knowledge discovery from structured data by beam-wise graphbased induction. Lecture Notes in Computer Science, 2417:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170255.
                                        pdf.

Tung:2002:BBP

[78] W. L. Tung and C. Quek. Back-POLE: Back propagation based on objective learning errors. Lecture Notes in Computer Science, 2417:265– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170265.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170265.
                                        pdf.

Lee:2002:MIO

[79] Tae-Seung Lee, Ho-Jin Choi, Young-Kil Kwag, and Byong-Won Hwang. A method on improvement of the online mode error backpropagation algorithm for pattern recognition. Lecture Notes in Computer Science, 2417:275– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170275.
                                        pdf.

Inoue:2002:OMC

[80] Hirotaka Inoue and Hiroyuki Narihisa. Optimizing a multiple classifier system. Lecture Notes in Computer Science, 2417:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2417/24170285.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2417/24170285.
pdf.

Arif:2002:GIL

[81] M. Arif, T. Ishihara, and H. Inooka. Generalization of iterative learning control for multiple desired trajectories in robotic systems. Lecture Notes in Computer Science, 2417:295– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170295.
pdf.

Yairi:2002:LTM

[82] Takehisa Yairi, Masahito Togami, and Koichi Hori. Learning topological maps from sequential observation and action data under partially observable environment. Lecture Notes in Computer Science, 2417:305– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170305.
pdf.

Cheoi:2002:SMM

[83] Kyungjoo Cheoi and Yillbyung Lee. A saliency map model for active perception using color information and local competitive mechanism. Lecture Notes in Computer Science, 2417:315– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170315.
                                        pdf.

Endo:2002:GOB

[84] Ken Endo, Funinori Yamasaki, Takashi Maeno, and Hiroaki Kitano. Generation of optimal biped walking for humanoid robot by co-evolving morphology and controller. Lecture Notes in Computer Science, 2417:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170325.
                                        pdf.

Valk:2002:MAC

[85] Jeroen Valk and Cees Witteveen. Multi-agent coordination in planning. Lecture Notes in Computer Science, 2417:335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2417/24170335.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170335.
                                        pdf.

Lao:2002:MAB

[86] Yizhi Lao and Hon Wai Leong. A multi-agent based approach to the inventory routing problem. Lecture Notes in Computer Science, 2417:345– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170345.
pdf.

Zhang:2002:ABH

[87] Zili Zhang and Chengqi Zhang. An agent-based hybrid intelligent system for financial investment planning. Lecture Notes in Computer Science, 2417: 355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170355.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170355.
pdf.

Hexmoor:2002:SIA

[88] Henry Hexmoor and Xin Zhang. Socially intelligent aerial combat simulator. Lecture Notes in Computer Science, 2417:365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2417/24170365.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2417/24170365.pdf.

Hwang:2002:CLS

[89] Kyu-Baek Hwang, Jae Won Lee, Seung-Woo Chung, and Byoung-Tak Zhang. Construction of largescale Bayesian networks by local to global search. Lecture Notes in Computer Science, 2417:375–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170375.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170375.
                                        pdf.

Bang:2002:UBN

[90] Jung-Wook Bang and Duncan Gillies. Using Bayesian networks with hidden nodes to recognise neural cell morphology. Lecture Notes in Computer Science, 2417:385–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2417/24170385.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2417/24170385.pdf.

Huang:2002:RSU

[91] Hung-Ju Huang and Chun-Nan Hsu. Recognizing 100 speakers using homologous naive Bayes. Lecture Notes in Computer Science, 2417:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170395.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170395.
                                        pdf.

Aizawa:2002:AMC

[92] Akiko Aizawa. An approach to microscopic clustering of terms and documents. Lecture Notes in Computer Science, 2417:404–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2417/24170404.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2417/24170404.pdf.

Kim:2002:EMI

[93] Sang-Bum Kim, Hae-Chang Rim, DongSuk Yook, and Heui-Seok Lim. Effective methods for improving naive Bayes text classifiers. Lecture Notes in Computer Science, 2417:414–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170414.
pdf.

Pantel:2002:ECD

[94] Patrick Pantel and Dekang Lin. Efficiently clustering documents with committees. Lecture Notes in Computer Science, 2417:424–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2417/24170424.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2417/24170424.pdf.

Chang:2002:TET

[95] Jeong-Ho Chang, Jae Won Lee, Yuseop Kim, and Byoung-Tak Zhang. Topic extraction from text documents using multiple-cause networks. Lecture Notes in Computer Science, 2417:434– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170434.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170434.
                                        pdf.

Lee:2002:CSS

[96] Kang Hyuk Lee, Judy Kay, Byeong Ho Kang, and Uwe Rosebrock. A comparative study on statistical machine learning algorithms and thresholding strategies for automatic text categorization. Lecture Notes in Computer Science, 2417:444–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170444.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170444.
                                        pdf.

Lee:2002:PAI

[97] Jae Won Lee and Sung-Dong Kim. PATI: An approach for identifying and resolving ambiguities. Lecture Notes in Computer Science, 2417:454– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170454.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170454.
                                        pdf.

Ku:2002:EUP

[98] Sebon Ku, Bogju Lee, and Eunyong Ha. Extracting user profiles from E-mails using the setoriented classifier. Lecture Notes in Computer Science, 2417:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170463.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170463.
pdf.

Yang:2002:WGU

[99] Jaeyoung Yang, Jungsun Kim, Kyoung-Goo Doh, and Joongmin Choi. Wrapper generation by using XML-based domain knowledge for intelligent information extraction. Lecture Notes in Computer Science, 2417:472–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170472.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170472.
pdf.

Kitakami:2002:MPM

[100] Hajime Kitakami, Tomoki Kanbara, Yasuma Mori, Susumu Kuroki, and Yukiko Yamazaki. Modified PrefixSpan method for motif discovery in sequence databases. Lecture Notes in Computer Science, 2417:482–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170482.
pdf.

Karasavvas:2002:MAB

[101] Konstantinos Karasavvas, Albert Burger, and Richard Baldock. A multiagent bioinformatics integration system with adjustable autonomy. Lecture Notes in Computer Science, 2417: 492–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170492.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170492.
                                        pdf.

Reyes:2002:UCB

[102] Rhodora L. Reyes and Raymund C. Sison. Using case-based reasoning approach in planning instructional activities. Lecture Notes in Computer Science, 2417:502–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170502.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170502.
                                        pdf.

Tang:2002:FCS

[103] Tiffany Y. Tang and Keith C. Chan. Feature construction for student group forming based on their browsing behaviors in an E-learning system. Lecture Notes in Computer Science, 2417: 512–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170512.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2417/24170512.
pdf.

Woo:2002:WBT

[104] Chong woo Woo and Jin woo Choi. Web-based tutoring system for computer security. Lecture Notes in Computer Science, 2417:522–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170522.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170522.
pdf.

Ko:2002:SIN

[105] Jaepil Ko, Eunju Kim, and Hyeran Byun. A simple illumination normalization algorithm for face recognition. Lecture Notes in Computer Science, 2417:532–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170532.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170532.
pdf.

Robertson:2002:SAA

[106] Paul Robertson and Robert Laddaga. A self-adaptive architecture and its application to robust face identification. Lecture Notes in Computer Science, 2417:542–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2417/24170542.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2417/24170542.
                                        pdf.

Okuno:2002:RAV

[107] Hiroshi G. Okuno, Kazuhiro Nakadai, and Hiroaki Kitano. Realizing audiovisually triggered ELIZA-like nonverbal behaviors. Lecture Notes in Computer Science, 2417:552–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170552.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170552.
                                        pdf.

Lee:2002:AVC

[108] Soonkyu Lee and Dongsuk Yook. Audio-to-visual conversion using hidden Markov models. Lecture Notes in Computer Science, 2417:563–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170563.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170563.
                                        pdf.

Prendinger:2002:SBM

[109] Helmut Prendinger, Sylvain Descamps, and Mitsuru Ishizuka. Scripting the bodies and minds of lifelike characters. Lecture Notes in Computer Science, 2417:571–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170571.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2417/24170571.
pdf.

Baillie:2002:ADM

[110] Penny Baillie and Dickson Lukose. An affective decision making agent architecture using emotion appraisals. Lecture Notes in Computer Science, 2417: 581–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170581.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170581.
pdf.

Hayashi:2002:LPA

[111] Hisashi Hayashi, Kenta Cho, and Akihiko Ohsuga. Logic programming for agents. Lecture Notes in Computer Science, 2417:591–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170591.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170591.
pdf.

Piron:2002:EMA

[112] Marie Piron, Alain Cardon, and Christophe Cambier. Evolutionary multi-agents system for prediction of social behavior in large cities. Lecture Notes in Computer Science, 2417:592– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170592.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170592.
                                        pdf.

Shakshuki:2002:ABC

[113] Elhadi Shakshuki and Trang Dang. Agent-based cooperative distributed tutoring systems. Lecture Notes in Computer Science, 2417:594–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170594.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170594.
                                        pdf.

Oehlmann:2002:CMR

[114] Ruediger Oehlmann. A computational model of reasoning as sociallyconstructed process. Lecture Notes in Computer Science, 2417:595–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170595.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170595.
                                        pdf.

Taniguchi:2002:MIC

[115] Ken Taniguchi, Setsuya Kurahashi, and Takao Terano. Managing information complexity of supply chains via agent-based genetic programming. Lecture Notes in Computer Science, 2417: 596–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170596.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2417/24170596.
pdf.

Izumi:2002:SIB

[116] Noriaki Izumi and Takahira Yamaguchi. Semantic integration of E-business models based on multilayered repository. Lecture Notes in Computer Science, 2417:597–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170597.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170597.
pdf.

Kyburg:2002:RUI

[117] Henry E. Kyburg and Choh Man Teng. Randomization and uncertain inference. Lecture Notes in Computer Science, 2417:598–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170598.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170598.
pdf.

Kim:2002:CSE

[118] Kyung-Joong Kim and Sung-Bae Cho. Checkers strategy evolution with speciated neural networks. Lecture Notes in Computer Science, 2417:599–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170599.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170599.
                                        pdf.

Swiatnicki:2002:GOF

[119] Zbigniew Świa̧tnicki and Vladimír Olej. Generation and optimization of fuzzy neural network structure. Lecture Notes in Computer Science, 2417:600– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170600.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170600.
                                        pdf.

Tung:2002:FFR

[120] W. L. Tung and C. Quek. FuzzyDrive: A fuzzy rule-based auto pilot system. Lecture Notes in Computer Science, 2417:601–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2417/24170601.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170601.
                                        pdf.

Beydoun:2002:INU

[121] Ghassan Beydoun and Lina Al-Jadir. Implementing NRDR using OO database management system (OODBMS). Lecture Notes in Computer Science, 2417:602–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170602.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2417/24170602.
pdf.

Kwok:2002:URR

[122] Rex B. H. Kwok. Using ripple down rules for actions and planning. Lecture Notes in Computer Science, 2417: 604–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170604.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170604.
pdf.

Diaconescu:2002:GDG

[123] Stefan Diaconescu. A generative dependency grammar. Lecture Notes in Computer Science, 2417:605–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170605.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170605.
pdf.

Uliniansyah:2002:IMP

[124] Mohammad Teduh Uliniansyah, Shun Ishizaki, and Kiyoko Uchiyama. Indonesian morphological parser with minimum connectivity cost to solve ambiguities. Lecture Notes in Computer Science, 2417:606–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2417/24170606.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2417/24170606.pdf.

Kim:2002:TWS

[125] Yuseop Kim, Jeong-Ho Chang, and Byoung-Tak Zhang. Target word selection using WordNet and data-driven models in machine translation. Lecture Notes in Computer Science, 2417: 607–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170607.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170607.
                                        pdf.

Kobayashi:2002:SUN

[126] Ichiro Kobayashi, Michiaki Iwazume, Shino Iwashita, Toru Sugimoto, and Michio Sugeno. A study on using natural language as a computer communication protocol. Lecture Notes in Computer Science, 2417:608–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170608.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170608.
                                        pdf.

Matsumura:2002:AIB

[127] Naohiro Matsumura, Yukio Ohsawa, and Mitsuru Ishizuka. Automatic indexing based on term activities. Lecture Notes in Computer Science, 2417: 609–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170609.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2417/24170609.
pdf.

Lee:2002:AEF

[128] Kyung-Soon Lee, Jae-Ho Kim, and Key-Sun Choi. Answer extraction by flexible matching, filtering, and interpretation. Lecture Notes in Computer Science, 2417:610–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170610.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170610.
pdf.

Lee:2002:SIV

[129] Kwanyong Lee and Hyeyoung Park. A statistical identification and verification method for biometrics. Lecture Notes in Computer Science, 2417:611– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170611.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170611.
pdf.

Araki:2002:PMD

[130] Masahiro Araki, Kiyoshi Ueda, Masashi Akita, Takuya Nishimoto, and Yasuhisa Niimi. Proposal of a multimodal dialogue description language. Lecture Notes in Computer Science, 2417:612– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170612.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170612.
                                        pdf.

Yanai:2002:ICW

[131] Keiji Yanai. Image classification by Web images. Lecture Notes in Computer Science, 2417:613–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170613.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170613.
                                        pdf.

Lee:2002:RTF

[132] Chang-Woo Lee, Yeon-Chul Lee, Sang-Yong Bak, and Hang-Joon Kim. Realtime face detection and tracking using PCA and NN. Lecture Notes in Computer Science, 2417:615–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170615.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170615.
                                        pdf.

Bredeche:2002:WBA

[133] Nicolas Bredeche, Jean-Daniel Zucker, and Yann Chevaleyre. A wrapperbased approach to robot learning concepts from images. Lecture Notes in Computer Science, 2417:616–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2417/24170616.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2417/24170616.
pdf.

Park:2002:EHB

[134] Hyuk-Jang Park and Sung-Bae Cho. An effective HMM-based intrusion detection system with privilege change event modeling. Lecture Notes in Computer Science, 2417:617–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170617.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170617.
pdf.

Lee:2002:MIE

[135] Tae-Seung Lee, Ho-Jin Choi, Seung-Hoe Choi, and Byong-Won Hwang. A method on improving of enrolling speed for the MLP-based speaker verification system through reducing learning data. Lecture Notes in Computer Science, 2417:619–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2417/24170619.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2417/24170619.
pdf.

Meyer:2002:SRS

[136] Thomas Meyer, Aditya Ghose, and Samir Chopra. Syntactic representations of semantic merging operations. Lecture Notes in Computer Science, 2417:620–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2417/24170620.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2417/24170620.
                                        pdf.

Anonymous:2002:AIb

[137] Anonymous. Author index. Lecture Notes in Computer Science, 2417:621– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2417/2417auth.pdf.

Wan:2002:ESX

[138] Changxuan Wan and Yunsheng Liu. Efficient supporting XML query and keyword search in relational database systems. Lecture Notes in Computer Science, 2419:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190001.
                                        pdf.

Lv:2002:PED

[139] Jianhua Lv, Guoren Wang, Jeffrey X. Yu, Ge Yu, Hongjun Lu, and Bing Sun. Performance evaluation of a DOMbased XML database: Storage, indexing and query optimization. Lecture Notes in Computer Science, 2419:13– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190013.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2419/24190013.
pdf.

Zheng:2002:SMN

[140] Shihui Zheng, Aoying Zhou, Jeffrey Xu Yu, Long Zhang, and Han Tao. Structural map: A new index for efficient XML path expression processing. Lecture Notes in Computer Science, 2419: 25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190025.
pdf.

Tossebro:2002:ADM

[141] Erlend Tøssebro and Mads Nygård. An advanced discrete model for uncertain spatial data. Lecture Notes in Computer Science, 2419:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190037.
pdf.

Shim:2002:STR

[142] Choon-Bo Shim and Jae-Woo Chang. A spatio-temporal representation scheme for content- and semanticbased video retrieval on moving objects’ trajectories. Lecture Notes in Computer Science, 2419:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190052.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190052.
                                        pdf.

Kwon:2002:EAT

[143] Joon-Hee Kwon and Yong-Ik Yoon. Efficient access technique using levelized data in Web-based GIS. Lecture Notes in Computer Science, 2419:64– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190064.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190064.
                                        pdf.

Shi:2002:UPT

[144] Shengfei Shi, Jianzhong Li, and Chaokun Wang. Using PR-tree and HPIR to manage coherence of semantic cache for location dependent data in mobile database. Lecture Notes in Computer Science, 2419:72– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190072.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190072.
                                        pdf.

Ma:2002:TER

[145] Xiu li Ma, Shi wei Tang, Dong qing Yang, and Xiao ping Du. Towards efficient re-mining of frequent patterns upon threshold changes. Lecture Notes in Computer Science, 2419:80– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190080.
pdf.

Zuo:2002:MPA

[146] Jie Zuo, Changjie Tang, and Tianqing Zhang. Mining predicate association rule by gene expression programming. Lecture Notes in Computer Science, 2419:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190092.
pdf.

Chen:2002:LCH

[147] Jianzhong Chen, Mary Shapcott, Sally McClean, and Kenny Adamson. Learning with concept hierarchies in probabilistic relational data mining. Lecture Notes in Computer Science, 2419:104– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190104.
pdf.

Zhou:2002:ESL

[148] Aoying Zhou, Xiong Fang, and Weining Qian. An empirical study of learning-based Web search. Lecture Notes in Computer Science, 2419:116– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190116.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190116.
                                        pdf.

He:2002:ODI

[149] Zengyou He, Shengchun Deng, and Xiaofei Xu. Outlier detection integrating semantic knowledge. Lecture Notes in Computer Science, 2419:126– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190126.
                                        pdf.

Ma:2002:EIS

[150] Liping Ma, John Shepherd, and Yanchun Zhang. Extracting information from semistructured data. Lecture Notes in Computer Science, 2419:132– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190132.
                                        pdf.

Ng:2002:MCI

[151] Wilfred Ng. Maintaining consistency of integrated XML trees. Lecture Notes in Computer Science, 2419:145– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190145.
pdf.

Yuan:2002:EFP

[152] Qingqing Yuan, Yubo Lou, Haofeng Zhou, Wei Wang, and Baile Shi. Extract frequent pattern from simple graph data. Lecture Notes in Computer Science, 2419:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190158.
pdf.

Xu:2002:XBD

[153] Peng Xu, Wenjun Yang, and Kehong Wang. XML-based data rendering engine for content management system. Lecture Notes in Computer Science, 2419:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190170.
pdf.

Chen:2002:TDU

[154] Kang Chen, Weimin Zheng, Xiatie Deng, Haodi Feng, and Shanfeng Zhu. Text distinguishers used in an interactive meta search engine. Lecture Notes in Computer Science, 2419:181– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190181.
                                        pdf.

Cao:2002:TWO

[155] Jian Cao, Shensheng Zhang, and Xiaoguang Zhang. Team work oriented flexible workflow management system. Lecture Notes in Computer Science, 2419:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2419/24190189.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190189.
                                        pdf.

Kumar:2002:ERB

[156] Akhil Kumar and J. Leon Zhao. EROICA: A rule-based approach to organizational policy management in workflow systems. Lecture Notes in Computer Science, 2419:201–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190201.
                                        pdf.

Kim:2002:ATW

[157] Kwang-Hoon Kim and In-Soo Kim. The admon-time workflow client: Why do we need the third type of workflow client designated for administration and monitoring services? Lecture Notes in Computer Science, 2419:213– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190213.
pdf.

Wagner:2002:PCD

[158] Matthias Wagner, Werner Kießling, and Wolf-Tilo Balke. Progressive content delivery for mobile E-services. Lecture Notes in Computer Science, 2419: 225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190225.
pdf.

Tanoue:2002:GBS

[159] Junko Tanoue, Noboru Matoba, Masatoshi Yoshikawa, and Shunsuke Uemura. GeneAround: A browsing system for gene annotation using XML technologies. Lecture Notes in Computer Science, 2419:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190236.
pdf.

Choi:2002:WDH

[160] Chi-Hon Choi, Jeffrey Xu Yu, and Gang Gou. What difference heuristics make: Maintenance-cost viewselection revisited. Lecture Notes in Computer Science, 2419:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190247.
pdf.

Medina:2002:RCM

[161] Enrique Medina and Juan Trujillo. Representing conceptual multidimensional properties using the common warehouse metamodel (CWM). Lecture Notes in Computer Science, 2419: 259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190259.
pdf.

Feng:2002:OQP

[162] Jianhua Feng, Li Chao, Xudong Jiang, and Lizhu Zhou. OLAP query processing algorithm based on relational storage. Lecture Notes in Computer Science, 2419:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190271.
pdf.

Fu:2002:QCV

[163] Lixin Fu. Querying and clustering very large data sets using dynamic bucketing approach. Lecture Notes in Computer Science, 2419:279– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190279.
pdf.

Chang:2002:CCB

[164] Jae-Woo Chang. CBCM: A cellbased clustering method for data mining applications. Lecture Notes in Computer Science, 2419:291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190291.
pdf.

Chen:2002:CVT

[165] Hanxiong Chen, Jiyuan An, Kazutaka Furuse, and Nobuo Ohbo. C^{2} VA: Trim high dimensional indexes. Lecture Notes in Computer Science, 2419: 303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190303.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190303.
pdf.

Wang:2002:TED

[166] Xiaoyu Wang, Zhiguo Lu, and Aoying Zhou. Topic exploration and distillation for Web search by a similarity-based analysis. Lecture Notes in Computer Science, 2419:316– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190316.
pdf.

Guillaume:2002:ESE

[167] Jean-Loup Guillaume, Matthieu Latapy, and Laurent Viennot. Efficient and simple encodings for the Web graph. Lecture Notes in Computer Science, 2419:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190328.
pdf.

Hui:2002:OEB

[168] Song Hui, Zhang Ling, Ye Yunming, and Ma Fanyuan. Objectextraction-based hidden Web information retrieval. Lecture Notes in Computer Science, 2419:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190338.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2419/24190338.
pdf.

Ding:2002:IUB

[169] Chen Ding, Chi-Hung Chi, and Tiejian Luo. An improved usagebased ranking. Lecture Notes in Computer Science, 2419:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190346.
pdf.

Zhang:2002:LSB

[170] Qing Zhang and Xuemin Lin. On linear-spline based histograms. Lecture Notes in Computer Science, 2419:354– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190354.
pdf.

Dang:2002:GEA

[171] Tran Khanh Dang, Josef Küng, and Roland Wagner. A general and efficient approach for solving nearest neighbor problem in the vague query system. Lecture Notes in Computer Science, 2419:367–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190367.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2419/24190367.
                                        pdf.

Lau:2002:CBW

[172] Kelvin Lau and Yiu-Kai Ng. A clientbased Web-cache management system. Lecture Notes in Computer Science, 2419:379–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2419/24190379.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190379.
                                        pdf.

Kim:2002:IPV

[173] Sang-Kyun Kim, Myungcheol Lee, and Kyu-Chul Lee. Immediate and partial validation mechanism for the conflict resolution of update operations in XML databases. Lecture Notes in Computer Science, 2419:387– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190387.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190387.
                                        pdf.

Liu:2002:RAC

[174] YunSheng Liu, GuoQiong Liao, Guo-Hui Li, and JiaLi Xia. Relaxed atomic commit for real-time transactions in mobile computing environment. Lecture Notes in Computer Science, 2419: 397–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190397.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2419/24190397.
pdf.

Xiang:2002:RTF

[175] Zhenggui Xiang, Tianxiang Yao, Jiang Li, and Keman Yu. Realtime facial patterns mining and emotion tracking. Lecture Notes in Computer Science, 2419:409–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190409.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190409.
pdf.

Guo:2002:DRN

[176] Gongde Guo, Hui Wang, and David Bell. Data reduction and noise filtering for predicting times series. Lecture Notes in Computer Science, 2419:421– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2419/24190421.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2419/24190421.
pdf.

Zhao:2002:EII

[177] Xianfeng Zhao, Weinong Wang, and Kefei Chen. Exploiting the intrinsic irreversibility of adaptive technologies to enhance the security of digital watermarking. Lecture Notes in Computer Science, 2419:430–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190430.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190430.
                                        pdf.

Asano:2002:WLV

[178] Yasuhito Asano, Hiroshi Imai, Masashi Toyoda, and Masaru Kitsuregawa. Web-linkage viewer: Finding graph structures in the Web. Lecture Notes in Computer Science, 2419:441–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190441.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190441.
                                        pdf.

Chen:2002:GWS

[179] JinShui Chen and Yan Tu. GDWFMS: A workflow system for submitting graphs and documents. Lecture Notes in Computer Science, 2419:443–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2419/24190443.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2419/24190443.
                                        pdf.

Anonymous:2002:AIc

[180] Anonymous. Author index. Lecture Notes in Computer Science, 2419:445– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2419/2419auth.pdf.

Aspinall:2002:SCC

[181] David Aspinall and Donald Sannella. From specifications to code in CASL. Lecture Notes in Computer Science, 2422:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2422/24220001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2422/24220001.
pdf.

Walukiewicz:2002:AGS

[182] Igor Walukiewicz. Automata and games for synthesis. Lecture Notes in Computer Science, 2422:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220015.
pdf.

Mosses:2002:PMS

[183] Peter D. Mosses. Pragmatics of modular SOS. Lecture Notes in Computer Science, 2422:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220021.
pdf.

Barthe:2002:TAS

[184] Gilles Barthe, Pierre Courtieu, Guillaume Dufay, and Simão Melo de Sousa. Tool-assisted specification and verification of the JavaCard platform. Lecture Notes in Computer Science, 2422:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220041.
pdf.

Miller:2002:HOQ

[185] Dale Miller. Higher-order quantification and proof search. Lecture Notes in Computer Science, 2422:60– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220060.
pdf.

Fiadeiro:2002:ASS

[186] José Luiz Fiadeiro. Algebraic support for service-oriented architecture (extended abstract). Lecture Notes in Computer Science, 2422:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220075.
pdf.

Haack:2002:FAA

[187] Christian Haack, Brian Howard, Allen Stoughton, and Joe B. Wells.

Fully automatic adaptation of software components based on semantic specifications. Lecture Notes in Computer Science, 2422:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220083.
pdf.

Schroder:2002:HTI

[188] Lutz Schröder and Till Mossakowski. HASCASL: Towards integrated specification and development of functional programs. Lecture Notes in Computer Science, 2422:99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220099.
pdf.

Alpuente:2002:RRA

[189] María Alpuente, Santiago Escobar, and Salvador Lucas. Removing redundant arguments of functions. Lecture Notes in Computer Science, 2422:117– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220117.
pdf.

Adelaide:2002:CDP

[190] Michaël Adélaïde and Olivier Roux. A class of decidable parametric hybrid systems. Lecture Notes in Computer Science, 2422:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220132.
                                        pdf.

Dong:2002:VCM

[191] Yifei Dong, Beata Sarna-Starosta, C. R. Ramakrishnan, and Scott A. Smolka. Vacuity checking in the modal mu-calculus. Lecture Notes in Computer Science, 2422:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220147.
                                        pdf.

Hornus:2002:STL

[192] Samuel Hornus and Philippe Schnoebelen. On solving temporal logic queries. Lecture Notes in Computer Science, 2422:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220163.
                                        pdf.

Guo:2002:MCB

[193] Guangyuan Guo and Ryszard Janicki. Modelling concurrent behaviours by commutativity and weak causality relations. Lecture Notes in Computer Science, 2422:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220178.
pdf.

Devillers:2002:ANS

[194] Raymond Devillers, Hanna Klaudel, Maciej Koutny, and Franck Pommereau. An algebra of nonsafe Petri boxes. Lecture Notes in Computer Science, 2422:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220192.
pdf.

Ulidowski:2002:RSI

[195] Irek Ulidowski. Refusal simulation and interactive games. Lecture Notes in Computer Science, 2422:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220208.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220208.
pdf.

Thati:2002:TMT

[196] Prasannaa Thati, Reza Ziaei, and Gul Agha. A theory of may testing for asynchronous calculi with locality and no name matching. Lecture Notes in Computer Science, 2422:223– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220223.
                                        pdf.

Aceto:2002:EAP

[197] Luca Aceto, Zoltán Ésik, and Anna Ingólfsdóttir. Equational axioms for probabilistic bisimilarity. Lecture Notes in Computer Science, 2422:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220239.
                                        pdf.

Baldan:2002:BU

[198] Paolo Baldan, Andrea Bracciali, and Roberto Bruni. Bisimulation by unification. Lecture Notes in Computer Science, 2422:254–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220254.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220254.
                                        pdf.

Bossi:2002:TPC

[199] Annalisa Bossi, Riccardo Focardi, Carla Piazza, and Sabina Rossi. Transforming processes to check and ensure information flow security. Lecture Notes in Computer Science, 2422:271– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220271.
pdf.

Borgstrom:2002:BSC

[200] Johannes Borgström and Uwe Nestmann. On bisimulations for the spi calculus. Lecture Notes in Computer Science, 2422:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2422/24220287.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2422/24220287.pdf.

Breunesse:2002:SVD

[201] Cees-Bart Breunesse, Bart Jacobs, and Joachim van den Berg. Specifying and verifying a decimal representation in Java for Smart Cards. Lecture Notes in Computer Science, 2422:304– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220304.
pdf.

Haneberg:2002:MSS

[202] Dominik Haneberg, Wolfgang Reif, and Kurt Stenzel. A method for secure Smartcard applications. Lecture Notes in Computer Science, 2422:319– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220319.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220319.
                                        pdf.

Anonymous:2002:AId

[203] Anonymous. Author index. Lecture Notes in Computer Science, 2422:321– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2422/2422ind.pdf.

Trentelman:2002:EJS

[204] Kerry Trentelman and Marieke Huisman. Extending JML specifications with temporal logic. Lecture Notes in Computer Science, 2422:334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220334.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220334.
                                        pdf.

Giegerich:2002:ADP

[205] Robert Giegerich and Carsten Meyer. Algebraic dynamic programming. Lecture Notes in Computer Science, 2422: 349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220349.
pdf.

Simon:2002:ASB

[206] Axel Simon and Andy King. Analyzing string buffers in C. Lecture Notes in Computer Science, 2422:365– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220365.
pdf.

Hill:2002:FEA

[207] Patricia M. Hill and Fausto Spoto. A foundation of escape analysis. Lecture Notes in Computer Science, 2422:380– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220380.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220380.
pdf.

Stell:2002:FOS

[208] John G. Stell. A framework for order-sorted algebra. Lecture Notes in Computer Science, 2422:396–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220396.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220396.
                                        pdf.

Pavlovic:2002:GTE

[209] Dusko Pavlovic and Douglas R. Smith. Guarded transitions in evolving specifications. Lecture Notes in Computer Science, 2422:411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220411.
                                        pdf.

Lopes:2002:RCA

[210] Antónia Lopes and José Luiz Fiadeiro. Revisiting the categorical approach to systems. Lecture Notes in Computer Science, 2422:426–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220426.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220426.
                                        pdf.

Schairer:2002:PTE

[211] Axel Schairer and Dieter Hutter. Proof transformations for evolutionary formal software development. Lecture Notes in Computer Science, 2422:441– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2422/24220441.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2422/24220441.
pdf.

Skoglund:2002:SOR

[212] Mats Skoglund. Sharing objects by read-only references. Lecture Notes in Computer Science, 2422:457–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220457.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220457.
pdf.

Reus:2002:CBV

[213] Bernhard Reus. Class-based versus object-based: A denotational comparison. Lecture Notes in Computer Science, 2422:473–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2422/24220473.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2422/24220473.pdf.

Rybina:2002:BBR

[214] Tatiana Rybina and Andrei Voronkov. BRAIN : Backward reachability analysis with integers. Lecture Notes in Computer Science, 2422:489–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2422/24220489.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2422/24220489.
pdf.

Autexier:2002:DGM

[215] Serge Autexier, Dieter Hutter, Till Mossakowski, and Axel Schairer. The development graph manager MAYA. Lecture Notes in Computer Science, 2422:495–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2422/24220495.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2422/24220495.
                                        pdf.

Smith:2002:RSI

[216] Elisa Barney Smith and Xiaohui Qiu. Relating statistical image differences and degradation features. Lecture Notes in Computer Science, 2423:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230001.
                                        pdf.

Dhanya:2002:SIP

[217] D. Dhanya and A. G. Ramakrishnan. Script identification in printed bilingual documents. Lecture Notes in Computer Science, 2423:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230013.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230013.
                                        pdf.

Dhanya:2002:OFE

[218] D. Dhanya and A. G. Ramakrishnan. Optimal feature extraction for bilingual OCR. Lecture Notes in Computer Science, 2423:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230025.
pdf.

Kumar:2002:MRP

[219] B. Vijay Kumar and A. G. Ramakrishnan. Machine recognition of printed kannada text. Lecture Notes in Computer Science, 2423:37– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230037.
pdf.

Vezzosi:2002:ISA

[220] Stefano Vezzosi, Luigi Bedini, and Anna Tonazzini. An integrated system for the analysis and the recognition of characters in ancient documents. Lecture Notes in Computer Science, 2423: 49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230049.
pdf.

Aparna:2002:CTO

[221] K. G. Aparna and A. G. Ramakrishnan. A complete tamil optical character recognition system. Lecture Notes in Computer Science, 2423:53– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230053.
pdf.

Santos:2002:DBH

[222] José Eduardo Bastos Dos Santos, Bernard Dubuisson, and Flávio Bortolozzi. Distinguishing between handwritten and machine printed text in bank cheque images. Lecture Notes in Computer Science, 2423:58– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230058.
pdf.

Ueda:2002:MES

[223] Katsuhiko Ueda and Ken’ichi Matsuo. Multi-expert seal imprint verification system for bankcheck processing. Lecture Notes in Computer Science, 2423: 62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230062.
pdf.

Murshed:2002:ART

[224] Nabeel Murshed. Automatic reading of traffic tickets. Lecture Notes in Computer Science, 2423:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230066.
pdf.

Xue:2002:SMC

[225] Hanhong Xue and Venu Govindaraju. A stochastic model combining discrete symbols and continuous attributes and its application to handwriting recognition. Lecture Notes in Computer Science, 2423:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230070.
pdf.

Ishidera:2002:TLW

[226] Eiki Ishidera, Simon M. Lucas, and Andrew C. Downton. Top-down likelihood word image generation model for holistic word recognition. Lecture Notes in Computer Science, 2423:82– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230082.
pdf.

Zheng:2002:SIH

[227] Yefeng Zheng, Huiping Li, and David Doermann. The segmentation and identification of handwriting in noisy document images. Lecture Notes in Computer Science, 2423:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230095.
                                        pdf.

Velek:2002:ILT

[228] Ondrej Velek and Masaki Nakagawa. The impact of large training sets on the recognition rate of offline Japanese Kanji character classifiers. Lecture Notes in Computer Science, 2423:106– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230106.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230106.
                                        pdf.

Ryu:2002:ACK

[229] Sungho Ryu and Jin-Hyung Kim. Automatic completion of Korean words for open vocabulary pen interface. Lecture Notes in Computer Science, 2423:111– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230111.
                                        pdf.

Velek:2002:USN

[230] Ondrej Velek and Masaki Nakagawa. Using stroke-number-characteristics for improving efficiency of combined online and offline Japanese character classifiers. Lecture Notes in Computer Science, 2423:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230115.
pdf.

Bodansky:2002:CGD

[231] Eugene Bodansky and Alexander Gribov. Closing gaps of discontinuous lines: A new criterion for choosing the best prolongation. Lecture Notes in Computer Science, 2423:119– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230119.
pdf.

Veeramachaneni:2002:CAN

[232] Sriharsha Veeramachaneni and George Nagy. Classifier adaptation with nonrepresentative training data. Lecture Notes in Computer Science, 2423:123– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230123.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2423/24230123.
                                        pdf.

Takebe:2002:LPB

[233] Hiroaki Takebe, Koji Kurokawa, Yutaka Katsuyama, and Satoshi Naoi. A learning pseudo Bayes discriminant method based on difference distribution of feature vectors. Lecture Notes in Computer Science, 2423:134– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230134.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230134.
                                        pdf.

Bovino:2002:INC

[234] L. Bovino, G. Dimauro, S. Impedovo, G. Pirlo, and A. Salzo. Increasing the number of classifiers in multiclassifier systems: A complementaritybased analysis. Lecture Notes in Computer Science, 2423:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230145.
                                        pdf.

Dimauro:2002:DRD

[235] G. Dimauro, S. Impedovo, M. G. Lucchese, G. Pirlo, and A. Salzo. Discovering rules for dynamic configuration of multi-classifier systems. Lecture Notes in Computer Science, 2423: 157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230157.
pdf.

Rahman:2002:MCC

[236] A. F. R. Rahman, H. Alam, and M. C. Fairhurst. Multiple classifier combination for character recognition: Revisiting the majority voting system and its variations. Lecture Notes in Computer Science, 2423:167– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230167.
pdf.

Spitz:2002:CVS

[237] A. Lawrence Spitz. Correcting for variable skew. Lecture Notes in Computer Science, 2423:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230179.
pdf.

Breuel:2002:TGA

[238] Thomas M. Breuel. Two geometric algorithms for layout analysis. Lecture Notes in Computer Science, 2423:188– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230188.
                                        pdf.

Tombre:2002:TGS

[239] Karl Tombre, Salvatore Tabbone, Loïc Pélissier, Bart Lamiroy, and Philippe Dosch. Text/graphics separation revisited. Lecture Notes in Computer Science, 2423:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230200.
                                        pdf.

Wang:2002:SDZ

[240] Yalin Wang, Ihsin T. Phillips, and Robert M. Haralick. A study on the document zone content classification problem. Lecture Notes in Computer Science, 2423:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230212.
                                        pdf.

Liang:2002:LLD

[241] Jian Liang and David Doermann. Logical labeling of document images using layout graph matching with adaptive learning. Lecture Notes in Computer Science, 2423:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230224.
pdf.

Antonacopoulos:2002:GTT

[242] A. Antonacopoulos and H. Meng. A ground-truthing tool for layout analysis performance evaluation. Lecture Notes in Computer Science, 2423:236– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230236.
pdf.

Suvichakorn:2002:SLS

[243] Aimamorn Suvichakorn, Sarin Watcharabusaracum, and Wasin Sinthupinyo. Simple layout segmentation of gray-scale document images. Lecture Notes in Computer Science, 2423:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230245.
pdf.

Wang:2002:DTH

[244] Yalin Wang and Jianying Hu. Detecting tables in HTML documents. Lecture Notes in Computer Science, 2423: 249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230249.
                                        pdf.

Sako:2002:DFI

[245] Hiroshi Sako, Naohiro Furukawa, Masakazu Fujio, and Shigeru Watanabe. Document-form identification using constellation matching of keywords abstracted by character recognition. Lecture Notes in Computer Science, 2423:261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230261.
                                        pdf.

Wang:2002:TDP

[246] Yalin Wang, Ihsin T. Phillips, and Robert M. Haralick. Table detection via probability optimization. Lecture Notes in Computer Science, 2423:272– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230272.
                                        pdf.

Amano:2002:CTF

[247] Akira Amano and Naoki Asada. Complex table form analysis using graph grammar. Lecture Notes in Computer Science, 2423:283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230283.
pdf.

Alrashed:2002:DAT

[248] Saleh Alrashed and W. A. Gray. Detection approaches for table semantics in text. Lecture Notes in Computer Science, 2423:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230287.
pdf.

Wasserman:2002:TFM

[249] Howard Wasserman, Keitaro Yukawa, Bon Sy, Kui-Lam Kwok, and Ihsin T. Phillips. A theoretical foundation and a method for document table structure extraction and decompositon. Lecture Notes in Computer Science, 2423:291– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230291.
pdf.

Antonacopoulos:2002:FSC

[250] A. Antonacopoulos and D. Karatzas. Fuzzy segmentation of characters in Web images based on human colour perception. Lecture Notes in Computer Science, 2423:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230295.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230295.
                                        pdf.

Loo:2002:WSE

[251] Poh Kok Loo and Chew Lim Tan. Word and sentence extraction using irregular pyramid. Lecture Notes in Computer Science, 2423:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230307.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230307.
                                        pdf.

Lu:2002:WSD

[252] Yue Lu and Chew Lim Tan. Word searching in document images using word portion matching. Lecture Notes in Computer Science, 2423:319– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230319.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230319.
                                        pdf.

Byun:2002:STE

[253] Hye-Ran Byun, Myung-Cheol Roh, Kil-Cheon Kim, Yeong-Woo Choi, and Seong-Whan Lee. Scene text extraction in complex images. Lecture Notes in Computer Science, 2423:329– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230329.
pdf.

Byun:2002:TED

[254] Hyeran Byun, Inyoung Jang, and Yeongwoo Choi. Text extraction in digital news video using morphology. Lecture Notes in Computer Science, 2423: 341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230341.
pdf.

Cesarini:2002:RLS

[255] Francesca Cesarini, Simone Marinai, and Giovanni Soda. Retrieval by layout similarity of documents represented with MXY trees. Lecture Notes in Computer Science, 2423:353– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230353.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230353.
pdf.

Liu:2002:AIN

[256] Qing H. Liu and Chew L. Tan. Automatic indexing of newspaper microfilm images. Lecture Notes in Computer Science, 2423:365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230365.
pdf.

Klink:2002:IDR

[257] Stefan Klink, Armin Hust, Markus Junker, and Andreas Dengel. Improving document retrieval by automatic query expansion using collaborative learning of term-based concepts. Lecture Notes in Computer Science, 2423: 376–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230376.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230376.
pdf.

Kise:2002:SWR

[258] Koichi Kise, Masaaki Tsujino, and Keinosuke Matsumoto. Spotting where to read on pages — retrieval of relevant parts from page images. Lecture Notes in Computer Science, 2423:388– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230388.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230388.
pdf.

Schmidt:2002:MDC

[259] Andreas Schmidt and Markus Junker. Mining documents for complex semantic relations by the use of context classification. Lecture Notes in Computer Science, 2423:400–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230400.
pdf.

Taghva:2002:HSE

[260] Kazem Taghva and Jeffrey Coombs. Hairetes: A search engine for OCR documents. Lecture Notes in Computer Science, 2423:412–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230412.
pdf.

Thoma:2002:TVA

[261] George R. Thoma, Glenn Ford, Daniel Le, and Zhirong Li. Text verification in an automated system for the extraction of bibliographic data. Lecture Notes in Computer Science, 2423:423– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230423.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230423.
pdf.

Dengel:2002:SRD

[262] Andreas R. Dengel and Bertin Klein. smartFIX : A requirementsdriven system for document analysis and understanding. Lecture Notes in Computer Science, 2423:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230433.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230433.
                                        pdf.

Bagdanov:2002:MSD

[263] Andrew D. Bagdanov and Marcel Worring. Multi-scale document description using rectangular granulometries. Lecture Notes in Computer Science, 2423: 445–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230445.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230445.
                                        pdf.

Wnek:2002:MLG

[264] Janusz Wnek. Machine learning of generalized document templates for data extraction. Lecture Notes in Computer Science, 2423:457–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230457.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230457.
                                        pdf.

Hadjar:2002:CRM

[265] Karim Hadjar, Oliver Hitz, Lyse Robadey, and Rolf Ingold. Configuration REcognition model for complex reverse engineering methods: 2(CREM). Lecture Notes in Computer Science, 2423:469–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230469.
pdf.

Mikheev:2002:EDP

[266] Artem Mikheev, Luc Vincent, Mike Hawrylycz, and Léon Bottou. Electronic document publishing using DjVu. Lecture Notes in Computer Science, 2423:480–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2423/24230480.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2423/24230480.pdf.

Cinque:2002:DAS

[267] L. Cinque, S. Levialdi, A. Malizia, and F. De Rosa. DAN: An automatic segmentation and classification engine for paper documents. Lecture Notes in Computer Science, 2423:491– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230491.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2423/24230491.
                                        pdf.

Lee:2002:DRE

[268] Kyong-Ho Lee, Yoon-Chul Choy, Sung-Bae Cho, Xiao Tang, and Victor Mc-Crary. Document reverse engineering: From paper to XML. Lecture Notes in Computer Science, 2423:503– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230503.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230503.
                                        pdf.

Baird:2002:HIP

[269] Henry S. Baird and Kris Popat. Human interactive proofs and document image analysis. Lecture Notes in Computer Science, 2423:507–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230507.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230507.
                                        pdf.

Todoran:2002:DGC

[270] Leon Todoran, Marcel Worring, and Arnold W. M. Smeulders. Data GroundTruth, complexity, and evaluation measures for color document analysis. Lecture Notes in Computer Science, 2423:519–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2423/24230519.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230519.
pdf.

Lopresti:2002:EWR

[271] Daniel Lopresti. Exploiting WWW resources in experimental document analysis research. Lecture Notes in Computer Science, 2423:532–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230532.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230532.
pdf.

Antonacopoulos:2002:ATC

[272] A. Antonacopoulos and D. P. Kennedy. An automated tachograph chart analysis system. Lecture Notes in Computer Science, 2423:544–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2423/24230544.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2423/24230544.
pdf.

Morde:2002:MSA

[273] Ashutosh Morde, Ramanujan S. Kashi, Michael K. Brown, Deborah Grove, and James L. Flanagan. A multimodal system for accessing driving directions. Lecture Notes in Computer Science, 2423:556– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2423/24230556.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2423/24230556.
                                        pdf.

Anonymous:2002:AIe

[274] Anonymous. Author index. Lecture Notes in Computer Science, 2423:569– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2423/2423auth.pdf.

Dieng:2002:CSW

[275] Rose Dieng. Corporate Semantic Webs. Lecture Notes in Computer Science, 2425:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2425/24250001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250001.
                                        pdf.

Pigneur:2002:FDB

[276] Yves Pigneur. A framework for defining E-business models. Lecture Notes in Computer Science, 2425:2– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250002.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250002.
                                        pdf.

Sabloniere:2002:GB

[277] Pierre Sablonière. GRID in E-business. Lecture Notes in Computer Science, 2425:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2425/24250004.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2425/24250004.
pdf.

Wang:2002:LDC

[278] Yingxu Wang. The latest development on cognitive informatics. Lecture Notes in Computer Science, 2425: 5–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250005.
pdf.

Loverdos:2002:PST

[279] Christos K. K. Loverdos, Kostas Saidis, Anya Sotiropoulou, and Dimitrios Theotokis. Pluggable services for tailorable E-content delivery. Lecture Notes in Computer Science, 2425:6– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250006.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250006.
pdf.

Gouscos:2002:OOA

[280] Dimitris Gouscos, Stathis Rouvas, Costas Vassilakis, and Panagiotis Georgiadis. An object-oriented approach for designing administrative E-forms and transactional E-services. Lecture Notes in Computer Science, 2425:19– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250019.
                                        pdf.

Calcanis:2002:TOE

[281] Costas Calcanis, Dilip Patel, and Shushma Patel. Trust objects in electronic commerce transactions. Lecture Notes in Computer Science, 2425:31– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250031.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250031.
                                        pdf.

Habela:2002:OMS

[282] Piotr Habela and Kazimierz Subieta. OODBMS metamodel supporting configuration management of large applications. Lecture Notes in Computer Science, 2425:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250040.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250040.
                                        pdf.

Roantree:2002:GAO

[283] Mark Roantree and Kazimierz Subieta. Generic applications for objectoriented databases. Lecture Notes in Computer Science, 2425:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250053.
pdf.

Sampaio:2002:VCM

[284] Sandra de F. Mendes Sampaio, Norman W. Paton, Jim Smith, and Paul Watson. Validated cost models for parallel OQL query processing. Lecture Notes in Computer Science, 2425:60– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250060.
pdf.

Martinez:2002:OOS

[285] José Martinez, Rania Lutfi, and Marc Gelgon. An object-oriented schema for querying audio. Lecture Notes in Computer Science, 2425:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250076.
pdf.

Haire:2002:SWD

[286] Brendan Haire, David Lowe, and Brian Henderson-Sellers. Supporting Web development in the OPEN process: Additional roles and techniques. Lecture Notes in Computer Science, 2425:82– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250082.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250082.
                                        pdf.

Kim:2002:SIH

[287] Hyon Hee Kim and Seung Soo Park. Semantic integration of heterogeneous XML data sources. Lecture Notes in Computer Science, 2425:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250095.
                                        pdf.

Al-Jadir:2002:FXS

[288] Lina Al-Jadir and Fatmé El-Moukaddem. F2/XML: Storing XML documents in object databases. Lecture Notes in Computer Science, 2425:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250108.
                                        pdf.

Cappi:2002:CPN

[289] Juan Cappi, Gustavo Rossi, and Andres Fortier. Customization policies need more than rule objects. Lecture Notes in Computer Science, 2425:117– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250117.
pdf.

Deneckere:2002:UMP

[290] Rébecca Deneckère. Using metapatterns to construct patterns. Lecture Notes in Computer Science, 2425: 124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250124.
pdf.

Conte:2002:TFD

[291] Agnès Conte, Mounia Fredj, Ibtissem Hassine, Jean-Pierre Giraudin, and Dominique Rieu. A tool and a formalism to design and apply patterns. Lecture Notes in Computer Science, 2425:135– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250135.
pdf.

Talens:2002:SSR

[292] Guilaine Talens, Isabelle Dedun, and Danielle Boulanger. A support system for reuse knowledge components. Lecture Notes in Computer Science, 2425:

147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250147.
                                        pdf.

Garzas:2002:OOD

[293] Javier Garzás and Mario Piattini. Object oriented design knowledge: Ontology and measurement of impact. Lecture Notes in Computer Science, 2425: 153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250153.
                                        pdf.

Kuzniarz:2002:GDM

[294] Ludwik Kuzniarz and Miroslaw Staron. Generating domain models from ontologies. Lecture Notes in Computer Science, 2425:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2425/24250160.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2425/24250160.pdf.

Sala:2002:FRC

[295] Michel Sala, Pierre Pompidor, and Danièle Hérin. A framework to review complex experimental knowledge. Lecture Notes in Computer Science, 2425: 167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250167.
pdf.

Sanchez:2002:FTU

[296] Pedro Sánchez, Patricio Letelier, Juan A. Pastor, and Juan A. Ortega. A framework to translate UML class generalization into Java code. Lecture Notes in Computer Science, 2425:173– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250173.
pdf.

Georg:2002:UAS

[297] Geri Georg and Robert France. UML aspect specification using role models. Lecture Notes in Computer Science, 2425:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2425/24250186.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2425/24250186.
pdf.

Kovse:2002:GXB

[298] Jernej Kovse and Theo Härder. Generic XMI-based UML model transformations. Lecture Notes in Computer Science, 2425:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250192.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250192.
                                        pdf.

Uhl:2002:UVM

[299] Axel Uhl and Horst Lichter. A UML variant for modeling system searchability. Lecture Notes in Computer Science, 2425:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250199.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250199.
                                        pdf.

Salinesi:2002:MFU

[300] Camille Salinesi and Jaana Wäyrynen. A methodological framework for understanding IS adaptation through enterprise change1. Lecture Notes in Computer Science, 2425:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250211.
                                        pdf.

Mirbel:2002:AAD

[301] Isabelle Mirbel and Violaine de Rivieres. Adapting analysis and design to software context: The JECKO approach. Lecture Notes in Computer Science, 2425:223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250223.
pdf.

Serour:2002:OTO

[302] M. K. Serour, B. Henderson-Sellers, J. Hughes, D. Winder, and L. Chow. Organizational transition to object technology: Theory and practice. Lecture Notes in Computer Science, 2425: 229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250229.
pdf.

Cazzola:2002:RAD

[303] Walter Cazzola, Ahmed Ghoneim, and Gunter Saake. Reflective analysis and design for adapting object run-time behavior. Lecture Notes in Computer Science, 2425:242–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250242.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250242.
pdf.

Liang:2002:GOM

[304] Ying Liang. Generation of object models for information systems from business system models. Lecture Notes in Computer Science, 2425:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250255.
                                        pdf.

Escalona:2002:RCW

[305] M. J. Escalona, J. Torres, and M. Mejías. Requirements capture workflow in global information systems. Lecture Notes in Computer Science, 2425: 267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250267.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250267.
                                        pdf.

Fabry:2002:SDE

[306] Johan Fabry. Supporting development of enterprise JavaBeans through declarative meta programming. Lecture Notes in Computer Science, 2425: 280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250280.
                                        pdf.

Osthoff:2002:EDC

[307] Carla Osthoff, Cristiana Bentes, Daniel Ariosto, Marta Mattoso, and Claudio L. Amorim. Evaluating the DSMIO cache-coherence algorithm in clusterbased parallel ODBMS. Lecture Notes in Computer Science, 2425:286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250286.
pdf.

Washizaki:2002:RTS

[308] Hironori Washizaki and Yoshiaki Fukazawa. A retrieval technique for software components using directed replaceability similarity. Lecture Notes in Computer Science, 2425:298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250298.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250298.
pdf.

Bielkowicz:2002:EIS

[309] Peter Bielkowicz, Preeti Patel, and Thein Than Tun. Evaluating information systems development methods: A new framework. Lecture Notes in Computer Science, 2425:311– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250311.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250311.
pdf.

Hagimont:2002:NFC

[310] Daniel Hagimont and Noël De Palma. Non-functional capability-based access control in the Java environment. Lecture Notes in Computer Science, 2425: 323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250323.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250323.
                                        pdf.

Wang:2002:ECA

[311] Yingxu Wang and Graham King. A European COTS architecture with built-in tests. Lecture Notes in Computer Science, 2425:336–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250336.
                                        pdf.

Benattou:2002:AOC

[312] Mohammed Benattou and Jean-Michel Bruel. Active objects for coordination in distributed testing. Lecture Notes in Computer Science, 2425:348– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250348.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250348.
                                        pdf.

Kristensen:2002:AMP

[313] Bent Bruun Kristensen. Associative modeling and programming. Lecture Notes in Computer Science, 2425:358– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250358.
pdf.

Genero:2002:CEV

[314] Marcela Genero, Luis Jiménez, and Mario Piattini. A controlled experiment for validating class diagram structural complexity metrics. Lecture Notes in Computer Science, 2425:372– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250372.
pdf.

Goedicke:2002:DSR

[315] Michael Goedicke, Klaus Pohl, and Uwe Zdun. Domain-specific runtime variability in product line architectures. Lecture Notes in Computer Science, 2425:384–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2425/24250384.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2425/24250384.pdf.

Romero:2002:MAS

[316] José Romero, Oscar Pastor, and Jorge Belenguer. Methodological approach to software quality assurance through high-level object-oriented metrics. Lecture Notes in Computer Science, 2425:

397–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250397.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250397.
                                        pdf.

Henderson-Sellers:2002:SUC

[317] B. Henderson-Sellers, D. Zowghi, T. Klemola, and S. Parasuram. Sizing use cases: How to create a standard metrical approach. Lecture Notes in Computer Science, 2425:409–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250409.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250409.
                                        pdf.

Villanova-Oliver:2002:PAS

[318] Marlène Villanova-Oliver, Jérôme Gensel, and Hervé Martin. Progressive access: A step towards adaptability in Web-based information systems. Lecture Notes in Computer Science, 2425: 422–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250422.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250422.
                                        pdf.

Amous:2002:CMD

[319] Ikram Amous, Anis Jedidi, and Florence Sèdes. A contribution to multimedia document modeling and organizing. Lecture Notes in Computer Science, 2425:434–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250434.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250434.
pdf.

Lonchamp:2002:OOC

[320] Jacques Lonchamp. An object oriented collaboration flow management system for virtual team support. Lecture Notes in Computer Science, 2425: 445–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250445.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250445.
pdf.

Traverson:2002:CCC

[321] Bruno Traverson and Nesrine Yahiaoui. Connectors for CORBA components. Lecture Notes in Computer Science, 2425:458–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2425/24250458.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2425/24250458.
pdf.

Marangozova:2002:NFR

[322] Vania Marangozova and Daniel Hagimont. Non-functional replication management in the Corba component model. Lecture Notes in Computer Science, 2425:464–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250464.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250464.
                                        pdf.

Pais:2002:RCA

[323] Ana Paula V. Pais, Bárbara O. B. Corrêa, Carlo E. T. Oliveira, and Gilson Tavares. A responsive client architecture with local object behavior deployment. Lecture Notes in Computer Science, 2425:470–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250470.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250470.
                                        pdf.

Myllymaki:2002:SPL

[324] Tommi Myllymäki, Kai Koskimies, and Tommi Mikkonen. Structuring product-lines: A layered architectural style. Lecture Notes in Computer Science, 2425:482–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250482.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250482.
                                        pdf.

Kogiomtzis:2002:IHC

[325] George Kogiomtzis and Drakoulis Martakos. Integrating heterogeneous communication and messaging systems in an object-oriented middleware framework. Lecture Notes in Computer Science, 2425:488–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250488.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250488.
pdf.

Sekharaiah:2002:OSP

[326] K. Chandra Sekharaiah and D. Janaki Ram. Object schizophrenia problem in object role system design. Lecture Notes in Computer Science, 2425:494– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250494.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250494.
pdf.

Hanenberg:2002:RAS

[327] Stefan Hanenberg and Rainer Unland. Roles and aspects: Similarities, differences, and synergetic potential. Lecture Notes in Computer Science, 2425: 507–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2425/24250507.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2425/24250507.
pdf.

Pfeifer:2002:FOO

[328] Daniel Pfeifer. Flexible object-oriented views using method propagation. Lecture Notes in Computer Science, 2425: 521–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250521.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250521.
                                        pdf.

Cherfi:2002:TAR

[329] Samira Si-Said Cherfi and Nadira Lammari. Towards an assisted reorganization of Is_{A} hierarchies. Lecture Notes in Computer Science, 2425:536– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2425/24250536.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2425/24250536.
                                        pdf.

Anonymous:2002:AIf

[330] Anonymous. Author index. Lecture Notes in Computer Science, 2425:549– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2425/2425ind.pdf.

Hermanns:2002:I

[331] Holger Hermanns. 1. introduction. Lecture Notes in Computer Science, 2428: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2428/24280001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2428/24280001.
pdf.

Hermanns:2002:IP

[332] Holger Hermanns. 2. interactive processes. Lecture Notes in Computer Science, 2428:7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2428/24280007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2428/24280007.
pdf.

Hermanns:2002:MC

[333] Holger Hermanns. 3. Markov chains. Lecture Notes in Computer Science, 2428:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2428/24280035.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2428/24280035.
pdf.

Hermanns:2002:IMCa

[334] Holger Hermanns. 4. interactive Markov chains. Lecture Notes in Computer Science, 2428:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2428/24280057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2428/24280057.
pdf.

Hermanns:2002:AIM

[335] Holger Hermanns. 5. algebra of interactive Markov chains. Lecture Notes in Computer Science, 2428:89– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2428/24280089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2428/24280089.
                                        pdf.

Hermanns:2002:IMCb

[336] Holger Hermanns. 6. interactive Markov chains in practice. Lecture Notes in Computer Science, 2428:129– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2428/24280129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2428/24280129.
                                        pdf.

Hermanns:2002:C

[337] Holger Hermanns. 7. conclusions. Lecture Notes in Computer Science, 2428: 155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2428/24280155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2428/24280155.
                                        pdf.

Hermanns:2002:APC

[338] Holger Hermanns. Appendix A: Proofs for chapters 3 and 4. Lecture Notes in Computer Science, 2428:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2428/24280161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2428/24280161.
pdf.

Hermanns:2002:ABP

[339] Holger Hermanns. Appendix B: Proofs for chapter 5. Lecture Notes in Computer Science, 2428:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2428/24280175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2428/24280175.
pdf.

Anonymous:2002:B

[340] Anonymous. Bibliography. Lecture Notes in Computer Science, 2428:207– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2428/24280207.pdf.

Bergmann:2002:I

[341] Ralph Bergmann. 1. introduction. Lecture Notes in Computer Science, 2432: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2432/24320001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2432/24320001.
pdf.

Bergmann:2002:EM

[342] Ralph Bergmann. 2. experience management. Lecture Notes in Computer Science, 2432:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2432/24320025.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2432/24320025.pdf.

Bergmann:2002:RE

[343] Ralph Bergmann. 3. representing experience. Lecture Notes in Computer Science, 2432:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2432/24320047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2432/24320047.
                                        pdf.

Bergmann:2002:AEU

[344] Ralph Bergmann. 4. assessing experience utility. Lecture Notes in Computer Science, 2432:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2432/24320093.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2432/24320093.
                                        pdf.

Bergmann:2002:RKA

[345] Ralph Bergmann. 5. representing knowledge for adaptation. Lecture Notes in Computer Science, 2432:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2432/24320141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2432/24320141.
pdf.

Bergmann:2002:UC

[346] Ralph Bergmann. 6. user communication. Lecture Notes in Computer Science, 2432:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2432/24320155.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2432/24320155.pdf.

Bergmann:2002:ER

[347] Ralph Bergmann. 7. experience retrieval. Lecture Notes in Computer Science, 2432:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2432/24320187.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2432/24320187.pdf.

Bergmann:2002:EA

[348] Ralph Bergmann. 8. experience adaptation. Lecture Notes in Computer Science, 2432:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2432/24320219.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2432/24320219.pdf.

Bergmann:2002:DME

[349] Ralph Bergmann. 9. developing and maintaining experience management applications. Lecture Notes in Computer Science, 2432:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2432/24320241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2432/24320241.
                                        pdf.

Bergmann:2002:EMEa

[350] Ralph Bergmann. 10. experience management for electronic commerce. Lecture Notes in Computer Science, 2432: 281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2432/24320281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2432/24320281.
                                        pdf.

Bergmann:2002:EMS

[351] Ralph Bergmann. 11. experience management for self-service and help-desk support. Lecture Notes in Computer Science, 2432:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2432/24320315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2432/24320315.
                                        pdf.

Bergmann:2002:EMEb

[352] Ralph Bergmann. 12. experience management for electronic design reuse. Lecture Notes in Computer Science, 2432:347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2432/24320347.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2432/24320347.
pdf.

Bergmann:2002:LS

[353] Ralph Bergmann. List of symbols. Lecture Notes in Computer Science, 2432: 365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2432/24320365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2432/24320365.
pdf.

Anonymous:2002:R

[354] Anonymous. References. Lecture Notes in Computer Science, 2432:367– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2432/2432refe.pdf.

Anonymous:2002:I

[355] Anonymous. Index. Lecture Notes in Computer Science, 2432:391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2432/2432indx.pdf.

Jang:2002:RTI

[356] Heejin Jang and Sangwook Kim. Realtime intruder tracing through selfreplication. Lecture Notes in Computer Science, 2433:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2433/24330001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2433/24330001.pdf.

Kwon:2002:DPP

[357] Taekyoung Kwon. On the difficulty of protecting private keys in software. Lecture Notes in Computer Science, 2433:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2433/24330017.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2433/24330017.
pdf.

Baras:2002:IDS

[358] John S. Baras and Maben Rabi. Intrusion detection with support vector machines and generative models. Lecture Notes in Computer Science, 2433: 32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330032.
pdf.

Satoh:2002:SHS

[359] Akashi Satoh and Sumio Morioka. Small and high-speed hardware architectures for the 3GPP standard cipher KASUMI. Lecture Notes in Computer Science, 2433:48–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330048.
pdf.

Lipmaa:2002:FSI

[360] Helger Lipmaa. Fast software implementations of SC2000. Lecture Notes in Computer Science, 2433:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330063.
pdf.

Grembowski:2002:CAH

[361] Tim Grembowski, Roar Lien, Kris Gaj, Nghi Nguyen, Peter Bellows, Jaroslav Flidr, Tom Lehman, and Brian Schott. Comparative analysis of the hardware implementations of hash functions SHA-1 and SHA-512. Lecture Notes in Computer Science, 2433:75– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330075.
pdf.

Jallad:2002:ICC

[362] Kahil Jallad, Jonathan Katz, and Bruce Schneier. Implementation of chosen-ciphertext attacks against PGP and GnuPG. Lecture Notes in Computer Science, 2433:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330090.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330090.
                                        pdf;   http://www.counterpane.com/
                                        pgp-attack.html.

Navarro:2002:RBA

[363] G. Navarro, S. Robles, and J. Borrell. Role-based access control for E-commerce sea-of-data applications. Lecture Notes in Computer Science, 2433:102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2433/24330102.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330102.
                                        pdf.

Gabillon:2002:ACM

[364] Alban Gabillon, Manuel Munier, Jean-Jacques Bascou, Laurent Gallon, and Emmanuel Bruno. An access control model for tree data structures. Lecture Notes in Computer Science, 2433: 117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330117.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2433/24330117.
pdf.

Dawson:2002:NDP

[365] Ed Dawson, Javier Lopez, Jose A. Montenegro, and Eiji Okamoto. A new design of privilege management infrastructure for organizations using outsourced PKI. Lecture Notes in Computer Science, 2433:136–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330136.
pdf.

Zhu:2002:PAK

[366] Feng Zhu, Duncan S. Wong, Agnes H. Chan, and Robbie Ye. Password authenticated key exchange based on RSA for imbalanced wireless networks. Lecture Notes in Computer Science, 2433:150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2433/24330150.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2433/24330150.
pdf.

Hsu:2002:QPL

[367] Tsan sheng Hsu, Churn-Jung Liau, Da-Wei Wang, and Jeremy K.-P. Chen. Quantifying privacy leakage through answering database queries. Lecture Notes in Computer Science, 2433:162– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330162.
                                        pdf.

Qiu:2002:NOP

[368] Weidong Qiu, Kefei Chen, and Dawu Gu. A new offline privacy protecting E-cash system with revokable anonymity. Lecture Notes in Computer Science, 2433:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330177.
                                        pdf.

Abe:2002:RFS

[369] Masayuki Abe and Koutarou Suzuki. Receipt-free sealed-bid auction. Lecture Notes in Computer Science, 2433: 191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330191.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330191.
                                        pdf.

Gonzalez-Deleito:2002:EFM

[370] Nicolás González-Deleito and Olivier Markowitch. Exclusion-freeness in multi-party exchange protocols. Lecture Notes in Computer Science, 2433: 200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2433/24330200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330200.
pdf.

Ferrer-Gomila:2002:RPM

[371] Josep Lluís Ferrer-Gomila, Magdalena Payeras-Capellà, and Llorenç Huguet-Rotger. A realistic protocol for multiparty certified electronic mail. Lecture Notes in Computer Science, 2433:210– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330210.
pdf.

Araki:2002:NRS

[372] Shunsuke Araki. A Nyberg-Rueppel signature for multiple messages and its batch verification. Lecture Notes in Computer Science, 2433:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330220.
pdf.

Martinez-Nadal:2002:CUM

[373] Apol·lònia Martínez-Nadal and Josep Lluís Ferrer-Gomila. Comments to the UNCITRAL model law on electronic signatures. Lecture Notes in Computer Science, 2433:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330229.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330229.
                                        pdf.

Shigetomi:2002:ALS

[374] Rie Shigetomi, Akira Otsuka, Takahide Ogawa, and Hideki Imai. An anonymous loan system based on group signature scheme. Lecture Notes in Computer Science, 2433:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330244.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330244.
                                        pdf.

Yonezawa:2002:TSS

[375] Shoko Yonezawa, Goichiro Hanaoka, Junji Shikata, and Hideki Imai. Traceability schemes for signed documents. Lecture Notes in Computer Science, 2433:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2433/24330257.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330257.
                                        pdf.

Bresson:2002:PKN

[376] Emmanuel Bresson and Jacques Stern. Proofs of knowledge for non-monotone discrete-log formulae and applications. Lecture Notes in Computer Science, 2433:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330272.
pdf.

Kim:2002:IDS

[377] Nam-Yeun Kim, Dae-Ghon Kho, and Kee-Young Yoo. Inversion/division systolic architecture for public-key cryptosystems in GF(2^{m} ). Lecture Notes in Computer Science, 2433:289– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330289.
pdf.

Kwon:2002:EBS

[378] Soonhak Kwon and Heuisu Ryu. Efficient bit serial multiplication using optimal normal bases of type II in GF (2^{m}). Lecture Notes in Computer Science, 2433:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330300.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330300.
pdf.

Watanabe:2002:CCD

[379] Yuji Watanabe and Masayuki Numao. Conditional cryptographic delegation for P2P data sharing. Lecture Notes in Computer Science, 2433:309– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330309.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330309.
                                        pdf.

Chen:2002:CPK

[380] L. Chen, K. Harrison, A. Moss, D. Soldera, and N. P. Smart. Certification of public keys within an identity based system. Lecture Notes in Computer Science, 2433:322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330322.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330322.
                                        pdf.

Zheng:2002:NPK

[381] Jiande Zheng. A new public key cryptosystem for constrained hardware. Lecture Notes in Computer Science, 2433:334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2433/24330334.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330334.
                                        pdf.

Daza:2002:DCS

[382] Vanesa Daza, Javier Herranz, Carles Padró, and Germán Sáez. A distributed and computationally secure key distribution scheme. Lecture Notes in Computer Science, 2433:342– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330342.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330342.
pdf.

Lipmaa:2002:OHT

[383] Helger Lipmaa. On optimal hash tree traversal for interval time-stamping. Lecture Notes in Computer Science, 2433:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2433/24330357.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2433/24330357.
pdf.

Goodrich:2002:EDD

[384] Michael T. Goodrich, Roberto Tamassia, and Jasminka Hasić. An efficient dynamic and distributed cryptographic accumulator. Lecture Notes in Computer Science, 2433:372–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330372.
pdf.

Okeya:2002:SOD

[385] Katsuyuki Okeya and Kouichi Sakurai. A second-order DPA attack breaks a window-method based countermeasure against side channel attacks. Lecture Notes in Computer Science, 2433:389– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330389.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330389.
                                        pdf.

Moller:2002:PEC

[386] Bodo Möller. Parallelizable elliptic curve point multiplication method with resistance against sidechannel attacks. Lecture Notes in Computer Science, 2433:402–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330402.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330402.
                                        pdf.

Martinelli:2002:AAS

[387] Fabio Martinelli, Marinella Petrocchi, and Anna Vaccarelli. Automated analysis of some security mechanisms of SCEP. Lecture Notes in Computer Science, 2433:414–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330414.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330414.
                                        pdf.

Monteiro:2002:APC

[388] José R. M. Monteiro and Ricardo Dahab. An attack on a protocol for certified delivery. Lecture Notes in Computer Science, 2433:428–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330428.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330428.
pdf.

Kikuchi:2002:OCM

[389] Hiroaki Kikuchi. Oblivious counter and majority protocol. Lecture Notes in Computer Science, 2433:437–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330437.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330437.
pdf.

Soo:2002:EMC

[390] Wai Han Soo, Azman Samsudin, and Alwyn Goh. Efficient mental card shuffling via optimised arbitrary-sized Benes permutation network. Lecture Notes in Computer Science, 2433:446– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2433/24330446.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2433/24330446.
pdf.

Fernandez:2002:FCC

[391] M. Fernandez and M. Soriano. Fingerprinting concatenated codes with efficient identification. Lecture Notes in Computer Science, 2433:459–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330459.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330459.
                                        pdf.

Domingo-Ferrer:2002:PSA

[392] Josep Domingo-Ferrer. A provably secure additive and multiplicative privacy homomorphism. Lecture Notes in Computer Science, 2433:471–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330471.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330471.
                                        pdf.

Sakai:2002:AES

[393] Yasuyuki Sakai and Kouichi Sakurai. Algorithms for efficient simultaneous elliptic scalar multiplication with reduced joint Hamming weight representation of scalars. Lecture Notes in Computer Science, 2433:484– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2433/24330484.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2433/24330484.
                                        pdf.

Anonymous:2002:AIg

[394] Anonymous. Author index. Lecture Notes in Computer Science, 2433:501– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2433/2433auth.pdf.

Atzeni:2002:TCW

[395] Paolo Atzeni. Time: A coordinate for Web site modelling. Lecture Notes in Computer Science, 2435:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350001.
pdf.

Boyens:2002:TEP

[396] Claus Boyens and Oliver Günther. Trust is not enough: Privacy and security in ASP and Web service environments. Lecture Notes in Computer Science, 2435:8–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350008.
pdf.

Schek:2002:IIS

[397] Hans-Jörg Schek, Heiko Schuldt, Christoph Schuler, and Roger Weber. Infrastructure for information spaces. Lecture Notes in Computer Science, 2435:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2435/24350023.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2435/24350023.
                                        pdf.

Giannella:2002:AAD

[398] Chris Giannella. An axiomatic approach to defining approximation measures for functional dependencies. Lecture Notes in Computer Science, 2435: 37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350037.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350037.
                                        pdf.

Koval:2002:ISI

[399] Robert Koval and Pavol Návrat. Intelligent support for information retrieval in WWW environment. Lecture Notes in Computer Science, 2435:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350051.
                                        pdf.

Zhou:2002:AIT

[400] Shuigeng Zhou and Jihong Guan. An approach to improve text classification efficiency. Lecture Notes in Computer Science, 2435:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350065.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2435/24350065.
pdf.

Polcicova:2002:SSC

[401] Gabriela Polčicová and Pavol Návrat. Semantic similarity in content-based filtering. Lecture Notes in Computer Science, 2435:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2435/24350080.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2435/24350080.pdf.

Wojciechowski:2002:DAP

[402] Marek Wojciechowski and Maciej Zakrzewicz. Data access paths for frequent itemsets discovery. Lecture Notes in Computer Science, 2435:86– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350086.
pdf.

Ilarri:2002:MCL

[403] Sergio Ilarri, Eduardo Mena, and Arantza Illarramendi. Monitoring continuous location queries using mobile agents. Lecture Notes in Computer Science, 2435:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350092.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2435/24350092.
pdf.

Lee:2002:OCC

[404] Ukhyun Lee and Buhyun Hwang. Optimistic concurrency control based on timestamp interval for broadcast environment. Lecture Notes in Computer Science, 2435:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350106.
pdf.

Samaras:2002:FPA

[405] George Samaras and Christoforos Panayiotou. A flexible personalization architecture for wireless Internet based on mobile agents. Lecture Notes in Computer Science, 2435:120– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350120.
pdf.

Shigiltchoff:2002:MDB

[406] Oleg Shigiltchoff, Panos K. Chrysanthis, and Evaggelia Pitoura. Multiversion data broadcast organizations. Lecture Notes in Computer Science, 2435: 135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350135.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2435/24350135.
pdf.

Brakatsoulas:2002:RRT

[407] Sotiris Brakatsoulas, Dieter Pfoser, and Yannis Theodoridis. Revisiting Rtree construction principles. Lecture Notes in Computer Science, 2435:149– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350149.
pdf.

Corral:2002:AAD

[408] Antonio Corral, Joaquin Cañadas, and Michael Vassilakopoulos. Approximate algorithms for distancebased queries in high-dimensional data spaces using R-trees. Lecture Notes in Computer Science, 2435:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350163.
pdf.

Jurado:2002:ESS

[409] Elena Jurado and Manuel Barrena. Efficient similarity search in feature spaces with the Q-tree. Lecture Notes in Computer Science, 2435:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2435/24350177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350177.
                                        pdf.

El-Geresy:2002:STG

[410] Baher A. El-Geresy, Alia I. Abdelmoty, and Christopher B. Jones. Spatiotemporal geographic information systems: A causal perspective. Lecture Notes in Computer Science, 2435:191– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350191.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350191.
                                        pdf.

Kwon:2002:AMI

[411] Joon-Hee Kwon and Yong-Ik Yoon. An access method for integrating multiscale geometric data. Lecture Notes in Computer Science, 2435:204–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350204.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350204.
                                        pdf.

Akal:2002:OQE

[412] Fuat Akal, Klemens Böhm, and Hans-Jörg Schek. OLAP query evaluation in a database cluster: A performance study on intra-query parallelism. Lecture Notes in Computer Science, 2435: 218–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2435/24350218.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350218.
pdf.

Medina:2002:SRM

[413] Enrique Medina and Juan Trujillo. A standard for representing multidimensional properties: The common warehouse metamodel (CWM). Lecture Notes in Computer Science, 2435:232– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350232.
pdf.

Caplinskas:2002:FAE

[414] Albertas Caplinskas, Audrone Lupeikiene, and Olegas Vasilecas. A framework to analyse and evaluate information systems specification languages. Lecture Notes in Computer Science, 2435:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350248.
pdf.

Habela:2002:FMO

[415] Piotr Habela, Mark Roantree, and Kazimierz Subieta. Flattening the metamodel for object databases. Lecture Notes in Computer Science, 2435:263– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350263.
                                        pdf.

Parama:2002:SQO

[416] José R. Paramá, Nieves R. Brisaboa, Miguel R. Penabad, and Ángeles S. Places. A semantic query optimization approach to optimize linear datalog programs. Lecture Notes in Computer Science, 2435:277–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350277.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350277.
                                        pdf.

Zamulin:2002:OAO

[417] Alexandre Zamulin. An object algebra for the ODMG standard. Lecture Notes in Computer Science, 2435:291– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350291.
                                        pdf.

Feyer:2002:MDS

[418] Thomas Feyer and Bernhard Thalheim. Many-dimensional schema modeling. Lecture Notes in Computer Science, 2435:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2435/24350305.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2435/24350305.
pdf.

Konovalov:2002:OOD

[419] Alexandre Konovalov. Object-oriented data model for data warehouse. Lecture Notes in Computer Science, 2435: 319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350319.
pdf.

Eder:2002:MMS

[420] Johann Eder and Wolfgang Gruber. A meta model for structured workflows supporting workflow transformations. Lecture Notes in Computer Science, 2435:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350326.
pdf.

Grinev:2002:TES

[421] Maxim Grinev and Sergey Kuznetsov. Towards an exhaustive set of rewriting rules for XQuery optimization: BizQuery experience. Lecture Notes in Computer Science, 2435:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350340.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350340.
                                        pdf.

Marathe:2002:ABQ

[422] Mona Marathe and Hemalatha Diwakar. Architecture of a blended-query and result-visualization mechanism for Web-accessible databases and associated implementation issues. Lecture Notes in Computer Science, 2435:346– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350346.
                                        pdf.

Stavrakas:2002:ACS

[423] Yannis Stavrakas, Manolis Gergatsoulis, Christos Doulkeridis, and Vassilis Zafeiris. Accommodating changes in semistructured databases using multidimensional OEM. Lecture Notes in Computer Science, 2435:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2435/24350360.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2435/24350360.
                                        pdf.

Liu:2002:DWE

[424] Jixue Liu and Chengfei Liu. A declarative way of extracting XML data in XSL. Lecture Notes in Computer Science, 2435:374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350374.
pdf.

Dolog:2002:TVM

[425] Peter Dolog and Mária Bieliková. Towards variability modelling for reuse in hypermedia engineering. Lecture Notes in Computer Science, 2435:388– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350388.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350388.
pdf.

Harada:2002:CTP

[426] Lilian Harada. Complex temporal patterns detection over continuous data streams. Lecture Notes in Computer Science, 2435:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2435/24350401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2435/24350401.
pdf.

Anonymous:2002:AIh

[427] Anonymous. Author index. Lecture Notes in Computer Science, 2435:415– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2435/24350415.pdf.

Chang:2002:GBD

[428] Shi-Kuo Chang. A growing book for distance learning. Lecture Notes in Computer Science, 2436:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360003.
                                        pdf.

Shen:2002:OFS

[429] Ruimin Shen, Peng Han, Fan Yang, Qiang Yang, and Joshua Zhexe Huang. An open framework for smart and personalized distance learning. Lecture Notes in Computer Science, 2436:19– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360019.
                                        pdf.

Song:2002:MFD

[430] William W. Song. A metadata framework for description of learning objects. Lecture Notes in Computer Science, 2436:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2436/24360031.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2436/24360031.
pdf.

Jun:2002:WBM

[431] Woochun Jun, Le Gruenwald, Jungmoon Park, and Suk ki Hong. A Webbased motivation-supporting model for effective teaching-learning. Lecture Notes in Computer Science, 2436:44– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360044.
pdf.

Jung:2002:DIW

[432] Hyosook Jung, Woochun Jun, Le Gruenwald, and Suk ki Hong. The design and implementation of a Webbased teaching-learning model for information communication technology application education. Lecture Notes in Computer Science, 2436:56– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360056.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360056.
pdf.

Wang:2002:AMS

[433] Feng-Hsu Wang. On analysis and modeling of student browsing behavior in Web-based asynchronous learning environments. Lecture Notes in Computer Science, 2436:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360069.
                                        pdf.

Luo:2002:CNB

[434] Sanding Luo, Sha Sha, Deyao Shen, and Weijia Jia. Conceptual network based courseware navigation and Web presentation mechanisms. Lecture Notes in Computer Science, 2436:81– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360081.
                                        pdf.

Albermani:2002:WBK

[435] Faris Albermani and Kwok wing Chau. Web-based knowledge-based system on liquid retaining structure design as instructional tool. Lecture Notes in Computer Science, 2436:95– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360095.
                                        pdf.

Gamper:2002:WBL

[436] Johann Gamper and Judith Knapp. A Web-based language learning system. Lecture Notes in Computer Science, 2436:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2436/24360106.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2436/24360106.
pdf.

Liu:2002:WWB

[437] Chunnian Liu, Lei Zheng, Junzhong Ji, Dong Jia, Chuan Jiang, and Wensheng Yang. WebMath: A Webbased ITS system. Lecture Notes in Computer Science, 2436:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360119.
pdf.

Shi:2002:SRC

[438] Yuanchun Shi, Weikai Xie, and Guangyou Xu. Smart remote classroom: Creating a revolutionary real-time interactive distance learning system. Lecture Notes in Computer Science, 2436:130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360130.
pdf.

Wang:2002:CSF

[439] Hao Wang, Heng Xu, Hock Chuan Chan, and Ling Chen. Critical success factors for Web-based organizational IT training systems. Lecture Notes in Computer Science, 2436:142– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360142.
                                        pdf.

Kwan:2002:AQA

[440] Reggie Kwan and Jimmy S. F. Chan. Application of quality assurance in Web-based laboratories. Lecture Notes in Computer Science, 2436:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360154.
                                        pdf.

Qu:2002:TOS

[441] Changtao Qu and Wolfgang Nejdl. Towards open standards: The evolution of an XML/JSP/WebDAV based collaborative courseware generating system. Lecture Notes in Computer Science, 2436:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360167.
                                        pdf.

Bouras:2002:ECU

[442] Christos Bouras, Dimitrios Psaltoulis, Christos Psaroudis, and Thrasyvoulos Tsiatsos. An educational community using collaborative virtual environments. Lecture Notes in Computer Science, 2436:180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2436/24360180.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2436/24360180.pdf.

Zhuge:2002:KKG

[443] Hai Zhuge, Yanyan Li, Jia Bi, and To yat Cheung. KGCL: A knowledge-grid-based cooperative learning environment. Lecture Notes in Computer Science, 2436:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360192.
pdf.

Lee:2002:CBA

[444] Yugyung Lee and James Geller. A component-based architecture for adaptive, collaborative Web-based learning. Lecture Notes in Computer Science, 2436:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2436/24360203.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2436/24360203.pdf.

Cao:2002:VPL

[445] Jiannong Cao, Alvin Chan, Weidong Cao, and Cassidy Yeung. Virtual programming lab for online distance learning. Lecture Notes in Computer Science, 2436:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360216.
pdf.

Xu:2002:VGG

[446] Zhiwei Xu, Wei Li, Hongguang Fu, and Zhenbing Zeng. The vega grid and grid-based education. Lecture Notes in Computer Science, 2436:228– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360228.
pdf.

Chin:2002:ERF

[447] Francis Y. L. Chin, C. K. M. Wong, and V. Mak. Experiences in running a flexible, Web-based, and self-paced course. Lecture Notes in Computer Science, 2436:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360241.
pdf.

Kurbel:2002:VTL

[448] Karl Kurbel. Virtual teaching and learning based on multimedia techniques: Experiences from a master program delivered over the Internet. Lecture Notes in Computer Science, 2436: 252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360252.
pdf.

Barrett:2002:EDL

[449] John Barrett, Nadia Hing-Wah Ng, and Kathy Edwards. Experiences in developing a large-scale school IT project: Processes, outcomes, and issues. Lecture Notes in Computer Science, 2436: 263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360263.
pdf.

Klamma:2002:ECC

[450] Ralf Klamma, Matthias Jarke, Elisabeth Hollender, and Dagmar Börner-Klein. Enabling communities by constructed media: The case of a Web-based study environment for a talmudic tractate. Lecture Notes in Computer Science, 2436:275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360275.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360275.
                                        pdf.

Cheung:2002:ITT

[451] Ronnie C. T. Cheung. Innovative teaching through the cyber university. Lecture Notes in Computer Science, 2436:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360287.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360287.
                                        pdf.

Tang:2002:SPW

[452] Changjie Tang, Qing Li, Rynson W. H. Lau, and Xiaodong Huang. Supporting practices in Web-based learning. Lecture Notes in Computer Science, 2436: 300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360300.
                                        pdf.

Okada:2002:WBL

[453] Akira Okada and Yahiko Kambayashi. A Web-based lecture video database system with flexible indexing method using action logs. Lecture Notes in Computer Science, 2436:313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360313.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2436/24360313.
pdf.

Park:2002:LBT

[454] Chanjung Park and Hwakyung Rim. A locking-based transaction scheduling algorithm for supporting Web-based classes. Lecture Notes in Computer Science, 2436:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2436/24360325.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2436/24360325.pdf.

Zhang:2002:DMS

[455] Jianting Zhang, Le Gruenwald, Chris Candler, Gary McNutt, and Wei Shung Chung. Database and metadata support of a Web-based multimedia digital library for medical education. Lecture Notes in Computer Science, 2436: 339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360339.
pdf.

Sumitomo:2002:DEL

[456] Chisa Sumitomo, Akira Okada, and Yahiko Kambayashi. Design and evaluation of lecture support functions for question databases. Lecture Notes in Computer Science, 2436:351– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360351.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360351.
                                        pdf.

Wang:2002:UPC

[457] Daling Wang, Yubin Bao, Ge Yu, and Guoren Wang. Using page classification and association rule mining for personalized recommendation in distance learning. Lecture Notes in Computer Science, 2436:363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360363.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360363.
                                        pdf.

So:2002:WBV

[458] Simon So. Web-based video indexing and retrieval for teaching and learning. Lecture Notes in Computer Science, 2436:377–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360377.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360377.
                                        pdf.

Huang:2002:AMA

[459] Yueh-Min Huang, Chin-Chiang Chang, Pin-Ying Chiu, Shin-Chang Huang, and William Cheng-Chung Chu. Applying multimedia authoring tool and XML techniques to standardized knowledge management for Webbased learning. Lecture Notes in Computer Science, 2436:386–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360386.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360386.
pdf.

Fong:2002:FNF

[460] Joseph Fong, Irene Kwan, Margaret Ng, and K. L. Lau. Fifth normal form made easy with novel Webbased CAI HCI. Lecture Notes in Computer Science, 2436:398–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2436/24360398.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2436/24360398.
pdf.

Zhuang:2002:MKE

[461] Yueting Zhuang and Xiang Liu. Multimedia knowledge exploitation for Elearning: Some enabling techniques. Lecture Notes in Computer Science, 2436:411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2436/24360411.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2436/24360411.
pdf.

Yang:2002:WBI

[462] Li Yang and Mustafa Sanver. Webbased interactive 3D visualization for computer graphics education. Lecture Notes in Computer Science, 2436:423– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2436/24360423.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2436/24360423.
                                        pdf.

Anonymous:2002:AIi

[463] Anonymous. Author index. Lecture Notes in Computer Science, 2436:433– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2436/24360433.pdf.

Master:2002:AAC

[464] Paul Master. The age of adaptive computing is here. Lecture Notes in Computer Science, 2438:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380001.
                                        pdf.

Hartenstein:2002:DTD

[465] Reiner W. Hartenstein. Disruptive trends by data-stream-based computing. Lecture Notes in Computer Science, 2438:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2438/24380004.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2438/24380004.pdf.

Brebner:2002:MLC

[466] Gordon Brebner. Multithreading for logic-centric systems. Lecture Notes in Computer Science, 2438:5– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380005.
pdf.

Sarmadi:2002:FPC

[467] Siavash Bayat Sarmadi, Seyed Ghassem Miremadi, Ghazanfar Asadi, and Ali Reza Ejlali. Fast prototyping with co-operation of simulation and emulation. Lecture Notes in Computer Science, 2438:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380015.
pdf.

Krupnova:2002:HFR

[468] Helena Krupnova, Veronique Meurou, Christophe Barnichon, Carlos Serra, and Farid Morsi. How fast is rapid FPGA-based prototyping: Lessons and challenges from the digital TV design prototyping project. Lecture Notes in Computer Science, 2438:26– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380026.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380026.
                                        pdf.

Ho:2002:IAC

[469] Quoc Thai Ho, Jean-Baptiste Rigaud, Laurent Fesquet, Marc Renaudin, and Robin Rolland. Implementing asynchronous circuits on LUT based FPGAs. Lecture Notes in Computer Science, 2438:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380036.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380036.
                                        pdf.

Martino:2002:TFS

[470] Beniamino Di Martino, Nicola Mazzocca, Giacinto Paolo Saggese, and Antonio G. M. Strollo. A technique for FPGA synthesis driven by automatic source code analysis and transformations. Lecture Notes in Computer Science, 2438:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380047.
                                        pdf.

Compton:2002:FRA

[471] Katherine Compton, Akshay Sharma, Shawn Phillips, and Scott Hauck. Flexible routing architecture generation for domain-specific reconfigurable subsystems. Lecture Notes in Computer Science, 2438:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380059.
pdf.

Khan:2002:IVP

[472] Jawad Khan, Manish Handa, and Ranga Vemuri. iPACE-V1: A portable adaptive computing engine for real time applications. Lecture Notes in Computer Science, 2438:69– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380069.
pdf.

Sima:2002:FPC

[473] Mihai Sima, Stamatis Vassiliadis, Sorin Cotofana, Jos T. J. van Eijndhoven, and Kees Vissers. Fieldprogrammable custom computing machines — A taxonomy. Lecture Notes in Computer Science, 2438:79– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380079.
pdf.

Leijten-Nowak:2002:ERL

[474] Katarzyna Leijten-Nowak and Jef L. van Meerbergen. Embedded reconfigurable logic core for DSP applications. Lecture Notes in Computer Science, 2438:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2438/24380089.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380089.
                                        pdf.

Cardells-Tormo:2002:EFB

[475] Francisco Cardells-Tormo, Javier Valls-Coquillat, Vicenc Almenar-Terre, and Vicente Torres-Carot. Efficient FPGAbased QPSK demodulation loops: Application to the DVB standard. Lecture Notes in Computer Science, 2438: 102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380102.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380102.
                                        pdf.

Dick:2002:FQD

[476] Chris Dick and Fred Harris. FPGA QAM demodulator design. Lecture Notes in Computer Science, 2438:112– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380112.
                                        pdf.

Lemieux:2002:AFS

[477] Guy G. Lemieux and David M. Lewis. Analytical framework for switch block design. Lecture Notes in Computer Science, 2438:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380122.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380122.
pdf.

Koorapaty:2002:MFS

[478] Aneesh Koorapaty and Lawrence Pileggi. Modular, fabric-specific synthesis for programmable architectures. Lecture Notes in Computer Science, 2438: 132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380132.
pdf.

Fan:2002:ODU

[479] Hongbing Fan, Jiping Liu, Yu-Liang Wu, and Chak-Chung Cheung. On optimum designs of universal switch blocks. Lecture Notes in Computer Science, 2438:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380142.
pdf.

Robertson:2002:IFS

[480] Ian Robertson, James Irvine, Patrick Lysaght, and David Robinson. Improved functional simulation of dynamically reconfigurable logic. Lecture Notes in Computer Science, 2438:152– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380152.
pdf.

Lopez-Buedo:2002:RTR

[481] S. Lopez-Buedo, P. Riviere, P. Pernas, and E. Boemo. Run-time reconfiguration to check temperature in custom computers: An application of JBits technology. Lecture Notes in Computer Science, 2438:162– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380162.
pdf.

Smit:2002:DRM

[482] Gerard J. M. Smit, Paul J. M. Havinga, Lodewijk T. Smit, Paul M. Heysters, and Michel A. J. Rosien. Dynamic reconfiguration in mobile systems. Lecture Notes in Computer Science, 2438: 171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380171.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2438/24380171.
pdf.

Horta:2002:UPI

[483] Edson L. Horta, John W. Lockwood, and Sérgio T. Kofuji. Using PAR-BIT to implement partial run-time reconfigurable systems. Lecture Notes in Computer Science, 2438:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380182.
pdf.

Turner:2002:MLR

[484] R. H. Turner, R. Woods, and T. Courtney. Multiplier-less realization of a poly-phase filter using LUT-based FPGAs. Lecture Notes in Computer Science, 2438:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380192.
pdf.

Melnikoff:2002:SRF

[485] Stephen J. Melnikoff, Steven F. Quigley, and Martin J. Russell. Speech recognition on an FPGA using discrete and continuous hidden Markov models. Lecture Notes in Computer Science, 2438:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380202.
                                        pdf.

Jamin:2002:FIW

[486] Antony Jamin and Petri Mähönen. FPGA implementation of the wavelet packet transform for high speed communications. Lecture Notes in Computer Science, 2438:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380212.
                                        pdf.

Carreira:2002:MIB

[487] A. Carreira, T. W. Fox, and L. E. Turner. A method for implementing bit-serial finite impulse response digital filters in FPGAs using JBits^{TM}. Lecture Notes in Computer Science, 2438: 222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380222.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380222.
                                        pdf.

Manohararajah:2002:API

[488] Valavan Manohararajah, Terry Borer, Stephen D. Brown, and Zvonko Vranesic. Automatic partitioning for improved placement and routing in complex programmable logic devices. Lecture Notes in Computer Science, 2438:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2438/24380232.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2438/24380232.
pdf.

Kannan:2002:RRR

[489] Parivallal Kannan, Shankar Balachandran, and Dinesh Bhatia. Rapid and reliable routability estimation for FP-GAs. Lecture Notes in Computer Science, 2438:242–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2438/24380242.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2438/24380242.pdf.

Danek:2002:IIA

[490] Martin Danek and Zdenek Muzikár. Integrated iterative approach to FPGA placement. Lecture Notes in Computer Science, 2438:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380253.
pdf.

Cabral:2002:TDM

[491] Lucídio A. F. Cabral, Júlio S. Aude, and Nelson Maculan. TDR: A distributed-memory parallel routing algorithm for FPGAs. Lecture Notes in Computer Science, 2438:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380263.
                                        pdf.

Kielbik:2002:HLP

[492] Rafal Kielbik, Juan Manuel Moreno, Andrzej Napieralski, Grzegorz Jablonski, and Tomasz Szymanski. Highlevel partitioning of digital systems based on dynamically reconfigurable devices. Lecture Notes in Computer Science, 2438:271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380271.
                                        pdf.

Yamaguchi:2002:HSH

[493] Yoshiki Yamaguchi, Yosuke Miyajima, Tsutomu Maruyama, and Akihiko Konagaya. High speed homology search using run-time reconfiguration. Lecture Notes in Computer Science, 2438:281– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380281.
                                        pdf.

Dyer:2002:PRC

[494] Matthias Dyer, Christian Plessl, and Marco Platzner. Partially reconfigurable cores for Xilinx virtex. Lecture Notes in Computer Science, 2438:292– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380292.
pdf.

Gericota:2002:LDR

[495] Manuel G. Gericota, Gustavo R. Alves, Miguel L. Silva, and José M. Ferreira. On-line defragmentation for run-time partially reconfigurable FPGAs. Lecture Notes in Computer Science, 2438: 302–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380302.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380302.
pdf.

Poon:2002:FPM

[496] Kara K. W. Poon, Andy Yan, and Steven J. E. Wilton. A flexible power model for FPGAs. Lecture Notes in Computer Science, 2438:312– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380312.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380312.
pdf.

Cadenas:2002:CTP

[497] Oswaldo Cadenas and Graham Megson. A clocking technique with power savings in virtex-based pipelined designs. Lecture Notes in Computer Science, 2438:322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2438/24380322.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2438/24380322.pdf.

Martina:2002:EER

[498] Maurizio Martina, Guido Masera, Gianluca Piccinini, Fabrizio Vacca, and Maurizio Zamboni. Energy evaluation on a reconfigurable, multimediaoriented wireless sensor. Lecture Notes in Computer Science, 2438:332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380332.
pdf.

Todorovich:2002:TAE

[499] E. Todorovich, M. Gilabert, G. Sutter, S. Lopez-Buedo, and E. Boemo. A tool for activity estimation in FPGAs. Lecture Notes in Computer Science, 2438: 340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380340.
pdf.

Sutter:2002:FDL

[500] Gustavo Sutter, Elias Todorovich, Sergio Lopez-Buedo, and Eduardo Boemo. FSM decomposition for low power in FPGA. Lecture Notes in Computer Science, 2438:350–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380350.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380350.
pdf.

Nam:2002:HRF

[501] Gi-Joon Nam, Karem Sakallah, and Rob Rutenbar. Hybrid routing for FP-GAs by integrating Boolean satisfiability with geometric search. Lecture Notes in Computer Science, 2438:360– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380360.
pdf.

Benkrid:2002:PBH

[502] K. Benkrid, D. Crookes, A. Benkrid, and S. Belkacemi. A Prolog-based hardware development environment. Lecture Notes in Computer Science, 2438:370–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2438/24380370.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2438/24380370.
pdf.

Ho:2002:FMH

[503] C. H. Ho, P. H. W. Leong, K. H. Tsoi, R. Ludewig, P. Zipf, A. G. Ortiz, and M. Glesner. Fly — A modifiable hardware compiler. Lecture Notes in Computer Science, 2438:381– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380381.
                                        pdf.

Bolsens:2002:COF

[504] Ivo Bolsens. Challenges and opportunities for FPGA platforms. Lecture Notes in Computer Science, 2438:391– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380391.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380391.
                                        pdf.

Kirimura:2002:DIF

[505] Masayuki Kirimura, Yoshifumi Takamoto, Takanori Mori, Keiichi Yasumoto, Akio Nakata, and Teruo Higashino. Design and implementation of FPGA circuits for high speed network monitors. Lecture Notes in Computer Science, 2438: 393–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380393.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380393.
                                        pdf.

Gokhale:2002:GTG

[506] Maya Gokhale, Dave Dubois, Andy Dubois, Mike Boorman, Steve Poole, and Vic Hogsett. Granidt: Towards gigabit rate network intrusion detection technology. Lecture Notes in Computer Science, 2438:404–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380404.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380404.
pdf.

Zhou:2002:FSH

[507] Kuan Zhou, Channakeshav, Jong-Ru Guo, Chao You, Bryan S. Goda, Russell P. Kraft, and John F. Mc-Donald. Fast SiGe HBT BiCMOS FPGAs with new architecture and power saving techniques. Lecture Notes in Computer Science, 2438:414– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380414.
pdf.

Hall:2002:FPA

[508] Tyson S. Hall, Paul Hasler, and David V. Anderson. Field-programmable analog arrays: A floating-gate approach. Lecture Notes in Computer Science, 2438:424–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2438/24380424.

                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2438/24380424.pdf.

Tanigawa:2002:GEM

[509] Kazuya Tanigawa, Tetsuo Hironaka, Akira Kojima, and Noriyoshi Yoshida. A generalized execution model for programming on reconfigurable architectures and an architecture supporting the model. Lecture Notes in Computer Science, 2438:434–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380434.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380434.
                                        pdf.

Pionteck:2002:FTR

[510] T. Pionteck, P. Zipf, L. D. Kabulepa, and M. Glesner. A framework for teaching (re)configurable architectures in student projects. Lecture Notes in Computer Science, 2438: 444–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380444.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380444.
                                        pdf.

Cho:2002:SHD

[511] Young H. Cho, Shiva Navab, and William H. Mangione-Smith. Specialized hardware for deep network packet filtering. Lecture Notes in Computer Science, 2438:452–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380452.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380452.
pdf.

Buerner:2002:ISE

[512] Thomas Buerner. Implementation of a successive erasure BCH(16,7,6) decoder and performance simulation by rapid prototyping. Lecture Notes in Computer Science, 2438:462–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380462.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380462.
pdf.

Ramirez:2002:FRF

[513] J. Ramírez, A. García, U. Meyer-Baese, and A. Lloris. Fast RNS FPL-based communications receiver design and implementation. Lecture Notes in Computer Science, 2438:472– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380472.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380472.
pdf.

Haynes:2002:URA

[514] Simon D. Haynes, Henry G. Epsom, Richard J. Cooper, and Paul L. McAlpine. UltraSONIC: A reconfigurable architecture for video image processing. Lecture Notes in Computer Science, 2438:482–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380482.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380482.
                                        pdf.

Fox:2002:IDC

[515] Trevor W. Fox and Laurence E. Turner. Implementing the discrete cosine transform using the Xilinx virtex FPGA. Lecture Notes in Computer Science, 2438:492–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380492.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380492.
                                        pdf.

Staller:2002:IJS

[516] Alexander Staller, Peter Dillinger, and Reinhard Männer. Implementation of the JPEG 2000 standard on a virtex 1000 FPGA. Lecture Notes in Computer Science, 2438:503– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380503.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380503.
                                        pdf.

Beuchat:2002:SMB

[517] Jean-Luc Beuchat and Arnaud Tisserand. Small multiplier-based multiplication and division operators for virtex-II devices. Lecture Notes in Computer Science, 2438:513–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380513.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380513.
pdf.

Gaffar:2002:ACF

[518] Altaf Abdul Gaffar, Wayne Luk, Peter Y. K. Cheung, Nabeel Shirazi, and James Hwang. Automating customisation of floating-point designs. Lecture Notes in Computer Science, 2438: 523–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380523.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380523.
pdf.

Jang:2002:EEM

[519] Ju wook Jang, Seonil Choi, and Viktor K. Prasanna. Energy-efficient matrix multiplication on FPGAs. Lecture Notes in Computer Science, 2438:534– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380534.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380534.
pdf.

Seng:2002:RTA

[520] Shay Seng, Wayne Luk, and Peter Y. K. Cheung. Run-time adaptive flexible instruction processors. Lecture Notes in Computer Science, 2438:545– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380545.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380545.
                                        pdf.

deSousa:2002:DDA

[521] José T. de Sousa, Fernando M. Gonçalves, Nuno Barreiro, and João Moura. DARP — A digital audio reconfigurable processor. Lecture Notes in Computer Science, 2438:556– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380556.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380556.
                                        pdf.

Charlwood:2002:SLM

[522] Stephen Charlwood, Jonathan Mangnall, and Steven Quigley. System-level modelling for performance estimation of reconfigurable coprocessors. Lecture Notes in Computer Science, 2438:567– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380567.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380567.
                                        pdf.

Ting:2002:FBS

[523] Kurt K. Ting, Steve C. L. Yuen, K. H. Lee, and Philip H. W. Leong. An FPGA based SHA-256 processor. Lecture Notes in Computer Science, 2438: 577–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380577.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380577.
pdf.

Zipf:2002:HFF

[524] P. Zipf, M. Glesner, C. Bauer, and H. Wojtkowiak. Handling FPGA faults and configuration sequencing using a hardware extension. Lecture Notes in Computer Science, 2438:586– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380586.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380586.
pdf.

Krasniewski:2002:STP

[525] Andrzej Krasniewski. On the set of target path delay faults in sequential subcircuits of LUT-based FPGAs. Lecture Notes in Computer Science, 2438: 596–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380596.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380596.
pdf.

Rebaudengo:2002:SBA

[526] M. Rebaudengo, M. Sonza Reorda, and M. Violante. Simulationbased analysis of SEU effects on SRAM-based FPGAs. Lecture Notes in Computer Science, 2438:607–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380607.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380607.
pdf.

Krasniewski:2002:ERE

[527] Andrzej Krasniewski. Exploiting reconfigurability for effective testing of delay faults in sequential subcircuits of LUT-based FPGAs. Lecture Notes in Computer Science, 2438:616–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380616.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380616.
pdf.

Matousek:2002:LNS

[528] Rudolf Matousek, Milan Tichý, Zdenek Pohl, Jirí Kadlec, Chris Softley, and Nick Coleman. Logarithmic number system and floating-point arithmetics on FPGA. Lecture Notes in Computer Science, 2438:627–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380627.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2438/24380627.
pdf.

Roesler:2002:NOH

[529] Eric Roesler and Brent Nelson. Novel optimizations for hardware floating-point units in a modern FPGA architecture. Lecture Notes in Computer Science, 2438:637–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380637.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380637.
pdf.

Chiricescu:2002:MM

[530] Silviu Chiricescu, Michael Schuette, Robin Glinton, and Herman Schmit. Morphable multipliers. Lecture Notes in Computer Science, 2438:647–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380647.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380647.
pdf.

Belanovic:2002:LPF

[531] Pavle Belanovic and Miriam Leeser. A library of parameterized floatingpoint modules and their use. Lecture Notes in Computer Science, 2438:657– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380657.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2438/24380657.
                                        pdf.

Stansfield:2002:WAP

[532] Tony Stansfield. Wordlength as an architectural parameter for reconfigurable computing devices. Lecture Notes in Computer Science, 2438:667– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380667.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380667.
                                        pdf.

Baleani:2002:EPF

[533] Massimo Baleani, Massimo Conti, Alberto Ferrari, Valerio Frascolla, and Alberto Sangiovanni-Vincentelli. An enhanced POLIS framework for fast exploration and implementation of I/O subsystems on CSoC platforms. Lecture Notes in Computer Science, 2438: 677–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380677.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380677.
                                        pdf.

Wigley:2002:IRO

[534] Grant B. Wigley, David A. Kearney, and David Warren. Introducing ReConfigME: An operating system for reconfigurable computing. Lecture Notes in Computer Science, 2438:687– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2438/24380687.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380687.
pdf.

Sidhu:2002:EMU

[535] Reetinder Sidhu and Viktor K. Prasanna. Efficient metacomputation using self-reconfiguration. Lecture Notes in Computer Science, 2438:698– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380698.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380698.
pdf.

Arias-Estrada:2002:FCP

[536] Miguel Arias-Estrada and Eduardo Rodríguez-Palacios. An FPGA coprocessor for real-time visual tracking. Lecture Notes in Computer Science, 2438:710–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2438/24380710.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2438/24380710.
pdf.

Fischer:2002:IDA

[537] Viktor Fischer, Milos Drutarovský, and Rastislav Lukac. Implementation of 3-D adaptive LUM smoother in reconfigurable hardware. Lecture Notes in Computer Science, 2438:720– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380720.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380720.
                                        pdf.

Amira:2002:CCB

[538] A. Amira, A. Bouridane, P. Milligan, and F. Bensaali. Custom coprocessor based matrix algorithms for image and signal processing. Lecture Notes in Computer Science, 2438:730– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380730.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380730.
                                        pdf.

Aranki:2002:PFI

[539] Nazeeh Aranki, Alex Moopenn, and Raoul Tawel. Parallel FPGA implementation of the split and merge discrete wavelet transform. Lecture Notes in Computer Science, 2438:740– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380740.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380740.
                                        pdf.

Kerins:2002:FPE

[540] Tim Kerins, Emanuel Popovici, William Marnane, and Patrick Fitzpatrick. Fully parameterizable elliptic curve cryptography processor over GF(2). Lecture Notes in Computer Science, 2438:750–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380750.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380750.
pdf.

Hamalainen:2002:GPS

[541] Antti Hämäläinen, Matti Tommiska, and Jorma Skyttä. 6.78 gigabits per second implementation of the IDEA cryptographic algorithm. Lecture Notes in Computer Science, 2438:760–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380760.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380760.
pdf.

Moreira:2002:RCE

[542] Emmanuel A. Moreira, Paul L. McAlpine, and Simon D. Haynes. Rijndael cryptographic engine on the Ultra-SONIC reconfigurable platform. Lecture Notes in Computer Science, 2438: 770–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380770.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380770.
pdf.

Quisquater:2002:CTM

[543] Jean-Jacques Quisquater, François-Xavier Standaert, Gael Rouvroy, Jean-Pierre David, and Jean-Didier Legat.

A cryptanalytic time-memory tradeoff: First FPGA implementation. Lecture Notes in Computer Science, 2438:780– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380780.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380780.
                                        pdf.

Lauwereins:2002:CWS

[544] Rudy Lauwereins. Creating a world of smart re-configurable devices. Lecture Notes in Computer Science, 2438: 790–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380790.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380790.
                                        pdf.

Marescaux:2002:INE

[545] Théodore Marescaux, Andrei Bartic, Dideriek Verkest, Serge Vernalde, and Rudy Lauwereins. Interconnection networks enable fine-grain dynamic multi-tasking on FPGAs. Lecture Notes in Computer Science, 2438:795– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380795.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380795.
                                        pdf.

Landaker:2002:MHS

[546] Wesley J. Landaker, Michael J. Wirthlin, and Brad L. Hutchings. Multitasking hardware on the SLAAC1-V reconfigurable computing system. Lecture Notes in Computer Science, 2438:806– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380806.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380806.
pdf.

Valtonen:2002:CFG

[547] Tuomas Valtonen, Jouni Isoaho, and Hannu Tenhunen. The case for fine-grained re-configurable architectures: An analysis of conceived performance. Lecture Notes in Computer Science, 2438:816–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380816.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380816.
pdf.

Kokosinski:2002:FIM

[548] Zbigniew Kokosinski and Wojciech Sikora. An FPGA implementation of a multi-comparand multi-search associative processor. Lecture Notes in Computer Science, 2438:826–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380826.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380826.
pdf.

Labbe:2002:AIF

[549] Anna Labbé and Annie Pérez. AES implementation on FPGA: Time — flexibility tradeoff. Lecture Notes in Computer Science, 2438:836–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380836.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380836.
                                        pdf.

Koeune:2002:FIL

[550] François Koeune, Gael Rouvroy, François-Xavier Standaert, Jean-Jacques Quisquater, Jean-Pierre David, and Jean-Didier Legat. An FPGA implementation of the linear cryptanalysis. Lecture Notes in Computer Science, 2438:845–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2438/24380845.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380845.
                                        pdf.

Budiu:2002:CAS

[551] Mihai Budiu and Seth Copen Goldstein. Compiling application-specific hardware. Lecture Notes in Computer Science, 2438:853–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2438/24380853.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2438/24380853.pdf.

Cardoso:2002:XVC

[552] João M. P. Cardoso and Markus Weinhardt. XPP-VC: A C compiler with temporal partitioning for the PACT-XPP architecture. Lecture Notes in Computer Science, 2438:864– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380864.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380864.
pdf.

Tripp:2002:SCS

[553] Justin L. Tripp, Preston A. Jackson, and Brad L. Hutchings. Sea cucumber: A synthesizing compiler for FPGAs. Lecture Notes in Computer Science, 2438:875–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380875.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380875.
pdf.

Carletta:2002:PCS

[554] J. E. Carletta and M. D. Rayman. Practical considerations in the synthesis of high performance digital filters for implementation on FPGAs. Lecture Notes in Computer Science, 2438: 886–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380886.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2438/24380886.
                                        pdf.

Meyer-Baese:2002:LPH

[555] U. Meyer-Baese, J. Ramírez, and A. García. Low power high speed algebraic integer frequency sampling filters using FPLDs. Lecture Notes in Computer Science, 2438:897–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380897.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380897.
                                        pdf.

Cardells-Tormo:2002:HPQ

[556] Francisco Cardells-Tormo and Javier Valls-Coquillat. High performance quadrature digital mixers for FPGAs. Lecture Notes in Computer Science, 2438:905–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2438/24380905.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380905.
                                        pdf.

Mencer:2002:HEM

[557] Oskar Mencer, Zhining Huang, and Lorenz Huelsbergen. HAGAR: Efficient multi-context graph processors. Lecture Notes in Computer Science, 2438: 915–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380915.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2438/24380915.
pdf.

Schafer:2002:SID

[558] Benjamin Carrión Schäfer, S. F. Quigley, and A. H. C. Chan. Scalable implementation of the discrete element method on a reconfigurable computing platform. Lecture Notes in Computer Science, 2438:925–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380925.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380925.
pdf.

Lam:2002:CTC

[559] Kai-Pui Lam and Sui-Tung Mak. On computing transitive-closure equivalence sets using a hybrid GA-DP approach. Lecture Notes in Computer Science, 2438:935–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2438/24380935.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2438/24380935.pdf.

Salcic:2002:RPC

[560] Zoran Salcic, Partha Roop, Morteza Biglari-Abhari, and Abbas Bigdeli. REFLIX: A processor core for reactive embedded applications. Lecture Notes in Computer Science, 2438:945– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2438/24380945.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380945.
                                        pdf.

Venkataramani:2002:FIP

[561] Girish Venkataramani, Suraj Sudhir, Mihai Budiu, and Seth Copen Goldstein. Factors influencing the performance of a CPU-RFU hybrid architecture. Lecture Notes in Computer Science, 2438:955–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380955.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380955.
                                        pdf.

Sanz:2002:ICF

[562] A. Sanz, J. I. García-Nicolás, and I. Urriza. Implementing converters in FPLD. Lecture Notes in Computer Science, 2438:966–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24380966.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24380966.
                                        pdf.

Benitez:2002:QUP

[563] Domingo Benitez. A quantitative understanding of the performance of reconfigurable coprocessors. Lecture Notes in Computer Science, 2438:976– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2438/24380976.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380976.
pdf.

Buchenrieder:2002:IRH

[564] Klaus Buchenrieder, Ulrich Nageldinger, Andreas Pyttel, and Alexander Sedlmeier. Integration of reconfigurable hardware into system-level design. Lecture Notes in Computer Science, 2438:987– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380987.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380987.
pdf.

Wolz:2002:RMG

[565] Frank Wolz and Reiner Kolla. A retargetable macro generation method for the evaluation of repetitive configurable architectures. Lecture Notes in Computer Science, 2438:997–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24380997.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24380997.
pdf.

Ramaswamy:2002:ISH

[566] Ramaswamy Ramaswamy and Russell Tessier. The integration of SystemC and hardware-assisted verification. Lecture Notes in Computer Science, 2438:1007–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2438/24381007.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381007.
                                        pdf.

Kaviani:2002:UDH

[567] Alireza S. Kaviani. Using design hierarchy to improve quality of results in FPGAs. Lecture Notes in Computer Science, 2438:1017–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381017.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381017.
                                        pdf.

Koutroumpezis:2002:ADR

[568] G. Koutroumpezis, K. Tatas, D. Soudris, S. Blionas, K. Masselos, and A. Thanailakis. Architecture design of a reconfigurable multiplier for flexible coarsegrain implementations. Lecture Notes in Computer Science, 2438:1027–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381027.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381027.
                                        pdf.

Kaneko:2002:GHD

[569] Naoto Kaneko and Hideharu Amano. A general hardware design model for multicontext FPGAs. Lecture Notes in Computer Science, 2438:1037–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381037.
pdf.

Porrmann:2002:DRH

[570] Mario Porrmann, Ulf Witkowski, Heiko Kalte, and Ulrich Rückert. Dynamically reconfigurable hardware — A new perspective for neural network implementations. Lecture Notes in Computer Science, 2438:1048–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381048.
pdf.

David:2002:CFD

[571] Raphael David, Daniel Chillet, Sebastien Pillement, and Olivier Sentieys. A compilation framework for a dynamically reconfigurable architecture. Lecture Notes in Computer Science, 2438:1058–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381058.
pdf.

Ichikawa:2002:DDC

[572] Shuichi Ichikawa and Shoji Yamamoto. Data dependent circuit for subgraph isomorphism problem. Lecture Notes in Computer Science, 2438:1068–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381068.
                                        pdf.

Schmidt:2002:EDS

[573] Jan Schmidt, Martin Novotný, Martin Jäger, Milos Becvár, and Michal Jáchim. Exploration of design space in ECDSA. Lecture Notes in Computer Science, 2438:1072–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381072.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381072.
                                        pdf.

Damaj:2002:CGA

[574] Issam Damaj, Suhaib Majzoub, and Hassan Diab. 2D and 3D computer graphics algorithms under MORPHOSYS. Lecture Notes in Computer Science, 2438:1076–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381076.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381076.
                                        pdf.

Blionas:2002:HIR

[575] S. Blionas, K. Masselos, C. Dre, C. Drosos, F. Ieromnimon, T. Pagonis, A. Pneymatikakis, A. Tatsaki, T. Trimis, A. Vontzalidis, and D. Metafas. A HIPER-LAN/2 — IEEE 802.11a reconfigurable system-on-chip. Lecture Notes in Computer Science, 2438:1080–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381080.
pdf.

McBader:2002:SFI

[576] Stephanie McBader, Luca Clementel, Alvise Sartori, Andrea Boni, and Peter Lee. SoftTOTEM: An FPGA implementation of the TOTEM parallel processor. Lecture Notes in Computer Science, 2438:1084–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381084.
pdf.

Yokota:2002:RTM

[577] Takashi Yokota, Masamichi Nagafuchi, Yoshito Mekada, Tsutomu Yoshinaga, Kanemitsu Ootsu, and Takanobu Baba. Real-time medical diagnosis on a multiple FPGA-based system. Lecture Notes in Computer Science, 2438:1088– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2438/24381088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381088.
                                        pdf.

Aoyama:2002:TEB

[578] Kazuo Aoyama and Hiroshi Sawada. Threshold element-based symmetric function generators and their functional extension. Lecture Notes in Computer Science, 2438:1092–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381092.
                                        pdf.

Schlecker:2002:HIM

[579] Wolfgang Schlecker, Achim Engelhart, Werner G. Teich, and Hans-Jörg Pfleiderer. Hardware implementation of a multiuser detection scheme based on recurrent neural networks. Lecture Notes in Computer Science, 2438:1097– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381097.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381097.
                                        pdf.

Hwang:2002:BCF

[580] James Hwang and Jonathan Ballagh. Building custom FIR filters using system generator. Lecture Notes in Computer Science, 2438:1101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381101.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381101.
pdf.

Feske:2002:SBL

[581] Klaus Feske, Georg Heinrich, Berndt Fritzsche, and Mark Langer. SoC based low cost design of digital audio broadcasting transport network applications. Lecture Notes in Computer Science, 2438:1105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381105.
pdf.

Jamro:2002:DCC

[582] Ernest Jamro and Kazimierz Wiatr. Dynamic constant coefficient convolvers implemented in FPGAs. Lecture Notes in Computer Science, 2438: 1110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2438/24381110.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2438/24381110.
pdf.

Kanus:2002:VIF

[583] Urs Kanus, Gregor Wetekam, Johannes Hirche, and Michael Meißner. VIZARD II: An FPGA-based interactive volume rendering system. Lecture Notes in Computer Science, 2438:1114– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381114.
                                        pdf.

Izu:2002:RNR

[584] Naoyuki Izu, Tomonori Yokoyama, Junichiro Tsuchiya, Konosuke Watanabe, and Hideharu Amano. RHiNET/ NI: A reconfigurable network interface for cluster computing. Lecture Notes in Computer Science, 2438:1118– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381118.
                                        pdf.

Miletic:2002:GPP

[585] Filip Miletic, Rene van Leuken, and Alexander de Graaf. General purpose prototyping platform for dataprocessor research and development. Lecture Notes in Computer Science, 2438:1122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2438/24381122.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381122.
                                        pdf.

Kobori:2002:HSC

[586] Tomoyoshi Kobori and Tsutomu Maruyama. High speed computation of three dimensional cellular automata with FPGA. Lecture Notes in Computer Science, 2438:1126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381126.
pdf.

Poussier:2002:SBE

[587] Sylvain Poussier, Hassan Rabah, and Serge Weber. SOPC-based embedded smart strain gage sensor. Lecture Notes in Computer Science, 2438:1131– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381131.
pdf.

Ha:2002:AHS

[588] Yajun Ha, Radovan Hipik, Serge Vernalde, Diederik Verkest, Marc Engels, Rudy Lauwereins, and Hugo De Man. Adding hardware support to the HotSpot virtual machine for domain specific applications. Lecture Notes in Computer Science, 2438:1135– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381135.
pdf.

Gaudino:2002:FBN

[589] Roberto Gaudino, Vito De Feo, M. Chiaberge, and C. Sansoè. An FPGA-based node controller for a high capacity WDM optical packet network. Lecture Notes in Computer Science, 2438:1139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381139.
                                        pdf.

Calmon:2002:FMF

[590] F. Calmon, M. Fathallah, P. J. Viverge, C. Gontrand, J. Carrabina, and P. Foussier. FPGA and mixed FPGA-DSP implementations of electrical drive algorithms. Lecture Notes in Computer Science, 2438:1144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381144.
                                        pdf.

Melis:2002:IRR

[591] Wim J. C. Melis, Peter Y. K. Cheung, and Wayne Luk. Image registration of real-time broadcast video using the UltraSONIC reconfigurable computer. Lecture Notes in Computer Science, 2438:1148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2438/24381148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381148.
pdf.

Carline:2002:NWT

[592] Dylan Carline and Paul Coulton. A novel watermarking technique for LUT based FPGA designs. Lecture Notes in Computer Science, 2438:1152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381152.
pdf.

Henz:2002:ICL

[593] Martin Henz, Edgar Tan, and Roland H. C. Yap. Implementing CSAT local search on FPGAs. Lecture Notes in Computer Science, 2438:1156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381156.
pdf.

Niyonkuru:2002:RPA

[594] Adronis Niyonkuru, Göran Eggers, and Hans Christoph Zeidler. A reconfigurable processor architecture. Lecture Notes in Computer Science, 2438:1160– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381160.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381160.
                                        pdf.

Friebe:2002:RSC

[595] Sebastian Friebe, Steffen Köhler, Rainer G. Spallek, Henrik Juhr, and Klaus Künanz. A reconfigurable system-on-chip-based fast EDM process monitor. Lecture Notes in Computer Science, 2438:1164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381164.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381164.
                                        pdf.

Guccione:2002:GMU

[596] Steven A. Guccione and Eric Keller. Gene matching using JBits. Lecture Notes in Computer Science, 2438:1168– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2438/24381168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2438/24381168.
                                        pdf.

Saab:2002:MPR

[597] Daniel G. Saab, Fatih Kocan, and Jacob A. Abraham. Massively parallel/reconfigurable emulation model for the D-algorithm. Lecture Notes in Computer Science, 2438:1172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2438/24381172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381172.
pdf.

Miyashita:2002:PRA

[598] Akira Miyashita, Toshihito Fujiwara, and Tsutomu Maruyama. A placement/routing approach for FPGA accelerators. Lecture Notes in Computer Science, 2438:1177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2438/24381177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2438/24381177.
pdf.

Anonymous:2002:AIj

[599] Anonymous. Author index. Lecture Notes in Computer Science, 2438:1183– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2438/2438indx.pdf.

Neuwirth:2002:GTT

[600] Christine M. Neuwirth. Groupware and text technologies. Lecture Notes in Computer Science, 2440:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2440/24400003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2440/24400003.
pdf.

Stahl:2002:GGS

[601] Gerry Stahl. Groupware goes to school. Lecture Notes in Computer Science, 2440:7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2440/24400007.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2440/24400007.
pdf.

Ferreira:2002:CLW

[602] J. M. Martins Ferreira, Gustavo R. C. Alves, Ricardo Costa, and Nick Hine. Collaborative learning in a Web-accessible workbench. Lecture Notes in Computer Science, 2440:25– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2440/24400025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2440/24400025.
pdf.

Baloian:2002:DMC

[603] Nelson Baloian, Alexander Berges, Stephan Buschmann, Katrin Gaßner, Jens Hardings, Heinz Ulrich Hoppe, and Wolfram Luther. Document management in a computer-integrated classroom. Lecture Notes in Computer Science, 2440:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2440/24400035.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2440/24400035.pdf.

Antunes:2002:HCM

[604] Pedro Antunes and Carlos J. Costa. Handheld CSCW in the meeting environment. Lecture Notes in Computer Science, 2440:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2440/24400047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2440/24400047.
pdf.

Lopez:2002:ECM

[605] Juan Carlos Leyva López and Dina Esperanza López Elizalde. Effect of the coordination modes in supporting group multiple criteria decision making in a distributed and asynchronous environment. Lecture Notes in Computer Science, 2440:61– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2440/24400061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2440/24400061.
pdf.

Wang:2002:CVH

[606] Weigang Wang, Jörg M. Haake, and Jessica Rubart. A cooperative visual hypermedia approach to planning and conducting virtual meetings. Lecture Notes in Computer Science, 2440:70– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2440/24400070.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2440/24400070.
                                        pdf.

Rubart:2002:TUG

[607] Jessica Rubart and Peter Dawabi. Towards UML-G: A UML profile for modeling groupware. Lecture Notes in Computer Science, 2440:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2440/24400093.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2440/24400093.
                                        pdf.

Ochoa:2002:DCI

[608] Sergio F. Ochoa, Luis A. Guerrero, David A. Fuller, and Oriel Herrera. Designing the communications infrastructure of groupware systems. Lecture Notes in Computer Science, 2440: 114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2440/24400114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2440/24400114.
                                        pdf.

Blois:2002:CBA

[609] Ana P. T. Bacelo Blois and Karin Becker. A component-based architecture to support collaborative application design. Lecture Notes in Computer Science, 2440:134–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2440/24400134.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2440/24400134.
pdf.

Moran:2002:BGT

[610] Alberto L. Morán, Jesus Favela, Ana M. Martínez-Enríquez, and Dominique Decouchant. Before getting there: Potential and actual collaboration. Lecture Notes in Computer Science, 2440:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2440/24400147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2440/24400147.
pdf.

Alarcon:2002:IAS

[611] Rosa Alarcón and David Fuller. Intelligent awareness in support of collaborative virtual work groups. Lecture Notes in Computer Science, 2440:168– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2440/24400168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2440/24400168.
pdf.

Leiva-Lobos:2002:OFP

[612] Edmundo P. Leiva-Lobos and Eliana Covarrubias. The 3-ontology: A framework to place cooperative awareness. Lecture Notes in Computer Science, 2440:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2440/24400189.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2440/24400189.
                                        pdf.

Collazos:2002:ECL

[613] César A. Collazos, Luis A. Guerrero, José A. Pino, and Sergio F. Ochoa. Evaluating collaborative learning processes. Lecture Notes in Computer Science, 2440:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2440/24400203.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2440/24400203.
                                        pdf.

deAraujo:2002:CLG

[614] Renata Mendes de Araujo, Flavia M. Santoro, and Marcos R. S. Borges. The CSCW lab for groupware evaluation. Lecture Notes in Computer Science, 2440:222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2440/24400222.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2440/24400222.
                                        pdf.

Fernandez:2002:TGW

[615] Alejandro Fernández, Jörg M. Haake, and Adele Goldberg. Tailoring group work. Lecture Notes in Computer Science, 2440:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2440/24400232.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2440/24400232.pdf.

Khlifi:2002:BGD

[616] Hechmi Khlifi, Jocelyn Desbiens, and Mohamed Cheriet. Building groupwares over duplicated object systems. Lecture Notes in Computer Science, 2440:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2440/24400245.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2440/24400245.
pdf.

Lukosch:2002:ATD

[617] Stephan Lukosch. Adaptive and transparent data distribution support for synchronous groupware. Lecture Notes in Computer Science, 2440:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2440/24400255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2440/24400255.
pdf.

David:2002:CWI

[618] José Maria N. David and Marcos R. S. Borges. COPSE-Web: An infrastructure for developing Web-based groupware applications. Lecture Notes in Computer Science, 2440:275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2440/24400275.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2440/24400275.
                                        pdf.

Anonymous:2002:AIk

[619] Anonymous. Author index. Lecture Notes in Computer Science, 2440:285– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2440/2440auth.pdf.

Murphy:2002:EAS

[620] Sean Murphy and Matthew J. B. Robshaw. Essential algebraic structure within the AES. Lecture Notes in Computer Science, 2442:1–16, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        URL http://www.isg.rhul.ac.uk/~
                                        mrobshaw/aes-crypto.pdf.

Joux:2002:BAA

[621] Antoine Joux, Gwenaëlle Martinet, and Frédéric Valette. Blockwiseadaptive attackers: Revisiting the (in)security of some provably secure encryption modes: CBC, GEM, IACBC. Lecture Notes in Computer Science, 2442:17–30, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420017.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420017.pdf.

Liskov:2002:TBC

[622] Moses Liskov, Ronald L. Rivest, and David Wagner. Tweakable block ciphers. Lecture Notes in Computer Science, 2442:31–46, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420031.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420031.pdf.

Halevy:2002:LBE

[623] Dani Halevy and Adi Shamir. The LSD broadcast encryption scheme. Lecture Notes in Computer Science, 2442:47–60, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420047.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420047.pdf.

Camenisch:2002:DAA

[624] Jan Camenisch and Anna Lysyanskaya. Dynamic accumulators and application to efficient revocation of anonymous credentials. Lecture Notes in Computer Science, 2442:61–76, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420143.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420143.pdf.

Hopper:2002:PSS

[625] Nicholas J. Hopper, John Langford, and Luis von Ahn. Provably secure steganography: (extended abstract). Lecture Notes in Computer Science, 2442:77–92, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420077.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420077.pdf.

Stern:2002:FAP

[626] Jacques Stern, David Pointcheval, John Malone-Lee, and Nigel P. Smart. Flaws in applying proof methodologies to signature schemes. Lecture Notes in Computer Science, 2442:93–110, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420093.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420093.pdf.

Nielsen:2002:SRO

[627] Jesper Buus Nielsen. Separating random oracle proofs from complexity theoretic proofs: The noncommitting encryption case. Lecture Notes in Computer Science, 2442: 111–126, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2442/24420111.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2442/24420111.
                                        pdf.

Jonsson:2002:SRE

[628] Jakob Jonsson and Burton S. Kaliski, Jr. On the security of RSA encryption in TLS. Lecture Notes in Computer Science, 2442:127–142, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420017.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420017.pdf.

Canetti:2002:SAI

[629] Ran Canetti and Hugo Krawczyk. Security analysis of IKE’s signaturebased key-exchange protocol. Lecture Notes in Computer Science, 2442: 143–161, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2442/24420143.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2442/24420143.
pdf.

Bellare:2002:GSI

[630] Mihir Bellare and Adriana Palacio. GQ and Schnorr identification schemes: Proofs of security against impersonation under active and concurrent attacks. Lecture Notes in Computer Science, 2442:162–177, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2501/25010397.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2501/25010397.pdf.

Gennaro:2002:RSM

[631] Rosario Gennaro, Yuval Ishai, Eyal Kushilevitz, and Tal Rabin. On 2-round secure multiparty computation. Lecture Notes in Computer Science, 2442:178–193, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2501/25010548.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2501/25010548.pdf.

Blaser:2002:PCC

[632] Markus Bläser, Andreas Jakoby, Maciej Liskiewicz, and Bodo Siebert. Private computation — k-connected versus 1-connected networks. Lecture Notes in Computer Science, 2442: 194–209, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2420/24200519.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2420/24200519.
                                        pdf.

Nguyen:2002:AIN

[633] Phong Q. Nguyen and David Pointcheval. Analysis and improvements of NTRU encryption paddings. Lecture Notes in Computer Science, 2442:210– 225, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2442/24420210.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2442/24420210.
                                        pdf.

Coron:2002:UPS

[634] Jean-Sébastien Coron, Marc Joye, David Naccache, and Pascal Paillier. Universal padding schemes for RSA. Lecture Notes in Computer Science, 2442:226–241, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

May:2002:CUR

[635] Alexander May. Cryptanalysis of unbalanced RSA with small CRTexponent. Lecture Notes in Computer Science, 2442:242–256, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420242.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420242.pdf.

Lu:2002:HEA

[636] Chi-Jen Lu. Hyper-encryption against space-bounded adversaries from online strong extractors. Lecture Notes in Computer Science, 2442: 257–271, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2442/24420257.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2442/24420257.
pdf.

Cramer:2002:OBB

[637] Ronald Cramer and Serge Fehr. Optimal black-box secret sharing over arbitrary Abelian groups. Lecture Notes in Computer Science, 2442: 272–287, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2442/24420272.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2442/24420272.
pdf.

Wagner:2002:GBP

[638] David Wagner. A generalized birthday problem: (extended abstract). Lecture Notes in Computer Science, 2442: 288–303, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/23/20/85/
                                        34/32/abstract.html.

Mironov:2002:RSR

[639] Ilya Mironov. (not so) random shuffles of RC4. Lecture Notes in Computer Science, 2442:304–319, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Black:2002:BBA

[640] John Black, Phillip Rogaway, and Thomas Shrimpton. Black-box analysis of the block-cipher-based hashfunction constructions from PGV. Lecture Notes in Computer Science, 2442:320–335, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420320.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420320.pdf.

Rubin:2002:SAV

[641] Karl Rubin and Alice Silverberg. Supersingular Abelian varieties in cryptology. Lecture Notes in Computer Science, 2442:336–353, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420336.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2442/24420336.pdf.

Barreto:2002:EAP

[642] Paulo S. L. M. Barreto, Hae Y. Kim, Ben Lynn, and Michael Scott. Efficient algorithms for pairing-based cryptosystems. Lecture Notes in Computer Science, 2442:354–368, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420354.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420354.pdf.

Vercauteren:2002:CZF

[643] Frederik Vercauteren. Computing zeta functions of hyperelliptic curves over finite fields of characteristic 2. Lecture Notes in Computer Science, 2442:369–384, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420369.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420369.pdf.

MacKenzie:2002:TPA

[644] Philip MacKenzie, Thomas Shrimpton, and Markus Jakobsson. Threshold password-authenticated key exchange: (extended abstract). Lecture Notes in Computer Science, 2442:385– 400, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550224.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2455/24550224.
                                        pdf.

Nielsen:2002:TPF

[645] Jesper Buus Nielsen. A threshold pseudorandom function construction and its applications. Lecture Notes in Computer Science, 2442:401–416, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Algesheimer:2002:ECM

[646] Joy Algesheimer, Jan Camenisch, and Victor Shoup. Efficient computation modulo a shared secret with application to the generation of shared safe-prime products. Lecture Notes in Computer Science, 2442:417–432, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420417.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420417.pdf.

Li:2002:HNP

[647] Wen-Ching W. Li, Mats Näslund, and Igor E. Shparlinski. Hidden number problem with the trace and bit security of XTR and LUC. Lecture Notes in Computer Science, 2442:433–448, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420031.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420031.pdf.

Damgaard:2002:EPF

[648] Ivan Damgård and Jesper Buus Nielsen. Expanding pseudorandom functions; or: From known-plaintext security to chosen-plaintext security. Lecture Notes in Computer Science, 2442:449–464, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420565.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420565.pdf.

Bresson:2002:TRS

[649] Emmanuel Bresson, Jacques Stern, and Michael Szydlo. Threshold ring signatures and applications to ad-hoc groups. Lecture Notes in Computer Science, 2442:465–480, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420465.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420465.pdf.

Naor:2002:DRA

[650] Moni Naor. Deniable ring authentication. Lecture Notes in Computer Science, 2442:481–498, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420210.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420210.pdf.

Itkis:2002:SSB

[651] Gene Itkis and Leonid Reyzin. SiBIR: Signer-Base Intrusion-Resilient signatures. Lecture Notes in Computer Science, 2442:499–514, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420499.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420499.pdf.

Coppersmith:2002:CSC

[652] Don Coppersmith, Shai Halevi, and Charanjit Jutla. Cryptanalysis of stream ciphers with linear masking. Lecture Notes in Computer Science, 2442:515–532, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420515.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420515.pdf.

Sarkar:2002:FCM

[653] Palash Sarkar. The filter-combiner model for memoryless synchronous stream ciphers. Lecture Notes in Computer Science, 2442:533– 548, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2442/24420533.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2442/24420533.
                                        pdf.

Carlet:2002:LCC

[654] Claude Carlet. A larger class of cryptographic Boolean functions via a study of the Maiorana–McFarland construction. Lecture Notes in Computer Science, 2442:549–564, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420549.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420549.pdf.

Fehr:2002:LVD

[655] Serge Fehr and Ueli Maurer. Linear VSS and distributed commitments based on secret sharing and pairwise checks. Lecture Notes in Computer Science, 2442:565–580, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420565.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420565.pdf.

Damgaard:2002:PHP

[656] Ivan Damgård and Jesper Buus Nielsen. Perfect hiding and perfect binding universally composable commitment schemes with constant expansion factor. Lecture Notes in Computer Science, 2442:581–596, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2442/24420565.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2442/24420565.pdf.

Lysyanskaya:2002:USV

[657] Anna Lysyanskaya. Unique signatures and verifiable random functions from the DH–DDH separation. Lecture Notes in Computer Science, 2442:597–612, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420597.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420597.pdf.

Coron:2002:SPP

[658] Jean-Sébastien Coron. Security proof for partial-domain hash signature schemes. Lecture Notes in Computer Science, 2442:613–626, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2442/24420613.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2442/24420613.pdf.

House:2002:MSS

[659] David House and Björn Granström. Multimodal speech synthesis: Improving information flow in dialogue systems using 3D talking heads. Lecture Notes in Computer Science, 2443:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2443/24430001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2443/24430001.
                                        pdf.

Popelinsky:2002:ERL

[660] Luboš Popelínský. Efficient relational learning from sparse data. Lecture Notes in Computer Science, 2443:11– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430011.
pdf.

Smirnov:2002:EIR

[661] E. N. Smirnov, I. G. Sprinkhuizen-Kuyper, and H. J. van den Herik. Efficient instance retraction. Lecture Notes in Computer Science, 2443:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430021.
pdf.

Osmani:2002:LPM

[662] A. Osmani. Learning patterns in multidimensional space using interval algebra. Lecture Notes in Computer Science, 2443:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430031.
pdf.

Monsifrot:2002:MLA

[663] Antoine Monsifrot, François Bodin, and René Quiniou. A machine learning approach to automatic production of compiler heuristics. Lecture Notes in Computer Science, 2443:41– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430041.
pdf.

Gregoire:2002:UFL

[664] Éric Grégoire, Bertrand Mazure, and Lakhdar Saïs. Using failed local search for SAT as an oracle for tackling harder A.I. problems more efficiently. Lecture Notes in Computer Science, 2443:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430051.
pdf.

Vrakas:2002:HPB

[665] Dimitris Vrakas and Ioannis Vlahavas. A heuristic for planning based on action evaluation. Lecture Notes in Computer Science, 2443:61– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430061.
pdf.

Lopez:2002:WWB

[666] Natalia López, Manuel Núñez, Ismael Rodríguez, and Fernando Rubio. WHAT: Web-based Haskell adaptive tutor. Lecture Notes in Computer Science, 2443:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430071.
pdf.

Magro:2002:DDC

[667] Diego Magro and Pietro Torasso. Decomposing and distributing configuration problems. Lecture Notes in Computer Science, 2443:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430081.
pdf.

Baruch:2002:DAN

[668] Ieroham Baruch, Alfredo del Carmen Martinez Q., Ruben Garrido, and Boyka Nenkova. Direct adaptive neural control with integralplus-state action. Lecture Notes in Computer Science, 2443:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430091.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2443/24430091.
                                        pdf.

Boytcheva:2002:ITF

[669] Svetla Boytcheva. ILP techniques for free-text input processing. Lecture Notes in Computer Science, 2443:101– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2443/24430101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2443/24430101.
                                        pdf.

Choenni:2002:DSA

[670] Sunil Choenni and Henk Blanken. A Dempster-Shafer approach to physical database design. Lecture Notes in Computer Science, 2443:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2443/24430111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2443/24430111.
                                        pdf.

Miliaev:2002:TDI

[671] Nestor Miliaev, Alison Cawsey, and Greg Michaelson. Technical documentation: An integrated architecture for supporting the author in generation and resource editing. Lecture Notes in Computer Science, 2443:122– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2443/24430122.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2443/24430122.
pdf.

Pacholczyk:2002:OVP

[672] Daniel Pacholczyk, Mohamed Quafafou, and Laurent Garcia. Optimistic vs. pessimistic interpretation of linguistic negation. Lecture Notes in Computer Science, 2443:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430132.
pdf.

Carrillo:2002:LCP

[673] Vicente Carrillo and Víctor J. Díaz. Left corner parser for tree insertion grammars. Lecture Notes in Computer Science, 2443:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430142.
pdf.

Stoykova:2002:BND

[674] Velislava Stoykova. Bulgarian noun — definite article in DATR. Lecture Notes in Computer Science, 2443:152– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430152.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2443/24430152.
pdf.

Simon:2002:MSO

[675] Gaële Simon, Marianne Flouret, and Bruno Mermet. A methodology to solve optimisation problems with MAS application to the graph colouring problem. Lecture Notes in Computer Science, 2443:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430162.
pdf.

Casas:2002:OOM

[676] Feliciano Manzano Casas and L. A. García. OCOA: An open, modular, ontology based autonomous robotic agent architecture. Lecture Notes in Computer Science, 2443:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430173.
pdf.

Boella:2002:RAR

[677] Guido Boella and Rossana Damiano. A replanning algorithm for a reactive agent architecture. Lecture Notes in Computer Science, 2443:183– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2443/24430183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430183.
pdf.

Lin:2002:BAM

[678] S. L. M. Lin. A broker approach for multi-agent scheduling. Lecture Notes in Computer Science, 2443:193– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430193.
pdf.

Kovacevic:2002:RCA

[679] Miloš Kovačević, Michelangelo Dilligenti, Marco Gori, and Veljko Milutinović. Recognition of common areas in a Web page using a visualization approach. Lecture Notes in Computer Science, 2443:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430203.
pdf.

Medina:2002:NAA

[680] Jesús Medina, Enrique Mérida-Casermeiro, and Manuel Ojeda-Aciego. A neural approach to abductive multiadjoint reasoning. Lecture Notes in Computer Science, 2443:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2443/24430213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2443/24430213.
                                        pdf.

deGuzman:2002:RTM

[681] I. P. de Guzmán, Manuel Ojeda-Aciego, and A. Valverde. Restricted Δ-trees in multiple-valued logics. Lecture Notes in Computer Science, 2443:223– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2443/24430223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2443/24430223.
                                        pdf.

Pop:2002:KME

[682] Daniel Pop and Viorel Negru. Knowledge management in expert system creator. Lecture Notes in Computer Science, 2443:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2443/24430233.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2443/24430233.pdf.

Choi:2002:TSG

[683] Seungyeob Choi. Towards semantic goal-directed forward reasoning in resolution. Lecture Notes in Computer Science, 2443:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2443/24430243.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2443/24430243.
pdf.

Dichev:2002:DWW

[684] Christo Dichev. Do what we do to find what you want. Lecture Notes in Computer Science, 2443:253– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430253.
pdf.

Maynard:2002:ARM

[685] Diana Maynard, Hamish Cunningham, Kalina Bontcheva, and Marin Dimitrov. Adapting a robust multigenre NE system for automatic content extraction. Lecture Notes in Computer Science, 2443:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2443/24430264.
pdf.

Goble:2002:SWK

[686] Carole Goble. The Semantic Web: A killer app for AI? Lecture Notes in Computer Science, 2443:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2443/24430274.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2443/24430274.
                                        pdf.

Anonymous:2002:AIl

[687] Anonymous. Author index. Lecture Notes in Computer Science, 2443:279– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2443/2443auth.pdf.

Mohan:2002:DBT

[688] C. Mohan. Dynamic E-business: Trends in Web services. Lecture Notes in Computer Science, 2444:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2444/24440001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2444/24440001.
                                        pdf.

Ronstrom:2002:TDS

[689] Mikael Ronström and Vinay P. Joosery. Telecom databases for the E-services industry. Lecture Notes in Computer Science, 2444:6– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2444/24440006.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2444/24440006.
                                        pdf.

Dogac:2002:IFU

[690] Asuman Dogac, Ibrahim Cingil, Gokce Laleci, and Yildiray Kabak. Improving the functionality of UDDI registries through Web service semantics. Lecture Notes in Computer Science, 2444: 9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2444/24440009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2444/24440009.
pdf.

Bussler:2002:PPI

[691] Christoph Bussler. Public process inheritance for business-to-business integration. Lecture Notes in Computer Science, 2444:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2444/24440019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2444/24440019.
pdf.

Valikov:2002:MTA

[692] Alexey Valikov, Alexei Akhounov, and Andreas Schmidt. A modeltransformers architecture for Web applications. Lecture Notes in Computer Science, 2444:29–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2444/24440029.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2444/24440029.pdf.

Mecella:2002:MSO

[693] Massimo Mecella, Francesco Parisi Presicce, and Barbara Pernici. Modeling E-service orchestration through Petri nets. Lecture Notes in Computer Science, 2444:38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2444/24440038.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2444/24440038.pdf.

Nori:2002:CAP

[694] Anil K. Nori and Rajiv Jain. Composite applications: Process based application development. Lecture Notes in Computer Science, 2444:48– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2444/24440048.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2444/24440048.
                                        pdf.

Papazoglou:2002:DMW

[695] Mike P. Papazoglou and Jian Yang. Design methodology for Web services and business processes. Lecture Notes in Computer Science, 2444:54– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2444/24440054.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2444/24440054.
                                        pdf.

Helal:2002:SBI

[696] Abdelsalam Helal and Jingting Lu. E-service based information fusion: A user-level information integration framework. Lecture Notes in Computer Science, 2444:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2444/24440065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2444/24440065.
pdf.

Aiello:2002:RLW

[697] M. Aiello, Mike P. Papazoglou, Jian Yang, M. Carman, M. Pistore, L. Serafini, and P. Traverso. A request language for Web-services based on planning and constraint satisfaction. Lecture Notes in Computer Science, 2444: 76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2444/24440076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2444/24440076.
pdf.

Ordille:2002:CFE

[698] Joann J. Ordille and Thomas Petsche. Communication flow expressions in a notification and response system. Lecture Notes in Computer Science, 2444: 86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2444/24440086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2444/24440086.
pdf.

Abrahams:2002:CDM

[699] Alan S. Abrahams, David M. Eyers, and Jean M. Bacon. A coveragedetermination mechanism for checking business contracts against organizational policies. Lecture Notes in Computer Science, 2444:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2444/24440097.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2444/24440097.
                                        pdf.

Li:2002:MBR

[700] Haifei Li, Jun-Jang Jeng, and Henry Chang. Managing business relationships in E-services using business commitments. Lecture Notes in Computer Science, 2444:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2444/24440107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2444/24440107.
                                        pdf.

Popovici:2002:AHT

[701] Andrei Popovici and Gustavo Alonso. Ad-hoc transactions for mobile services. Lecture Notes in Computer Science, 2444:118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2444/24440118.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2444/24440118.pdf.

Li:2002:AWS

[702] Jing Li, Xin Zhang, and Zhong Tian.

Advanced Web session provider for suspensible E-services. Lecture Notes in Computer Science, 2444:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2444/24440131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2444/24440131.
pdf.

Zeng:2002:PDB

[703] Liangzhao Zeng, David Flaxer, Henry Chang, and Jun-Jang Jeng. PLM_{flow}- dynamic business process composition and execution by rule inference. Lecture Notes in Computer Science, 2444: 141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2444/24440141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2444/24440141.
pdf.

Yang:2002:TBS

[704] Seokwon Yang, Herman Lam, and Stanley Y. W. Su. Trust-based security model and enforcement mechanism for Web service technology. Lecture Notes in Computer Science, 2444:151– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2444/24440151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2444/24440151.
pdf.

Ito:2002:FEU

[705] Chihiro Ito, Mizuho Iwaihara, and Yahiko Kambayashi. Fair exchange under limited trust. Lecture Notes in Computer Science, 2444:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2444/24440161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2444/24440161.
                                        pdf.

Anonymous:2002:AIm

[706] Anonymous. Author index. Lecture Notes in Computer Science, 2444:171– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2444/2444auth.pdf.

Balaban:2002:SIM

[707] Mira Balaban. Structure and interpretation of music concepts: Music from a computational perspective. Lecture Notes in Computer Science, 2445:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2445/24450001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2445/24450001.
                                        pdf.

Camurri:2002:EG

[708] Antonio Camurri. Expressive gesture. Lecture Notes in Computer Science, 2445:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2445/24450004.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2445/24450004.
pdf.

Bod:2002:GPM

[709] Rens Bod. A general parsing model for music and language. Lecture Notes in Computer Science, 2445:5– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2445/24450005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2445/24450005.
pdf.

Chew:2002:SAA

[710] Elaine Chew. The spiral array: An algorithm for determining key boundaries. Lecture Notes in Computer Science, 2445:18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2445/24450018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2445/24450018.
pdf.

Conklin:2002:RDV

[711] Darrell Conklin. Representation and discovery of vertical patterns in music. Lecture Notes in Computer Science, 2445:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2445/24450032.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2445/24450032.
                                        pdf.

Dannenberg:2002:DMS

[712] Roger B. Dannenberg and Ning Hu. Discovering musical structure in audio recordings. Lecture Notes in Computer Science, 2445:43–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2445/24450043.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2445/24450043.
                                        pdf.

Dixon:2002:RTT

[713] Simon Dixon, Werner Goebl, and Gerhard Widmer. Real time tracking and visualisation of musical expression. Lecture Notes in Computer Science, 2445:58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2445/24450058.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2445/24450058.
                                        pdf.

Herrera:2002:ACD

[714] Perfecto Herrera, Alexandre Yeterian, and Fabien Gouyon. Automatic classification of drum sounds: A comparison of feature selection methods and classification techniques. Lecture Notes in Computer Science, 2445:69– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2445/24450069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2445/24450069.
pdf.

Horton:2002:SFP

[715] Tim Horton. Some formal problems with schenkerian representations of tonal structure. Lecture Notes in Computer Science, 2445:81– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2445/24450081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2445/24450081.
pdf.

Igarashi:2002:RRM

[716] Soh Igarashi, Tomonobu Ozaki, and Koichi Furukawa. Respiration reflecting musical expression: Analysis of respiration during musical performance by inductive logic programming. Lecture Notes in Computer Science, 2445:94– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2445/24450094.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2445/24450094.
pdf.

Miranda:2002:MDI

[717] Eduardo Reck Miranda. Mimetic development of intonation. Lecture Notes in Computer Science, 2445:107– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2445/24450107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2445/24450107.
                                        pdf.

Pachet:2002:IML

[718] François Pachet. Interacting with a musical learning system: The continuator. Lecture Notes in Computer Science, 2445:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2445/24450119.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2445/24450119.
                                        pdf.

Pikrakis:2002:RIM

[719] Aggelos Pikrakis, Sergios Theodoridis, and Dimitris Kamarotos. Recognition of isolated musical patterns using hidden Markov models. Lecture Notes in Computer Science, 2445:133– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2445/24450133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2445/24450133.
                                        pdf.

Povel:2002:MPT

[720] Dirk-Jan Povel. A model for the perception of tonal melodies. Lecture Notes in Computer Science, 2445:144– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2445/24450144.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2445/24450144.
pdf.

Phon-Amnuaisuk:2002:CLH

[721] Somnuk Phon-Amnuaisuk. Control language for harmonisation process. Lecture Notes in Computer Science, 2445:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2445/24450155.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2445/24450155.
pdf.

Spevak:2002:EMS

[722] Christian Spevak, Belinda Thom, and Karin Höthker. Evaluating melodic segmentation. Lecture Notes in Computer Science, 2445:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2445/24450168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2445/24450168.
pdf.

Spiro:2002:CGB

[723] Neta Spiro. Combining grammar-based and memory-based models of perception of time signature and phase. Lecture Notes in Computer Science, 2445: 183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2445/24450183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2445/24450183.
pdf.

Temperley:2002:BAK

[724] David Temperley. A Bayesian approach to key-finding. Lecture Notes in Computer Science, 2445:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2445/24450195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2445/24450195.
                                        pdf.

Anonymous:2002:AIn

[725] Anonymous. Author index. Lecture Notes in Computer Science, 2445:207– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2445/2445auth.pdf.

Baeza-Yates:2002:ACW

[726] Ricardo Baeza-Yates and José Miguel Piquer. Agents, crawlers, and Web retrieval. Lecture Notes in Computer Science, 2446:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2446/24460001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2446/24460001.pdf.

Mase:2002:III

[727] Kenji Mase. Intelligent interfaces for information agents: Systems, experiences, future challenges. Lecture Notes in Computer Science, 2446:10– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460010.
pdf.

Noriega:2002:EIF

[728] Pablo Noriega and Carles Sierra. Electronic institutions: Future trends and challenges. Lecture Notes in Computer Science, 2446:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460014.
pdf.

Giunchiglia:2002:MPD

[729] Fausto Giunchiglia and Ilya Zaihrayeu. Making peer databases interact — A vision for an architecture supporting data coordination. Lecture Notes in Computer Science, 2446:18– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460018.
pdf.

Laukkanen:2002:TM

[730] Mikko Laukkanen, Heikki Helin, and Heimo Laamanen. Tourists on the move. Lecture Notes in Computer Science, 2446:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2446/24460036.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2446/24460036.
                                        pdf.

Mena:2002:ABA

[731] E. Mena, J. A. Royo, A. Illarramendi, and A. Goñi. An agent-based approach for helping users of hand-held devices to browse software catalogs. Lecture Notes in Computer Science, 2446:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2446/24460051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2446/24460051.
                                        pdf.

Iwao:2002:LSP

[732] Tadashige Iwao, Makoto Okada, Kazuya Kawashima, Satoko Matsumura, Hajime Kanda, Susumu Sakamoto, Tatsuya Kainuma, and Makoto Amamiya. Large scale peer-to-peer experiments with virtual private community (VPC) framework. Lecture Notes in Computer Science, 2446:66– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2446/24460066.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2446/24460066.
                                        pdf.

Garcia-Serrano:2002:CAD

[733] Ana García-Serrano and Javier Calle-Gómez. A cognitive architecture for the design of an interaction agent. Lecture Notes in Computer Science, 2446: 82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460082.
pdf.

Stefano:2002:MAR

[734] Antonella Di Stefano, Giuseppe Pappalardo, Corrado Santoro, and Emiliano Tramontana. A multi-agent reflective architecture for user assistance and its application to E-commerce. Lecture Notes in Computer Science, 2446:90– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460090.
pdf.

Simoff:2002:CN

[735] Simeon Simoff and John Debenham. Curious negotiator. Lecture Notes in Computer Science, 2446:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460104.
pdf.

Grimsley:2002:MBB

[736] Michael Grimsley and Anthony Meehan. Measuring behaviour-based trust between negotiating agents. Lecture Notes in Computer Science, 2446:112– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2446/24460112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2446/24460112.
                                        pdf.

Azoulay-Schwartz:2002:AOA

[737] Rina Azoulay-Schwartz and Sarit Kraus. Acquiring an optimal amount of information for choosing from alternatives. Lecture Notes in Computer Science, 2446:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2446/24460123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2446/24460123.
                                        pdf.

Kitamura:2002:CIR

[738] Yasuhiko Kitamura, Toshiki Sakamoto, and Shoji Tatsumi. A competitive information recommendation system and its behavior. Lecture Notes in Computer Science, 2446:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2446/24460138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2446/24460138.
                                        pdf.

Uchyigit:2002:AML

[739] Gulden Uchyigit and Keith Clark.

Agents that model and learn user interests for dynamic collaborative filtering. Lecture Notes in Computer Science, 2446:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2446/24460152.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2446/24460152.
pdf.

Montaner:2002:OBF

[740] Miquel Montaner, Beatriz López, and Josep Lluís de la Rosa. Opinionbased filtering through trust. Lecture Notes in Computer Science, 2446:164– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460164.
pdf.

Koubarakis:2002:DML

[741] M. Koubarakis, C. Tryfonopoulos, P. Raftopoulou, and T. Koutris. Data models and languages for agent-based textual information dissemination. Lecture Notes in Computer Science, 2446: 179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460179.
pdf.

Cost:2002:IDI

[742] R. Scott Cost, Srikanth Kallurkar, Hemali Majithia, Charles Nicholas, and Yongmei Shi. Integrating distributed information sources with CARROT II. Lecture Notes in Computer Science, 2446:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2446/24460194.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2446/24460194.
pdf.

Gandon:2002:DAI

[743] Fabien Gandon and Rose Dieng-Kuntz. Distributed artificial intelligence for distributed corporate knowledge management. Lecture Notes in Computer Science, 2446:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460202.
pdf.

Macho-Gonzalez:2002:IIG

[744] Santiago Macho-Gonzalez and Boi Faltings. Integrating information gathering and problem-solving in open environments. Lecture Notes in Computer Science, 2446:218–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460218.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460218.
pdf.

Chalmers:2002:SVO

[745] Stuart Chalmers, Peter M. D. Gray, and Alun Preece. Supporting virtual organisations using BDI agents and constraints. Lecture Notes in Computer Science, 2446:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460226.
pdf.

Serrano:2002:AAC

[746] Juan M. Serrano and Sascha Ossowski. An approach to agent communication based on organisational roles. Lecture Notes in Computer Science, 2446:241– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460241.
pdf.

Stuckenschmidt:2002:EPS

[747] Heiner Stuckenschmidt. Exploiting partially shared ontologies for multiagent communication. Lecture Notes in Computer Science, 2446:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460249.
pdf.

Ben-Ami:2002:EDC

[748] David Ben-Ami and Onn Shehory. Evaluation of distributed and centralized agent location mechanisms. Lecture Notes in Computer Science, 2446: 264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2446/24460264.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2446/24460264.
                                        pdf.

Hadad:2002:ECT

[749] Meirav Hadad and Sarit Kraus. Exchanging and combining temporal information in a cooperative environment. Lecture Notes in Computer Science, 2446:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2446/24460279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2446/24460279.
                                        pdf.

Montanari:2002:PAM

[750] Rebecca Montanari, Gianluca Tonti, and Cesare Stefanelli. Programming agent mobility. Lecture Notes in Computer Science, 2446:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2446/24460287.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2446/24460287.
                                        pdf.

Cubaleska:2002:MPM

[751] Biljana Cubaleska and Markus Schneider. A method for protecting mobile agents against denial of service attacks. Lecture Notes in Computer Science, 2446:297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460297.
pdf.

Morawek:2002:TRO

[752] Roman Morawek. Threshold route optimization algorithm for information retrieving mobile agents. Lecture Notes in Computer Science, 2446:312– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2446/24460312.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2446/24460312.
pdf.

Anonymous:2002:AIo

[753] Anonymous. Author index. Lecture Notes in Computer Science, 2446:321– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2446/2446auth.pdf.

Hand:2002:PDD

[754] David J. Hand. Pattern detection and discovery. Lecture Notes in Computer Science, 2447:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2447/24470001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2447/24470001.
pdf.

Morik:2002:DII

[755] Katharina Morik. Detecting interesting instances. Lecture Notes in Computer Science, 2447:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2447/24470013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2447/24470013.
pdf.

Siebes:2002:CDM

[756] Arno Siebes and Zbyszek Struzik. Complex data: Mining using patterns. Lecture Notes in Computer Science, 2447:24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2447/24470024.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2447/24470024.
pdf.

Bolton:2002:DHR

[757] Richard J. Bolton, David J. Hand, and Niall M. Adams. Determining hit rate in pattern search. Lecture Notes in Computer Science, 2447:36– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2447/24470036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2447/24470036.
pdf.

Cohen:2002:UAS

[758] Paul Cohen, Brent Heeringa, and Niall M. Adams. An unsupervised algorithm for segmenting categorical timeseries into episodes. Lecture Notes in Computer Science, 2447:49– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2447/24470049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2447/24470049.
pdf.

Unwin:2002:IYC

[759] Antony Unwin. If you can’t see the pattern, is it there? Lecture Notes in Computer Science, 2447:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2447/24470063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2447/24470063.
pdf.

Wojciechowski:2002:DFT

[760] Marek Wojciechowski and Maciej Zakrzewicz. Dataset filtering techniques in constraint-based frequent pattern mining. Lecture Notes in Computer Science, 2447:77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2447/24470077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2447/24470077.
                                        pdf.

Kryszkiewicz:2002:CRA

[761] Marzena Kryszkiewicz. Concise representations of association rules. Lecture Notes in Computer Science, 2447:92– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2447/24470092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2447/24470092.
                                        pdf.

Jeudy:2002:CBD

[762] Baptiste Jeudy and Jean-François Boulicaut. Constraint-based discovery and inductive queries: Application to association rule mining. Lecture Notes in Computer Science, 2447: 110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2447/24470110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2447/24470110.
                                        pdf.

Goethals:2002:RAR

[763] Bart Goethals and Jan Van den Bussche. Relational association rules: Getting WARMeR. Lecture Notes in Computer Science, 2447:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2447/24470125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2447/24470125.
pdf.

Delgado:2002:MTD

[764] M. Delgado, M. J. Martín-Bautista, D. Sánchez, and M. A. Vila. Mining text data: Special features and patterns. Lecture Notes in Computer Science, 2447:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2447/24470140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2447/24470140.
pdf.

Spiliopoulou:2002:MIB

[765] Myra Spiliopoulou and Carsten Pohle. Modelling and incorporating background knowledge in the Web mining process. Lecture Notes in Computer Science, 2447:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2447/24470154.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2447/24470154.pdf.

Mladenic:2002:MIT

[766] Dunja Mladenić. Modeling information in textual data combining labeled and unlabeled data. Lecture Notes in Computer Science, 2447:170– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2447/24470170.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2447/24470170.
                                        pdf.

Ahonen-Myka:2002:DFW

[767] Helena Ahonen-Myka. Discovery of frequent word sequences in text. Lecture Notes in Computer Science, 2447:180– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2447/24470180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2447/24470180.
                                        pdf.

Rolland:2002:PDD

[768] Pierre-Yves Rolland and Jean-Gabriel Ganascia. Pattern detection and discovery: The case of music data mining. Lecture Notes in Computer Science, 2447:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2447/24470190.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2447/24470190.
                                        pdf.

Hoppner:2002:DCE

[769] Frank Höppner. Discovery of core episodes from sequences. Lecture Notes in Computer Science, 2447:199– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2447/24470199.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2447/24470199.
pdf.

Gather:2002:PDD

[770] Ursula Gather, Roland Fried, Michael Imhoff, and Claudia Becker. Patterns of dependencies in dynamic multivariate data. Lecture Notes in Computer Science, 2447:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2447/24470214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2447/24470214.
pdf.

Anonymous:2002:AIp

[771] Anonymous. Author index. Lecture Notes in Computer Science, 2447:227– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2447/2447auth.pdf.

Grana:2002:CST

[772] Jorge Graña, Miguel A. Alonso, and Manuel Vilares. A common solution for tokenization and partof-speech tagging. Lecture Notes in Computer Science, 2448:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480003.
pdf.

Al-Sughaiyer:2002:RPA

[773] Imad A. Al-Sughaiyer and Ibrahim A. Al-Kharashi. Rule parser for Arabic stemmer. Lecture Notes in Computer Science, 2448:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480011.
pdf.

Kveton:2002:AAC

[774] Pavel Květoň and Karel Oliva. Achieving an almost correct PoStagged corpus. Lecture Notes in Computer Science, 2448:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480019.
pdf.

Oguro:2002:EJS

[775] Rei Oguro, Hiromi Sekiya, Yuhei Morooka, Kazuyuki Takagi, and Kazuhiko Ozeki. Evaluation of a Japanese sentence compression method based on phrase significance and interphrase dependency. Lecture Notes in Computer Science, 2448:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480027.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2448/24480027.
pdf.

Boldasov:2002:UQU

[776] Michael V. Boldasov, Elena G. Sokolova, and Michael G. Malkovsky. User query understanding by the In-BASE system as a source for a multilingual NL generation module. Lecture Notes in Computer Science, 2448:33– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480033.
pdf.

Montoyo:2002:RWM

[777] Andrés Montoyo, Rafael Romero, Sonia Vázquez, Carmen Calle, and Susana Soler. The role of WSD for multilingual natural language applications. Lecture Notes in Computer Science, 2448:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2448/24480041.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2448/24480041.
pdf.

Rozenknop:2002:GCF

[778] Antoine Rozenknop. Gibbsian contextfree grammar for parsing. Lecture Notes in Computer Science, 2448:49– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480049.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480049.
                                        pdf.

Oard:2002:CLA

[779] Douglas W. Oard, Dina Demner-Fushman, Jan Hajič, Bhuvana Ramabhadran, Samuel Gustman, William J. Byrne, Dagobert Soergel, Bonnie Dorr, Philip Resnik, and Michael Picheny. Cross-language access to recorded speech in the MALACH project. Lecture Notes in Computer Science, 2448: 57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480057.
                                        pdf.

Trabalka:2002:USW

[780] Marek Trabalka and Mária Bieliková. Using salient words to perform categorization of Web sites. Lecture Notes in Computer Science, 2448:65– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480065.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480065.
                                        pdf.

Alberti:2002:DSA

[781] Gábor Alberti and Helga M. Szabó. Discourse-semantic analysis of Hungarian sign language. Lecture Notes in Computer Science, 2448:73–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480073.
pdf.

Holan:2002:DAC

[782] Tomáš Holan. Dependency analyser configurable by measures. Lecture Notes in Computer Science, 2448:81– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480081.
pdf.

Hartrumpf:2002:GUL

[783] Sven Hartrumpf and Hermann Helbig. The generation and use of layer information in multilayered extended semantic networks. Lecture Notes in Computer Science, 2448:89– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480089.
pdf.

Zizka:2002:FLN

[784] Jan Žižka and Aleš Bourek. Filtering of large numbers of unstructured text documents by the developed tool TEA. Lecture Notes in Computer Science, 2448:99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480099.
                                        pdf.

Widdows:2002:VTA

[785] Dominic Widdows, Scott Cederberg, and Beate Dorow. Visualisation techniques for analysing meaning. Lecture Notes in Computer Science, 2448:107– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480107.
                                        pdf.

Huang:2002:SPS

[786] Liang Huang, Yinan Peng, Huan Wang, and Zhenyu Wu. Statistical part-of-speech tagging for classical chinese. Lecture Notes in Computer Science, 2448:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480115.
                                        pdf.

Rangel:2002:SNL

[787] Rodolfo A. Pazos Rangel, Alexander Gelbukh, J. Javier González Barbosa, Erika Alarcón Ruiz, Alejandro Mendoza Mejía, and A. Patricia Domínguez Sánchez. Spanish natural language interface for a relational database querying system. Lecture Notes in Computer Science, 2448:123– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480123.
pdf.

Suarez:2002:WSV

[788] Armando Suárez and Manuel Palomar. Word sense vs. word domain disambiguation: A maximum entropy approach. Lecture Notes in Computer Science, 2448:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480131.
pdf.

Sadat:2002:ETH

[789] Fatiha Sadat, Masatoshi Yoshikawa, and Shunsuke Uemura. Exploiting thesauri and hierarchical categories in cross-language information retrieval. Lecture Notes in Computer Science, 2448:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2448/24480139.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2448/24480139.
pdf.

Lopatkova:2002:VLC

[790] Markéta Lopatková, Veronika Řezníčková, and Zdeněk Žabokrtský. Valency lexicon for Czech: From verbs to nouns. Lecture Notes in Computer Science, 2448:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480147.
                                        pdf.

Nenadic:2002:TCU

[791] Goran Nenadić, Irena Spasić, and Sophia Ananiadou. Term clustering using a corpus-based similarity measure. Lecture Notes in Computer Science, 2448:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480151.
                                        pdf.

Kral:2002:WSD

[792] Robert Král. Word sense discrimination for Czech. Lecture Notes in Computer Science, 2448:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480155.
                                        pdf.

Bartuskova:2002:TSA

[793] Dita Bartušková and Radek Sedláček.

Tools for semi-automatic assignment of Czech nouns to declination patterns. Lecture Notes in Computer Science, 2448:159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480159.
pdf.

Sef:2002:ALS

[794] Tomaž Šef, Maja Škrjanc, and Matjaž Gams. Automatic lexical stress assignment of unknown words for highly inflected Slovenian language. Lecture Notes in Computer Science, 2448:165– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480165.
pdf.

Matousek:2002:GCS

[795] Jindřich Matoušek, Daniel Tihelka, Josef Psutka, and Jana Hesová. German and Czech speech synthesis using HMM-based speech segment database. Lecture Notes in Computer Science, 2448:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480173.
pdf.

Stemmer:2002:CCC

[796] Georg Stemmer, Stefan Steidl, Elmar Nöth, Heinrich Niemann, and Anton Batliner. Comparison and combination of confidence measures. Lecture Notes in Computer Science, 2448:181– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480181.
                                        pdf.

Nouza:2002:SDR

[797] Jan Nouza. Strategies for developing a real-time continuous speech recognition system for Czech language. Lecture Notes in Computer Science, 2448: 189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480189.
                                        pdf.

Nejedlova:2002:CSB

[798] Dana Nejedlová. Comparative study on bigram language models for spoken Czech recognition. Lecture Notes in Computer Science, 2448:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480197.
                                        pdf.

Wiggers:2002:ISR

[799] Pascal Wiggers and Leon J. M. Rothkrantz. Integration of speech recognition and automatic lip-reading. Lecture Notes in Computer Science, 2448:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2448/24480205.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2448/24480205.
pdf.

Bouwman:2002:UVB

[800] Gies Bouwman and Lou Boves. Utterance verification based on the likelihood distance to alternative paths. Lecture Notes in Computer Science, 2448: 213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480213.
pdf.

Bartos:2002:RTB

[801] Tomáš Bartoš and Luděk Müller. Rejection technique based on the mumble model. Lecture Notes in Computer Science, 2448:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480221.
pdf.

Motlicek:2002:ENE

[802] Petr Motlíček and Lukáš Burget. Efficient noise estimation and its application for robust speech recognition. Lecture Notes in Computer Science, 2448: 229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480229.
pdf.

Secujski:2002:ASS

[803] Milan Sečujski, Radovan Obradović, Darko Pekar, Ljubomir Jovanov, and Vlado Delić. AlfaNum system for speech synthesis in serbian language. Lecture Notes in Computer Science, 2448:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480237.
pdf.

Zibert:2002:SFE

[804] Janez Žibert, France Mihelič, and Nikola Pavešić. Speech features extraction using cone-shaped kernel distribution. Lecture Notes in Computer Science, 2448:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480245.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2448/24480245.
pdf.

Psutka:2002:ATC

[805] Josef Psutka, Pavel Ircing, Josef V. Psutka, Vlasta Radová, William J. Byrne, Jan Hajič, Samuel Gustman, and Bhuvana Ramabhadran. Automatic transcription of Czech language oral history in the MALACH project: Resources and initial experiments. Lecture Notes in Computer Science, 2448: 253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480253.
pdf.

Zervas:2002:FGT

[806] P. Zervas, I. Potamitis, N. Fakotakis, and G. Kokkinakis. On the first greek-TTS based on festival speech synthesis. Lecture Notes in Computer Science, 2448:261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480261.
pdf.

Batusek:2002:ALD

[807] Robert Batušek. An analysis of limited domains for speech synthesis. Lecture Notes in Computer Science, 2448: 265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480265.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480265.
                                        pdf.

Baudoin:2002:AVL

[808] Geneviève Baudoin, François Capman, Jan Černocký, Fadi El Chami, Maurice Charbit, Gérard Chollet, and Dijana Petrovska-Delacrétaz. Advances in very low bit rate speech coding using recognition and synthesis techniques. Lecture Notes in Computer Science, 2448:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480269.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480269.
                                        pdf.

Demuynck:2002:CDA

[809] Kris Demuynck and Tom Laureys. A comparison of different approaches to automatic speech segmentation. Lecture Notes in Computer Science, 2448: 277–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480277.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480277.
                                        pdf.

Ayed:2002:KSU

[810] Yassine Ben Ayed, Dominique Fohr, Jean Paul Haton, and Gérard Chollet. Keyword spotting using support vector machines. Lecture Notes in Computer Science, 2448:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480285.
pdf.

Aversano:2002:APE

[811] Guido Aversano and Anna Esposito. Automatic parameter estimation for a context-independent speech segmentation algorithm. Lecture Notes in Computer Science, 2448:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480293.
pdf.

Nocera:2002:PLB

[812] Pascal Nocera, Georges Linares, Dominique Massonié, and Loïc Lefort. Phoneme lattice based A* search algorithm for speech recognition. Lecture Notes in Computer Science, 2448: 301–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480301.
pdf.

Prcin:2002:HSM

[813] Michal Prcín and Luděk Müller. Heuristic and statistical methods for speech/non-speech detector design. Lecture Notes in Computer Science, 2448:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2448/24480309.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480309.
                                        pdf.

Karagjosova:2002:ACR

[814] Elena Karagjosova and Ivana Kruijff-Korbayová. An analysis of conditional responses in dialogue. Lecture Notes in Computer Science, 2448:317– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480317.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480317.
                                        pdf.

Matejka:2002:SLI

[815] Pavel Matějka, Petr Schwarz, Martin Karafiát, and Jan Černocký. Some like it Gaussian …. Lecture Notes in Computer Science, 2448:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480321.
                                        pdf.

Kocsor:2002:KSD

[816] András Kocsor and Kornél Kovács. Kernel springy discriminant analysis and its application to a phonological awareness teaching system. Lecture Notes in Computer Science, 2448:325– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480325.
pdf.

Rotovnik:2002:LVS

[817] Tomaž Rotovnik, Mirjam Sepesy Maučec, Bogomir Horvat, and Zdravko Kačič. Large vocabulary speech recognition of Slovenian language using data-driven morphological models. Lecture Notes in Computer Science, 2448: 329–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480329.
pdf.

Zgank:2002:USR

[818] Andrej Žgank, Tomaž Rotovnik, Zdravko Kačič, and Bogomir Horvat. Uniform speech recognition platform for evaluation of new algorithms. Lecture Notes in Computer Science, 2448: 333–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480333.
pdf.

Potamitis:2002:SEU

[819] Ilyas Potamitis, Nikos Fakotakis, Nikos Liolios, and George Kokkinakis. Speech enhancement using mixtures of Gaussians for speech and noise. Lecture Notes in Computer Science, 2448:337– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480337.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480337.
                                        pdf.

Hecht:2002:FGG

[820] Robert Hecht, Jürgen Riedler, and Gerhard Backfried. Fitting German into N-gram language models. Lecture Notes in Computer Science, 2448:341– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480341.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480341.
                                        pdf.

Lublinskaya:2002:ACE

[821] Marina Lublinskaya and Tatiana Sherstinova. Audio collections of endangered Arctic languages in the Russian Federation. Lecture Notes in Computer Science, 2448:347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480347.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480347.
                                        pdf.

Batliner:2002:PCO

[822] Anton Batliner, Viktor Zeißler, Elmar Nöth, and Heinrich Niemann. Prosodic classification of offtalk: First experiments. Lecture Notes in Computer Science, 2448:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480357.
pdf.

Maragoudakis:2002:SDM

[823] Manolis Maragoudakis, Aristomenis Thanopoulos, and Nikos Fakotakis. Statistical decision making from text and dialogue corpora for effective plan recognition. Lecture Notes in Computer Science, 2448:365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480365.
pdf.

Gatius:2002:NLG

[824] Marta Gatius and Horacio Rodríguez. Natural language guided dialogues for accessing the Web. Lecture Notes in Computer Science, 2448:373–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480373.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2448/24480373.
                                        pdf.

Martinez-Hinarejos:2002:EPD

[825] Carlos D. Martínez-Hinarejos and Francisco Casacuberta. Evaluating a probabilistic dialogue model for a railway information task. Lecture Notes in Computer Science, 2448:381– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480381.
                                        pdf.

Garay-Vitoria:2002:EPMa

[826] Nestor Garay-Vitoria, Julio Abascal, and Luis Gardeazabal. Evaluation of prediction methods applied to an inflected language. Lecture Notes in Computer Science, 2448:389– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480389.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480389.
                                        pdf.

Garay-Vitoria:2002:EPMb

[827] Nestor Garay-Vitoria, Julio Abascal, and Luis Gardeazabal. Evaluation of prediction methods applied to an inflected language. Lecture Notes in Computer Science, 2448:397– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2448/24480397.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480397.
pdf.

Yang:2002:KBS

[828] C. K. Yang and L. J. M. Rothkrantz. Knowledge based speech interfacing for handhelds. Lecture Notes in Computer Science, 2448:405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480405.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480405.
pdf.

Mast:2002:DAB

[829] Marion Mast, Thomas Roß, Henrik Schulz, and Heli Harrikari. Different approaches to build multilingual conversational systems. Lecture Notes in Computer Science, 2448:413– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480413.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480413.
pdf.

Arranz:2002:SOP

[830] Victoria Arranz, Núria Castell, and Montserrat Civit. Strategies to overcome problematic input in a Spanish dialogue system. Lecture Notes in Computer Science, 2448:421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480421.
                                        pdf.

Camilleri:2002:DSP

[831] Guy Camilleri. Dialogue systems and planning. Lecture Notes in Computer Science, 2448:429–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480429.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480429.
                                        pdf.

Cenek:2002:FFE

[832] Pavel Cenek. A flexible framework for evaluation of new algorithms for dialogue systems. Lecture Notes in Computer Science, 2448:437–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480437.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480437.
                                        pdf.

Ferreras:2002:HVF

[833] César González Ferreras, David Escudero Mancebo, and Valentín Cardeñoso Payo. From HTML to VoiceXML: A first approach. Lecture Notes in Computer Science, 2448:441–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2448/24480441.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480441.
pdf.

Nouza:2002:VCV

[834] Jan Nouza, Petr Kolář, and Josef Chaloupka. Voice chat with a virtual character: The Good Soldier Švejk case project. Lecture Notes in Computer Science, 2448:445–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480445.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480445.
pdf.

Azzini:2002:ASD

[835] I. Azzini, T. Giorgino, D. Falavigna, and R. Gretter. Application of spoken dialogue technology in a medical domain. Lecture Notes in Computer Science, 2448:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2448/24480449.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2448/24480449.
pdf.

Dobrisek:2002:VDW

[836] Simon Dobrišek, Jerneja Gros, Boštjan Vesnicer, France Mihelič, and Nikola Pavešić. A voice-driven Web browser for blind people. Lecture Notes in Computer Science, 2448:453–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480453.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480453.
                                        pdf.

Horak:2002:EBA

[837] Aleš Horák, Vladimír Kadlec, and Pavel Smrž. Enhancing best analysis selection and parser comparison. Lecture Notes in Computer Science, 2448: 461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2448/24480461.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2448/24480461.
                                        pdf.

Anonymous:2002:AIq

[838] Anonymous. Author index. Lecture Notes in Computer Science, 2448:467– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2448/2448auth.pdf.

Anonymous:2002:SI

[839] Anonymous. Subject index. Lecture Notes in Computer Science, 2448:471– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2448/2448subj.pdf.

Piguet:2002:FQE

[840] Christian Piguet. The first quartz electronic watch. Lecture Notes in Computer Science, 2451:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510001.
pdf.

Helms:2002:IPM

[841] D. Helms, E. Schmidt, A. Schulz, A. Stammermann, and W. Nebel. An improved power macro-model for arithmetic datapath components. Lecture Notes in Computer Science, 2451:16– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510016.
pdf.

Dao:2002:PCV

[842] Hoang Q. Dao and Vojin G. Oklobdzija. Performance comparison of VLSI adders using logical effort. Lecture Notes in Computer Science, 2451:25– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510025.
pdf.

Pessolano:2002:MHP

[843] F. Pessolano, J. Kessels, and A. Peeters. MDSP: A high-performance low-power DSP architecture. Lecture Notes in Computer Science, 2451:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510035.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510035.
                                        pdf.

Carballo:2002:ITP

[844] Juan-Antonio Carballo and Sani R. Nassif. Impact of technology in power-grid-induced noise. Lecture Notes in Computer Science, 2451:45– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510045.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510045.
                                        pdf.

Windschiegl:2002:EML

[845] Armin Windschiegl, Paul Zuber, and Walter Stechele. Exploiting metal layer characteristics for low-power routing. Lecture Notes in Computer Science, 2451:55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510055.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510055.
                                        pdf.

Picot:2002:CMT

[846] F. Picot, P. Coll, and D. Auvergne. Crosstalk measurement technique for CMOS ICs. Lecture Notes in Computer Science, 2451:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510065.
pdf.

Nikolaidis:2002:ISI

[847] S. Nikolaidis, N. Kavvadias, P. Neofotistos, K. Kosmatopoulos, T. Laopoulos, and L. Bisdounis. Instrumentation set-up for instruction level power modeling. Lecture Notes in Computer Science, 2451:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510071.
pdf.

Allier:2002:LPA

[848] Emmanuel Allier, Laurent Fesquet, Marc Renaudin, and Gilles Sicard. Low-power asynchronous A/D conversion. Lecture Notes in Computer Science, 2451:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2451/24510081.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2451/24510081.pdf.

Lemberski:2002:OTL

[849] Igor Lemberski and Mark Josephs. Optimal two-level delay — insensitive implementation of logic functions. Lecture Notes in Computer Science, 2451: 92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510092.
                                        pdf.

Saas:2002:RMC

[850] Christoph Saas and Josef A. Nossek. Resonant multistage charging of dominant capacitances. Lecture Notes in Computer Science, 2451:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510101.
                                        pdf.

Garnica:2002:NMD

[851] Oscar Garnica, Juan Lanchares, and Román Hermida. A new methodology to design low-power asynchronous circuits. Lecture Notes in Computer Science, 2451:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510108.
                                        pdf.

Blotti:2002:DCL

[852] Antonio Blotti, Maurizio Castellucci, and Roberto Saletti. Designing carry look-ahead adders with an adiabatic logic standard-cell library. Lecture Notes in Computer Science, 2451:118– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510118.
pdf.

Oklobdzija:2002:CCS

[853] Vojin G. Oklobdzija. Clocking and clocked storage elements in multi-GHz environment. Lecture Notes in Computer Science, 2451:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510128.
pdf.

Mahnke:2002:DSV

[854] Torsten Mahnke, Walter Stechele, and Wolfgang Hoeld. Dual supply voltage scaling in a conventional power-driven logic synthesis environment. Lecture Notes in Computer Science, 2451:146– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510146.
pdf.

Landrault:2002:TLS

[855] A. Landrault, L. Pellier, A. Richard, C. Jay, M. Robert, and D. Auvergne.

Transistor level synthesis dedicated to fast I.P. prototyping. Lecture Notes in Computer Science, 2451:156– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510156.
pdf.

Aloul:2002:RSB

[856] Fadi A. Aloul, Soha Hassoun, Karem A. Sakallah, and David Blaauw. Robust SAT-based search algorithm for leakage power reduction. Lecture Notes in Computer Science, 2451:167– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510167.
pdf.

Choi:2002:PZP

[857] Kyu won Choi and Abhijit Chatterjee. PA-ZSA (power-aware zero-slack algorithm): A graph-based timing analysis for ultra-low power CMOS VLSI. Lecture Notes in Computer Science, 2451: 178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510178.
pdf.

Gonzalez:2002:NME

[858] Daniel González, Antonio García, Graham A. Jullien, Javier Ramírez, Luis Parrilla, and Antonio Lloris. A new methodology for efficient synchronization of RNS-based VLSI systems. Lecture Notes in Computer Science, 2451: 188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510188.
pdf.

Casu:2002:CDN

[859] M. R. Casu, M. Graziano, G. Masera, G. Piccinini, M. M. Prono, and M. Zamboni. Clock distribution network optimization under self-heating and timing constraints. Lecture Notes in Computer Science, 2451:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510198.
pdf.

Jimenez:2002:TGC

[860] Raúl Jiménez, Pilar Parra, Pedro Sanmartín, and Antonio Acosta. A technique to generate CMOS VLSI flipflops based on differential latches. Lecture Notes in Computer Science, 2451: 209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510209.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510209.
                                        pdf.

Rossello:2002:CCB

[861] José Luis Rossello and Jaume Segura. A compact charge-based propagation delay model for submicronic CMOS buffers. Lecture Notes in Computer Science, 2451:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510219.
                                        pdf.

Nikolaidis:2002:OWE

[862] S. Nikolaidis, H. Pournara, and A. Chatzigeorgiou. Output waveform evaluation of basic pass transistor structure. Lecture Notes in Computer Science, 2451:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510229.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510229.
                                        pdf.

Alioto:2002:AEC

[863] M. Alioto, G. Palumbo, and M. Poli. An approach to energy consumption modeling in RC ladder circuits. Lecture Notes in Computer Science, 2451: 239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2451/24510239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510239.
pdf.

Maurine:2002:SIR

[864] P. Maurine, N. Azemard, and D. Auvergne. Structure independent representation of output transition time for CMOS library. Lecture Notes in Computer Science, 2451:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510247.
pdf.

Jayapala:2002:LEC

[865] Murali Jayapala, Francisco Barat, Pieter Op de Beeck, Francky Catthoor, Geert Deconinck, and Henk Corporaal. A low energy clustered instruction memory hierarchy for long instruction word processors. Lecture Notes in Computer Science, 2451:258– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510258.
pdf.

Zhao:2002:DRL

[866] Xue mei Zhao and Yi zheng Ye. Design and realization of a low power register file using energy model. Lecture Notes in Computer Science, 2451:268– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510268.
                                        pdf.

Takamura:2002:RFE

[867] Hiroshi Takamura, Koji Inoue, and Vasily G. Moshnyaga. Register file energy reduction by operand data reuse. Lecture Notes in Computer Science, 2451:278–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2451/24510278.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510278.
                                        pdf.

Ponomarev:2002:EED

[868] Dmitry Ponomarev, Gurhan Kucuk, and Kanad Ghose. Energy-efficient design of the reorder buffer. Lecture Notes in Computer Science, 2451:289– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510289.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510289.
                                        pdf.

Itoh:2002:TUV

[869] Kiyoo Itoh. Trends in ultralowvoltage RAM technology. Lecture Notes in Computer Science, 2451:300– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2451/24510300.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510300.
pdf.

Benini:2002:ODP

[870] Luca Benini, Alberto Macii, and Enrico Macii. Offline data profiling techniques to enhance memory compression in embedded systems. Lecture Notes in Computer Science, 2451:314– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510314.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510314.
pdf.

Zervas:2002:PPC

[871] N. D. Zervas, G. Pagkless, M. Dasigenis, and D. Soudris. Performance and power comparative study of discrete wavelet transform on programmable processors. Lecture Notes in Computer Science, 2451:323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510323.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510323.
pdf.

Senn:2002:PCE

[872] Eric Senn, Nathalie Julien, Johann Laurent, and Eric Martin. Power consumption estimation of a C program for data-intensive applications. Lecture Notes in Computer Science, 2451:332– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510332.
                                        pdf.

Kretzschmar:2002:LOA

[873] Claudia Kretzschmar, Robert Siegmund, and Dietmar Müller. A low overhead auto-optimizing bus encoding scheme for low power data transmission. Lecture Notes in Computer Science, 2451:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510342.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510342.
                                        pdf.

Baena:2002:MSA

[874] C. Baena, J. Juan-Chico, M. J. Bellido, P. Ruiz de Clavijo, C. J. Jiménez, and M. Valencia. Measurement of the switching activity of CMOS digital circuits at the gate level. Lecture Notes in Computer Science, 2451:353– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510353.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510353.
                                        pdf.

Sutter:2002:LPF

[875] G. Sutter, E. Todorovich, S. Lopez-Buedo, and E. Boemo. Lowpower FSMs in FPGA: Encoding alternatives. Lecture Notes in Computer Science, 2451:363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510363.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510363.
pdf.

Linares-Barranco:2002:SGE

[876] Alejandro Linares-Barranco, Gabriel Jiménez, Antón Civit, and Bernabé Linares-Barranco. Synthetic generation of events for address-event-representation communications. Lecture Notes in Computer Science, 2451: 371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510371.
pdf.

Sato:2002:REC

[877] Toshinori Sato and Itsujiro Arita. Reducing energy consumption via lowcost value prediction. Lecture Notes in Computer Science, 2451:380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510380.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510380.
pdf.

Salhiene:2002:DVS

[878] Mohammed Es Salhiene, Laurent Fesquet, and Marc Renaudin. Dynamic voltage scheduling for real time asynchronous systems. Lecture Notes in Computer Science, 2451:390–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510390.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510390.
                                        pdf.

Ruiz-de-Clavijo:2002:EFC

[879] Paulino Ruiz de Clavijo, Jorge Juan, Manuel J. Bellido, Alejandro Millán, and David Guerrero. Efficient and fast current curve estimation of CMOS digital circuits at the logic level. Lecture Notes in Computer Science, 2451:400– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510400.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510400.
                                        pdf.

Masselos:2002:PEV

[880] Kostas Masselos, Panagiotis Merakos, and Costas E. Goutis. Power efficient vector quantization design using pixel truncation. Lecture Notes in Computer Science, 2451:409–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510409.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2451/24510409.
pdf.

Wroblewski:2002:MSS

[881] Artur Wróblewski, Florian Auernhammer, and Josef A. Nossek. Minimizing spurious switching activities in CMOS circuits. Lecture Notes in Computer Science, 2451:419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510419.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510419.
pdf.

Alioto:2002:MPD

[882] M. Alioto and G. Palumbo. Modeling propagation delay of MUX, XOR, and D-latch source-coupled logic gates. Lecture Notes in Computer Science, 2451: 429–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510429.
pdf.

Cappuccino:2002:ORM

[883] Gregorio Cappuccino and Giuseppe Cocorullo. Operating region modelling and timing analysis of CMOS gates driving transmission lines. Lecture Notes in Computer Science, 2451:438– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510438.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510438.
                                        pdf.

Parra:2002:SCG

[884] Pilar Parra, Antonio Acosta, and Manuel Valencia. Selective clockgating for low power/low noise synchronous counters. Lecture Notes in Computer Science, 2451:448–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510448.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510448.
                                        pdf.

Freimann:2002:PPE

[885] Achim Freimann. Probabilistic power estimation for digital signal processing architectures. Lecture Notes in Computer Science, 2451:458–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510458.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2451/24510458.
                                        pdf.

Mita:2002:MPD

[886] Rosario Mita and Gaetano Palumbo. Modeling of propagation delay of a first order circuit with a ramp input. Lecture Notes in Computer Science, 2451: 468–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2451/24510468.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2451/24510468.
pdf.

Millan:2002:CNP

[887] Alejandro Millán, Jorge Juan, Manuel J. Bellido, Paulino Ruiz de Clavijo, and David Guerrero. Characterization of normal propagation delay for delay degradation model (DDM). Lecture Notes in Computer Science, 2451:477– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510477.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510477.
pdf.

Ionita:2002:ADM

[888] Razvan Ionita, Andrei Vladimirescu, and Paul Jespers. Automated design methodology for CMOS analog circuit blocks in complex systems. Lecture Notes in Computer Science, 2451:487– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2451/24510487.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2451/24510487.
pdf.

Anonymous:2002:AIr

[889] Anonymous. Author index. Lecture Notes in Computer Science, 2451:495– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2451/2451auth.pdf.

Abiteboul:2002:IMW

[890] S. Abiteboul. Issues in monitoring Web data. Lecture Notes in Computer Science, 2453:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530001.
                                        pdf.

Vassilakis:2002:CCD

[891] C. Vassilakis and G. Lepouras. Controlled caching of dynamic WWW pages. Lecture Notes in Computer Science, 2453:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2453/24530009.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2453/24530009.pdf.

Dhyani:2002:DVS

[892] D. Dhyani, S. S. Bhowmick, and W. K. Ng. Deriving and verifying statistical distribution of a hyperlink-based Web page quality metric. Lecture Notes in Computer Science, 2453:19– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530019.
                                        pdf.

Shikata:2002:MRA

[893] M. Shikata, A. Nadamoto, K. Sumiya, and K. Tanaka. MWM: Retrieval and automatic presentation of manual data for mobile terminals. Lecture Notes in Computer Science, 2453:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530029.
pdf.

Younas:2002:ERS

[894] M. Younas and B. Eaglestone. Ensuring recovery for SACReD Web transactions in the E-commerce applications. Lecture Notes in Computer Science, 2453:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2453/24530040.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2453/24530040.
pdf.

Zhao:2002:WSA

[895] X. Zhao. Workflow simulation across multiple workflow domains. Lecture Notes in Computer Science, 2453:50– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530050.
pdf.

Grefen:2002:TWW

[896] P. Grefen. Transactional workflows or workflow transactions? Lecture Notes in Computer Science, 2453:60– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530060.
                                        pdf.

Lonchamp:2002:CFM

[897] J. Lonchamp. Collaboration flow management: A new paradigm for virtual team support. Lecture Notes in Computer Science, 2453:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530070.
                                        pdf.

Chun:2002:DKB

[898] S. A. Chun, V. Atluri, and N. R. Adam. Domain knowledge-based automatic workflow generation. Lecture Notes in Computer Science, 2453:81– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530081.
                                        pdf.

Cai:2002:CAE

[899] M. Cai, S. Ghandeharizadeh, R. Schmidt, and S. Song. A comparison of alternative encoding mechanisms for Web services. Lecture Notes in Computer Science, 2453:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530093.
pdf.

Liu:2002:CSH

[900] M. Liu. Capturing semantics in HTML documents. Lecture Notes in Computer Science, 2453:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530103.
pdf.

Ladner:2002:ASD

[901] R. Ladner and F. E. Petry. Assessment of spatial data mining tools for integration into an objectoriented GIS (GIDB). Lecture Notes in Computer Science, 2453:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530113.
pdf.

Naggar:2002:MTS

[902] P. Naggar, L. Pontieri, M. Pupo, G. Terracina, and E. Virardi. A model and a toolkit for supporting incremental data warehouse construction. Lecture Notes in Computer Science, 2453: 123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530123.
                                        pdf.

Kaya:2002:EAM

[903] M. Kaya, R. Alhajj, F. Polat, and A. Arslan. Efficient automated mining of fuzzy association rules. Lecture Notes in Computer Science, 2453:133– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530133.
                                        pdf.

Kim:2002:TPS

[904] S.-D. Kim, J. W. Lee, J. Lee, and J. Chae. A two-phase stock trading system using distributional differences. Lecture Notes in Computer Science, 2453:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2453/24530143.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530143.
                                        pdf.

Lee:2002:MAQ

[905] J. W. Lee and J. O. A multiagent Q-learning framework for optimizing stock trading systems. Lecture Notes in Computer Science, 2453:153– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530153.
pdf.

Kroha:2002:UAP

[906] P. Kroha and S. Bürgel. Using adaptive programming in legal advisory systems. Lecture Notes in Computer Science, 2453:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530163.
pdf.

Matsumoto:2002:WIR

[907] C. Matsumoto, Q. Ma, and K. Tanaka. Web information retrieval based on the localness degree. Lecture Notes in Computer Science, 2453:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530172.
pdf.

Goda:2002:RTL

[908] K. Goda, T. Tamura, M. Oguchi, and M. Kitsuregawa. Run-time load balancing system on SAN-connected PC cluster for dynamic injection of CPU and disk resource – A case study of data mining application –. Lecture Notes in Computer Science, 2453:182– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530182.
pdf.

Ling:2002:VSP

[909] T. W. Ling, W. C. Low, Z. W. Luo, S. Y. Lee, and H. G. Li. Variable sized partitions for range query algorithms. Lecture Notes in Computer Science, 2453:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530193.
pdf.

List:2002:CDW

[910] B. List, R. M. Bruckner, K. Machaczek, and J. Schiefer. A comparison of data warehouse development methodologies case study of the process warehouse. Lecture Notes in Computer Science, 2453:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2453/24530203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530203.
pdf.

Wookey:2002:SMD

[911] L. Wookey, Y. Hwang, S.-H. Kang, S. Kim, C. Kim, and Y. Lee. Selfmaintainable data warehouse views using differential files. Lecture Notes in Computer Science, 2453:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530216.
pdf.

Dube:2002:UER

[912] K. Dube, B. Wu, and J. B. Grimson. Using ECA rules in database systems to support clinical protocols. Lecture Notes in Computer Science, 2453:226– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530226.
pdf.

Quirchmayr:2002:EVE

[913] G. Quirchmayr, Z. Milosevic, R. Tagg, J. Cole, and S. Kulkarni. Establishment of virtual enterprise contracts. Lecture Notes in Computer Science, 2453:236– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2453/24530236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530236.
                                        pdf.

Dzemydiene:2002:MRA

[914] D. Dzemydiene and V. Rudzkiene. Multiple regression analysis in crime pattern warehouse for decision support. Lecture Notes in Computer Science, 2453:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2453/24530249.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530249.
                                        pdf.

Wouters:2002:PWO

[915] C. Wouters, T. Dillon, W. Rahayu, and E. Chang. A practical walkthrough of the ontology derivation rules. Lecture Notes in Computer Science, 2453:259– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530259.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530259.
                                        pdf.

Arenas:2002:WHA

[916] M. Arenas, W. Fan, and L. Libkin. What’s hard about XML schema constraints? Lecture Notes in Computer Science, 2453:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2453/24530269.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2453/24530269.pdf.

Kha:2002:ARP

[917] D. D. Kha, M. Yoshikawa, and S. Uemura. Application of rUID in processing XML queries on structure and keyword. Lecture Notes in Computer Science, 2453:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530279.
pdf.

Chatvichienchai:2002:TAA

[918] S. Chatvichienchai, M. Iwaihara, and Y. Kambayashi. Translating access authorizations for transformed XML documents. Lecture Notes in Computer Science, 2453:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530290.
pdf.

Perez:2002:XXB

[919] J. M. Pérez, M. J. Aramburu, and R. Berlanga. XRL: A XML-based query language for advanced services in digital libraries. Lecture Notes in Computer Science, 2453:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2453/24530300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530300.
                                        pdf.

Palmirani:2002:NSL

[920] M. Palmirani and R. Brighi. Normasystem: A legal document system for managing consolidated acts. Lecture Notes in Computer Science, 2453:310– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530310.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530310.
                                        pdf.

Lora:2002:EMP

[921] A. Troncoso Lora, José Riquelme Santos, Jesús Riquelme Santos, J. L. Martínez Ramos, and A. Gómez Expósito. Electricity market price forecasting: Neural networks versus weighted-distance k nearest neighbours. Lecture Notes in Computer Science, 2453:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2453/24530321.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2453/24530321.pdf.

Chung:2002:TPM

[922] J. D. Chung, O. H. Paek, J. W. Lee, and K. H. Ryu. Temporal pattern mining of moving objects for location-based service. Lecture Notes in Computer Science, 2453:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530331.
pdf.

Varde:2002:MCV

[923] A. S. Varde and E. A. Rundensteiner. MEDWRAP: Consistent view maintenance over distributed multi-relation sources. Lecture Notes in Computer Science, 2453:341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530341.
pdf.

Ruberg:2002:ECP

[924] G. Ruberg, F. Baião, and M. Mattoso. Estimating costs of path expression evaluation in distributed object databases. Lecture Notes in Computer Science, 2453:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530351.
pdf.

Perich:2002:PDD

[925] F. Perich, S. Avancha, D. Chakraborty, A. Joshi, and Yelena Yesha. Profile driven data management for pervasive environments. Lecture Notes in Computer Science, 2453:361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530361.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530361.
                                        pdf.

Gobel:2002:IUI

[926] S. Göbel, J. Haist, H. Reiterer, and F. Müller. INVISIP: Usage of information visualization techniques to access geospatial data archives. Lecture Notes in Computer Science, 2453:371– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530371.
                                        pdf.

Manco:2002:ETI

[927] G. Manco, C. Pizzuti, and D. Talia. Eureka!: A tool for interactive knowledge discovery. Lecture Notes in Computer Science, 2453:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530381.
                                        pdf.

Kuramitsu:2002:ETC

[928] K. Kuramitsu and K. Sakamura. Electronic tickets on contactless Smartcard database. Lecture Notes in Computer Science, 2453:392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530392.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530392.
pdf.

Ye:2002:MAD

[929] D. Y. Ye, M. C. Lee, and T. I. Wang. Mobile agents for distributed transactions of a distributed heterogeneous database system. Lecture Notes in Computer Science, 2453:403– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530403.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530403.
pdf.

Ou:2002:ADB

[930] D.-X. Ou, K.-Y. Lam, and D.-C. Dong. An adaptive direction-based location update scheme for next generation PCS networks. Lecture Notes in Computer Science, 2453:413–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530413.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530413.
pdf.

Vargas-Solar:2002:AAE

[931] G. Vargas-Solar and C. Collet. ADEES: An adaptable and extensible event based infrastructure. Lecture Notes in Computer Science, 2453:423–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530423.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530423.
                                        pdf.

Bramer:2002:UJP

[932] M. Bramer. Using J-pruning to reduce overfitting of classification rules in noisy domains. Lecture Notes in Computer Science, 2453:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530433.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530433.
                                        pdf.

Hajji:2002:FRO

[933] H. Hajji and E. Terzi. A framework for reasoning on objects with geometric constraints. Lecture Notes in Computer Science, 2453:443–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530443.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530443.
                                        pdf.

Bettini:2002:DAV

[934] C. Bettini and S. Ruffini. Deriving abstract views of multi-granularity temporal constraint networks. Lecture Notes in Computer Science, 2453:454– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530454.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530454.
pdf.

Fu:2002:RCB

[935] G. Fu, J. Shao, S. M. Embury, and W. A. Gray. Representing constraint business rules extracted from legacy systems. Lecture Notes in Computer Science, 2453:464–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530464.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530464.
pdf.

Debenham:2002:AIP

[936] J. Debenham. Agents for industry process management. Lecture Notes in Computer Science, 2453:474– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530474.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530474.
pdf.

Zhao:2002:NED

[937] L. Zhao, S. S. Yuan, S. Peng, and L. T. Wang. A new efficient data cleansing method. Lecture Notes in Computer Science, 2453:484–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530484.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530484.
                                        pdf.

OMahony:2002:PRA

[938] M. P. O’Mahony, N. J. Hurley, and G. C. M. Silvestre. Promoting recommendations: An attack on collaborative filtering. Lecture Notes in Computer Science, 2453:494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530494.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530494.
                                        pdf.

Roger:2002:BTD

[939] M. Roger, A. Simonet, and M. Simonet. Bringing together description logics and database in an object oriented model. Lecture Notes in Computer Science, 2453:504–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530504.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530504.
                                        pdf.

Jodlowski:2002:ORS

[940] A. Jodłowski, P. Habela, J. Płodzien, and K. Subieta. Objects and roles in the stack-based approach. Lecture Notes in Computer Science, 2453:514– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530514.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530514.
pdf.

Galante:2002:DSE

[941] R. de Matos Galante, A. Bueno da Silva Roma, A. Jantsch, N. Edelweiss, and C. Saraiva dos Santos. Dynamic schema evolution management using version in temporal objectoriented databases. Lecture Notes in Computer Science, 2453:524–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530524.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530524.
pdf.

Stumme:2002:EDM

[942] G. Stumme. Efficient data mining based on formal concept analysis. Lecture Notes in Computer Science, 2453: 534–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530534.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530534.
pdf.

Galton:2002:TAE

[943] A. Galton and J. C. Augusto. Two approaches to event definition. Lecture Notes in Computer Science, 2453:547– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530547.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530547.
                                        pdf.

Artelsmair:2002:CAT

[944] C. Artelsmair, W. Essmayr, P. Lang, R. Wagner, and E. Weippl. CoSMo: An approach towards Co nceptual S ecurity Mo deling. Lecture Notes in Computer Science, 2453:557–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530557.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530557.
                                        pdf.

Harr:2002:ECG

[945] R. Harr. Exploring the concept of group interaction through action in a mobile context. Lecture Notes in Computer Science, 2453:567–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530567.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530567.
                                        pdf.

Guan:2002:IST

[946] W. Guan and R. K. Wong. Intersystem triggering for heterogeneous database systems. Lecture Notes in Computer Science, 2453:577–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530577.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530577.
pdf.

Lowden:2002:CIR

[947] B. G. T. Lowden and J. Robinson. Constructing inter-relational rules for semantic query optimisation. Lecture Notes in Computer Science, 2453:587– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530587.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530587.
pdf.

Bosc:2002:ASJ

[948] P. Bosc, L. Duval, and O. Pivert. About selections and joins in possibilistic queries addressed to possibilistic databases. Lecture Notes in Computer Science, 2453:597–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530597.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530597.
pdf.

Khan:2002:LSC

[949] S. Khan and P. L. Mott. LeedsCQ: A scalable continual queries system. Lecture Notes in Computer Science, 2453:

607–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530607.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530607.
                                        pdf.

Moro:2002:TTV

[950] M. Moura Moro, N. Edelweiss, A. Pereira Zaupa, and C. Saraiva dos Santos. TVQL — temporal versioned query language. Lecture Notes in Computer Science, 2453:618–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530618.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530618.
                                        pdf.

Ramadour:2002:AMB

[951] P. Ramadour and C. Cauvet. Approach and model for business components specification. Lecture Notes in Computer Science, 2453:628–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530628.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530628.
                                        pdf.

Gams:2002:TFP

[952] E. Gams, T. Berka, and S. Reich. The TrailTRECer framework — A platform for trail-enabled recommender applications. Lecture Notes in Computer Science, 2453:638–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530638.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530638.
pdf.

Loh:2002:AMR

[953] Z. X. Loh, T. W. Ling, C. H. Ang, and S. Y. Lee. Adaptive method for range top- k queries in OLAP data cubes. Lecture Notes in Computer Science, 2453:648–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530648.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530648.
pdf.

Tok:2002:DSD

[954] W. H. Tok and S. Bressan. dbRouter — A scaleable and distributed query optimization and processing framework. Lecture Notes in Computer Science, 2453:658–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530658.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530658.
pdf.

Frikken:2002:OPR

[955] K. Frikken, M. Atallah, S. Prabhakar, and R. Safavi-Naini. Optimal parallel I/O for range queries through replication. Lecture Notes in Computer Science, 2453:669–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530669.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530669.
pdf.

Werstein:2002:DRU

[956] P. Werstein and J. R. McDonald. A database for repetitive, unpredictable moving objects. Lecture Notes in Computer Science, 2453:679– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530679.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530679.
pdf.

Reiner:2002:HSS

[957] B. Reiner, K. Hahn, G. Höfling, and P. Baumann. Hierarchical storage support and management for largescale multidimensional array database management systems. Lecture Notes in Computer Science, 2453:689–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530689.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530689.
pdf.

Daassi:2002:MVR

[958] C. Daassi, M.-C. Fauvet, and L. Nigay. Multiple visual representation of temporal data. Lecture Notes in Computer Science, 2453:701–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530701.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530701.
pdf.

Gryz:2002:QOE

[959] J. Gryz and D. Liang. Query optimization via empty joins. Lecture Notes in Computer Science, 2453:710– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530710.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530710.
pdf.

Chen:2002:EPX

[960] Y. Chen, S. Madria, K. Passi, and S. Bhowmick. Efficient processing of XPath queries using indexes. Lecture Notes in Computer Science, 2453:721– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530721.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530721.
pdf.

Kalashnikov:2002:EEC

[961] D. Kalashnikov, S. Prabhakar, S. Hambrusch, and W. Aref. Efficient evaluation of continuous range queries on moving objects. Lecture Notes in Computer Science, 2453:731–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530731.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530731.
pdf.

Shin:2002:PBS

[962] H. Shin, B. Moon, and S. Lee. Partition-based similarity join in high dimensional data spaces. Lecture Notes in Computer Science, 2453:741– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530741.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530741.
pdf.

Bouzeghoub:2002:HDS

[963] M. Bouzeghoub, B. Farias Lóscio, Z. Kedad, and A. Soukane. Heterogeneous data source integration and evolution (extended abstract). Lecture Notes in Computer Science, 2453:751– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530751.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530751.
pdf.

Hatano:2002:IRS

[964] K. Hatano, H. Kinutani, M. Yoshikawa, and S. Uemura. Information retrieval system for XML documents. Lecture Notes in Computer Science, 2453:758– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530758.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530758.
pdf.

Ueda:2002:SRD

[965] T. Ueda, T. Amagasa, M. Yoshikawa, and S. Uemura. A system for retrieval and digest creation of video data based on geographic objects. Lecture Notes in Computer Science, 2453:768– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530768.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530768.
pdf.

Tada:2002:SIR

[966] M. Tada, T. Kato, and I. Shinohara. Similarity image retrieval system using hierarchical classification. Lecture Notes in Computer Science, 2453:779– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530779.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530779.
pdf.

Hemnani:2002:IET

[967] A. Hemnani and S. Bressan. Information extraction — tree alignment approach to pattern discovery in Web documents. Lecture Notes in Computer Science, 2453:789–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530789.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530789.
pdf.

Vossough:2002:PCQ

[968] E. Vossough and J. R. Getta. Processing of continuous queries over unlimited data streams. Lecture Notes in Computer Science, 2453:799–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530799.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530799.
pdf.

Dang:2002:IIH

[969] T. K. Dang, J. Küng, and R. Wagner. ISA — an incremental hypersphere approach for efficiently solving complex vague queries. Lecture Notes in Computer Science, 2453:810– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530810.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530810.
pdf.

Hahn:2002:PQS

[970] K. Hahn, B. Reiner, G. Höfling, and P. Baumann. Parallel query support for multidimensional data: Interobject parallelism. Lecture Notes in Computer Science, 2453:820–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530820.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530820.
pdf.

Guan:2002:PTC

[971] J. Guan and S. Zhou. Pruning training corpus to speedup text classification. Lecture Notes in Computer Science, 2453:831–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530831.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530831.
pdf.

Park:2002:TRV

[972] Y. Park, P. Kim, W. Kim, J.-J. Song, and S. Panchanathan. Towards retrieval of visual information based on the semantic models. Lecture Notes in Computer Science, 2453: 841–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530841.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530841.
pdf.

Song:2002:ORS

[973] J.-J. Song and F. Golshani. 3D object retrieval by shape similarity. Lecture Notes in Computer Science, 2453: 851–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530851.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530851.
                                        pdf.

Buccafurri:2002:BTH

[974] F. Buccafurri, F. Furfaro, G. Lax, and D. Saccà. Binary-tree histograms with tree indices. Lecture Notes in Computer Science, 2453:861–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530861.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530861.
                                        pdf.

Ong:2002:IGD

[975] T.-H. Ong, K.-L. Tan, and H. Wang. Indexing genomic databases for fast homology searching. Lecture Notes in Computer Science, 2453:871–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530871.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530871.
                                        pdf.

Stockinger:2002:BIS

[976] K. Stockinger. Bitmap indices for speeding up high-dimensional data analysis. Lecture Notes in Computer Science, 2453:881–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2453/24530881.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2453/24530881.pdf.

Sugiyama:2002:MIF

[977] K. Sugiyama, K. Hatano, M. Yoshikawa, and S. Uemura. A method of improving feature vector for Web pages reflecting the contents of their out-linked pages. Lecture Notes in Computer Science, 2453:891–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530891.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530891.
pdf.

Wang:2002:CLC

[978] Y. Wang and M. Kitsuregawa. On combining link and contents information for Web page clustering. Lecture Notes in Computer Science, 2453:902– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530902.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530902.
pdf.

Cappelli:2002:MTH

[979] R. Cappelli, A. Lumini, and D. Maio. MKL-Tree: A hierarchical data structure for indexing multidimensional data. Lecture Notes in Computer Science, 2453:914–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2453/24530914.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2453/24530914.pdf.

Thomas:2002:BBP

[980] M. Thomas and J. M. Hellerstein. Boolean bounding predicates for spatial access methods. Lecture Notes in Computer Science, 2453:925–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530925.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530925.
                                        pdf.

Cha:2002:CIT

[981] G.-H. Cha and Y.-I. Yoon. Clustered indexing technique for multidimensional index structures. Lecture Notes in Computer Science, 2453:935– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2453/24530935.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2453/24530935.
                                        pdf.

Tjoa:2002:DES

[982] A. M. Tjoa and R. R. Wagner. Database and expert systems 2002 “Quo vadis”? Lecture Notes in Computer Science, 2453:945–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2453/24530945.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2453/24530945.
pdf.

Anonymous:2002:AIs

[983] Anonymous. Author index. Lecture Notes in Computer Science, 2453:949– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2453/2453auth.pdf.

Botta:2002:CBQ

[984] Marco Botta, Jean-François Boulicaut, Cyrille Masson, and Rosa Meo. A comparison between query languages for the extraction of association rules. Lecture Notes in Computer Science, 2454: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540001.
pdf.

Teng:2002:LD

[985] Choh Man Teng. Learning from dissociations. Lecture Notes in Computer Science, 2454:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2454/24540011.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2454/24540011.pdf.

Braga:2002:MAR

[986] Daniele Braga, Alessandro Campi, Mika Klemettinen, and PierLuca Lanzi. Mining association rules from XML data. Lecture Notes in Computer Science, 2454:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540021.
pdf.

Li:2002:EJP

[987] Tao Li, Shenghuo Zhu, Mitsunori Ogihara, and Yinhe Cheng. Estimating joint probabilities from marginal ones. Lecture Notes in Computer Science, 2454:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540031.
pdf.

Yun:2002:STC

[988] Ching-Huang Yun, Kun-Ta Chuang, and Ming-Syan Chen. Self-tuning clustering: An adaptive clustering method for transaction data. Lecture Notes in Computer Science, 2454:42– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540042.
pdf.

Zhu:2002:CAN

[989] Shenghuo Zhu, Tao Li, and Mitsuonri Ogihara. CoFD: An algorithm for non-distance based clustering in high dimensional spaces. Lecture Notes in Computer Science, 2454:52– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540052.
pdf.

Chu:2002:EMB

[990] Shu-Chuan Chu, John F. Roddick, and J. S. Pan. An efficient Kmedoids-based algorithm using previous medoid index, triangular inequality elimination criteria, and partial distance search. Lecture Notes in Computer Science, 2454:63–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540063.
pdf.

Jespersen:2002:HAW

[991] Søren E. Jespersen, Jesper Thorhauge, and Torben Bach Pedersen. A hybrid approach to Web usage mining. Lecture Notes in Computer Science, 2454: 73–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540073.
                                        pdf.

Pohle:2002:BEA

[992] Carsten Pohle and Myra Spiliopoulou. Building and exploiting ad hoc concept hierarchies for Web log analysis. Lecture Notes in Computer Science, 2454: 83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540083.
                                        pdf.

Bhargava:2002:ABE

[993] Bharat Bhargava and Yuhui Zhong. Authorization based on evidence and trust. Lecture Notes in Computer Science, 2454:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2454/24540094.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2454/24540094.pdf.

Terzi:2002:ABU

[994] Evimaria Terzi, Yuhui Zhong, Bharat Bhargava, Pankaj, and Sanjay Madria. An algorithm for building user-role profiles in a trust environment. Lecture Notes in Computer Science, 2454:104– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540104.
pdf.

Lee:2002:NBA

[995] Yue-Shi Lee and Show-Jane Yen. Neural-based approaches for improving the accuracy of decision trees. Lecture Notes in Computer Science, 2454:114– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540114.
pdf.

Angiulli:2002:ACP

[996] Fabrizio Angiulli and Clara Pizzuti. Approximate k-closest-pairs with space filling curves. Lecture Notes in Computer Science, 2454:124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540124.
pdf.

Bohm:2002:ODO

[997] Christian Böhm, Florian Krebs, and Hans-Peter Kriegel. Optimal dimension order: A generic technique for the similarity join. Lecture Notes in Computer Science, 2454:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540135.
                                        pdf.

Lin:2002:FDS

[998] Ming-Yen Lin and Suh-Yin Lee. Fast discovery of sequential patterns by memory indexing. Lecture Notes in Computer Science, 2454:150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540150.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540150.
                                        pdf.

Zhao:2002:DSF

[999] Li Zhao, Sung Sam Yuan, Qi Xiao Yang, and Sun Peng. Dynamic similarity for fields with NULL values. Lecture Notes in Computer Science, 2454: 161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540161.
                                        pdf.

Hawkins:2002:ODU

[1000] Simon Hawkins, Hongxing He, Graham J. Williams, and Rohan Baxter. Outlier detection using replicator neural networks. Lecture Notes in Computer Science, 2454:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540170.
pdf.

Luong:2002:CKB

[1001] Viet Phan Luong. The closed keys base of frequent itemsets. Lecture Notes in Computer Science, 2454:181– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540181.
pdf.

Lee:2002:NRA

[1002] Yujin Lee, Wootaek Kim, and Kyungsook Han. New representation and algorithm for drawing RNA structure with pseudoknots. Lecture Notes in Computer Science, 2454:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540191.
pdf.

Viaene:2002:BNB

[1003] Stijn Viaene, Richard Derrig, and Guido Dedene. Boosting naive Bayes for claim fraud diagnosis. Lecture Notes in Computer Science, 2454:202– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540202.
                                        pdf.

Ko:2002:OAW

[1004] Su-Jeong Ko and Jung-Hyun Lee. Optimization of association word knowledge base through genetic algorithm. Lecture Notes in Computer Science, 2454:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2454/24540212.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540212.
                                        pdf.

Lin:2002:MTP

[1005] Weiqiang Lin, Mehmet A. Orgun, and Graham J. Williams. Mining temporal patterns from health care data. Lecture Notes in Computer Science, 2454: 222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540222.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540222.
                                        pdf.

Bizarro:2002:APO

[1006] Pedro Bizarro and Henrique Madeira.

Adding a performance-oriented perspective to data warehouse design. Lecture Notes in Computer Science, 2454: 232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540232.
pdf.

Pedersen:2002:CME

[1007] Dennis Pedersen, Karsten Riis, and Torben Bach Pedersen. Cost modeling and estimation for OLAP-XML federations. Lecture Notes in Computer Science, 2454:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540245.
pdf.

Bhowmick:2002:CFJ

[1008] Sourav S. Bhowmick, Wee Keong Ng, Sanjay Madria, and Mukesh Mohania. Constraint-free join processing on hyperlinked Web data. Lecture Notes in Computer Science, 2454:255– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540255.
pdf.

Ciferri:2002:FDD

[1009] Cristina Dutra de Aguiar Ciferri and Fernando da Fonseca de Souza. Focusing on data distribution in the WebD^{2} W system. Lecture Notes in Computer Science, 2454:265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540265.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540265.
                                        pdf.

Hummer:2002:DMM

[1010] W. Hümmer, W. Lehner, A. Bauer, and L. Schlesinger. A decathlon in multidimensional modeling: Open issues and some solutions. Lecture Notes in Computer Science, 2454:275– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540275.
                                        pdf.

Wu:2002:MIL

[1011] Xintao Wu and Daniel Barbará. Modeling and imputation of large incomplete multidimensional datasets. Lecture Notes in Computer Science, 2454: 286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540286.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540286.
                                        pdf.

Bellatreche:2002:PES

[1012] Ladjel Bellatreche, Michel Schneider, Mukesh Mohania, and Bharat Bhargava. PartJoin: An efficient storage and query execution for data warehouses. Lecture Notes in Computer Science, 2454:296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540296.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540296.
pdf.

Liu:2002:TAP

[1013] Bin Liu, Songting Chen, and Elke A.

Rundensteiner. A transactional approach to parallel data warehouse maintenance. Lecture Notes in Computer Science, 2454:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2454/24540307.
pdf.

Bruckner:2002:STN

[1014] Robert M. Bruckner, Beate List, and Josef Schiefer. Striving towards near real-time data integration for data warehouses. Lecture Notes in Computer Science, 2454:317– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2454/24540317.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2454/24540317.
                                        pdf.

Furtado:2002:TIS

[1015] Pedro Furtado and João Pedro Costa. Time-interval sampling for improved estimations in data warehouses. Lecture Notes in Computer Science, 2454: 327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2454/24540327.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2454/24540327.
                                        pdf.

Anonymous:2002:AIt

[1016] Anonymous. Author index. Lecture Notes in Computer Science, 2454:339– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2454/2454auth.pdf.

Hawryszkiewycz:2002:EVE

[1017] I. T. Hawryszkiewycz. Enabling virtual enterprises: A case for multidisciplinary research. Lecture Notes in Computer Science, 2455:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550001.
                                        pdf.

Kothari:2002:IRR

[1018] Ravi Kothari, Mukesh Mohania, and Yahiko Kambayashi. Increasing realized revenue in a Web based Dutch auction. Lecture Notes in Computer Science, 2455:7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550007.
pdf.

Aknine:2002:SBA

[1019] Samir Aknine. Strategies and behaviours of agents in multi-phased negotiations. Lecture Notes in Computer Science, 2455:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2455/24550017.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2455/24550017.pdf.

Kameshwaran:2002:NAD

[1020] S. Kameshwaran and Y. Narahari.

A new approach to the design of electronic exchanges. Lecture Notes in Computer Science, 2455:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550027.
pdf.

Bichler:2002:WDA

[1021] Martin Bichler, Jayant Kalagnanam, Ho Soo Lee, and Juhnyoung Lee. Winner determination algorithms for electronic auctions: A framework design. Lecture Notes in Computer Science, 2455:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2455/24550037.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2455/24550037.
pdf.

Garcia:2002:WBC

[1022] F. J. García, A. B. Gil, N. Moreno, and B. Curto. A Web-based Ecommerce facilitator intermediary for small and medium enterprises: A B2B/B2C hybrid proposal. Lecture Notes in Computer Science, 2455:47– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550047.
pdf.

Puhretmair:2002:EDM

[1023] Franz Pühretmair, Hildegard Rumetshofer, and Erwin Schaumlechner. Extended decision making in tourism information systems. Lecture Notes in Computer Science, 2455:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550057.
pdf.

Debenham:2002:IAO

[1024] John Debenham. Identifying arbitrage opportunities in E-markets. Lecture Notes in Computer Science, 2455:67– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550067.
pdf.

Lu:2002:ABU

[1025] Jin Lu and Lihui Chen. An architecture for building user-driven Web tasks via Web services. Lecture Notes in Computer Science, 2455:77– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550077.
pdf.

Nambiar:2002:EXD

[1026] Ullas Nambiar, Zo’e Lacroix, Stéphane Bressan, Mong Li Lee, and Ying Guang Li. Efficient XML data management: An analysis. Lecture Notes in Computer Science, 2455:87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550087.
pdf.

Maurino:2002:CTD

[1027] Andrea Maurino and Piero Fraternali. Commercial tools for the development of personalized Web applications: A survey. Lecture Notes in Computer Science, 2455:99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550099.
pdf.

Buccafurri:2002:ABH

[1028] F. Buccafurri, D. Rosaci, G. M. L. Sarnè, and D. Ursino. An agentbased hierarchical clustering approach for E-commerce environments. Lecture Notes in Computer Science, 2455:109– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550109.
pdf.

Younas:2002:MAA

[1029] M. Younas, N. H. Shah, and K.- M. Chao. A multi-agent approach to SACReD transactions for Ecommerce applications. Lecture Notes in Computer Science, 2455:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550119.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2455/24550119.
pdf.

Wang:2002:PDM

[1030] Yan Wang, Kian-Lee Tan, and Xiaolin Pang. A parallel dispatch model with secure and robust routes for mobile agents. Lecture Notes in Computer Science, 2455:129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550129.
pdf.

Kneer:2002:ABF

[1031] Helmut Kneer, Henrik Stormer, Harald Häuschen, and Burkhard Stiller. An agent-based framework for monitoring service contracts. Lecture Notes in Computer Science, 2455:139– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550139.
pdf.

Kozawa:2002:CSC

[1032] Masayuki Kozawa, Mizuho Iwaihara, and Yahiko Kambayashi. Constraint search for comparing multipleincentive merchandises. Lecture Notes in Computer Science, 2455:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2455/24550152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550152.
                                        pdf.

Park:2002:XQP

[1033] Sangwon Park, Yoonra Choi, and Hyoung-Joo Kim. XML query processing using signature and DTD. Lecture Notes in Computer Science, 2455:162– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550162.
                                        pdf.

Dhyani:2002:MPW

[1034] Devanshu Dhyani, Sourav S. Bhowmick, and Wee Keong Ng. Modelling and predicting Web page accesses using burrell’s model. Lecture Notes in Computer Science, 2455:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550172.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550172.
                                        pdf.

Karagiannis:2002:MP

[1035] Dimitris Karagiannis and Harald Kühn. Metamodelling platforms. Lecture Notes in Computer Science, 2455: 182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550182.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2455/24550182.
pdf.

Woss:2002:HEB

[1036] Wolfram Wöß and Anton Dunzendorfer. Homogeneous EDI between heterogeneous Web-based tourism information systems. Lecture Notes in Computer Science, 2455:183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550183.
pdf.

Passi:2002:MXS

[1037] Kalpdrum Passi, Louise Lane, Sanjay Madria, Bipin C. Sakamuri, Mukesh Mohania, and Sourav Bhowmick. A model for XML schema integration. Lecture Notes in Computer Science, 2455:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2455/24550193.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2455/24550193.
pdf.

Lopez:2002:XBD

[1038] Javier López, Antonio Maña, and Mariemma I. Yagüe. XML-based distributed access control system. Lecture Notes in Computer Science, 2455: 203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2455/24550203.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550203.
                                        pdf.

Fahrenholtz:2002:TSD

[1039] Dietrich Fahrenholtz and Winfried Lamersdorf. Transactional security for a distributed reputation management system. Lecture Notes in Computer Science, 2455:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550214.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550214.
                                        pdf.

Mana:2002:PMD

[1040] Antonio Maña and Sonia Matamoros. Practical mobile digital signatures. Lecture Notes in Computer Science, 2455:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550224.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550224.
                                        pdf.

Au:2002:SCA

[1041] Richard Au, Ming Yao, Mark Looi, and Paul Ashley. Secure client agent environment (SCAE) for World Wide Web. Lecture Notes in Computer Science, 2455:234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2455/24550234.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550234.
pdf.

Ko:2002:UPM

[1042] SuJeong Ko and JungHyun Lee. User preference mining through collaborative filtering and content based filtering in recommender system. Lecture Notes in Computer Science, 2455:244– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550244.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550244.
pdf.

Kim:2002:IRA

[1043] Taek-Hun Kim, Young-Suk Ryu, Seok-In Park, and Sung-Bong Yang. An improved recommendation algorithm in collaborative filtering. Lecture Notes in Computer Science, 2455:254– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550254.
pdf.

Modani:2002:SDT

[1044] Natwar Modani, Parul A. Mittal, Amit A. Nanavati, and Biplav Srivastava. Series of dynamic targeted recommendations. Lecture Notes in Computer Science, 2455:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550262.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550262.
                                        pdf.

Ko:2002:SCC

[1045] Seung-Kyu Ko, Won-Sung Sohn, Kyong-Ho Lee, Soon-Bum Lim, and Yoon-Chul Choy. A study of content conversion between eBook standards. Lecture Notes in Computer Science, 2455:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2455/24550273.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550273.
                                        pdf.

Nikolaidou:2002:WBS

[1046] Mara Nikolaidou and Dimosthenis Anagnostopoulos. Web-based system configuration and performance evaluation using a knowledgebased methodology. Lecture Notes in Computer Science, 2455:284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550284.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550284.
                                        pdf.

Sonneck:2002:MUW

[1047] Georg Sonneck, Renate Motschnig, and Thomas Mueck. Mapping UML Web navigation stereotypes to XML data skeletons. Lecture Notes in Computer Science, 2455:294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550294.
pdf.

Bozsak:2002:KTL

[1048] Erol Bozsak, Marc Ehrig, Siegfried Handschuh, Andreas Hotho, Alexander Maedche, Boris Motik, Daniel Oberle, Christoph Schmitz, Steffen Staab, Ljiljana Stojanovic, Nenad Stojanovic, Rudi Studer, Gerd Stumme, York Sure, Julien Tane, Raphael Volz, and Valentin Zacharias. KAON — towards a large scale Semantic Web. Lecture Notes in Computer Science, 2455: 304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550304.
pdf.

Enzmann:2002:PPT

[1049] Matthias Enzmann, Thomas Kunz, and Markus Schneider. Privacy protection through unlinkability of customer activities in business processes using mobile agents. Lecture Notes in Computer Science, 2455:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550314.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550314.
                                        pdf.

Lee:2002:SEC

[1050] Hyung-Woo Lee, Sung-Min Lee, and Im-Yeong Lee. Secure electronic copyright distribution with public key based traitor tracing. Lecture Notes in Computer Science, 2455:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550324.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550324.
                                        pdf.

Wang:2002:WEM

[1051] Shiuh-Jeng Wang and Kai-Sheng Yang. Watermark embedding mechanism using modulus-based for intellectual property protection on image data. Lecture Notes in Computer Science, 2455:333–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2455/24550333.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550333.
                                        pdf.

Ferrer-Gomila:2002:JVC

[1052] Josep Lluís Ferrer-Gomila, Apol·lònia Martínez-Nadal, Magdalena Payeras-Capellà, and Llorenç Huguet-Rotger. A juridical validation of a contract signing protocol. Lecture Notes in Computer Science, 2455:343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550343.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550343.
pdf.

Schmid:2002:DBP

[1053] Hans Albrecht Schmid and Gustavo Rossi. Designing business processes in E-commerce applications. Lecture Notes in Computer Science, 2455:353– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550353.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550353.
pdf.

Ward:2002:GSS

[1054] Christopher Ward, Melissa J. Buco, Rong N. Chang, and Laura Z. Luan. A generic SLA semantic model for the execution management of E-business outsourcing contracts. Lecture Notes in Computer Science, 2455:363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550363.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550363.
pdf.

Ouyang:2002:FES

[1055] Chun Ouyang, Lars Michael Kristensen, and Jonathan Billington. A formal and executable specification of the Internet Open Trading Protocol. Lecture Notes in Computer Science, 2455:

377–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550377.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550377.
                                        pdf.

Khu-Smith:2002:UEC

[1056] Vorapranee Khu-Smith and Chris J. Mitchell. Using EMV cards to protect E-commerce transactions. Lecture Notes in Computer Science, 2455:388– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550388.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550388.
                                        pdf.

Kreyer:2002:SPP

[1057] Nina Kreyer, Key Pousttchi, and Klaus Turowski. Standardized payment procedures as key enabling factor for mobile commerce. Lecture Notes in Computer Science, 2455:400– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2455/24550400.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2455/24550400.
                                        pdf.

Tjoa:2002:STR

[1058] A. Min Tjoa and Gerald Quirchmayr. Some thoughts on research perspectives in E-business. Lecture Notes in Computer Science, 2455:410– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2455/24550410.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2455/24550410.
pdf.

Anonymous:2002:AIu

[1059] Anonymous. Author index. Lecture Notes in Computer Science, 2455:413– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2455/2455auth.pdf.

Gelenbe:2002:GNM

[1060] Erol Gelenbe. G-networks: Multiple classes of positive customers, signals, and product form results. Lecture Notes in Computer Science, 2459: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590001.
pdf.

Mitrani:2002:SES

[1061] Isi Mitrani. Spectral expansion solutions for Markov-modulated queues. Lecture Notes in Computer Science, 2459:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2459/24590017.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2459/24590017.
                                        pdf.

Riska:2002:MGT

[1062] Alma Riska and Evgenia Smirni. M/G/1-type Markov processes: A tutorial. Lecture Notes in Computer Science, 2459:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2459/24590036.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2459/24590036.pdf.

Fourneau:2002:AAS

[1063] J. M. Fourneau and N. Pekergin. An algorithmic approach to stochastic bounds. Lecture Notes in Computer Science, 2459:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2459/24590064.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2459/24590064.
                                        pdf.

Yao:2002:DSP

[1064] David D. Yao. Dynamic scheduling via polymatroid optimization. Lecture Notes in Computer Science, 2459:89– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2459/24590089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2459/24590089.
                                        pdf.

Feitelson:2002:WMP

[1065] Dror G. Feitelson. Workload modeling for performance evaluation. Lecture Notes in Computer Science, 2459: 114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590114.
pdf.

Almeida:2002:CPW

[1066] Virgilio A. F. Almeida. Capacity planning for Web services. Lecture Notes in Computer Science, 2459:142– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590142.
pdf.

Cremonesi:2002:EEP

[1067] Paolo Cremonesi and Giuseppe Serazzi. End-to-end performance of Web services. Lecture Notes in Computer Science, 2459:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2459/24590158.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2459/24590158.pdf.

Weicker:2002:B

[1068] Reinhold Weicker. Benchmarking. Lecture Notes in Computer Science, 2459:

179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590179.
pdf.

Andreolini:2002:BMT

[1069] Mauro Andreolini, Valeria Cardellini, and Michele Colajanni. Benchmarking models and tools for distributed Web-server systems. Lecture Notes in Computer Science, 2459:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590208.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590208.
pdf.

Bernardo:2002:SPA

[1070] Marco Bernardo, Lorenzo Donatiello, and Paolo Ciancarini. Stochastic process algebra: From an algebraic formalism to an architectural description language. Lecture Notes in Computer Science, 2459:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590236.
pdf.

Baier:2002:APD

[1071] Christel Baier, Boudewijn Haverkort, Holger Hermanns, and Joost-Pieter Katoen. Automated performance and dependability evaluation using model checking. Lecture Notes in Computer Science, 2459:261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590261.
pdf.

Iyer:2002:MBA

[1072] Ravishankar K. Iyer and Zbigniew Kalbarczyk. Measurement-based analysis of system dependability using fault injection and field failure data. Lecture Notes in Computer Science, 2459: 290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590290.
pdf.

Trivedi:2002:SRR

[1073] Kishor S. Trivedi and Kalyanaraman Vaidyanathan. Software reliability and rejuvenation: Modeling and analysis. Lecture Notes in Computer Science, 2459:318–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590318.
pdf.

Grassi:2002:PVM

[1074] Vincenzo Grassi, Vittorio Cortellessa, and Raffaela Mirandola. Performance validation of mobile software architectures. Lecture Notes in Computer Science, 2459:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2459/24590346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2459/24590346.
                                        pdf.

deSouzaeSilva:2002:PIM

[1075] Edmundo de Souza e Silva, Rosa M. M. Leão, Berthier Ribeiro-Neto, and Sérgio Campos. Performance issues of multimedia applications. Lecture Notes in Computer Science, 2459: 374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2459/24590374.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2459/24590374.
                                        pdf.

Horvath:2002:MMR

[1076] András Horváth and Miklós Telek. Markovian modeling of real data traffic: Heuristic phase type and MAP fitting of heavy tailed and fractal like samples. Lecture Notes in Computer Science, 2459:405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2459/24590405.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2459/24590405.
pdf.

Conti:2002:OBE

[1077] Marco Conti and Enrico Gregori.

Optimization of bandwidth and energy consumption in wireless local area networks. Lecture Notes in Computer Science, 2459:435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590435.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590435.
pdf.

Rolia:2002:SCC

[1078] Jerry Rolia, Rich Friedrich, and Chandrakant Patel. Service centric computing — next generation Internet computing. Lecture Notes in Computer Science, 2459:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2459/24590463.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2459/24590463.
pdf.

Gagliardi:2002:EDP

[1079] Fabrizio Gagliardi, Bob Jones, Mario Reale, and Stephen Burke. European DataGrid project: Experiences of deploying a large scale testbed for E-science applications. Lecture Notes in Computer Science, 2459:480– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2459/24590480.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2459/24590480.
                                        pdf.

Anonymous:2002:AIv

[1080] Anonymous. Author index. Lecture Notes in Computer Science, 2459:501– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2459/2459auth.pdf.

Cook:2002:STS

[1081] William Cook. Solving traveling salesman problems. Lecture Notes in Computer Science, 2461:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610001.
                                        pdf.

Dey:2002:CSP

[1082] Tamal K. Dey. Computing shapes from point cloud data. Lecture Notes in Computer Science, 2461:2– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610002.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610002.
                                        pdf.

Karlin:2002:MDF

[1083] Anna R. Karlin. Mechanism design for fun and profit. Lecture Notes in Computer Science, 2461:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610003.
pdf.

Thorup:2002:DOR

[1084] Mikkel Thorup. On distance oracles and routing in graphs. Lecture Notes in Computer Science, 2461:4– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610004.
pdf.

Agarwal:2002:KMK

[1085] Pankaj K. Agarwal, Jie Gao, and Leonidas J. Guibas. Kinetic medians and kd-trees. Lecture Notes in Computer Science, 2461:5–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610005.
pdf.

Agarwal:2002:RSC

[1086] Pankaj K. Agarwal, Sathish Govindarajan, and S. Muthukrishnan. Range searching in categorical data: Colored range searching on grid. Lecture Notes in Computer Science, 2461:17– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610017.
pdf.

Agarwal:2002:NLT

[1087] Pankaj K. Agarwal, Sariel Har-Peled, Nabil H. Mustafa, and Yusu Wang. Near-linear time approximation algorithms for curve simplification. Lecture Notes in Computer Science, 2461:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610029.
pdf.

Agarwal:2002:TPO

[1088] Pankaj K. Agarwal, Torben Hagerup, Rahul Ray, Micha Sharir, Michiel Smid, and Emo Welzl. Translating a planar object to maximize point containment. Lecture Notes in Computer Science, 2461:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610042.
pdf.

Agarwal:2002:AAL

[1089] Pankaj K. Agarwal, Cecilia M. Procopiuc, and Kasturi R. Varadarajan. Approximation algorithms for k-line center. Lecture Notes in Computer Science, 2461:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610054.
pdf.

Ahr:2002:NHL

[1090] Dino Ahr and Gerhard Reinelt. New heuristics and lower bounds for the min-max k-Chinese postman problem. Lecture Notes in Computer Science, 2461:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2461/24610064.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2461/24610064.
pdf.

Althaus:2002:SSC

[1091] Ernst Althaus, Alexander Bockmayr, Matthias Elf, Michael Jünger, Thomas Kasper, and Kurt Mehlhorn. SCIL — symbolic constraints in integer linear programming. Lecture Notes in Computer Science, 2461:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610075.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2461/24610075.
                                        pdf.

Arge:2002:IED

[1092] Lars Arge, Octavian Procopiuc, and Jeffrey Scott Vitter. Implementing I/O-efficient data structures using TPIE. Lecture Notes in Computer Science, 2461:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610088.
                                        pdf.

Baier:2002:SFP

[1093] Georg Baier, Ekkehard Köhler, and Martin Skutella. On the ksplittable flow problem. Lecture Notes in Computer Science, 2461:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610101.
                                        pdf.

Barkan:2002:PAT

[1094] Arye Barkan and Haim Kaplan. Partial alphabetic trees. Lecture Notes in Computer Science, 2461:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610114.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2461/24610114.
pdf.

Barrett:2002:CCS

[1095] Chris Barrett, Keith Bisset, Riko Jacob, Goran Konjevod, and Madhav Marathe. Classical and contemporary shortest path problems in road networks: Implementation and experimental analysis of the TRANSIMS router. Lecture Notes in Computer Science, 2461:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610126.
pdf.

Bender:2002:STM

[1096] Michael A. Bender, Richard Cole, Erik D. Demaine, and Martin Farach-Colton. Scanning and traversing: Maintaining data for traversals in a memory hierarchy. Lecture Notes in Computer Science, 2461:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610139.
pdf.

Bender:2002:TSA

[1097] Michael A. Bender, Richard Cole, Erik D. Demaine, Martin Farach-Colton, and Jack Zito. Two simplified algorithms for maintaining order in a list. Lecture Notes in Computer Science, 2461:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610152.
pdf.

Bender:2002:ETL

[1098] Michael A. Bender, Erik D. Demaine, and Martin Farach-Colton. Efficient tree layout in a multilevel memory hierarchy. Lecture Notes in Computer Science, 2461:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610165.
pdf.

Berberich:2002:CBC

[1099] Eric Berberich, Arno Eigenwillig, Michael Hemmer, Susan Hert, Kurt Mehlhorn, and Elmar Schömer. A computational basis for conic arcs and Boolean operations on conic polygons. Lecture Notes in Computer Science, 2461:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610174.
pdf.

deBerg:2002:TNV

[1100] Mark de Berg, Joachim Gudmundsson, Matthew J. Katz, Christos Levcopoulos, Mark H. Overmars, and A. Frank van der Stappen. TSP with neighborhoods of varying size. Lecture Notes in Computer Science, 2461:187– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610187.
pdf.

Berman:2002:AAS

[1101] Piotr Berman, Sridhar Hannenhalli, and Marek Karpinski. 1.375approximation algorithm for sorting by reversals. Lecture Notes in Computer Science, 2461:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610200.
pdf.

Bodlaender:2002:RLP

[1102] Hans L. Bodlaender, Hajo Broersma, Fedor V. Fomin, Artem V. Pyatkin, and Gerhard J. Woeginger. Radio labeling with preassigned frequencies. Lecture Notes in Computer Science, 2461:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610211.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610211.
                                        pdf.

Bontridder:2002:BBA

[1103] Koen M. J. De Bontridder, B. J. Lageweg, Jan K. Lenstra, James B. Orlin, and Leen Stougie. Branchand-bound algorithms for the test cover problem. Lecture Notes in Computer Science, 2461:223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610223.
                                        pdf.

Bose:2002:CPS

[1104] Prosenjit Bose, Joachim Gudmundsson, and Michiel Smid. Constructing plane spanners of bounded degree and low weight. Lecture Notes in Computer Science, 2461:234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610234.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610234.
                                        pdf.

Brandes:2002:ESO

[1105] Ulrik Brandes. Eager st-ordering. Lecture Notes in Computer Science, 2461: 247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610247.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2461/24610247.
pdf.

Buchsbaum:2002:TDL

[1106] Adam L. Buchsbaum and Michael T.

Goodrich. Three-dimensional layers of maxima. Lecture Notes in Computer Science, 2461:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610257.
pdf.

Chen:2002:OTC

[1107] Danny Z. Chen, Xiaobo S. Hu, Shuang Luan, Xiaodong Wu, and Cedric X. Yu. Optimal terrain construction problems and applications in intensitymodulated radiation therapy. Lecture Notes in Computer Science, 2461:270– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610270.
pdf.

Chen:2002:GAD

[1108] Danny Z. Chen, Michiel Smid, and Bin Xu. Geometric algorithms for density-based data clustering. Lecture Notes in Computer Science, 2461:284– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2461/24610284.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610284.
                                        pdf.

Cohen:2002:BRA

[1109] Edith Cohen and Haim Kaplan. Balanced-replication algorithms for distribution trees. Lecture Notes in Computer Science, 2461:297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610297.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610297.
                                        pdf.

Datar:2002:BPP

[1110] Mayur Datar. Butterflies and peerto-peer networks. Lecture Notes in Computer Science, 2461:310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610310.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610310.
                                        pdf.

Datar:2002:ERS

[1111] Mayur Datar and S. Muthukrishnan. Estimating rarity and similarity over data stream windows. Lecture Notes in Computer Science, 2461:323– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610323.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2461/24610323.
pdf.

Bonis:2002:ECG

[1112] Annalisa De Bonis and Ugo Vaccaro.

Efficient constructions of generalized superimposed codes with applications to group testing and conflict resolution in multiple access channels. Lecture Notes in Computer Science, 2461: 335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610335.
pdf.

Demaine:2002:FEI

[1113] Erik D. Demaine, Alejandro López-Ortiz, and J. Ian Munro. Frequency estimation of Internet packet streams with limited space. Lecture Notes in Computer Science, 2461:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610348.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610348.
pdf.

Deshmukh:2002:TCD

[1114] Kaustubh Deshmukh, Andrew V.

Goldberg, Jason D. Hartline, and Anna R. Karlin. Truthful and competitive double auctions. Lecture Notes in Computer Science, 2461:361– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610361.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610361.
                                        pdf.

Dessmark:2002:OGE

[1115] Anders Dessmark and Andrzej Pelc. Optimal graph exploration without good maps. Lecture Notes in Computer Science, 2461:374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610374.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610374.
                                        pdf.

Dey:2002:AMA

[1116] Tamal K. Dey and Wulue Zhao. Approximating the medial axis from the Voronoi diagram with a convergence guarantee. Lecture Notes in Computer Science, 2461:387–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610387.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610387.
                                        pdf.

Doerr:2002:NIR

[1117] Benjamin Doerr and Henning Schnieder. Non-independent randomized rounding and an application to digital halftoning. Lecture Notes in Computer Science, 2461:399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610399.
pdf.

Efrat:2002:CHS

[1118] Alon Efrat, Stephen G. Kobourov, and Anna Lubiw. Computing homotopic shortest paths efficiently. Lecture Notes in Computer Science, 2461:411– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610411.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610411.
pdf.

Elbassioni:2002:ADP

[1119] Khaled M. Elbassioni. An algorithm for dualization in products of lattices and its applications. Lecture Notes in Computer Science, 2461:424– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610424.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610424.
pdf.

Enosh:2002:DSC

[1120] Angela Enosh, Klara Kedem, and Joel Bernstein. Determining similarity of conformational polymorphs. Lecture Notes in Computer Science, 2461:436– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610436.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610436.
                                        pdf.

Epstein:2002:MMS

[1121] Leah Epstein and Rob van Stee. Minimizing the maximum starting time on-line. Lecture Notes in Computer Science, 2461:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610449.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610449.
                                        pdf.

Epstein:2002:VAP

[1122] Leah Epstein and Tamir Tassa. Vector assignment problems: A general framework. Lecture Notes in Computer Science, 2461:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610461.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610461.
                                        pdf.

Ezra:2002:SIC

[1123] Eti Ezra, Dan Halperin, and Micha Sharir. Speeding up the incremental construction of the union of geometric objects in practice. Lecture Notes in Computer Science, 2461:473– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2461/24610473.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610473.
pdf.

Fahle:2002:SFI

[1124] Torsten Fahle. Simple and fast:

Improving a branch-and-bound algorithm for maximum clique. Lecture Notes in Computer Science, 2461:485– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610485.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610485.
pdf.

Fiat:2002:OCC

[1125] Amos Fiat, Manor Mendel, and Steven S. Seiden. Online companion caching. Lecture Notes in Computer Science, 2461:499–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610499.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610499.
pdf.

Gasieniec:2002:DCR

[1126] Leszek Gasieniec, Aris Pagourtzis, and Igor Potapov. Deterministic communication in radio networks with large labels. Lecture Notes in Computer Science, 2461:512–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610512.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610512.
                                        pdf.

Gentile:2002:PAS

[1127] Claudio Gentile, Utz-Uwe Haus, Matthias Köppe, Giovanni Rinaldi, and Robert Weismantel. A primal approach to the stable set problem. Lecture Notes in Computer Science, 2461: 525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610525.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610525.
                                        pdf.

Haxell:2002:WSN

[1128] Penny Haxell, April Rasala, Gordon Wilfong, and Peter Winkler. Wide-sense nonblocking WDM crossconnects. Lecture Notes in Computer Science, 2461:538–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2461/24610538.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2461/24610538.pdf.

Heggernes:2002:EIM

[1129] Pinar Heggernes and Yngve Villanger. Efficient implementation of a minimal triangulation algorithm. Lecture Notes in Computer Science, 2461:550– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2461/24610550.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610550.
pdf.

Jansen:2002:SMP

[1130] Klaus Jansen. Scheduling malleable parallel tasks: An asymptotic fully polynomial-time approximation scheme. Lecture Notes in Computer Science, 2461:562–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610562.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610562.
pdf.

Kaporis:2002:PAG

[1131] Alexis C. Kaporis, Lefteris M. Kirousis, and Efthimios G. Lalas. The probabilistic analysis of a greedy satisfiability algorithm. Lecture Notes in Computer Science, 2461:574–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610574.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610574.
pdf.

Karavelas:2002:DAW

[1132] Menelaos I. Karavelas and Mariette Yvinec. Dynamic additively weighted Voronoi diagrams in 2D. Lecture Notes in Computer Science, 2461:586– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610586.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610586.
                                        pdf.

Kohler:2002:TEG

[1133] Ekkehard Köhler, Katharina Langkau, and Martin Skutella. Time-expanded graphs for flow-dependent transit times. Lecture Notes in Computer Science, 2461:599–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2461/24610599.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2461/24610599.pdf.

Kolliopoulos:2002:POK

[1134] Stavros G. Kolliopoulos and George Steiner. Partially-ordered knapsack and applications to scheduling. Lecture Notes in Computer Science, 2461: 612–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610612.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610612.
                                        pdf.

Konstantinou:2002:SLE

[1135] Elisavet Konstantinou, Yiannis Stamatiou, and Christos Zaroliagis. A software library for elliptic curve cryptography. Lecture Notes in Computer Science, 2461:625–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2461/24610625.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2461/24610625.pdf.

Krumke:2002:RTD

[1136] Sven O. Krumke, Jörg Rambau, and Luis M. Torres. Real-time dispatching of guided and unguided automobile service units with soft time windows. Lecture Notes in Computer Science, 2461: 637–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610637.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610637.
pdf.

Laber:2002:RAA

[1137] Eduardo Laber, Ojas Parekh, and R. Ravi. Randomized approximation algorithms for query optimization problems on two processors. Lecture Notes in Computer Science, 2461:649– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610649.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610649.
pdf.

Langerman:2002:CTT

[1138] Stefan Langerman and Pat Morin.

Covering things with things. Lecture Notes in Computer Science, 2461:662– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610662.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610662.
                                        pdf.

Lipmann:2002:LDR

[1139] Maarten Lipmann, X. Lu, Willem E. de Paepe, Rene A. Sitters, and Leen Stougie. On-line dial-a-ride problems under a restricted information model. Lecture Notes in Computer Science, 2461:674–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610674.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610674.
                                        pdf.

Lorys:2002:AAM

[1140] Krzysztof Lorys and Grazyna Zwozniak. Approximation algorithm for the maximum leaf spanning tree problem for cubic graphs. Lecture Notes in Computer Science, 2461:686–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610686.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610686.
                                        pdf.

Manzini:2002:ELS

[1141] Giovanni Manzini and Paolo Ferragina. Engineering a lightweight suffix array construction algorithm (extended abstract). Lecture Notes in Computer Science, 2461:698–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610698.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610698.
pdf.

Maxova:2002:CCD

[1142] Jana Maxová and Jaroslav Nešetril.

Complexity of compatible decompositions of Eulerian graphs and their transformations. Lecture Notes in Computer Science, 2461:711–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610711.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610711.
pdf.

Mehlhorn:2002:EMB

[1143] Kurt Mehlhorn and Ulrich Meyer.

External-memory breadth-first search with sublinear I/O. Lecture Notes in Computer Science, 2461:723–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610723.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610723.
pdf.

Molloy:2002:FCA

[1144] Michael Molloy and Mohammad R.

Salavatipour. Frequency channel assignment on planar networks. Lecture Notes in Computer Science, 2461:736– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610736.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610736.
                                        pdf.

Naher:2002:DIE

[1145] Stefan Näher and Oliver Zlotowski. Design and implementation of efficient data types for static graphs. Lecture Notes in Computer Science, 2461:748– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610748.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610748.
                                        pdf.

Nielsen:2002:EAU

[1146] Benny K. Nielsen, Pawel Winter, and Martin Zachariasen. An exact algorithm for the uniformly-oriented Steiner tree problem. Lecture Notes in Computer Science, 2461:760–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610760.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610760.
                                        pdf.

Ohta:2002:FAS

[1147] Kenichiro Ohta, Kunihiko Sadakane, Akiyoshi Shioura, and Takeshi Tokuyama. A fast, accurate and simple method for pricing european-asian and savingasian options. Lecture Notes in Computer Science, 2461:772–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2461/24610772.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2461/24610772.pdf.

Peczarski:2002:SER

[1148] Marcin Peczarski. Sorting 13 elements requires 34 comparisons. Lecture Notes in Computer Science, 2461: 785–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610785.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610785.
pdf.

Polzin:2002:ERT

[1149] Tobias Polzin and Siavash Vahdati Daneshmand. Extending reduction techniques for the steiner tree problem. Lecture Notes in Computer Science, 2461:795–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2461/24610795.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2461/24610795.
pdf.

Pruhs:2002:CMP

[1150] Kirk Pruhs and Patchrawat Uthaisombut. A comparison of multicast pull models. Lecture Notes in Computer Science, 2461:808–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610808.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610808.
                                        pdf.

Racke:2002:OSS

[1151] Harald Räcke, Christian Sohler, and Matthias Westermann. Online scheduling for sorting buffers. Lecture Notes in Computer Science, 2461:820–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610820.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610820.
                                        pdf.

Schurr:2002:FST

[1152] Ingo Schurr and Tibor Szabó. Finding the sink takes some time. Lecture Notes in Computer Science, 2461:833– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610833.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610833.
                                        pdf.

Sellmann:2002:LCC

[1153] Meinolf Sellmann, Georg Kliewer, and Achim Koberstein. Lagrangian cardinality cuts and variable fixing for capacitated network design. Lecture Notes in Computer Science, 2461:845– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610845.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610845.
pdf.

Shachnai:2002:MMP

[1154] Hadas Shachnai, Tami Tamir, and Gerhard J. Woeginger. Minimizing makespan and preemption costs on a system of uniform machines. Lecture Notes in Computer Science, 2461:859– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610859.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610859.
pdf.

vanStee:2002:MTC

[1155] Rob van Stee and Han La Poutré.

Minimizing the total completion time on-line on a single machine, using restarts. Lecture Notes in Computer Science, 2461:872–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2461/24610872.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2461/24610872.
pdf.

Wein:2002:HLF

[1156] Ron Wein. High-level filtering for arrangements of conic arcs (extended abstract). Lecture Notes in Computer Science, 2461:884–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610884.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610884.
                                        pdf.

Woeginger:2002:ASC

[1157] Gerhard J. Woeginger. An approximation scheme for cake division with a linear number of cuts. Lecture Notes in Computer Science, 2461:896– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610896.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610896.
                                        pdf.

Zhang:2002:SLT

[1158] Huaming Zhang and Xin He. A simple linear time algorithm for finding even triangulations of 2-connected bipartite plane graphs. Lecture Notes in Computer Science, 2461:902–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2461/24610902.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2461/24610902.
                                        pdf.

Anonymous:2002:AIw

[1159] Anonymous. Author index. Lecture Notes in Computer Science, 2461:915– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
papers/2461/2461ind.pdf.

Anonymous:2002:AIx

[1160] Anonymous. Author index. Lecture Notes in Computer Science, 2463:000– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2463/2463auth.pdf.

Socha:2002:ASU

[1161] Krzysztof Socha, Joshua Knowles, and Michael Sampels. A MAX − MIN ant system for the university course timetabling problem. Lecture Notes in Computer Science, 2463:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630001.
pdf.

Blum:2002:AAG

[1162] Christian Blum. ACO applied to group shop scheduling: A case study on intensification and diversification. Lecture Notes in Computer Science, 2463: 14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630014.
pdf.

Nouyan:2002:ABA

[1163] Shervin Nouyan. Agent-based approach to dynamic task allocation. Lecture Notes in Computer Science, 2463: 28–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630028.
pdf.

Shmygelska:2002:ACO

[1164] Alena Shmygelska, Rosalía Aguirre-Hernández, and Holger H. Hoos. An ant colony optimization algorithm for the 2D HP protein folding problem. Lecture Notes in Computer Science, 2463:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630040.
pdf.

Ellabib:2002:ESS

[1165] Ismail Ellabib, Otman A. Basir, and Paul Calamai. An experimental study of a simple ant colony system for the vehicle routing problem with time windows. Lecture Notes in Computer Science, 2463:53– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630053.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2463/24630053.
pdf.

Bautista:2002:AAA

[1166] Joaquín Bautista and Jordi Pereira.

Ant algorithms for assembly line balancing. Lecture Notes in Computer Science, 2463:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2463/24630065.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2463/24630065.pdf.

Silva:2002:ACL

[1167] Carlos A. Silva, Thomas A. Runkler, João M. Sousa, and Rainer Palm. Ant colonies as logistic processes optimizers. Lecture Notes in Computer Science, 2463:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630076.
pdf.

Eyckelhof:2002:ASD

[1168] Casper Joost Eyckelhof and Marko Snoek. Ant systems for a dynamic TSP. Lecture Notes in Computer Science, 2463:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2463/24630088.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2463/24630088.
pdf.

Montgomery:2002:APT

[1169] James Montgomery and Marcus Randall. Anti-pheromone as a tool for better exploration of search space. Lecture Notes in Computer Science, 2463: 100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630100.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630100.
                                        pdf.

Guntsch:2002:APB

[1170] Michael Guntsch and Martin Middendorf. Applying population based ACO to dynamic optimization problems. Lecture Notes in Computer Science, 2463:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2463/24630111.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630111.
                                        pdf.

Wittner:2002:CEG

[1171] Otto Wittner and Bjarne E. Helvik. Cross-entropy guided ant-like agents finding cyclic paths in scarcely meshed networks. Lecture Notes in Computer Science, 2463:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630123.
                                        pdf.

Reimann:2002:IBA

[1172] Marc Reimann, Karl Doerner, and Richard F. Hartl. Insertion based ants for vehicle routing problems with backhauls and time windows. Lecture Notes in Computer Science, 2463:135– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630135.
pdf.

Merkle:2002:MAC

[1173] Daniel Merkle and Martin Middendorf. Modelling ACO: Composed permutation problems. Lecture Notes in Computer Science, 2463:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630149.
pdf.

Buhl:2002:SON

[1174] Jérôme Buhl, Jean-Louis Deneubourg, and Guy Theraulaz. Self-organized networks of galleries in the ant messor sancta. Lecture Notes in Computer Science, 2463:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630163.
pdf.

Bianchi:2002:SHP

[1175] Leonora Bianchi, Luca Maria Gambardella, and Marco Dorigo. Solving the homogeneous probabilistic traveling salesman problem by the ACO metaheuristic. Lecture Notes in Computer Science, 2463:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630176.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630176.
                                        pdf.

Birattari:2002:TFF

[1176] Mauro Birattari, Gianni Di Caro, and Marco Dorigo. Toward the formal foundation of ant programming. Lecture Notes in Computer Science, 2463:188– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630188.
                                        pdf.

Svennebring:2002:TBT

[1177] Jonas Svennebring and Sven Koenig. Towards building terrain-covering ant robots. Lecture Notes in Computer Science, 2463:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2463/24630202.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2463/24630202.pdf.

Dreo:2002:NAC

[1178] Johann Dréo and Patrick Siarry.

A new ant colony algorithm using the heterarchical concept aimed at optimization of multiminima continuous functions. Lecture Notes in Computer Science, 2463:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630216.
pdf.

Maniezzo:2002:ABF

[1179] Vittorio Maniezzo and Matteo Milandri. An ant-based framework for very strongly constrained problems. Lecture Notes in Computer Science, 2463:222– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630222.
pdf.

Cordon:2002:ABW

[1180] Oscar Cordón, Iñaki Fernández de Viana, and Francisco Herrera. Analysis of the best-worst ant system and its variants on the QAP. Lecture Notes in Computer Science, 2463:228– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630228.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2463/24630228.
                                        pdf.

Canright:2002:AL

[1181] Geoffrey Canright. Ants and loops. Lecture Notes in Computer Science, 2463:235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2463/24630235.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630235.
                                        pdf.

Randall:2002:CSS

[1182] Marcus Randall and James Montgomery. Candidate set strategies for ant colony optimisation. Lecture Notes in Computer Science, 2463:243– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630243.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630243.
                                        pdf.

Garlick:2002:DWR

[1183] Ryan M. Garlick and Richard S. Barr. Dynamic wavelength routing in WDM networks via ant colony optimization. Lecture Notes in Computer Science, 2463:250–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630250.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630250.
                                        pdf.

Hoe:2002:HAW

[1184] Kok Meng Hoe, Weng Kin Lai, and Tracy S. Y. Tai. Homogeneous ants for Web document similarity modeling and categorization. Lecture Notes in Computer Science, 2463:256– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630256.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630256.
pdf.

Rahoual:2002:PAS

[1185] Malek Rahoual, Riad Hadji, and Vincent Bachelet. Parallel ant system for the set covering problem. Lecture Notes in Computer Science, 2463:262– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630262.
pdf.

Vogel:2002:RWS

[1186] Andre Vogel, Marco Fischer, Hendrik Jaehn, and Tobias Teich. Realworld shop floor scheduling by ant colony optimization. Lecture Notes in Computer Science, 2463:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630268.
pdf.

Sahin:2002:SNA

[1187] Erol Sahin and Nigel R. Franks. Simulation of nest assessment behavior by ant scouts. Lecture Notes in Computer Science, 2463:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630274.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630274.
                                        pdf.

Pilat:2002:UGA

[1188] Marcin L. Pilat and Tony White. Using genetic algorithms to optimize ACS-TSP. Lecture Notes in Computer Science, 2463:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630282.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630282.
                                        pdf.

Ling:2002:MSO

[1189] Chen Ling, Sheng Jie, Qin Ling, and Chen Hongjian. A method for solving optimization problems in continuous space using ant colony algorithm. Lecture Notes in Computer Science, 2463: 288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630288.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630288.
                                        pdf.

Wolf:2002:NLT

[1190] Tom De Wolf, Liesbeth Jaco, Tom Holvoet, and Elke Steegmans. A nested layered threshold model for dynamic task allocation. Lecture Notes in Computer Science, 2463:290– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630290.
pdf.

Fidanova:2002:AAA

[1191] Stefka Fidanova. ACO algorithm with additional reinforcement. Lecture Notes in Computer Science, 2463:292– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630292.
pdf.

Ouadfel:2002:ACS

[1192] Salima Ouadfel, Mohamed Batouche, and Catherine Garbay. Ant colony system for image segmentation using Markov random field. Lecture Notes in Computer Science, 2463:294– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630294.
pdf.

Vallone:2002:BSP

[1193] Ugo Vallone. Bidimensional shapes polygonalization by ACO. Lecture Notes in Computer Science, 2463:296– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630296.
                                        pdf.

Branke:2002:CAA

[1194] Jürgen Branke, Michael Decker, Daniel Merkle, and Hartmut Schmeck. Coevolutionary ant algorithms playing games. Lecture Notes in Computer Science, 2463:298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2463/24630298.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2463/24630298.pdf.

Acan:2002:GGA

[1195] Adnan Acan. GAACO: A GA + ACO hybrid for faster and better search capability. Lecture Notes in Computer Science, 2463:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2463/24630300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2463/24630300.
                                        pdf.

Saleh:2002:GPN

[1196] Hussain Aziz Saleh. GPS positioning networks design: An application of the ant colony system. Lecture Notes in Computer Science, 2463:302– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2463/24630302.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2463/24630302.
pdf.

Alonso:2002:UES

[1197] Miguel A. Alonso, Jesús Vilares, and Víctor M. Darriba. On the usefulness of extracting syntactic dependencies for text indexing. Lecture Notes in Computer Science, 2464:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640003.
pdf.

Buckeridge:2002:ULS

[1198] Alan M. Buckeridge and Richard F. E. Sutcliffe. Using latent semantic indexing as a measure of conceptual association for noun compound disambiguation. Lecture Notes in Computer Science, 2464:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640012.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2464/24640012.
                                        pdf.

Crean:2002:RFA

[1199] Brian P. Crean and Diarmuid O’Donoghue. RADAR: Finding analogies using attributes of structure. Lecture Notes in Computer Science, 2464:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640020.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640020.
                                        pdf.

Cummins:2002:CLB

[1200] Fred Cummins. Classifying languages based on speech rhythm. Lecture Notes in Computer Science, 2464:28– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640028.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640028.
                                        pdf.

Doran:2002:FAT

[1201] Jim Doran. Finding agents in a twodimensional Boolean STaM. Lecture Notes in Computer Science, 2464:36– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640036.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640036.
                                        pdf.

Fdez-Riverola:2002:NSS

[1202] Florentino Fdez-Riverola, Juan M.

Corchado, and Jesús M. Torres. Neuro-symbolic system for forecasting red tides. Lecture Notes in Computer Science, 2464:45–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640045.
pdf.

Keller:2002:ILH

[1203] Bill Keller and Rudi Lutz. Improved learning for hidden Markov models using penalized training. Lecture Notes in Computer Science, 2464:53– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640053.
pdf.

Lutz:2002:RHL

[1204] Rudi Lutz. Recovering high-level structure of software systems using a minimum description length principle. Lecture Notes in Computer Science, 2464: 61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640061.
pdf.

Dermott:2002:SMA

[1205] Paula Mc Dermott and Colm O’Riordan. A system for multi-agent information retrieval. Lecture Notes in Computer Science, 2464:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640070.
                                        pdf.

Narayanan:2002:ATM

[1206] Ajit Narayanan. All there is to the mind is to have the right genes, or, consciousness as a form of genetic engineering. Lecture Notes in Computer Science, 2464:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640078.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640078.
                                        pdf.

OMahony:2002:TRC

[1207] Michael P. O’Mahony, Neil J. Hurley, and Guenole C. M. Silvestre. Towards robust collaborative filtering. Lecture Notes in Computer Science, 2464:87– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640087.
                                        pdf.

Pereira:2002:GNG

[1208] Francisco B. Pereira, Jorge Tavares, Penousal Machado, and Ernesto Costa. GVR: A new genetic representation for the vehicle routing problem. Lecture Notes in Computer Science, 2464:95– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640095.
pdf.

Silva:2002:ECP

[1209] Arlindo Silva, Ana Neves, and Ernesto Costa. An empirical comparison of particle swarm and predator prey optimisation. Lecture Notes in Computer Science, 2464:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640103.
pdf.

Smyth:2002:DMS

[1210] Barry Smyth, David Wilson, and Derry O’Sullivan. Data mining support for case-based collaborative recommendation. Lecture Notes in Computer Science, 2464:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640111.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2464/24640111.
                                        pdf.

Sutcliffe:2002:FML

[1211] Richard F. E. Sutcliffe and Kieran White. The feasibility of machine learning for query answering — an experiment in two domains. Lecture Notes in Computer Science, 2464:119– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640119.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640119.
                                        pdf.

Veale:2002:MKA

[1212] Tony Veale. Meta-knowledge annotation for efficient natural-language question-answering. Lecture Notes in Computer Science, 2464:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640127.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640127.
                                        pdf.

Brabazon:2002:FTS

[1213] Anthony Brabazon. Financial time series modelling using neural networks: An assessment of the utility of a stacking methodology. Lecture Notes in Computer Science, 2464:137– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640137.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2464/24640137.
pdf.

Bridge:2002:ESR

[1214] Derek Bridge and Jerome Kelleher. Experiments in sparsity reduction: Using clustering in collaborative recommenders. Lecture Notes in Computer Science, 2464:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640144.
pdf.

Corchado:2002:IVF

[1215] Emilio Corchado and Colin Fyfe.

Identification of visual features using a neural version of exploratory projection pursuit. Lecture Notes in Computer Science, 2464:150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640150.
pdf.

Costello:2002:HPC

[1216] Fintan Costello. How people compare an item’s placement in two alternative categories. Lecture Notes in Computer Science, 2464:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2464/24640158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640158.
                                        pdf.

Dempsey:2002:IMI

[1217] Ian Dempsey, Michael O’Neill, and Anthony Brabazon. Investigations into market index trading models using evolutionary automatic programming. Lecture Notes in Computer Science, 2464:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640165.
                                        pdf.

Fairclough:2002:ISE

[1218] Chris Fairclough and Pádraig Cunningham. An interactive story engine. Lecture Notes in Computer Science, 2464: 171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640171.
                                        pdf.

Glass:2002:CEB

[1219] David H. Glass. Coherence, explanation, and Bayesian networks. Lecture Notes in Computer Science, 2464:177– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640177.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2464/24640177.
pdf.

Gomes:2002:CCB

[1220] Paulo Gomes, Francisco C. Pereira, Nuno Seco, Paulo Paiva, Paulo Carreiro, José L. Ferreira, and Carlos Bento. Combining case-based reasoning and analogical reasoning in software design. Lecture Notes in Computer Science, 2464:183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640183.
pdf.

Jones:2002:CMI

[1221] Gareth J. F. Jones and Adenike M.

Lam-Adesina. Combination methods for improving the reliability of machine translation based cross-language information retrieval. Lecture Notes in Computer Science, 2464:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640190.
pdf.

Lahart:2002:SMI

[1222] Orla Lahart and Colm O’Riordan. A system for music information retrieval. Lecture Notes in Computer Science, 2464:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2464/24640197.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640197.
                                        pdf.

Madden:2002:NBN

[1223] Michael G. Madden. A new Bayesian network structure for classification tasks. Lecture Notes in Computer Science, 2464:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640203.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640203.
                                        pdf.

Ginty:2002:EPB

[1224] Lorraine Mc Ginty and Barry Smyth. Evaluating preference-based feedback in recommender systems. Lecture Notes in Computer Science, 2464:209– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640209.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640209.
                                        pdf.

McGrath:2002:DMI

[1225] Declan McGrath. Design of a musical instrument classifier system based on mel scaled cepstral coefficient supervectors and a supervised two-layer feedforward neural network. Lecture Notes in Computer Science, 2464:215– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640215.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640215.
pdf.

McSherry:2002:ILE

[1226] David McSherry. An interactive learning environment for knowledge engineering. Lecture Notes in Computer Science, 2464:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640221.
pdf.

Medori:2002:CCI

[1227] Julia Medori, Eric Atwell, Paul Gent, and Clive Souter. Customising a copying-identifier for biomedical science student reports: Comparing simple and smart analyses. Lecture Notes in Computer Science, 2464:228– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2464/24640228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2464/24640228.
pdf.

Panadero:2002:HGS

[1228] R. Panadero and J. L. Fernández-Villacañas. A hybridised GA for the steiner minimal tree problem. Lecture Notes in Computer Science, 2464:234– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640234.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640234.
                                        pdf.

Ross:2002:SAI

[1229] Robert J. Ross, Bryan McEleney, Robert Kelly, Tarek Abu-Amer, Michael Walsh, Julie Carson-Berndsen, and Gregory M. P. O’Hare. Speaking autonomous intelligent devices. Lecture Notes in Computer Science, 2464: 240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2464/24640240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2464/24640240.
                                        pdf.

Anonymous:2002:AIy

[1230] Anonymous. Author index. Lecture Notes in Computer Science, 2464:247– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2464/2464auth.pdf.

Milano:2002:RCB

[1231] Michela Milano and Willem J. van Hoeve. Reduced cost-based ranking for generating promising subproblems. Lecture Notes in Computer Science, 2470:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2470/24700001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2470/24700001.
pdf.

Appa:2002:ICI

[1232] Gautam Appa, Ioannis Mourtos, and Dimitris Magos. Integrating constraint and integer programming for the orthogonal Latin squares problem. Lecture Notes in Computer Science, 2470: 17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700017.
pdf.

Amaral:2002:OCI

[1233] Paula Amaral and Pedro Barahona.

On optimal correction of inconsistent linear constraints. Lecture Notes in Computer Science, 2470:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700033.
pdf.

Dimopoulos:2002:TPT

[1234] Yannis Dimopoulos and Alfonso Gerevini. Temporal planning through mixed integer programming: A preliminary report. Lecture Notes in Computer Science, 2470:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700047.
                                        pdf.

Beldiceanu:2002:NMR

[1235] Nicolas Beldiceanu and Mats Carlsson. A new multi-resource cumulatives constraint with negative heights. Lecture Notes in Computer Science, 2470:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700063.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700063.
                                        pdf.

Yunes:2002:SCR

[1236] Tallys H. Yunes. On the sum constraint:relaxation and applications. Lecture Notes in Computer Science, 2470:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2470/24700080.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700080.
                                        pdf.

Frisch:2002:GCL

[1237] Alan Frisch, Brahim Hnich, Zeynep Kiziltan, Ian Miguel, and Toby Walsh. Global constraints for lexicographic orderings. Lecture Notes in Computer Science, 2470:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2470/24700093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700093.
pdf.

Lebbah:2002:GFA

[1238] Yahia Lebbah, Michel Rueher, and Claude Michel. A global filtering algorithm for handling systems of quadratic equations and inequations. Lecture Notes in Computer Science, 2470:109– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700109.
pdf.

Cicirello:2002:ASP

[1239] Vincent A. Cicirello and Stephen F.

Smith. Amplification of search performance through randomization of heuristics. Lecture Notes in Computer Science, 2470:124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2470/24700124.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2470/24700124.pdf.

Muscettola:2002:CES

[1240] Nicola Muscettola. Computing the envelope for stepwise-constant resource allocations. Lecture Notes in Computer Science, 2470:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2470/24700139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700139.
                                        pdf.

Kamarainen:2002:LPA

[1241] Olli Kamarainen and Hani El Sakkout. Local probing applied to scheduling. Lecture Notes in Computer Science, 2470:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2470/24700155.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700155.
                                        pdf.

Habet:2002:HAS

[1242] Djamal Habet, Chu Min Li, Laure Devendeville, and Michel Vasquez. A hybrid approach for SAT. Lecture Notes in Computer Science, 2470:172– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700172.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700172.
                                        pdf.

Ostrowski:2002:RES

[1243] Richard Ostrowski, Éric Grégoire, Bertrand Mazure, and Lakhdar Saïs. Recovering and exploiting structural knowledge from CNF formulas. Lecture Notes in Computer Science, 2470: 185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700185.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2470/24700185.
pdf.

Zhang:2002:TST

[1244] Lintao Zhang and Sharad Malik. Towards a symmetric treatment of satisfaction and conflicts in quantified Boolean formula evaluation. Lecture Notes in Computer Science, 2470:200– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700200.
pdf.

Wei:2002:ARW

[1245] Wei Wei and Bart Selman. Accelerating random walks. Lecture Notes in Computer Science, 2470:216– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700216.
pdf.

Hutter:2002:SPS

[1246] Frank Hutter, Dave A. D. Tompkins, and Holger H. Hoos. Scaling and probabilistic smoothing: Efficient dynamic local search for SAT. Lecture Notes in Computer Science, 2470:233– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700233.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700233.
                                        pdf.

Rossi:2002:LSS

[1247] Francesca Rossi, Alessandro Sperduti, Kristen B. Venable, Lina Khatib, Paul Morris, and Robert Morris. Learning and solving soft temporal constraints: An experimental study. Lecture Notes in Computer Science, 2470: 249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700249.
                                        pdf.

Meseguer:2002:OSR

[1248] Pedro Meseguer, Martí Sánchez, and Gérard Verfaillie. Opportunistic specialization in Russian doll search. Lecture Notes in Computer Science, 2470: 264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700264.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700264.
                                        pdf.

Petit:2002:RBA

[1249] Thierry Petit, Jean-Charles Régin, and Christian Bessière. Range-based algorithm for max-CSP. Lecture Notes in Computer Science, 2470:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2470/24700280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700280.
pdf.

Mitchell:2002:RCR

[1250] David G. Mitchell. Resolution complexity of random constraints. Lecture Notes in Computer Science, 2470:295– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700295.
pdf.

Dalmau:2002:CSB

[1251] Víctor Dalmau, Phokion G. Kolaitis, and Moshe Y. Vardi. Constraint satisfaction, bounded treewidth, and finite-variable logics. Lecture Notes in Computer Science, 2470:310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700310.
pdf.

Angelsmark:2002:DNS

[1252] Ola Angelsmark, Peter Jonsson, Svante Linusson, and Johan Thapper. Determining the number of solutions to binary CSP instances. Lecture Notes in Computer Science, 2470:327– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2470/24700327.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700327.
                                        pdf.

Skiadopoulos:2002:CCQ

[1253] Spiros Skiadopoulos and Manolis Koubarakis. Consistency checking for qualitative spatial reasoning with cardinal directions. Lecture Notes in Computer Science, 2470:341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700341.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700341.
                                        pdf.

Faltings:2002:OCS

[1254] Boi Faltings and Santiago Macho-Gonzalez. Open constraint satisfaction. Lecture Notes in Computer Science, 2470:356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2470/24700356.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700356.
                                        pdf.

Bordeaux:2002:BNA

[1255] Lucas Bordeaux and Eric Monfroy. Beyond NP: Arc-consistency for quantified constraints. Lecture Notes in Computer Science, 2470:371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700371.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2470/24700371.
pdf.

Yokoo:2002:SDC

[1256] Makoto Yokoo, Koutarou Suzuki, and Katsutoshi Hirayama. Secure distributed constraint satisfaction: Reaching agreement without revealing private information. Lecture Notes in Computer Science, 2470:387–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700387.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700387.
pdf.

Smith:2002:DGT

[1257] Barbara M. Smith. A dual graph translation of a problem in ‘life’. Lecture Notes in Computer Science, 2470:402– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700402.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700402.
pdf.

Gent:2002:GCS

[1258] Ian P. Gent, Warwick Harvey, and Tom Kelsey. Groups and constraints: Symmetry breaking during search. Lecture Notes in Computer Science, 2470:415– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700415.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700415.
                                        pdf.

McDonald:2002:PSBa

[1259] Iain McDonald and Barbara Smith. Partial symmetry breaking. Lecture Notes in Computer Science, 2470:431– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700431.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700431.
                                        pdf.

Puget:2002:SBR

[1260] Jean-François Puget. Symmetry breaking revisited. Lecture Notes in Computer Science, 2470:446–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700446.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700446.
                                        pdf.

Flener:2002:BRC

[1261] Pierre Flener, Alan M. Frisch, Brahim Hnich, Zeynep Kiziltan, Ian Miguel, Justin Pearson, and Toby Walsh. Breaking row and column symmetries in matrix models. Lecture Notes in Computer Science, 2470:462–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700462.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2470/24700462.
pdf.

Barnier:2002:SKS

[1262] Nicolas Barnier and Pascal Brisset.

Solving the Kirkman’s schoolgirl problem in a few seconds. Lecture Notes in Computer Science, 2470:477–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700477.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700477.
pdf.

Lesaint:2002:ICT

[1263] David Lesaint. Inferring constraint types in constraint programming. Lecture Notes in Computer Science, 2470: 492–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700492.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700492.
pdf.

Williams:2002:MBP

[1264] Brian C. Williams and Michel D. Ingham. Model-based programming: Controlling embedded systems by reasoning about hidden state. Lecture Notes in Computer Science, 2470:508– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700508.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2470/24700508.
pdf.

Epstein:2002:ACE

[1265] Susan L. Epstein, Eugene C. Freuder, Richard Wallace, Anton Morozov, and Bruce Samuels. The adaptive constraint engine. Lecture Notes in Computer Science, 2470:525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700525.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700525.
pdf.

Dao:2002:IBS

[1266] Thi Bich Hanh Dao, Arnaud Lallouet, Andrei Legtchenko, and Lionel Martin. Indexical-based solver learning. Lecture Notes in Computer Science, 2470: 541–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700541.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700541.
pdf.

Leyton-Brown:2002:LEH

[1267] Kevin Leyton-Brown, Eugene Nudelman, and Yoav Shoham. Learning the empirical hardness of optimization problems. Lecture Notes in Computer Science, 2470:556–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700556.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2470/24700556.
pdf.

Ruan:2002:RPD

[1268] Yongshao Ruan, Eric Horvitz, and Henry Kautz. Restart policies with dependence among runs: A dynamic programming approach. Lecture Notes in Computer Science, 2470:573–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700573.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700573.
pdf.

Bartak:2002:VSE

[1269] Roman Barták. Visopt ShopFloor: On the edge of planning and scheduling. Lecture Notes in Computer Science, 2470:587–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2470/24700587.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2470/24700587.
pdf.

Benoist:2002:CPC

[1270] Thierry Benoist, Etienne Gaudin, and Benoit Rottembourg. Constraint programming contribution to Benders decomposition: A case study. Lecture Notes in Computer Science, 2470:603– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700603.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700603.
                                        pdf.

Christie:2002:MCC

[1271] Marc Christie, Éric Languénou, and Laurent Granvilliers. Modeling camera control with constrained hypertubes. Lecture Notes in Computer Science, 2470:618–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700618.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700618.
                                        pdf.

Pape:2002:RPS

[1272] Claude Le Pape, Laurent Perron, Jean-Charles Régin, and Paul Shaw. Robust and parallel solving of a network design problem. Lecture Notes in Computer Science, 2470:633–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700633.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700633.
                                        pdf.

Lauvergne:2002:CRR

[1273] Muriel Lauvergne, Philippe David, and Patrice Boizumault. Connections reservation with rerouting for ATM networks: A hybrid approach with constraints. Lecture Notes in Computer Science, 2470:649–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700649.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700649.
pdf.

Fernandez:2002:CCD

[1274] Cèsar Fernàndez, Ramón Béjar, Bhaskar Krishnamachari, and Carla Gomes. Communication and computation in distributed CSP algorithms. Lecture Notes in Computer Science, 2470:664–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2470/24700664.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2470/24700664.
pdf.

Ratschan:2002:CFO

[1275] Stefan Ratschan. Continuous firstorder constraint satisfactionwith equality and disequality constraints. Lecture Notes in Computer Science, 2470:680– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700680.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700680.
pdf.

Hooker:2002:RCC

[1276] John N. Hooker and Hong Yan. A relaxation of the cumulative constraint. Lecture Notes in Computer Science, 2470:686–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2470/24700686.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700686.
                                        pdf.

Stuckey:2002:IGU

[1277] Peter J. Stuckey and Lei Zheng. Improving GSAT using 2SAT. Lecture Notes in Computer Science, 2470:691– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700691.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700691.
                                        pdf.

Mauss:2002:RCS

[1278] Jakob Mauss, Frank Seelisch, and Mugur Tatar. A relational constraint solver for model-based engineering. Lecture Notes in Computer Science, 2470:696–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2470/24700696.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700696.
                                        pdf.

Elkhyari:2002:CBR

[1279] Abdallah Elkhyari, Christelle Guéret, and Narendra Jussien. Conflict-based repair techniques for solving dynamic scheduling problems. Lecture Notes in Computer Science, 2470:702–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2470/24700702.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700702.
pdf.

Parkes:2002:SPP

[1280] Andrew J. Parkes. Scaling properties of pure random walk on random 3-SAT. Lecture Notes in Computer Science, 2470:708–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700708.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700708.
pdf.

Roli:2002:CPS

[1281] Andrea Roli. Criticality and parallelism in structured SAT instances. Lecture Notes in Computer Science, 2470:714–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2470/24700714.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2470/24700714.
pdf.

Bennaceur:2002:CSP

[1282] Hachemi Bennaceur and Chu Min Li.

Characterizing SAT problems with the row convexity property. Lecture Notes in Computer Science, 2470:720– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700720.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2470/24700720.
                                        pdf.

Bistarelli:2002:ISC

[1283] Stefano Bistarelli, Boi Faltings, and Nicoleta Neagu. Interchangeability in soft CSPs. Lecture Notes in Computer Science, 2470:726–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700726.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700726.
                                        pdf.

Yorke-Smith:2002:CPIa

[1284] Neil Yorke-Smith and Carmen Gervet. On constraint problems with incompleteor erroneous data. Lecture Notes in Computer Science, 2470:732–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700732.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700732.
                                        pdf.

Sellmann:2002:HCP

[1285] Meinolf Sellmann and Warwick Harvey. Heuristic constraint propagation. Lecture Notes in Computer Science, 2470: 738–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700738.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700738.
                                        pdf.

Sellmann:2002:ACA

[1286] Meinolf Sellmann. An arc-consistency algorithm for the minimum weight all different constraint. Lecture Notes in Computer Science, 2470:744–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700744.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700744.
pdf.

Law:2002:APC

[1287] Y. C. Law and Jimmy H. M.

Lee. Algebraic properties of CSP model operators. Lecture Notes in Computer Science, 2470:750–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700750.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700750.
pdf.

vanDongen:2002:AEA

[1288] Marc R. C. van Dongen. AC-3_{d} an efficient arc-consistency algorithm with a low space-complexity. Lecture Notes in Computer Science, 2470:755– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700755.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700755.
pdf.

Ringwelski:2002:ISO

[1289] Georg Ringwelski. Integrating search objects in asynchronous constraint solving. Lecture Notes in Computer Science, 2470:761–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700761.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700761.
                                        pdf.

Schlenker:2002:DCB

[1290] Hans Schlenker. Distributed constraintbased railway simulation. Lecture Notes in Computer Science, 2470:762– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700762.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700762.
                                        pdf.

Petrie:2002:SBP

[1291] Karen E. Petrie. Symmetry breaking in peaceably coexisting armies of queens. Lecture Notes in Computer Science, 2470:763–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2470/24700763.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700763.
                                        pdf.

Vilim:2002:BPS

[1292] Petr Vilím. Batch processing with sequence dependent setup times. Lecture Notes in Computer Science, 2470:764– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700764.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700764.
pdf.

Muller:2002:IHS

[1293] Tomáš Müller. Interactive heuristic search algorithm. Lecture Notes in Computer Science, 2470:765–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700765.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700765.
pdf.

Yorke-Smith:2002:CPIb

[1294] Neil Yorke-Smith. On constraint problems with incomplete or erroneous data. Lecture Notes in Computer Science, 2470:766–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2470/24700766.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2470/24700766.pdf.

Roli:2002:DNM

[1295] Andrea Roli. Design of a new metaheuristic for MAXSAT problems. Lecture Notes in Computer Science, 2470: 767–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700767.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700767.
                                        pdf.

Salido:2002:DCC

[1296] Miguel A. Salido and Federico Barber. Disjunctive and continuous constraint satisfaction problems. Lecture Notes in Computer Science, 2470:768– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700768.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700768.
                                        pdf.

Lynce:2002:TRB

[1297] Inês Lynce and João Marques-Silva. Tuning randomization in backtrack search SAT algorithms. Lecture Notes in Computer Science, 2470:769–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700769.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700769.
                                        pdf.

Zhan:2002:CST

[1298] Yuan Zhan. Constraint solving in test-data generation. Lecture Notes in Computer Science, 2470:770–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700770.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2470/24700770.
pdf.

Bohlin:2002:ICC

[1299] Markus Bohlin. Improving cost calculations for global constraints in local search. Lecture Notes in Computer Science, 2470:772–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700772.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700772.
pdf.

Renker:2002:MFC

[1300] Gerrit Renker. A modeling framework for constraints. Lecture Notes in Computer Science, 2470:773–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700773.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700773.
pdf.

vanNorden:2002:LPB

[1301] Linda van Norden and Hans van Maaren. A linear programming based satisfiability solver using a new horndriven search tree design. Lecture Notes in Computer Science, 2470:775– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700775.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2470/24700775.
                                        pdf.

Quesada:2002:CCP

[1302] Luis Quesada and Peter Van Roy. A concurrent constraint programming approach for trajectory determination of autonomous vehicles. Lecture Notes in Computer Science, 2470:777– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700777.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700777.
                                        pdf.

Gronkvist:2002:UCP

[1303] Mattias Grönkvist. Using constraint propagation to accelerate column generation in aircraft scheduling. Lecture Notes in Computer Science, 2470:778– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700778.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700778.
                                        pdf.

Venable:2002:SLS

[1304] Kristen B. Venable. Solving and learning soft temporal constraints; ceteris paribus statements represented as soft constraints problems. Lecture Notes in Computer Science, 2470:779– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700779.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2470/24700779.
pdf.

Seelisch:2002:PSF

[1305] Frank Seelisch. A partially solved form for heterogeneous constraints in disjunctive normal form. Lecture Notes in Computer Science, 2470:780– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700780.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700780.
pdf.

Hnich:2002:MIP

[1306] Brahim Hnich and Toby Walsh. Models of injection problems. Lecture Notes in Computer Science, 2470:781– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700781.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700781.
pdf.

McDonald:2002:PSBb

[1307] Iain McDonald. Partial symmetry breaking. Lecture Notes in Computer Science, 2470:782–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700782.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700782.
pdf.

Drake:2002:AGI

[1308] Lyndon Drake, Alan Frisch, and Toby Walsh. Automatic generation of implied clauses for SAT. Lecture Notes in Computer Science, 2470:783–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700783.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700783.
                                        pdf.

Ansotegui:2002:BGB

[1309] Carlos Ansótegui and Felip Manyà. Bridging the gap between SAT and CSP. Lecture Notes in Computer Science, 2470:784–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2470/24700784.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2470/24700784.pdf.

Kiziltan:2002:RSM

[1310] Zeynep Kiziltan. Reducing symmetry in matrix models. Lecture Notes in Computer Science, 2470:786–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2470/24700786.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2470/24700786.
                                        pdf.

Neagu:2002:SIC

[1311] Nicoleta Neagu. Studying interchangeability in constraint satisfaction problems. Lecture Notes in Computer Science, 2470:787–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700787.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700787.
pdf.

Glaubius:2002:CMC

[1312] Robert Glaubius and Berthe Y.

Choueiry. Constraint modeling in the context of academic task assignment. Lecture Notes in Computer Science, 2470:789–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2470/24700789.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2470/24700789.
pdf.

Vetsikas:2002:DTA

[1313] Ioannis A. Vetsikas. Design tradeoffs for autonomous trading agents. Lecture Notes in Computer Science, 2470: 790–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2470/24700790.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2470/24700790.
pdf.

Anonymous:2002:AIz

[1314] Anonymous. Author index. Lecture Notes in Computer Science, 2470:793– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2470/2470ind.pdf.

Hayashi:2002:LCM

[1315] Susumu Hayashi and Yohji Akama. Limit-computable mathematics and its applications. Lecture Notes in Computer Science, 2471:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710001.
                                        pdf.

Neven:2002:ALX

[1316] Frank Neven. Automata, logic, and XML. Lecture Notes in Computer Science, 2471:2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2471/24710002.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2471/24710002.pdf.

Niwinski:2002:CGE

[1317] Damian Niwinski. μ-calculus via games (extended abstract). Lecture Notes in Computer Science, 2471:27– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710027.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710027.
                                        pdf.

Kanovich:2002:BBP

[1318] Max Kanovich. Bijections between partitions by two-directional rewriting techniques. Lecture Notes in Computer Science, 2471:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710044.
pdf.

Aehlig:2002:CN

[1319] Klaus Aehlig and Felix Joachimski.

On continuous normalization. Lecture Notes in Computer Science, 2471:59– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710059.
pdf.

Chernov:2002:VRP

[1320] Alexey V. Chernov, Dmitriy P.

Skvortsov, Elena Z. Skvortsova, and Nikolai K. Vereshchagin. Variants of realizability for propositional formulas and the logic of the weak law of excluded middle. Lecture Notes in Computer Science, 2471:74– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710074.
pdf.

Bridges:2002:CCC

[1321] Douglas Bridges, Hajime Ishihara, and Peter Schuster. Compactness and continuity, constructively revisited. Lecture Notes in Computer Science, 2471: 89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710089.
                                        pdf.

Nipkow:2002:HLR

[1322] Tobias Nipkow. Hoare logics for recursive procedures and unbounded nondeterminism. Lecture Notes in Computer Science, 2471:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710103.
                                        pdf.

Chen:2002:FTN

[1323] Yifeng Chen. A fixpoint theory for non-monotonic parallelism. Lecture Notes in Computer Science, 2471:120– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710120.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710120.
                                        pdf.

Esik:2002:GNF

[1324] Zoltán Ésik and Hans Leiß. Greibach normal form in algebraically complete semirings. Lecture Notes in Computer Science, 2471:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710135.
pdf.

Mairson:2002:PCS

[1325] Harry G. Mairson and Xavier Rival. Proofnets and context semantics for the additives. Lecture Notes in Computer Science, 2471:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710151.
pdf.

Hodas:2002:TFS

[1326] Joshua S. Hodas, Pablo López, Jeffrey Polakow, Lubomira Stoilova, and Ernesto Pimentel. A tagframe system of resource management for proof search in linearlogic programming. Lecture Notes in Computer Science, 2471:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710167.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2471/24710167.
                                        pdf.

Galmiche:2002:RTE

[1327] Didier Galmiche, Daniel Méry, and David Pym. Resource tableaux (extended abstract). Lecture Notes in Computer Science, 2471:183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710183.
                                        pdf.

Cenciarelli:2002:CT

[1328] Pietro Cenciarelli. Configuration theories. Lecture Notes in Computer Science, 2471:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2471/24710200.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2471/24710200.pdf.

Moshier:2002:LPS

[1329] M. Andrew Moshier and Achim Jung. A logic for probabilities in semantics. Lecture Notes in Computer Science, 2471:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710216.
                                        pdf.

Levy:2002:PWS

[1330] Paul Blain Levy. Possible world semantics for general storage in call-by-value. Lecture Notes in Computer Science, 2471:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2471/24710232.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2471/24710232.
pdf.

McCusker:2002:FAR

[1331] Guy McCusker. A fully abstract relational model of syntactic control of interference. Lecture Notes in Computer Science, 2471:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710247.
pdf.

Marcinkowski:2002:OCB

[1332] Jerzy Marcinkowski and Tomasz Truderung. Optimal complexity bounds for positive LTL games. Lecture Notes in Computer Science, 2471: 262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710262.
pdf.

Kucera:2002:SPR

[1333] Antonín Kucera and Jan Strejcek. The stuttering principle revisited: On the expressiveness of nested X and U operators in the logic LTL. Lecture Notes in Computer Science, 2471:276– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710276.
pdf.

Jurdzinski:2002:TPF

[1334] Marcin Jurdzinski, Orna Kupferman, and Thomas A. Henzinger. Trading probability for fairness. Lecture Notes in Computer Science, 2471:292– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710292.
pdf.

Beauquier:2002:LPD

[1335] Danièle Beauquier, Alexander Rabinovich, and Anatol Slissenko. A logic of probability with decidable model-checking. Lecture Notes in Computer Science, 2471:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710306.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710306.
pdf.

Cachat:2002:SPG

[1336] Thierry Cachat, Jacques Duparc, and Wolfgang Thomas. Solving pushdown games with a Σ_{3} winning condition. Lecture Notes in Computer Science, 2471:322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710322.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710322.
pdf.

Kreutzer:2002:PFP

[1337] Stephan Kreutzer. Partial fixed-point logic on infinite structures. Lecture Notes in Computer Science, 2471:337– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710337.
pdf.

Berwanger:2002:VHM

[1338] Dietmar Berwanger, Erich Grädel, and Giacomo Lenzi. On the variable hierarchy of the modal μ-calculus. Lecture Notes in Computer Science, 2471:352– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710352.
pdf.

Leivant:2002:ICC

[1339] Daniel Leivant. Implicit computational complexity for higher type functionals (extended abstract). Lecture Notes in Computer Science, 2471:367– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710367.
                                        pdf.

Baaz:2002:GST

[1340] Matthias Baaz and Georg Moser. On generalizations of semi-terms of particularly simple form. Lecture Notes in Computer Science, 2471:382–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710382.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710382.
                                        pdf.

Barbanchon:2002:LPP

[1341] Régis Barbanchon and Etienne Grandjean. Local problems, planar local problems and linear time. Lecture Notes in Computer Science, 2471:397– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710397.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710397.
                                        pdf.

Bohler:2002:EIB

[1342] Elmar Böhler, Edith Hemaspaandra, Steffen Reith, and Heribert Vollmer. Equivalence and isomorphism for Boolean constraint satisfaction. Lecture Notes in Computer Science, 2471: 412–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710412.
pdf.

Faggian:2002:TD

[1343] Claudia Faggian. Travelling on designs.

Lecture Notes in Computer Science, 2471:427–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2471/24710427.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2471/24710427.
pdf.

Faggian:2002:DDS

[1344] Claudia Faggian and Martin Hyland.

Designs, disputes and strategies. Lecture Notes in Computer Science, 2471: 442–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710442.
pdf.

Hasegawa:2002:CLL

[1345] Masahito Hasegawa. Classical linear logic of implications. Lecture Notes in Computer Science, 2471:458– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710458.
pdf.

Goubault-Larrecq:2002:HOP

[1346] Jean Goubault-Larrecq. Higher-order positive set constraints. Lecture Notes in Computer Science, 2471:473–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710473.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710473.
pdf.

Ogata:2002:PTA

[1347] Ichiro Ogata. A proof theoretical account of continuation passing style. Lecture Notes in Computer Science, 2471:490–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2471/24710490.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2471/24710490.
pdf.

Kakutani:2002:DBC

[1348] Yoshihiko Kakutani. Duality between call-by-name recursion and callby-value iteration. Lecture Notes in Computer Science, 2471:506–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710506.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710506.
pdf.

Schmidt-Schauss:2002:DBH

[1349] Manfred Schmidt-Schauß and Klaus U.

Schulz. Decidability of bounded higher-order unification. Lecture Notes in Computer Science, 2471:522–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710522.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710522.
pdf.

Geuvers:2002:OPO

[1350] Herman Geuvers and Gueorgui I. Jojgov. Open proofs and open terms: A basis for interactive logic. Lecture Notes in Computer Science, 2471:537– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2471/24710537.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710537.
pdf.

Goubault-Larrecq:2002:LRM

[1351] Jean Goubault-Larrecq, Slawomir Lasota, and David Nowak. Logical relations for monadic types. Lecture Notes in Computer Science, 2471:553– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710553.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710553.
                                        pdf.

Atserias:2002:ARR

[1352] Albert Atserias and María Luisa Bonet. On the automatizability of resolution and related propositional proof systems. Lecture Notes in Computer Science, 2471:569–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710569.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710569.
                                        pdf.

deNivelle:2002:EPC

[1353] Hans de Nivelle. Extraction of proofs from the clausal normal form transformation. Lecture Notes in Computer Science, 2471:584–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2471/24710584.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2471/24710584.
                                        pdf.

Beckmann:2002:RRP

[1354] Arnold Beckmann. Resolution refutations and propositional proofs with height-restrictions. Lecture Notes in Computer Science, 2471:599–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2471/24710599.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2471/24710599.
pdf.

Anonymous:2002:AIba

[1355] Anonymous. Author index. Lecture Notes in Computer Science, 2471:613– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2471/2471ind.pdf.

Reps:2002:SPA

[1356] Thomas Reps. Static program analysis via 3-valued logic. Lecture Notes in Computer Science, 2477:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770001.
pdf.

Hall:2002:OMM

[1357] Robert J. Hall. Open modeling in multi-stakeholder distributed systems: Research and tool challenges. Lecture Notes in Computer Science, 2477:2– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770002.
pdf.

Esparza:2002:AAS

[1358] Javier Esparza. An algebraic approach to the static analysis of concurrent software. Lecture Notes in Computer Science, 2477:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770003.
pdf.

Muller-Olm:2002:PCD

[1359] Markus Müller-Olm and Helmut Seidl. Polynomial constants are decidable. Lecture Notes in Computer Science, 2477:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2477/24770004.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2477/24770004.
pdf.

Nielson:2002:NHC

[1360] Flemming Nielson, Hanne Riis Nielson, and Helmut Seidl. Normalizable horn clauses, strongly recognizable relations, and spi. Lecture Notes in Computer Science, 2477:20– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770020.
pdf.

Venet:2002:NAA

[1361] Arnaud Venet. Nonuniform alias analysis of recursive data structures and arrays. Lecture Notes in Computer Science, 2477:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770036.
pdf.

Jeannet:2002:RAT

[1362] B. Jeannet. Representing and approximating transfer functions in abstract interpretation of hetereogeneous datatypes. Lecture Notes in Computer Science, 2477:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770052.
pdf.

Yavuz-Kahveci:2002:AVC

[1363] Tuba Yavuz-Kahveci and Tevfik Bultan. Automated verification of concurrent linked lists with counters. Lecture Notes in Computer Science, 2477:69– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770069.
pdf.

Lu:2002:BTI

[1364] Lunjin Lu and Andy King. Backward type inference generalises type checking. Lecture Notes in Computer Science, 2477:85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2477/24770085.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2477/24770085.pdf.

Vaucheret:2002:MPY

[1365] Claudio Vaucheret and Francisco Bueno. More precise yet efficient type inference for logic programs. Lecture Notes in Computer Science, 2477:102– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770102.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770102.
                                        pdf.

Mine:2002:FGB

[1366] Antoine Miné. A few graph-based relational numerical abstract domains. Lecture Notes in Computer Science, 2477: 117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770117.
                                        pdf.

Martel:2002:SAN

[1367] Matthieu Martel. Static analysis of the numerical stability of loops. Lecture Notes in Computer Science, 2477:133– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770133.
pdf.

Serebrenik:2002:TLP

[1368] Alexander Serebrenik and Danny De Schreye. On termination of logic programs with floating point computations. Lecture Notes in Computer Science, 2477:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2477/24770151.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2477/24770151.pdf.

Probst:2002:MCF

[1369] Christian W. Probst. Modular control flow analysis for libraries. Lecture Notes in Computer Science, 2477:165– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770165.
pdf.

Whaley:2002:EIB

[1370] John Whaley and Monica S. Lam. An efficient inclusion-based points-to analysis for strictly-typed languages. Lecture Notes in Computer Science, 2477: 180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770180.
                                        pdf.

Manevich:2002:CRF

[1371] R. Manevich, G. Ramalingam, J. Field, D. Goyal, and M. Sagiv. Compactly representing first-order structures for static analysis. Lecture Notes in Computer Science, 2477:196– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770196.
                                        pdf.

Bagnara:2002:PCC

[1372] Roberto Bagnara, Elisa Ricci, Enea Zaffanella, and Patricia M. Hill. Possibly not closed convex polyhedra and the parma polyhedra library. Lecture Notes in Computer Science, 2477:213– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770213.
                                        pdf.

Adams:2002:SDA

[1373] Stephen Adams, Thomas Ball, Manuvir Das, Sorin Lerner, Sriram K. Rajamani, Mark Seigle, and Westley Weimer. Speeding up dataflow analysis using flow-insensitive pointer analysis. Lecture Notes in Computer Science, 2477:230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2477/24770230.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2477/24770230.
pdf.

Ruf:2002:IPE

[1374] Erik Ruf. Improving the precision of equality-based dataflow analyses. Lecture Notes in Computer Science, 2477: 247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770247.
pdf.

Veldhuizen:2002:GOP

[1375] Todd L. Veldhuizen and Andrew Lumsdaine. Guaranteed optimization: Proving nullspace properties of compilers. Lecture Notes in Computer Science, 2477:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770263.
pdf.

Faxen:2002:RAC

[1376] Karl-Filip Faxén. Representation analysis for coercion placement. Lecture Notes in Computer Science, 2477:278– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770278.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770278.
                                        pdf.

Langenbach:2002:PMT

[1377] Marc Langenbach, Stephan Thesing, and Reinhold Heckmann. Pipeline modeling for timing analysis. Lecture Notes in Computer Science, 2477:294– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770294.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770294.
                                        pdf.

Pierro:2002:AAC

[1378] Alessandra Di Pierro, Chris Hankin, and Herbert Wiklicky. Analysing approximate confinement under uniform attacks. Lecture Notes in Computer Science, 2477:310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2477/24770310.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2477/24770310.pdf.

Corin:2002:ICB

[1379] Ricardo Corin and Sandro Etalle. An improved constraint-based system for the verification of security protocols. Lecture Notes in Computer Science, 2477:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770326.
pdf.

Blanchet:2002:SAS

[1380] Bruno Blanchet. From secrecy to authenticity in security protocols. Lecture Notes in Computer Science, 2477: 342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770342.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770342.
pdf.

Zanotti:2002:STA

[1381] Mirko Zanotti. Security typings by abstract interpretation. Lecture Notes in Computer Science, 2477:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770360.
pdf.

Sabelfeld:2002:SCC

[1382] Andrei Sabelfeld and Heiko Mantel. Securing communication in a concurrent language. Lecture Notes in Computer Science, 2477:376–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770376.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770376.
                                        pdf.

Gallardo:2002:RLF

[1383] María del Mar Gallardo, Pedro Merino, and Ernesto Pimentel. Refinement of LTL formulas for abstract model checking. Lecture Notes in Computer Science, 2477:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770395.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770395.
                                        pdf.

Ranzato:2002:MAM

[1384] Francesco Ranzato and Francesco Tapparo. Making abstract model checking strongly preserving. Lecture Notes in Computer Science, 2477:411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770411.
                                        pdf.

Masse:2002:SAI

[1385] Damien Massé. Semantics for abstract interpretation-based static analyzes of temporal properties. Lecture Notes in Computer Science, 2477:428– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2477/24770428.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770428.
pdf.

Hymans:2002:CSP

[1386] Charles Hymans. Checking safety properties of behavioral VHDL descriptions by abstract interpretation. Lecture Notes in Computer Science, 2477:444– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2477/24770444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2477/24770444.
pdf.

Giacobazzi:2002:SVT

[1387] Roberto Giacobazzi and Francesco Ranzato. States vs. traces in model checking by abstract interpretation. Lecture Notes in Computer Science, 2477:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2477/24770461.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2477/24770461.
pdf.

Bruynooghe:2002:RRT

[1388] Maurice Bruynooghe, Michael Codish, Samir Genaim, and Wim Vanhoof. Reuse of results in termination analysis of typed logic programs. Lecture Notes in Computer Science, 2477:477– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770477.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770477.
                                        pdf.

Lee:2002:FAP

[1389] Chin Soon Lee. Finiteness analysis in polynomial time. Lecture Notes in Computer Science, 2477:493– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2477/24770493.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770493.
                                        pdf.

Mesnard:2002:DOT

[1390] Fred Mesnard, Etienne Payet, and Ulrich Neumerkel. Detecting optimal termination conditions of logic programs. Lecture Notes in Computer Science, 2477:509–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2477/24770509.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2477/24770509.
                                        pdf.

Anonymous:2002:AIbb

[1391] Anonymous. Author index. Lecture Notes in Computer Science, 2477:527– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2477/2477auth.pdf.

Ciortuz:2002:LCL

[1392] Liviu Ciortuz. LIGHT — A constraint language and compiler system for typed-unification grammars. Lecture Notes in Computer Science, 2479: 3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790003.
pdf.

Zens:2002:PBS

[1393] Richard Zens, Franz Josef Och, and Hermann Ney. Phrase-based statistical machine translation. Lecture Notes in Computer Science, 2479:18– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790018.
pdf.

Bonzon:2002:CDA

[1394] Pierre Bonzon. Compiling dynamic agent conversations. Lecture Notes in Computer Science, 2479:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790033.
pdf.

Schwind:2002:DPI

[1395] Michael Schwind and Oliver Wendt. Dynamic pricing of information products based on reinforcement learning: A yield-management approach. Lecture Notes in Computer Science, 2479: 51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790051.
pdf.

Guetova:2002:IFD

[1396] Marina Guetova, Steffen Hölldobler, and Hans-Peter Störr. Incremental fuzzy decision trees. Lecture Notes in Computer Science, 2479:67– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790067.
pdf.

Borth:2002:LMB

[1397] Michael Borth. Learning from multiple Bayesian networks for the revision and refinement of expert systems. Lecture Notes in Computer Science, 2479: 82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790082.
pdf.

Baader:2002:PCS

[1398] Franz Baader and Anni-Yasmin Turhan. On the problem of computing small representations of least common subsumers. Lecture Notes in Computer Science, 2479:99–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790099.
pdf.

Stuckenschmidt:2002:AIF

[1399] Heiner Stuckenschmidt. Approximate information filtering on the Semantic Web. Lecture Notes in Computer Science, 2479:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790114.
pdf.

Bui:2002:PAP

[1400] The Duy Bui, Dirk Heylen, Mannes Poel, and Anton Nijholt. ParleE: An adaptive plan based event appraisal model of emotions. Lecture Notes in Computer Science, 2479:129– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790129.
pdf.

Schmid:2002:IFA

[1401] Ute Schmid, Marina Müller, and Fritz Wysotzki. Integrating function application in state-based planning. Lecture Notes in Computer Science, 2479:144– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2479/24790144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2479/24790144.
                                        pdf.

Jain:2002:FWT

[1402] Brijnesh J. Jain and Fritz Wysotzki. Fast winner-takes-all networks for the maximum clique problem. Lecture Notes in Computer Science, 2479:163– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2479/24790163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2479/24790163.
                                        pdf.

Skabar:2002:ASN

[1403] Andrew Skabar. Augmenting supervised neural classifier training using a corpus of unlabeled data. Lecture Notes in Computer Science, 2479:174– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2479/24790174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2479/24790174.
                                        pdf.

Geibel:2002:LCD

[1404] Peter Geibel, Kristina Schädler, and Fritz Wysotzki. Learning of class descriptions from class discriminations: A hybrid approach for relational objects. Lecture Notes in Computer Science, 2479:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790186.
pdf.

Hitzler:2002:WFS

[1405] Pascal Hitzler and Matthias Wendt.

The well-founded semantics is a stratified fitting semantics. Lecture Notes in Computer Science, 2479:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790205.
pdf.

Burghardt:2002:AFA

[1406] Jochen Burghardt. Axiomatization of finite algebras. Lecture Notes in Computer Science, 2479:222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790222.
pdf.

Gago:2002:AGC

[1407] M. Carmen Fernández Gago, Michael Fisher, and Clare Dixon. Algorithms for guiding clausal temporal resolution. Lecture Notes in Computer Science, 2479:235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790235.
pdf.

Wickel:2002:AMV

[1408] Jochen Wickel, Pablo Alvarado, Peter Dörfler, Thomas Krüger, and Karl-Friedrich Kraiss. Axiom — A modular visual object retrieval system. Lecture Notes in Computer Science, 2479:253– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790253.
pdf.

Arens:2002:RBK

[1409] Michael Arens and Hans-Hellmut Nagel. Representation of behavioral knowledge for planning and plan-recognition in a cognitive vision system. Lecture Notes in Computer Science, 2479:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790268.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2479/24790268.
pdf.

Stolzenburg:2002:QVB

[1410] Frieder Stolzenburg, Oliver Obst, and Jan Murray. Qualitative velocity and ball interception. Lecture Notes in Computer Science, 2479:283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790283.
pdf.

Gips:2002:SIL

[1411] Carsten Gips, Petra Hofstedt, and Fritz Wysotzki. Spatial inference — learning vs. constraint solving. Lecture Notes in Computer Science, 2479:299– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790299.
pdf.

Andre:2002:SDI

[1412] Elisabeth André. From simulated dialogues to interactive performances with virtual actors. Lecture Notes in Computer Science, 2479:317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790317.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2479/24790317.
pdf.

Wooldridge:2002:TKC

[1413] Michael Wooldridge. Time, knowledge, and cooperation: Alternatingtime temporal epistemic logic and its applications. Lecture Notes in Computer Science, 2479:318–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790318.
pdf.

Fensel:2002:SWE

[1414] Dieter Fensel. Semantic Web enabled Web services. Lecture Notes in Computer Science, 2479:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790319.
pdf.

Christaller:2002:DPP

[1415] Thomas Christaller. DFG priority program RoboCup (SPP-1125) cooperating teams of mobile robots in dynamic environments. Lecture Notes in Computer Science, 2479:323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790323.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2479/24790323.
pdf.

Kirn:2002:DPP

[1416] Stefan Kirn. DFG priority program (SPP-1083) intelligent agents and realistic commercial application scenarios. Lecture Notes in Computer Science, 2479:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790324.
pdf.

Malsch:2002:DPP

[1417] Thomas Malsch. DFG priority program (SPP-1077) socionics — investigating and modelling artificial societies. Lecture Notes in Computer Science, 2479: 325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2479/24790325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2479/24790325.
pdf.

Anonymous:2002:AIbc

[1418] Anonymous. Author index. Lecture Notes in Computer Science, 2479:327– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2479/2479auth.pdf.

Eder:2002:DWW

[1419] Johann Eder, Georg E. Olivotto, and Wolfgang Gruber. A data warehouse for workflow logs. Lecture Notes in Computer Science, 2480:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800001.
pdf.

Lai:2002:WKM

[1420] Jin Lai and Yushun Fan. Workflow and knowledge management: Approaching an integration. Lecture Notes in Computer Science, 2480:16– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800016.
pdf.

Qu:2002:LTI

[1421] Yang Qu, Chuang Lin, and Jiye Wang. Linear temporal inference of workflow management systems based on timed Petri nets models. Lecture Notes in Computer Science, 2480:30– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800030.
pdf.

vanderAalst:2002:DWP

[1422] W. M. P. van der Aalst and B. F. van Dongen. Discovering workflow performance models from timed logs. Lecture Notes in Computer Science, 2480:45– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800045.
pdf.

Lin:2002:PEA

[1423] Chuang Lin, Yang Qu, Fengyuan Ren, and Dan C. Marinescu. Performance equivalent analysis of workflow systems based on stochastic Petri net models. Lecture Notes in Computer Science, 2480:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800064.
pdf.

Liu:2002:AEF

[1424] Jianxun Liu, Shensheng Zhang, Jian Cao, and Jinmin Hu. An agent enhanced framework to support predispatching of tasks in workflow management systems. Lecture Notes in Computer Science, 2480:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800080.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2480/24800080.
                                        pdf.

Dong:2002:TNM

[1425] Ying Dong and Mingshu Li. TEMP-PLET: A new method for domainspecific ontology design. Lecture Notes in Computer Science, 2480:90– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800090.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800090.
                                        pdf.

Si:2002:EMD

[1426] Jinxin Si, Cungen Cao, Haitao Wang, Fang Gu, Qiangze Feng, Chunxia Zhang, Qingtian Zeng, Wen Tian, and Yufei Zheng. An environment for multi-domain ontology development and knowledge acquisition. Lecture Notes in Computer Science, 2480:104– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800104.
                                        pdf.

Yang:2002:AIR

[1427] Zhifeng Yang, Yue Liu, and Sujian Li. Applying information retrieval technology to incremental knowledge management. Lecture Notes in Computer Science, 2480:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2480/24800117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800117.
pdf.

Kim:2002:VDK

[1428] Hong-Gee Kim, Christian Fillies, Bob Smith, and Dietmar Wikarski. Visualizing a dynamic knowledge map using Semantic Web technology. Lecture Notes in Computer Science, 2480:130– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800130.
pdf.

Lee:2002:IRX

[1429] Sung Wan Kim Jaeho Lee and Hae Chull Lim. Indexing and retrieval of XML-encoded structured documents in dynamic environment. Lecture Notes in Computer Science, 2480:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800141.
pdf.

Ma:2002:KMS

[1430] Jun Ma and Matthias Hemmje. Knowledge management: System architectures, main functions, and implementing techniques. Lecture Notes in Computer Science, 2480:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800155.
                                        pdf.

Wang:2002:DMB

[1431] Fengjin Wang, Zhuofeng Zhao, and Yanbo Han. A dynamic matching and binding mechanism for business service integration. Lecture Notes in Computer Science, 2480:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800168.
                                        pdf.

Li:2002:UMA

[1432] Dongdong Li, Songlin Hu, and Shuo Bai. A uniform model for authorization and access control in enterprise information platform. Lecture Notes in Computer Science, 2480:180– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800180.
                                        pdf.

Hongwei:2002:CPM

[1433] Sun Hongwei, Zhang Shusheng, Zhou Jingtao, and Wang Jing. Constraintspreserving mapping algorithm from XML-schema to relational schema. Lecture Notes in Computer Science, 2480: 193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800193.
pdf.

Wang:2002:SSB

[1434] Dan Wang, Ge Yu, Baoyan Song, Derong Shen, and Guoren Wang. Study on SOAP-based mobile agent techniques. Lecture Notes in Computer Science, 2480:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2480/24800208.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2480/24800208.pdf.

Maxim:2002:SAB

[1435] Michael Maxim and Ashish Venugopal.

Securing agent based architectures. Lecture Notes in Computer Science, 2480:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2480/24800220.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2480/24800220.
pdf.

Yang:2002:SNM

[1436] Kun Yang, Alex Galis, Telma Mota, Xin Guo, and Chris Todd. Service and network management middleware for cooperative information systems through policies and mobile agents. Lecture Notes in Computer Science, 2480:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800232.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800232.
                                        pdf.

Hongxia:2002:REM

[1437] Xu Hongxia, Zhang Li, and Zhou Bosheng. Research on enterprise modeling of agile manufacturing. Lecture Notes in Computer Science, 2480:247– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800247.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800247.
                                        pdf.

Zhang:2002:HMD

[1438] Yan Zhang and Meilin Shi. HCM — A model describing cooperation of virtual enterprise. Lecture Notes in Computer Science, 2480:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800257.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800257.
                                        pdf.

Wiedeler:2002:DSS

[1439] Markus Wiedeler. A description for services supporting cooperations. Lecture Notes in Computer Science, 2480:267– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800267.
pdf.

Park:2002:AEP

[1440] Sung-Hoon Park. Analysis of an election problem for CSCW in asynchronous distributed systems. Lecture Notes in Computer Science, 2480:280– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800280.
pdf.

Woo:2002:LOB

[1441] Sung-Ho Woo and Sung-Bong Yang. A linear-order based access method for efficient network computations. Lecture Notes in Computer Science, 2480:289– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800289.
pdf.

Shen:2002:PNM

[1442] Jun Shen, Yun Yang, and Junzhou Luo. A Petri net model for session services. Lecture Notes in Computer Science, 2480:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800303.
                                        pdf.

Gruhn:2002:SPE

[1443] Volker Gruhn and Lothar Schöpe. Software processes for electronic commerce portal systems. Lecture Notes in Computer Science, 2480:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800315.
                                        pdf.

Han:2002:ASS

[1444] Jun Han and Pin Chen. Architecture support for system-of-systems evolution. Lecture Notes in Computer Science, 2480:332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800332.
                                        pdf.

Chen:2002:EGU

[1445] Jessica Chen. Expressing graphical user’s input for test specifications. Lecture Notes in Computer Science, 2480: 347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2480/24800347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800347.
pdf.

Cao:2002:IDS

[1446] Ying Cao, K. W. Chau, M. Anson, and Jianping Zhang. An intelligent decision support system in construction management by data warehousing technique. Lecture Notes in Computer Science, 2480:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800360.
pdf.

Lican:2002:DHI

[1447] Huang Lican and Wu Zhaohui. Distributed heterogeneous inspecting system and its implementation. Lecture Notes in Computer Science, 2480:370– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800370.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800370.
pdf.

Lin:2002:ADE

[1448] Yi Lin, XingShe Zhou, and XiaoAn Shi. Architecture for distributed embedded systems based on workflow and distributed resource management. Lecture Notes in Computer Science, 2480: 381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800381.
                                        pdf.

Cook:2002:KMC

[1449] David E. Cook, Lin Mellor, Guy Frost, and Reiner Creutzburg. Knowledge management and the control of duplication. Lecture Notes in Computer Science, 2480:396–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800396.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800396.
                                        pdf.

Jin:2002:ETM

[1450] Ying Jin, Susan D. Urban, Amy Sundermier, and Suzanne W. Dietrich. An execution and transaction model for active, rule-based component integration middleware. Lecture Notes in Computer Science, 2480:403– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800403.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800403.
                                        pdf.

Jiang:2002:IC

[1451] Jinlei Jiang and Meilin Shi. Implementation of CovaTM. Lecture Notes in Computer Science, 2480:418– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800418.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800418.
pdf.

Andreoli:2002:RBS

[1452] Jean-Marc Andreoli, Damián Arregui, François Pacull, and Jutta Willamowski. Resource-based scripting to stitch distributed components. Lecture Notes in Computer Science, 2480:429– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800429.
pdf.

Xu:2002:MAC

[1453] Ming Xu and Yi Zhuang. Multiagent coordination mechanism in distributed environment. Lecture Notes in Computer Science, 2480:444–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800444.
pdf.

Wei:2002:CAM

[1454] Yinxing Wei, Shensheng Zhang, and Jian Cao. Coordination among multi-agents using process calculus and ECA rule. Lecture Notes in Computer Science, 2480:456–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800456.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800456.
                                        pdf.

Zhang:2002:RBP

[1455] Yaying Zhang and Jinyuan You. An RBAC based policy enforcement coordination model in Internet environment. Lecture Notes in Computer Science, 2480:466–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800466.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800466.
                                        pdf.

Zhao:2002:CBN

[1456] Yongyi Zhao and Guangxing Wang. A CORBA-based negotiation strategy in E-commerce. Lecture Notes in Computer Science, 2480:478–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2480/24800478.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2480/24800478.
                                        pdf.

Sihvonen:2002:NFN

[1457] Markus Sihvonen and Jarkko Holappa. Negotiation framework for the next generation mobile middleware service environment. Lecture Notes in Computer Science, 2480:487– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800487.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800487.
pdf.

Dai:2002:IBC

[1458] Kaiyu Dai, Yinglin Wang, and Xiqing Xu. An Internet-based conference system for real-time distributed design evaluation. Lecture Notes in Computer Science, 2480:499–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800499.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800499.
pdf.

Xue:2002:IPM

[1459] Liyin Xue, Mehmet Orgun, and Kang Zhang. Intention preservation by multi-versioning in distributed realtime group editors. Lecture Notes in Computer Science, 2480:510–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800510.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800510.
pdf.

Hu:2002:SGA

[1460] Bin Hu, Andreas Kuhlenkamp, and Rolf Reinema. Supporting group awareness in Web based learning environments. Lecture Notes in Computer Science, 2480:525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800525.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800525.
pdf.

Matsubara:2002:RDO

[1461] Takashi Matsubara, Kozo Sugiyama, and Kazushi Nishimoto. Raison D’etre object: A cyber-hearth that catalyzes face-to-face informal communication. Lecture Notes in Computer Science, 2480:537–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800537.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800537.
pdf.

Barthelmess:2002:NPE

[1462] P. Barthelmess and C. A. Ellis. The neem platform: An extensible framework for the development of perceptual collaborative applications. Lecture Notes in Computer Science, 2480:547– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2480/24800547.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2480/24800547.
pdf.

Anonymous:2002:AIbd

[1463] Anonymous. Author index. Lecture Notes in Computer Science, 2480:563– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2480/2480auth.pdf.

Bar-Yossef:2002:CDE

[1464] Ziv Bar-Yossef, T. S. Jayram, Ravi Kumar, D. Sivakumar, and Luca Trevisan. Counting distinct elements in a data stream. Lecture Notes in Computer Science, 2483:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2483/24830001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2483/24830001.
pdf.

Parnas:2002:TCS

[1465] Michal Parnas, Dana Ron, and Ronitt Rubinfeld. On testing convexity and submodularity. Lecture Notes in Computer Science, 2483:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2483/24830011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2483/24830011.
pdf.

Chockler:2002:RLT

[1466] Hana Chockler and Orna Kupferman.

ω-regular languages are testable with a constant number of queries. Lecture Notes in Computer Science, 2483:26– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2483/24830026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2483/24830026.
pdf.

Obata:2002:OLB

[1467] Kenji Obata. Optimal lower bounds for 2-query locally decodable linear codes. Lecture Notes in Computer Science, 2483:39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2483/24830039.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2483/24830039.
pdf.

Dyer:2002:CSC

[1468] Martin Dyer, Leslie A. Goldberg, and Mark Jerrum. Counting and sampling H-colourings. Lecture Notes in Computer Science, 2483:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2483/24830051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2483/24830051.
pdf.

Dyer:2002:RMM

[1469] Martin Dyer, Mark Jerrum, and Eric Vigoda. Rapidly mixing Markov chains for dismantleable constraint graphs. Lecture Notes in Computer Science, 2483:68–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2483/24830068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2483/24830068.
pdf.

Achlioptas:2002:CRH

[1470] Dimitris Achlioptas and Cristopher Moore. On the 2-colorability of random hypergraphs. Lecture Notes in Computer Science, 2483:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2483/24830078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2483/24830078.
pdf.

Malon:2002:PFC

[1471] Christopher Malon and Igor Pak. Percolation on finite Cayley graphs. Lecture Notes in Computer Science, 2483: 91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2483/24830091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2483/24830091.
pdf.

Friedgut:2002:CGP

[1472] Ehud Friedgut, Jeff Kahn, and Avi Wigderson. Computing graph properties by randomized subcube partitions. Lecture Notes in Computer Science, 2483:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2483/24830105.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2483/24830105.
                                        pdf.

Diaz:2002:BRC

[1473] J. Díaz, N. Do, M. J. Serna, and N. C. Wormald. Bisection of random cubic graphs. Lecture Notes in Computer Science, 2483:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2483/24830114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2483/24830114.
                                        pdf.

Duckworth:2002:SDS

[1474] William Duckworth and Bernard Mans. Small k-dominating sets of regular graphs. Lecture Notes in Computer Science, 2483:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2483/24830126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2483/24830126.
                                        pdf.

Coja-Oghlan:2002:FSI

[1475] Amin Coja-Oghlan. Finding sparse induced subgraphs of semirandom graphs. Lecture Notes in Computer Science, 2483:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/

series/0558/bibs/2483/24830139.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2483/24830139.pdf.

Dyer:2002:MTS

[1476] Martin Dyer, Alistair Sinclair, Eric Vigoda, and Dror Weitz. Mixing in time and space for lattice spin systems: A combinatorial view. Lecture Notes in Computer Science, 2483:149– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2483/24830149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2483/24830149.
pdf.

Moore:2002:QWH

[1477] Cristopher Moore and Alexander Russell. Quantum walks on the hypercube. Lecture Notes in Computer Science, 2483:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2483/24830164.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2483/24830164.
pdf.

Santis:2002:ROC

[1478] Alfredo De Santis, Giovanni Di Crescenzo, and Giuseppe Persiano. Randomness-optimal characterization of two NP proof systems. Lecture Notes in Computer Science, 2483:179– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2483/24830179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2483/24830179.
                                        pdf.

Barak:2002:PTH

[1479] Boaz Barak. A probabilistic-time hierarchy theorem for “Slightly Nonuniform” algorithms. Lecture Notes in Computer Science, 2483:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2483/24830194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2483/24830194.
                                        pdf.

Goldreich:2002:DRW

[1480] Oded Goldreich and Avi Wigderson. Derandomization that is rarely wrong from short advice that is typically good. Lecture Notes in Computer Science, 2483:209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2483/24830209.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2483/24830209.
                                        pdf.

Engebretsen:2002:CST

[1481] Lars Engebretsen and Venkatesan Guruswami. Is constraint satisfaction over two variables always easy? Lecture Notes in Computer Science, 2483:224– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2483/24830224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2483/24830224.
pdf.

Magen:2002:DRP

[1482] Avner Magen. Dimensionality reductions that preserve volumes and distance to affine spaces, and their algorithmic applications. Lecture Notes in Computer Science, 2483:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2483/24830239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2483/24830239.
pdf.

Mihail:2002:EPL

[1483] Milena Mihail and Christos Papadimitriou. On the eigenvalue power law. Lecture Notes in Computer Science, 2483:254–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2483/24830254.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2483/24830254.
pdf.

Cooper:2002:CSI

[1484] Colin Cooper. Classifying special interest groups in Web graphs. Lecture Notes in Computer Science, 2483:263– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2483/24830263.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2483/24830263.
                                        pdf.

Anonymous:2002:AIbe

[1485] Anonymous. Author index. Lecture Notes in Computer Science, 2483:277– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2483/2483auth.pdf.

Antunes:2002:ISA

[1486] Cláudia M. Antunes and Arlindo L. Oliveira. Inference of sequential association rules guided by contextfree grammars. Lecture Notes in Computer Science, 2484:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840001.
                                        pdf.

Belz:2002:PLN

[1487] Anja Belz. PCFG learning by nonterminal partition search. Lecture Notes in Computer Science, 2484:14– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840014.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840014.
                                        pdf.

Cano:2002:ISR

[1488] Antonio Cano, José Ruiz, and Pedro García. Inferring subclasses of regular languages faster using RPNI and forbidden configurations. Lecture Notes in Computer Science, 2484:28– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840028.
pdf.

Cicchello:2002:BE

[1489] Orlando Cicchello and Stefan C. Kremer. Beyond EDSM. Lecture Notes in Computer Science, 2484:37– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840037.
pdf.

Florencio:2002:CIL

[1490] Christophe Costa Florêncio. Consistent identification in the limit of rigid grammars from strings is NP-hard. Lecture Notes in Computer Science, 2484:49– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840049.
pdf.

Denis:2002:SCR

[1491] François Denis, Aurélien Lemay, and Alain Terlutte. Some classes of regular languages identifiable in the limit from positive data. Lecture Notes in Computer Science, 2484:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840063.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840063.
                                        pdf.

Esposito:2002:LPR

[1492] Yann Esposito, Aurélien Lemay, François Denis, and Pierre Dupont. Learning probabilistic residual finite state automata. Lecture Notes in Computer Science, 2484:77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840077.
                                        pdf.

Fernau:2002:FEI

[1493] Henning Fernau. Fragmentation: Enhancing identifiability. Lecture Notes in Computer Science, 2484:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840092.
                                        pdf.

Foret:2002:LPS

[1494] Annie Foret and Yannick Le Nir. On limit points for some variants of rigid lambek grammars. Lecture Notes in Computer Science, 2484:106– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840106.
pdf.

Habrard:2002:GST

[1495] Amaury Habrard, Marc Bernard, and François Jacquenet. Generalized stochastic tree automata for multirelational data mining. Lecture Notes in Computer Science, 2484:120–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840120.
pdf.

delaHiguera:2002:SCI

[1496] Colin de la Higuera and Jose Oncina.

On sufficient conditions to identify in the limit classes of grammars from polynomial time and data. Lecture Notes in Computer Science, 2484:134– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840134.
pdf.

Kermorvant:2002:SGI

[1497] Christopher Kermorvant and Pierre Dupont. Stochastic grammatical inference with multinomial tests. Lecture Notes in Computer Science, 2484: 149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840149.
                                        pdf.

Kermorvant:2002:LLH

[1498] Christopher Kermorvant and Colin de la Higuera. Learning languages with help. Lecture Notes in Computer Science, 2484:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840161.
                                        pdf.

Nakamura:2002:ILC

[1499] Katsuhiko Nakamura and Masashi Matsumoto. Incremental learning of context free grammars. Lecture Notes in Computer Science, 2484:174– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840174.
                                        pdf.

Oates:2002:EGP

[1500] Tim Oates and Brent Heeringa. Estimating grammar parameters using bounded memory. Lecture Notes in Computer Science, 2484:185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840185.
pdf.

Rico-Juan:2002:STT

[1501] Juan Ramón Rico-Juan, Jorge Calera-Rubio, and Rafael C. Carrasco. Stochastic k-testable tree languages and applications. Lecture Notes in Computer Science, 2484:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840199.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840199.
pdf.

Seginer:2002:FLS

[1502] Yoav Seginer. Fast learning from strings of 2-letter rigid grammars. Lecture Notes in Computer Science, 2484: 213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840213.
pdf.

Sempere:2002:LLT

[1503] José M. Sempere and Pedro García.

Learning locally testable even linear languages from positive data. Lecture Notes in Computer Science, 2484:225– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840225.
                                        pdf.

Starkie:2002:IAG

[1504] Bradford Starkie. Inferring attribute grammars with structured data for natural language processing. Lecture Notes in Computer Science, 2484:237– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840237.
                                        pdf.

Tajima:2002:PLS

[1505] Yasuhiro Tajima and Matsuaki Terada. A PAC learnability of simple deterministic languages. Lecture Notes in Computer Science, 2484:249– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840249.
                                        pdf.

Terwijn:2002:LHM

[1506] Sebastiaan A. Terwijn. On the learnability of hidden Markov models. Lecture Notes in Computer Science, 2484: 261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840261.
pdf.

Thollard:2002:SPU

[1507] Franck Thollard and Alexander Clark.

Shallow parsing using probabilistic grammatical inference. Lecture Notes in Computer Science, 2484:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840269.
pdf.

Thomas:2002:LRB

[1508] D. G. Thomas, M. Humrosia Begam, K. G. Subramanian, and S. Gnanasekaran. Learning of regular bi-ω languages. Lecture Notes in Computer Science, 2484:283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2484/24840283.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2484/24840283.
pdf.

Adriaans:2002:EGI

[1509] Pieter Adriaans and Marco Vervoort. The EMILE 4.1 grammar induction toolbox. Lecture Notes in Computer Science, 2484:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840293.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840293.
                                        pdf.

Chalup:2002:SAR

[1510] Stephan K. Chalup and Alan D. Blair. Software for analysing recurrent neural nets that learn to predict non-regular languages. Lecture Notes in Computer Science, 2484:296– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840296.
                                        pdf.

Ciortuz:2002:FIL

[1511] Liviu Ciortuz. A framework for inductive learning of typed-unification grammars. Lecture Notes in Computer Science, 2484:299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2484/24840299.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2484/24840299.pdf.

Sofronie:2002:TLL

[1512] Daniela Dudau Sofronie, Isabelle Tellier, and Marc Tommasi. A tool for language learning based on categorial grammars and semantic information. Lecture Notes in Computer Science, 2484:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840303.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840303.
pdf.

Lievesley:2002:NAI

[1513] Sam Lievesley and Eric Atwell. ‘NAIL’:

Artificial intelligence software for learning natural language. Lecture Notes in Computer Science, 2484:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840306.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840306.
pdf.

Starkie:2002:LDS

[1514] Bradford Starkie, Greg Findlow, Khanh Ho, Alvaro Hui, Lawrence Law, Liron Lightwood, Simon Michnowicz, and Christian Walder. Lyrebird^{TM}: Developing spoken dialog systems using examples. Lecture Notes in Computer Science, 2484:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2484/24840309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2484/24840309.
pdf.

vanZaanen:2002:IAB

[1515] Menno van Zaanen. Implementing alignment-based learning. Lecture Notes in Computer Science, 2484:312– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2484/24840312.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2484/24840312.
                                        pdf.

Anonymous:2002:AIbf

[1516] Anonymous. Author index. Lecture Notes in Computer Science, 2484:315– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2484/2484indx.pdf.

Burrell:2002:CAC

[1517] Jenna Burrell, Geri K. Gay, Kiyo Kubo, and Nick Farina. Contextaware computing: A test case. Lecture Notes in Computer Science, 2498: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2498/24980001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2498/24980001.
                                        pdf.

Sumi:2002:CRI

[1518] Yasuyuki Sumi, Ryuuki Sakamoto, Keiko Nakao, and Kenji Mase. Comic-Diary: Representing individual experiences in a comics style. Lecture Notes in Computer Science, 2498:16– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2498/24980016.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2498/24980016.
pdf.

Goose:2002:MRP

[1519] Stuart Goose, Heiko Wanning, and Georg Schneider. Mobile reality: A PDA-based multimodal framework synchronizing a hybrid tracking solution with 3D graphics and locationsensitive speech interaction. Lecture Notes in Computer Science, 2498:33– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980033.
pdf.

Fleck:2002:RTC

[1520] Margaret Fleck, Marcos Frid, Tim Kindberg, Eamonn O’Brien-Strain, Rakhi Rajani, and Mirjana Spasojevic. Rememberer: A tool for capturing museum visits. Lecture Notes in Computer Science, 2498:48– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980048.
pdf.

Trevor:2002:IPS

[1521] Jonathan Trevor, David M. Hilbert, and Bill N. Schilit. Issues in personalizing shared ubiquitous devices. Lecture Notes in Computer Science, 2498: 56–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2498/24980056.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2498/24980056.
                                        pdf.

Consolvo:2002:UST

[1522] Sunny Consolvo, Larry Arnstein, and B. Robert Franza. User study techniques in the design and evaluation of a ubicomp environment. Lecture Notes in Computer Science, 2498:73– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2498/24980073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2498/24980073.
                                        pdf.

Intille:2002:CBI

[1523] Stephen S. Intille. Change blind information display for ubiquitous computing environments. Lecture Notes in Computer Science, 2498:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2498/24980091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2498/24980091.
                                        pdf.

Christensen:2002:SHA

[1524] Henrik Bærbak Christensen and Jakob E. Bardram. Supporting human activities – exploring activitycentered computing. Lecture Notes in Computer Science, 2498:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980107.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980107.
pdf.

Crowley:2002:PCC

[1525] James L. Crowley, Joëlle Coutaz, Gaeten Rey, and Patrick Reignier. Perceptual components for context aware computing. Lecture Notes in Computer Science, 2498:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980117.
pdf.

Darrell:2002:FRI

[1526] Trevor Darrell, Konrad Tollmar, Frank Bentley, Neal Checka, Loius-Phillipe Morency, Ali Rahimi, and Alice Oh. Face-responsive interfaces: From direct manipulation to perceptive presence. Lecture Notes in Computer Science, 2498:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980135.
pdf.

Nakanishi:2002:VBF

[1527] Yasuto Nakanishi, Takashi Fujii, Kotaro Kiatjima, Yoichi Sato, and Hideki Koike. Vision-based face tracking system for large displays. Lecture Notes in Computer Science, 2498:152– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980152.
pdf.

Ma:2002:FFR

[1528] Hongshen Ma and Joseph A. Paradiso. The FindIT flashlight: Responsive tagging based on optically triggered microprocessor wakeup. Lecture Notes in Computer Science, 2498:160– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980160.
pdf.

McElligott:2002:FFF

[1529] Lisa McElligott, Michelle Dillon, Krispin Leydon, Bruce Richardson, Mikael Fernström, and Joseph A. Paradiso. ‘ForSe FIElds’ — force sensors for interactive environments. Lecture Notes in Computer Science, 2498:168– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980168.
pdf.

Jiang:2002:AIF

[1530] Xiaodong Jiang, Jason I. Hong, and James A. Landay. Approximate information flows: Sociallybased modeling of privacy in ubiquitous computing. Lecture Notes in Computer Science, 2498:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980176.
pdf.

Want:2002:PSC

[1531] Roy Want, Trevor Pering, Gunner Danneels, Muthu Kumar, Murali Sundar, and John Light. The personal server: Changing the way we think about ubiquitous computing. Lecture Notes in Computer Science, 2498:194– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980194.
pdf.

Konomi:2002:QBI

[1532] Shin’ichi Konomi. QueryLens: Beyond ID-based information access. Lecture Notes in Computer Science, 2498:210– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980210.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2498/24980210.
                                        pdf.

vanLaerhoven:2002:PNO

[1533] Kristof van Laerhoven, Albrecht Schmidt, and Hans-Werner Gellersen. Pin&play: Networking objects through pins. Lecture Notes in Computer Science, 2498:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2498/24980219.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2498/24980219.pdf.

Russell:2002:SAU

[1534] Daniel M. Russell, Clemens Drews, and Alison Sue. Social aspects of using large public interactive displays for collaboration. Lecture Notes in Computer Science, 2498:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2498/24980229.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2498/24980229.
                                        pdf.

Langheinrich:2002:PAS

[1535] Marc Langheinrich. A privacy awareness system for ubiquitous computing environments. Lecture Notes in Computer Science, 2498:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2498/24980237.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2498/24980237.
pdf.

Jiang:2002:HLM

[1536] Changhao Jiang and Peter Steenkiste.

A hybrid location model with a computable location identifier for ubiquitous computing. Lecture Notes in Computer Science, 2498:246–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980246.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980246.
pdf.

Hazas:2002:NBU

[1537] Mike Hazas and Andy Ward. A novel broadband ultrasonic location system. Lecture Notes in Computer Science, 2498:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2498/24980264.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2498/24980264.
pdf.

Hoffmann:2002:LMD

[1538] Frank Hoffmann and James Scott.

Location of mobile devices using networked surfaces. Lecture Notes in Computer Science, 2498:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980281.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2498/24980281.
pdf.

Krumm:2002:SGI

[1539] John Krumm, Lyndsay Williams, and Greg Smith. SmartMoveX on a graph — an inexpensive active badge tracker. Lecture Notes in Computer Science, 2498:299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980299.
pdf.

Sanmugalingam:2002:GLE

[1540] Kumaresan Sanmugalingam and George Coulouris. A generic location event simulator. Lecture Notes in Computer Science, 2498:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980308.
pdf.

LaMarca:2002:PIP

[1541] Anthony LaMarca, Waylon Brunette, David Koizumi, Matthew Lease, Stefan B. Sigurdsson, Kevin Sikorski, Dieter Fox, and Gaetano Borriello. PlantCare: An investigation in practical ubiquitous systems. Lecture Notes in Computer Science, 2498:316– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2498/24980316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980316.
pdf.

Schmidt:2002:CAB

[1542] Albrecht Schmidt, Martin Strohbach, Kristof van Laerhoven, Adrian Friday, and Hans-Werner Gellersen. Context acquisition based on load sensing. Lecture Notes in Computer Science, 2498: 333–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980333.
pdf.

Antifakos:2002:PIF

[1543] Stavros Antifakos, Florian Michahelles, and Bernt Schiele. Proactive instructions for furniture assembly. Lecture Notes in Computer Science, 2498:351– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2498/24980351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2498/24980351.
pdf.

Lukowicz:2002:WDM

[1544] P. Lukowicz, H. Junker, M. Stäger, T. von Büren, and G. Tröster. Wear-NET: A distributed multi-sensor system for context aware wearables. Lecture Notes in Computer Science, 2498: 361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2498/24980361.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2498/24980361.
                                        pdf.

Mankoff:2002:ULC

[1545] Jennifer Mankoff, Gary Hsieh, Ho Chak Hung, Sharon Lee, and Elizabeth Nitao. Using low-cost sensing to support nutritional awareness. Lecture Notes in Computer Science, 2498:371– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2498/24980371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2498/24980371.
                                        pdf.

Anonymous:2002:AIbg

[1546] Anonymous. Author index. Lecture Notes in Computer Science, 2498:377– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2498/2498auth.pdf.